A Complete Bibliography of Publications in ACM SIGSOFT Software Engineering Notes: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 August 2018 Version 1.00

Title word cross-reference

# [dCJL06]. #15 [KS01].

$29.99 [Tra08]. $39.99 [Sau08]. < username > [KAKM05]. π [MO06, Oqu04a, Oqu04b, Oqu04c, Oqu06b, Oqu06a].

-AAL [MO06]. -ADL [Oqu04a, Oqu06b]. -ARL [Oqu04b, Oqu04c]. -calculus [Oqu04a]. -Method [Oqu06a].

0 [Bea05, Ber06a, Ber06b, Ber06c, Bre06, Che05a, Che05e, Che05c, Che05d, Els06a, Els06b, FV09a, Fen06, Fen07, Fra05, Law06, Law05, Mar05b, Pet06, Sau05b, Sau05c, Sau06a, Sau06b, Sau06c, Sch06c, Sch06d, Sch06e, Sch06f, WW05a, WW05b, WW06]. 0-13-02175-0 [Man00b]. 0-13-187250-8 [Sch06d]. 0-321-18612-5 [Bea05]. 0-321-18613-3 [Law05]. 0-321-19950-2 [Sau05c]. 0-321-20575-8 [WW05a]. 0-321-22391-8 [Sau05b]. 0-321-22847-2 [Che05b, Bre06]. 0-321-26838-5 [WW05b]. 0-321-26934-9 [Els06a]. 0-321-33205-9 [Fen06]. 0-321-33570-8 [WW06]. 0-321-33572-4 [Sau08]. 0-321-33662-0 [Sch06c]. 0-321-47789-8 [Tra08]. 0-470-12795-3 [Hen09a]. 0-471-20282-7 [Che05a, Els06b, FV09a]. 0-471-49892-0 [Law06]. 0-471-71051-2 [Sau06a]. 0-471-73849-2 [Ber06c]. 0-471-97047-6 [Fen07]. 0-521-52043-6 [Fra05]. 0-521-53983-8 [Che05c]. 0-521-54018-6 [Che05g]. 0-521-67595-2 [Sch09b].

0-521-68761-6 [FV09b]. 0-521-77477-2 [Pet06]. 0-521-785197 [Che05f]. 0-521-81733-1 [Ber06b]. 0-521-84347-2 [Sch06f]. 0-521-89308-9 [Che05e]. 0-7356-1965-4 [Che05d]. 0-7506-6507-6 [Ber06a]. 0-932633-45-5 [Mar05b]. 0-932633-46-3 [Sau06c]. 0-932633-59-5 [Sch06e]. 0-932633-62-5 [Ba05a]. 0-932633-64-1 [Sau06b]. ’08 [WWB09].

1 [Ano00b, Ber06b, Sau05d, Sau06b, Tra06a]. 1-4020-7869-2 [Tra06a]. 1-4184-9387-2 [Sau05d]. 1-4184-9388-0 [Sau05d]. 1984 [Ano07d]. 1993 [Ano07e]. 1st [CHR04, FFvdH01, GFF^{+}08].

2 [Ber06c, Bre06, Pet06, Sau05c, Sau05d, Sau06a, Sch06f, Tra06a]. 2-dimensional [XN09]. 2000 [Con00, Sau04b]. ’2001 [BOR02, Ano07b]. 2002 [Ano07c, Bea04, Enc02, Sau04a, WW04a]. 2003 [BM04a, BM04b, BM05a, BM05c, BM05d, Fra04a, Har04, Law04b, Law04a, Law04c, Pen04, Sau04e, Tra04e, Tra04f]. 2004 [BM04c, CGL^{+}04, EL04, Fra04b, PSS04, Sau04c, Sau04d, Tra04d, WW04b]. 2005 [Ano05a, BM05b, BM06, Ba05a, BK05b, GCL^{+}05a, Gla06, PA05, Tra04c]. 2006 [BM07a, BM07b, GGR^{+}06, Tra06b]. 2007 [BB07b, CEHK07]. 2009 [BFZ^{+}09, BB09, CPK09]. 21st [Ale02, Rho02]. 22nd [KS01]. 23rd [KS02]. 251pp. [Gve09b]. 295pp. [Gve09a]. 2nd [Che05f, DK09, Sau05d, Sau09a, EC01, Kim02, PSS05].

3 [Law05, Sau06c]. 3-540-70551-1 [Ebe09]. 334 [BM09a]. 3rd [Che05g, RPSS06].

4 [Che05d]. 403 [Hei09a]. 432 [BM08c]. 471 [BM08b]. 4th [BE05, CSSW01, DMS09, FFHL05, MZ01, PSSS07, WWB09].

5 [Bea05, Mar05b, Sch06e, WW05b]. 503 [Hei09b]. 5th [FFH^{+}05, PRRW04, WGM06].

6 [Ber06a, Fen07, Fra05]. 6th [CSSW04].

7 [Che05a, Els06b, FV09a]. 7-year

[WLCP09]. 7th [Con00].

8

[Che05c, Sau05b, Sch06d, WW05a, WW06]. 8th [AG01].

9 [Che05e, Els06a, Fen06]. 978-0-393-97950-3 [Sau09a]. 978-0-470-1763-6 [Gve09a]. 978-0-471-76867-8 [BM08c]. 978-0-521-69031-7 [Cho09]. 978-0-521-72596-5 [Gve09b]. 978-0-521-88038 [FV09a]. 978-0-521-88069-5 [Sch09a]. 978-00470-40129-3 [Hei09b]. 978-0470-14707-8 [BM09a]. 978-0521-71770-0 [Hei09a]. 978-3-527-40758-3 [Sau09b]. 987-0-521-87012-2 [Cho09]. 9th [CLS02].

A-MOST [DJP05]. AAL [MO06]. Aaron [Law05]. Abran [BM09a]. Abstract [Sch00b, CGS04, Cla00b, Dil00a, GGSV02, NFH05, Pin00, Wit00, Har04]. Abstraction [KH06, Bas08, CRH01]. abstractions [LS06, SHB05]. abstracts [ACM05a, ACM04b, ACM04a, ACM04d, ACM06a]. Abu [Hen09a]. Academic [Tra06a]. Academy [Win05d]. Accelerated [Fra01]. Accelerating [OOOiM05]. acceptance [US05]. access [Bow05, HCGGXFY00, HFH00, NC04, RDS05, VPDJ05, Zir06]. accreditation [LGA05]. Accumulation [CM05a]. Accumulative [YMI00]. accuracy [CMPB05]. ACE [Gup00]. Achieve [Tra04c]. Achieving [RMHL06, GZ05b]. ACID [Rom01]. ACM [Coo01b, Coo01a, DT06, Not07a, Not07b, Not07c, Not08a, Not08b, Omi01, Tra04c].

Acme [Ano00a]. AcmeStudio [Ano00a]. ACoM’07 [GGH^{+}07]. acquisition [CCHS09, HFA^{+}01, KY09, Sha02]. across [CSR06, KLS00]. ACT [Sch00a, NSAK09]. ACT-ONE [NSAK09]. action [Fra02b]. actions [Rom01, Sch05c, TGSP05]. Active [BRH04, CYD00, YFR00, ZXZ01]. ActiveAspect [CM05b]. ActiveScaffold [STB08]. activities [EH04, ZE00]. activity [Die07a, SM09a]. acts [dDLS02]. Ad [RM00b, JR02]. Adam [Sau05b]. adaptability [GB09, LE04]. adaptable [Hei00]. Adaptation [Sch00a, BK05a, FCSM05, GB09, MHB05, YYX06, YXP09, ZC05]. adaptations [DDMW05]. Adapting [DH00, LR01a]. Adaptive [ABNR05, DPT05, GFB05, LdSJR07, COWL02, GvdHT05, JAPM05, Lie00, MM09b, Wan05]. add [MN05, Win06b]. added [Dan01]. Addison [Bea04, Bea05, BM04a, BM05b, BM05d, BM07b, Bre06, Che05b, Fen06, Fra04a, Fra04b, Law04c, Law05, Pen04, Sau04c, Sau05c, Sau08, Sch06c, Tra04f, Tra08, WW04a, WW05a, WW05b, WW06]. Addison-Wesley [Bea04, BM07b, Bre06, Fen06, Fra04a, Fra04b, Sau04c, Sau05c, Tra04f, Tra08, WW05a, WW05b, WW06]. adjustment [AE05b]. ADL [Bra01, Oqu04a, Oqu06b]. Admission [Vog02a]. adoption [AR00, Cat09]. adoptions [PGL^{+}05]. Advanced [Boy02a, BH05, Pen01, Rom08]. Advances [CHR04, DJP05, PPV09, FLMW09]. advice [KFG04, Pen03a]. aerial [Jiu07]. affecting [Gil05, KM05a, SMC05b]. AFIS [Zir07]. after [OLHL02]. again [Sau05a]. against [MP05]. AGEDIS [HN04]. Agent [GCL^{+}05a, HS05, KGS09, Pan00, RM06a, BG09, BC05a, CGL^{+}04, DL05, Fas05, GCL^{+}05b, GGR^{+}06, GFB05, JR05a, Kei05, KG05, LMSF04, Lu04, Min05, OACL05, RM06b, SHRG05, TTOH01, TKA05, YYX06]. agent-based [BG09].

Agent-oriented [HS05, Pan00]. agents [ADD04, BG09, KWD05]. aggregation [CCWS02]. Agile [GMAC09, CBB^{+}07, CMV^{+}08, DGN05, HD07, LC05, NDG05, Pen04, Sil09, WLCP09, Che05a]. Agility [Bea05]. ahead [FBC06, GLB04]. Ahern [Law05]. aid [KD00]. aided [Alu00, Hen00]. aiming [WKP05]. air [MDC04]. Ákos [Sau04a]. Alain [BM09a]. Alan [Sau06b]. Albert [Hen09a]. algebra [Gog00]. algebraic [KJ05, RC01]. algebras [BCD00]. algorithm [LPH06, Nau02, SGAY09, TG06]. Algorithms [BE05, GK05a, WGM06, ASCP05a, BBM07b, BBG^{+}09, Law01b, PE04, Zir07]. All-Seeing [Ano00b]. Allen [Bea04, Gve09b]. allocation [Amr05]. also [Cho09]. alternative [BA00, NC00]. Ambler [Che05a, Che05f, Che05g]. Amman [FV09a]. amongst [CSTM06]. analyses [CHC05, Fel00, MRR02, Nau03]. analysing [PV06]. Analysis [EK01a, EL04, KWB05, KG05, Mar05b, Mor03, PTV01, Pin00, PA05, Rob05c, Tai00, Wan00, ASCP05c, Amm00, AMN^{+}06, Ano00b, Ano05b, Bas07, BM04c, BGK^{+}00, BM00b, BME^{+}08, BPSH05, Bow05, Boy01a, BCD06, BLS02, BC02, Che00, CKG06, Coo00, DSRJ04, DL04, DADY04, DPP05, Egy04, FV09b, FSN05, GL00, Gun00, HO05, HYD06, Hen08a, HP00, HSP06, IMP05, Jiu07, KD00, KB05, KJ05, KM06d, Lar06, Law01b, Lee09, LARSW00, LPH02, MSJ05, MSA^{+}04, MP05, Mas05, MO06, Mem05, MPCH05, Mye06, NMS09, NC04, Not03, Off00, Pan00, PGL^{+}05, PB05, PD06, Ram00a, RM06b, RSS^{+}04, bReWgH^{+}09, RSB04, RRU05, RKG04, RVR06, SK01a, SS09, Sce00, Sch08b, SK00, SGK09a, SS08, Sta00, Ste00, SMCH06, Sut05a, TG06, VTK02, VJ05, Vog02b, WT08]. analysis [YW04, ZYL04, ZLL04, Har04]. Analysis-based [Wan00]. analyst [HDS05]. analytical [MSC04]. analyze [Kei05].

Analyzer [Egy00]. Analyzing [Li02, SSG^{+}05a, Bul00, Egy00, EKHG01, FBR05]. anarchy [Min05]. Anchoring [AE05b]. Andersen [LPH06]. Andrew [Che05e]. Andy [Tra04d]. Anger [Tra04c]. ANN [SBS09a]. Anna [Cho09]. anomalies [Fas04]. anti [Joh00b]. anti-measurement [Joh00b]. anticipate [NPSS01]. Anticipatory [PSPS05]. antipatterns [BPSH05, Lon01]. Antonio [Sch09a]. any [dDLS02]. AOSD [CEHK07]. AOSQUAMO [KGK09]. API [DFMS09, Per06]. APIs [dSRC^{+}04]. appliances [Mar04]. Application [BBM07a, FFH^{+}05, GLM^{+}05, SKM09, ASD03, BB07a, BH05, BJL^{+}05, Kei05, LR01a, LSK^{+}05, MG00, Ram00b, SSG^{+}05a, SS05, TTOH01, Tan01, WSH05, YMI00, BM04c]. Applications [Car07, CHR04, ASCP05a, Ana02, BMSS09, BME^{+}08, BPSH05, CCN^{+}05, CDF^{+}00, DFW04, EV05, IT01, Jah01, MP05, MN05, MHB05, NC04, OG05b, PSPS05, RDS05, RM00b, RM04b, SHB05, SN05, Sta01, SCT04, ST00, Swi00, XXJ05]. applied [CMV^{+}08, Lak05, Rie05]. Applying [HKO^{+}05, NTS07, Sut05a, Die06b, LMSF04]. Approach [Sau06c, ASKS04, BTA06, BMK^{+}05, BBHK04, BBW05, BS00, BS01, BMSS09, CEM02, CC05, CdOBT05, DS05, DL05, Ehr05, FCSM05, GSV05, GR06, Ham05, HB05a, Jan01, KY09, KSWR06, KSM06, LKS00, LMSF04, LLT05, LvdBC00, MMvD05, MB05, Mei00, MLG05, NR00b, Ned00, OACL05, PJ05, RM01, RC01, RM06a, RTDR05, RM04b, SNEC06, SSM05, SABS05, SKG08, Smi00, Wen01, YYX06, ZGL05, ZD05, BM05a, BM07a, Sau05c, Sau09a, Fra04a]. Approaches [DS06, Ell07, KCM05, MBPRR01, Ric00a]. approximation [Mei04]. Approximations [Fas04]. apps [Win09a]. April [BM09a]. Arabica [TR00a]. Architecting [dLGR05, dLR05, BM00a, Far06, Tra04d].

Architectural

[AKL^{+}07, CEHK07, GvdHT05, LA06, LFW02, SN05, Sta00, WSH05, XGY02, dB06, AAAG^{+}05, ANGRB05, Ano00b, ALK09, BMMG05, BCD00, CNPD06, DDMW05, DYC^{+}05, GBRS06, GWH05, IMP05, JS00, JME05, LM08, Min00, MT01, PM03, PSK09, RM04a, RTDR05, Ros01, SvdVD06, Ter04, WLF01, XZRA05]. Architecture [BB08a, BB09, CPK09, KKS01, Oqu04a, AKF06, Aug09, BMK^{+}05, BC05a, BD03, BM09b, BM00b, CNS05, COWL02, DGP^{+}01, GGH05a, Hua07, JR05a, JRHD00, KJP05, LLT05, MO06, Mei00, MDR05, NPSS01, Oqu04b, Oqu06a, QXCMW05, Ran01, Ric00a, SBCO01, Smi00, Tai05, UCKM04, WK05, XYH04, YYX06]. architecture- [Ric00a]. architecture-based [KJP05, MDR05, QXCMW05, YYX06]. architecture-centric [Oqu06a]. Architectures [BKW04, KS01, KS02, BB01, Bar05a, CK01, EAG^{+}05, FBR05, FSN05, GMS01, MSR02, MO06, Oqu04a, Oqu04b, Oqu04c, Oqu06b, San09, SGY05]. ArchStudio [TR00b]. Arell [Ano07b]. Argo [TR00c]. Argo/UML [TR00c]. arguments [WDKM05]. ARGUS [Ano00b]. ARGUS-1 [Ano00b]. ARIES [HKKI05]. Arithmetic [RRR05]. ARL [Oqu04b, Oqu04c]. ARM [XMP08]. arrays [YCP04]. art [AH01, Ber08, CCHS09, Sch09d, Stu05]. Arthorne [WW05b]. Articulating [SAM05]. artifacts [OYH04]. artificial [Ben00]. ASM [WHW08]. ASM-based [WHW08]. Aspect [CPK09, KH01, AX06, BV08, DS06, GKK08, KY09, KFG04, KKG09, KGK09, MMvD05, RSB04, SGS^{+}05b]. Aspect-Oriented [CPK09, KH01, AX06, GKK08, KKG09, KGK09, RSB04, SGS^{+}05b]. aspectC [CKFS01]. aspects [BK09, CPK09, DB09, GBRS06, KMM05, MO04, Tho01b, ZMC06, CEHK07]. assertion [CR06b, ZGL05]. Assessing [Ano00c, CCK04, KPCD00, Sil09]. Assessment [Rot04, Bar05b, LKS00, Lee09, MM09b, MB09, MBM09, SDF00, SGK09b, SKT05, Wen01, GGH^{+}07]. asset [SDF00]. assigned [BBHK04]. assisted [JS00, PC05b]. association [STB08]. Assume [BGP06]. Assume-guarantee [BGP06]. assurance [HB05a, SYC^{+}04, WDKM05, You00, ZTSR05, ZHG05, ACM05b]. ASTEC [MB05]. asynchronous [SMK09]. atomic [Rom01, Sch05c]. atomicity [FFQ04]. attack [GW05a]. attacks [BP05, HO05, MP05]. Attitude [Sau06c]. attributed [FLY03]. Auerbach [BM04c, BM05a, BM05c, Sau04e]. Augustine [Sau05c]. authentication [RG06]. author [ACM04c]. AuthorHouse [Sau05d]. authorisation [SAM05]. authors [Jay06]. automata [BČVZ06, dAH01]. Automated [BOP00, CRL^{+}06, Dil00b, Eme00, GJJL00, GL00, Gun00, Mar01, Mei04, MPS00, PTV01, Raj06, Ste00, Var00, AX06, Ano01a, BS01, BJL^{+}05, BKM02, CSSW04, JM00, Ric00b, SSG05b, Say05, SSHP05]. Automatic [CRH01, Cla00a, IT01, NE02a, Rob05b, WML02, BRH04, Cle05, HBZ^{+}05, KLS00, Mar07, Smi00]. Automatically [Per06, LZ05a, SWC05]. Automating [DSRJ04, DL04, Jac00, MVT^{+}02]. Automation [DMS09, AMS05, BM05e, Gre00, JC09, MK04]. Automotive [PIÖ05, PSS05, PSSS07, RPSS06, BKKS05, EAG^{+}05, FSN05, PSS04, WBH^{+}05]. Autonomic [GLM^{+}05, ST05b, BMK^{+}05, CCN^{+}05, GZ05b, KAKM05, LLMY05, LWZ05, MN05, SMC05a, SL05, SMC05b, WSH05]. Availability [GWH05, HB05b]. available [Cho09, OYH04, ZRWS05]. avionics [Bra00]. avoid [Fra02a]. award [Bur00]. aware [DMNJ06, JR02, PSPS05, SM00, XC05, CC05]. awareness [ŚZZ05a].

B [BM05c, Gve09b, Jiu07, Sau06a, NRSS09, Sca05]. Back [ACM03, ACM04a]. Backmatter [ACM05a, ACM04b, ACM04c, ACM04d]. backwards [XDS04]. backwards-compatible [XDS04]. bad [CS09]. Bader [Law04b]. Bakar [Hen09a]. Balancing [Bea05]. banking [NTS07]. Barr [Sau05b]. Barriers [Cat09, LCS^{+}05]. Barry [Bea05]. Bart [BM05d]. base [GL06, RN00]. Based [BB07b, CSSW01, CSSW04, CRS^{+}05, Mar05b, SKK01b, TLG01, ASKS04, ASK04, ANGRB05, BV08, BNL05, Ban08, BG09, BLQ^{+}06, BMSS09, BM00a, Boe05a, BKM02, Bra01, BCD06, BP05, BČVZ06, BK05b, CNPD06, Car00, CTZ02, Cle00, CdOBT05, CLS02, DJP05, DeL05, DPP05, EK01a, ELC06, FFPT05, FLMW09, FK01, GBSO04, GKKW05, GG04a, GB08, GL06, GJJL00, GKS07, GR06, HCGGXFY00, HFH00, HR06, HN04, Hie04, HB05a, Jah01, JB05, Jiu07, Joh05, Kai00a, Kai00b, KS09c, KM09, KDJ07, KDKJ08a, KDKJ08b, KJP05, KBBL05, KSWR06, KM06a, KSM06, Lak05, LKMU05, Li02, LCS^{+}05, LWZ05, LCLP05, LLT05, LSK^{+}05, LYF^{+}05, Lu04, LR01b, MK01, Mao09, MMvD05, Mar07, Mar01, MZY01, MB09, MBPRR01, MDR05, NMS09, NC04, NS09, NR00b, NKB05, NAD02, OG05b, Oqu04a, PWZ^{+}04, Pan09b, Par05]. based [PSSA06, PP00, QXCMW05, RJH05, Raj06, RM01, RM04a, RRSW00, RM06a, RM06b, RCL05, RSS^{+}04, bReWgH^{+}09, RRCDHM05, Ric00a, Ric09, RRU05, SM09a, Say05, SABS05, Sce00, SMD^{+}05, SDF00, SKM07, SGK09a, SC05b, SS08, SW05, SPS^{+}08, SKR08, Sta00, SCFP00, ST00, Tra04d, TKA05, UCKM04, VTK02, VWW02, WHS01, Wan00, WZX06, WLF01, WLA05, Wit00, WHW08, YYX06, ZLX04, dH05, HIM00, BM09b, BS01, Bla00, FBHZ08, FT05, RM04b, Hei09b, Ano07b, FV09b]. Baselines [SHD05]. Basic [SV00, BHP^{+}06, XYH04]. Basic.NET [Che05c]. basis [DB05]. batch [KGS05]. batches [Ezi04]. BatchFlow [KGS05]. bazaar [FFvdH01]. Be [Win06a, Enc02, GHS^{+}06]. Beans [EK01b]. Beck [WW05a]. before [GG04b]. beginning [Lar01]. behavior [AMS05, Aug09, BRH04, SM09b, WT08]. Behavioral [FLF01, EKHG01, LBR06, Ric00c]. behavioural [MO06, RC07, UC04]. behaviours [LM02]. benefits [Dan01, Sut05b]. Bennatan [Sch06c]. Berlin [Ebe09]. Bernstein [BM06]. Best [Fen06, Sch06e, Bea04, CNS05, Pan09a]. Beth [BM04a]. Better [MW05, Boe05b, CCS00, Gil06]. between [DA05, RDS05, SC05a]. beyond [Eme00]. Bezem [Law04a]. B’Far [Ber06b]. bibliography [Kie01]. bidirectional [BM09b, Boo00]. Bijay [Sch06d]. Bilisoly [Gve09a]. Bill [Gla06]. binaries [CKG06]. biological [DS05]. Black [Ano07c, Sau04c, SK00, Mei04]. Black-box [SK00, Mei04]. Blanchet [Law06]. block [Mar01]. block-based [Mar01]. blues [Win06d]. Bob [DT06]. body [JMV04]. Boehm [Bea05]. Book [BM09a, Cho09, Ebe09, FV09a, FV09b, Gve09a, Gve09b, Hei09a, Hei09b, Hen09a, Hor01, Man00b, Sau05b, Sau09a, Sau09b, Sch09a, Sch09b, ACM05a]. Boolean [SCS06]. BOOT [Ric00a]. Börger [Har04]. boundaries [vdH01]. bounded [SYC^{+}04]. bounds [Bis02]. box [Mei04, SK00, TDSCdlR08]. branch [FD00]. breakpoints [PC05b]. Brede [Tra06a]. breeding [RRR04]. Brennan [BM07a]. brewing [CCS00]. Brian [Hei09a]. bridge [RDS05]. brief [XQZ^{+}05]. Broekman [BM05d]. Brooks [Hee03]. Bruce [Bea04]. buffer [ZLL04]. Bug [Sau05b, Lar06, LYF^{+}05]. Bugs [Pet00, Hol02, HSP06, JV00, OWB04, SZPZ06]. Build [Fen06]. Building [BK05a, Bre06, Che05b, DER04, Fra02a, Tan01, Vie05, Ben00, CBB^{+}07, CLS02, HDLC04, JMV04, PJ05, Pin00, Bea04]. built [Aug09, Ler04, Sau06c]. Bumgardner [Che05e]. Burt [Win06b]. Business [FSN05, NKB05, Ano05b, KJML05, Law04c, STG02, Sha02, Sta01, Väh05]. buying [Sha02].

C [Ano07e, BM04c, BM06, Che05e, Fra05, Gla06, Hen09a, Law04c, Pen01, Sau04e, Sau04a, Sau08, Sch06d, XDS04, Boy01b, CCN^{+}05, CYD00, FCSM05, Fra01, Fra02b, HBW00, MB05, Say05, SMA05, Sau09a]. C# [FV09b]. C-Cube [CCN^{+}05]. C3 [PM03]. Cairns [Cho09]. calculus [BBM07a, Bar05a, Oqu04a]. Calendar [Ano07a, ACM05a, ACM05d, Tra05, ACM04d, ACM05b, ACM06a, ACM04b, ACM04a]. Call [ACM04d, RKG04]. Calls [Ano07a]. Calvanese [Law04b]. Cambridge [BM04b, BM08b, Ber06b, Che05e, Che05c, Che05f, Che05g, Cho09, Fra05, Gla06, Gve09b, Hei09a, Law04b, Law04a, Pet06, Sch06f, Sch09a, Sch09b, Tra04e, WW04b]. Campbell [FV09b]. can [CMPB05, Chr08, Dek05, Mar04, ZBPD01, Zir07]. Capability [Bra00, Tra04d]. capable [Boo00]. Capture [SCFP00, OK05, SvdVD06]. Capture/Replay [SCFP00]. capturing [CNS05]. car [BHLRM05]. card [BJL^{+}05]. care [Win05b]. CAREER [Ros00]. Carlo [Sau04a]. Carol [BM07a]. Cascaded [BX01]. Case [BV00, Par05, BK09, Boy02a, Die06b, EAN01, FD00, Fra04a, FHK^{+}01, GT07, GGH05b, KS09b, KY09, LARSW00, MHB05, NWVO05, Oqu04c, PAGDLN05a, PIÖ05, Raj06, RN00, SM09a, SSG05b, SW05, SKR08, Ter04, WLA05, dFAM06, AR00]. case-base [RN00]. cases [ASK04, BM01, CP05, Die06b, EMR00, Fra04b, FSN05, KG01, NS09, OX04, Rie08c, ZXZ01, Zün01, dFAM06]. Catastrophe [Sch06c]. categorical [dCJL06, Sau04b]. categorization [PRSH07]. Categorizing [NB09, SSM05]. Causal [Ric09]. Causal-based [Ric09]. Cause [SPC09, Zel02]. cause-effect [Zel02]. caused [Mar04]. CBS [GT07]. centered [GKKW05]. Centric [Sau05c, Jah01, Oqu06a, RTDR05, XGY02]. century [Ale02, Rho02]. CEP [Luc00]. cepstrum [ARV05]. certain [DB02]. certificates [She05]. Certification [CSSW01, KS08b, LGA05, WBH^{+}05]. CHAIMS [WBSM00]. chains [RKG04, Zel02]. chair [ACM05b, ACM06a]. challenge [JW01]. Challenges [DMS09, And04, AKL^{+}07, BL04, CM05a, RB05c, Spr05, vdH01]. Change [Ba05a, CC05, GGT05, KWB05, NPSS01, Raj00, RRWB00, Win09c, Ano07b]. Change-based [Ano07b]. changeability [GKK08]. changes [ŚZZ05b, ZRWS05]. changing [NPSS01, WW00]. chaos [DA05, KPCD00]. characteristics [KB05, YEMT05]. characterization [Gil06, dCJL06, YCP04]. CHARMY [IMP05]. chart [AMN^{+}06, UKM01]. charts [SC02]. Cheaper [Sau05d]. checker [RMLMG05]. checkers [CEH03, FW06]. checking [Amm00, Bul00, CRH01, CR06b, Dil00a, Eme00, FBS04, GV02, GDC02, Hor00, NE02b, RRSW00, TDA07, ZGL05]. Chemical [Ebe09]. CHI [DFMS09]. Chimera [TR00d]. Choir [Tra05]. Choosing [MG00]. Chris [Sau05c, WW05b]. Chrissis [BM04a]. Christine [BM05c]. Clamp [AX06]. CLASP [Vie05]. class [Tee09, VJT08, Wor02, YWG04, YW04, YF00, ZLX04]. Classboxes [LS06]. Classes [BOP00, KS06, SKM09, TDA07, Ton04].

Classification

[KD00, BRH04, HKO^{+}05, NC00, RSB04]. Classifying [NC00]. classroom [WHS01]. Clayberg [Bre06, Che05b]. cleanroom [LP05]. ClearCase [Ben02]. Cleland [Sau04d]. Cleland-Huang [Sau04d]. client [JS01, MM08a, MM09c]. client-server [MM08a, MM09c]. client-side [JS01]. clone [HKKI05, KSNM05, KN05]. clones [KN05, YR08, WLK04]. Closing [ZBPD01]. clothes [dDLS02]. Clouse [Law05]. clues [SPT05]. clustering [AX06]. clusters [SSG^{+}05a]. Clymer [Hei09b]. CMM [LSK^{+}05]. CMMI [BM05a, Ber06c, Boy02b, BM04a, Law05]. CMMI-SW [Ber06c]. Co [DRL^{+}05, Mar05b, Dek05]. co-located [Dek05]. Co-Operative [DRL^{+}05]. Coaching [BM07b]. coaster [Win06f]. COBOL [Gla06]. Cocoa [Win07a]. Code [ASK04, KKB07, ASCP05b, BLS02, CL05, CKFS01, DNR06, Eng03, HKKI05, JC09, KCM05, KSNM05, KN05, KON05, LLL01, LZ05a, MR06, NFH05, SWC05, SGS03, TH02, VDM^{+}06, VJT08, YWA05a, YWA05b, YR08, ZYL04, ZRWS05, ZLL04, Zün01]. code-design [ZYL04]. CodeSurfer [Tei00]. Coding [Sau08, ARV05, GG04b]. coefficients [ARV05]. Cognitive [Fen07, KM06a, KSM06, KM06b, KM06c, KM06d, MM07, Mis07]. cohesion [CSTM06, WBG05, ZLX04]. Colin [FV09b]. Collaboration [CHC05, FFHL05, ABNR05, CCS00, FK01, OOOiM05, RK00, WW04a, dSRC^{+}04]. collaboration-based [FK01]. collaborations [Cap04, DOV^{+}01, Rie08a]. collaborative [OOOiM05, Ebe09]. collected [HKO^{+}05, Hor01]. collecting [CO04]. collection [MLG05]. colors [SGS03]. coloured [AMN^{+}06]. colouring [Bre05]. Column [BÁ09a, BÁ09b]. COM [EK01b, IT01, JS00]. COM/DCOM [IT01]. combination [MBPRR01]. Combinatorial [Bre05, KKN07, WW06, CSR06, LJD05]. Combining [Eng03, HO05, HBZ^{+}05, LQV01, WDKM05]. come [Ras00, SZPZ06]. comments [YWA05a, YWA05b]. commerce [Sca05, Sta01]. Commercial [Bre06, Che05b, FSN05, GGH05b, HBW00]. Commercial-Quality [Bre06]. commitment [Sch00a]. committee [ASCP05b]. commodities [ZMC06]. common [LZ05b, Sha00, VWW02]. communicate [AE05a]. communication [DeM00, JeMD04b, RMHL06, Tai05]. Communications [Sch06f]. communities [WGS07]. community [ELC^{+}02, HDLC04, SR05]. commutativity [DSRJ04]. companies [RB05b, Tra06a]. Company [Sau09a]. Comparative [GZES04, NH09, NSAK09, ZLX04, ZHG05]. comparing [BCD06]. Comparison [FD00, MM07, MSC04, Nam08, STB08, YWG04]. comparisons [Wil04]. compatible [XDS04]. Competitive [Ber06a]. compilation [PD06]. Compiler [BRBY00, Ezi04, Hin00, THL03]. compilers [UT05]. complementary [Mei00]. complete [Mun00]. completion [Bre05]. Complex [BG09, Hei09b, Luc00, Sti00, Bas08, Die07b, Die09, LR01a, RK00, Sch05a, Sch05b, WW00, YCP04]. Complexity [Bha05a, KS08c, Bha09, CG09, GG04a, GB08, KM06a, KM06b, KM06c, KM06d, MK06, MM07, MK07, Mis07, Pan09b, PH07, RR09, YK09]. Component [Bra01, BČVZ06, CSSW01, CLS02, CSSW04, CRS^{+}05, EK01b, GB09, MM09b, Stu05, WT08, BLQ^{+}06, BK05a, BMW01, Cle00, CO06, DYC^{+}05, FFPT05, GBSO04, GG04a, Gil06, GB08, GMS01, GKS07, GR06, JS01, Jah01, dCJL06, KS08a, KS09a, KS09b, KS09c, KS08c, KDJ07, KDKJ08a, KDKJ08b, LR01b, MZY01, MBPRR01, PSSA06, RRCDHM05, SKM07, SGK09a, SSH^{+}05, Sta00, SM09b, TK06, TvH00, WHS01, WZX06, WML02, Wit00]. Component-Based [CSSW01, CSSW04, CRS^{+}05, CLS02, BLQ^{+}06, BČVZ06, Cle00, GBSO04, GG04a, GB08, GKS07, GR06, Jah01, KDJ07, KDKJ08a, KDKJ08b, LR01b, MZY01, PSSA06, SGK09a, Sta00, WHS01, Wit00]. Component-interaction [BČVZ06]. component-oriented [KS08c]. Components [CFM05, DF04, Fel00, Ano05b, Bar05a, CLS02, GKS07, Ham00, Hei00, KS08b, KS09d, LB05, Lea00, PND01, PS00, RM01, Rie07a, SKG08, SGK09b, SC05b, SKK^{+}01a, Sre01, SS05, SD01, WLA05, XYH04, YKB02]. componentware [Lu04]. Composable [NAD02, COWL02]. Composing [HA00]. composite [Bul00]. Composition [WBSM00, BK09, KS09d, LdSJR07]. compositional [dCJL06, LS06, LvdBC00]. comprehension [KM06b, MBPRR01]. comprehensive [DL04, Vog02b]. compression [Zir06, Zir07]. computation [Law01a, Mai00]. Computational [Car09]. Computer [Alu00, BM06, BM08c, Ber06c, Cho09, CLS02, Ebe09, Fen07, Mar05b, Sau06a, Ano05b, Haz07, Hen00, Joh05, Sch09c, Tem09, WW02, Zel02]. Computer-aided [Alu00, Hen00]. Computer-Based [CLS02, Mar05b]. computers [Chr08, Neu00a, Neu00b, Neu00c, Neu00d, Neu01a, Neu01b, Neu01c, Neu01d, Neu02a, Neu02b, Neu02c, Neu02d, Neu03, Neu04a, Neu04b, Neu04c, Neu05e, Neu05f]. Computing [ASC02, Ano05a, Ber06b, Car07, TGSP05, Cam05, CEM02, HBR05, Kar00, Nau02, SMC05a, ST05b, SL05, SMC05b, XGY02, Hen09a]. Concept [Fra05, Bow05, MDC04, OG05a, TG06]. Concepts [Man00b, AH01, Ban09, BC05a, BHP^{+}06, Man01, Ran01, WHS01]. conceptual [Egy00]. Concern [HOST05, Mem05, UT05, LLT05, MMvD05, MR06, RM05]. Concerns [Rob05c, THO^{+}01a, GGH05a, Ham05, JME05, PAGDLN05b, SG05, SPT05, WBGM05]. concolic [SMA05]. concurrency [KvB04, Mis00, YKB02]. Concurrent [Cla00c, Ols00, Att00, Ban07, Ban09, BBM07c, BS00, Car00, CLS00, Cle00, DL04, Gog00, NR00a, Sis00, Tai00]. conditions [TvH00]. Conference [CLS02, KS01, KS02, ACM05d, TGF09, Tra05, GHP^{+}09, LRBH01, PHP08, SPH07, ACM05b, ACM06a]. confidence [JS01, RM01]. Configuration [Ben02, Fen06, vdH01, CCS00, CCWS02, DMNJ06, ELC^{+}02, GvdHT05, MZY01, PSPS05, TKA05, WLF01, YCP04]. configurations [CSR06]. conflict [CEM02, FFHL05]. Conformance [CO02, LM02]. connecting [Väh05]. Connector [SC05b, LLT05, MSR02, RM01, RTDR05]. Connector-based [SC05b, LLT05, RM01]. connector-centric [RTDR05]. connectors [IT01]. Connotations [Sri01]. conquering [WW00]. conservative [Nau02]. consider [Tho01b]. consideration [Zir06]. considerations [GG04a, GKS07, LLL01]. consistency [EKHG01, TDA07]. consistent [RG06]. conspiracy [Ban09]. Constraint [CGB06, HCGGXFY00, Van00a, BBW05, EK01a, EJTY04, GJJL00, JV00, Pon00, Ves00b]. constraint-based [GJJL00]. constraints [CRH01, CMV^{+}08, Jay00, TK06, Van00b]. constructing [KP02d, UT05]. construction [Bas08]. constructs [EAN01]. Containment [COWL02]. Contemporary [GGH^{+}07]. content [APS09, PSK09]. contents [ACM04e]. Context [XA05, Die06a, JR02, KG01, LPH06, UCKM04, WM00, XC05, Fra04b]. Context- [XA05]. context-aware [JR02, XC05]. context-sensitivity [LPH06]. contextaware [SHB05]. contexts [KG05, LS06]. continuing [DOV^{+}01].

Continuity [Ham02, Rie06]. Continuous [GLB04, OLHL02, SE04]. continuum [Rie05]. contract [CGB06]. contractors [Rot04]. contracts [BLS02, DPSJ05, FLF01]. Contributing [WW05a]. control [Bha05a, CMV^{+}08, FFHL05, HCGGXFY00, HFH00, HG05, Jiu07, JAPM05, LWZ05, MV01, NC04, RDS05, RVR06, Suk01, Tra04f, VPDJ05, Vog02a, WBH^{+}05, YKB02]. control-flow [RVR06]. controlled [NWVO05, TDSCdlR08]. controller [GKKW05]. Controlling [Ba05a]. controversy [dDLS02]. conversation [Rac01a]. Coombs [BM04b, Tra04e]. cooperating [PSK09]. Coordinated [Rom01, CF00, CL06, Sch05c]. Coordination [Amr05, PSSA06]. copyright [ACM04e]. CORBA [BRV01, EK01b]. Corps [Rug00]. corpus [FR05]. correctness [BBM07b, Fas04, Mei04]. Correlation [KAYE04]. COSMIC [ZTSR05]. Cost [Fen07, Tra08, Bin00, CMPB05, MSJ05, MPCH05]. Cost-Effective [Fen07]. costing [BM04b, Tra04e]. costs [Sut05b]. COTS [DF04, Kim02, Bal00, Bla00, DOV^{+}01, DYC^{+}05, FT05, HSGPSF05, KS09a, MRE05, RRCDHM05]. COTS-based [Bla00, FT05]. Countdown [Ba05a]. counter [HO05]. counterexample [TAC^{+}04]. counterexamples [CGS04]. countries [Gri00]. coupling [ASC02, HLL02, KKG09]. Coven [CCS00]. Cover [Cho09, ACM04e]. Coverage [MSP01, ASK04, BC05b, CL05, FHNS02, NMS09, Pet05, RMHL06, SCT04, TH02, VJ05]. Covering [YCP04]. Cox [Cho09]. CPAS [Bra00]. CRC [BM04c]. Creating [Hub01, Spr05]. creation [Ano05b]. Creativity [DBCP09]. crisis [Rom07]. criteria [MSP01, MK06, MK07, VWW02]. critical [CPDGP01, LGA05, MP09, MK05, PH05, Sch06a, Sau04c]. Criticism [Tra04c].

CROPS [CF00]. cross [LKF02, OOOiM05]. cross-cutting [LKF02]. cross-project [OOOiM05]. crosscutting [CM05b, MMvD05, PAGDLN05b, SPT05, ZGL05]. crossword [Ber06d]. CRUD [STB08]. CS [BM08c, Fra06]. CSCW [Sut05a]. CSE [Car09, HFA^{+}01]. CSEE&T [Let06]. CSP [NH09, NSAK09]. CTL [TK06]. Cube [CCN^{+}05]. Cultural [MHK05, Win09c]. cumulative [SPS^{+}08]. Cunningham [Els06a]. CUP [FV09a, FV09b]. Current [KJML05, Bha05c, TvH00]. Curricula [Ano05a]. Customer [GP05, vD01]. Customizable [Mil00, SHB05]. customization [CKFS01, MRE05]. CUTE [SMA05]. cutting [LKF02]. CVS [MLRW05]. cycle [KS09d, Ova04]. cycles [SA00]. cyclone [Mor03].

D [Hei09a]. Dale [Bea04]. damage [TGSP05]. Dan [Bre06, Che05b]. Daniel [Law04c]. Daniele [Law04b]. dark [Sch09d]. Daryl [Fra04b]. DAS-BOOT [Ric00a]. data [BRV01, Bha05a, BDE^{+}03, CRL^{+}06, CHC05, Coo00, DA01, EK01a, Eng03, FORG05, FD00, God01, HKO^{+}05, HBZ^{+}05, KGS09, Mad01, MSJ05, Mas05, MLG05, MACE02, Nau03, OX04, PC05b, RGB05, SGAY09, VJ05, Wil04, Zir06, Zir07, Sau04b]. data-model [BRV01]. database [ASD03, CDF^{+}00, DFW04, SCT04, TDSCdlR08]. databases [Zir06]. dataflow [FJM05, HYD06, Van00b]. datatypes [KJ05]. Datrix [LLL01]. David [Che05d, Hor01, Sch09b, Tra04c, Tra08]. Davis [Sau06b]. day [WHS01]. days [Sti00]. DCOM [IT01]. dead [ZBPD01]. Deadlock [KE01, IT01]. Dealing [AHW05, NvdBA05, Ter04]. Dean [Fra04a]. DEAS [GLM^{+}05]. debate [Bha05b]. Deborah [Law04b]. debugged [Sch08a]. debugger [Boo00, SM00]. Debugging [Hor00, Sch08a, AM03, BRBY00, Dev00, Dil00b, RB05c, THL03, Van00b]. December [Hen09a]. Decision [Ale02, MRE05, FBC06, KWD05, SPC09]. decision-making [KWD05]. decisions [CNPD06, GBRS06, ZM05]. declarative [OACL05, UK06, WV04]. Deducing [Bha09]. Deeper [Sch05a]. default [CR06a]. defect [Boe05b, Hen08a, KL05, LSH^{+}04, SNWV05, She05, WS05]. defect-detection [WS05]. Defects [Ba05a, Bag09, Hen08a, JeMD04a, Lev09, RR09]. Defense [Lak05]. Defining [Lev09, Egy00, WW04a]. Definition [BB08b, Bre05, Rey00]. Definitions [Rac01b, CCO00]. degradations [Lev09]. Delegation [YXP09]. delivered [FD00]. delivery [YFR00, Ba05a]. demands [ELC06]. Demeter [Lie00]. Demeter/adaptive [Lie00]. demographics [Rac01b]. demonstration [Sti00]. Dennis [Law05]. density [CCK04, SNWV05, She05]. dependability [AAAG^{+}05, FR05, KM05a, LG02, Rom08, XZRA05]. Dependable [dLGR05, dLR05, GvdHT05, RM00b]. Dependence [VTK02, NMS09]. Dependency [GB08, RR09, SGK09a]. Dependent [Bil03, WLA05]. DEPLOY [Rom08]. deployed [BOH03, LSH^{+}04]. deploying [Pin00]. deployment [CO06, OLHL02, Rom08]. depot [TR00g]. derivation [Nad00]. derived [OG05b]. Deriving [LvL02, GG04a]. describe [PM03]. Describing [KKS01, GCS05]. Description [Oqu04a, Law04b, Ric00d, CEHK07]. Design [Ebe09, FBC06, GLM^{+}05, Jay06, MM08a, PP01, Sti00, AK09, BV08, BNL05, BB07a, BG09, Bil03, Bla00, BME^{+}08, CS05, CNPD06, CGB06, Che02, CLS00, DS06, DBCP09, Dek05, DSRJ04, Dro09, DNR06, GBSO04, GKKW05, HLL02, Har04, HW05, KS09b, KM09, KG05, KPDdL00, LLMY05, Law02, LBR06, Lee09, LvL04, LMSF04, LCLP05, Man00a, MM09c, MK05, Nau03, NvdBA05, NPSS01, Ols00, Pan09a, Pan09b, Pen03b, PSSA06, Pet05, RM04a, RR09, Red00, RM09, Rey00, Ric00a, ST05a, San09, SABS05, Sch09d, SGCH01, SGS^{+}05b, Sut05a, Tee09, WSH05, WK05, YK09, YF00, ZMC06, ZM05, ZYL04, CPK09, Gve09b, Sch06d, Sch09a]. design- [Ric00a]. Designed [Sau06c]. Designing [Rie07a, RM00a, Sil09, Hei00, Jan01, KP02c, Pen03b, SMD^{+}05, Sha01]. designs [CS09, FK01, GW05a, RRR04, Ram00b]. details [Neu07a]. detect [MP05, RR09]. Detecting [BJ05, UKM01, LZ05a]. Detection [WLK04, CL05, GW05b, KE01, Lar06, Par05, PE04, Rei05b, WS05, XA05, XC05]. detectors [CJ04]. determine [HB05b]. Determining [HB05a, OCE^{+}00, SPS^{+}08]. Deugo [Pet06]. Developer [Els06b, RGB05, Enc02, Joh00b, Pen02]. developers [CSTM06, LKS00, Pin00]. Developing [Els06a, JB05, Ana02, CCN^{+}05, EAN01, Lan06, NPSS01, Sch05a, Sch05b]. Development [BM07b, DRL^{+}05, Sau05c, Sau06b, TLG01, Amr05, BV08, BS06, CR06b, CBB^{+}07, CMV^{+}08, DB02, DeM00, DOV^{+}01, Die07a, DK09, EAG^{+}05, Fen09, FT05, GCS05, GL00, HFA^{+}01, HDLC04, HD07, Hua07, Hub01, KS09d, Kei05, KDJ07, KDKJ08a, KDKJ08b, KPCD00, KM06a, KSM06, LG02, LCS^{+}05, LR01b, Mar05a, Mar07, MZY01, MBPRR01, Mun00, Pen04, RM06a, RCL05, RS07, Rob05a, RM00b, SE04, STG02, ST05a, Sca05, Spr05, ST02, SDD05, Sut05b, Tan02a, Väh05, VS08, VWW02, Zün01, ZHG05, dSRC^{+}04, BM05c, Fen07, Sau04d]. devices [BFN^{+}05, CBB^{+}07]. diagnosing [CGB06]. diagrams [LM02, NS09, RVR06, SM09a, VJT08, YWG04, YW04]. dialog [Tra02]. Diego [Law04b]. Dieter [Man00b]. differencing [JC09]. different [CL05, DH09, KS09d, KON05]. Dijkstra [BBM07a]. Dimensional [DA01, THO^{+}01a, BBG^{+}09, WBGM05, XN09]. Dingsøyr [Tra06a]. Dining [BBM07a]. direction [Bha05c]. Directions [GWS04, AG01, FD07, Hen00, LR02]. directive [Ros04a]. Discipline [Bea05, KM05b, Rho02]. DISCOA [DDMW05]. DiscoTect [SGY05]. discover [SPT05]. Discovering [Rec07a, SGY05]. discovery [dB06]. discrete [LKMU05]. discrete-time [LKMU05]. Discussion [KB05, BHP^{+}06, MM09c, Sre02, TWQ09]. Disentanglement [Sch06c]. disseminating [LCS^{+}05]. dissimilarity [Bha05b]. distance [YW04]. Distilled [Law05, Boy02b]. distinguishing [RMLMG05]. Distributed [DB09, Ebe09, HW00, KP05, TLG01, And00, AHW05, BD03, CO02, Dek05, Gog00, GMS01, JR05a, JRHD00, KM05b, Lon04, MSR02, MHB05, MT01, Mye06, PAGDLN05b, Pon00, Ram00b, Ros00, Sis00, SN05, TKA05, WSH05]. distribution [DLP01, FFPT05, KE01, LFW02, OW02]. DNA [BB07a]. do [Rac02, SZPZ06, ŚZZ05b]. Documentation [Ber06c, Mam00, GT07, KK06, Mar07, BM08c]. documented [Rie08a]. documents [BM08a, Mun00]. DoD [Tra04c]. dogs [BK01]. Doke [Gla06]. Domain [Sto02, BLQ^{+}06, Dev00, NTS07, PND01, RM06a, WAK00, Zir07]. domain-and-organization-specific [PND01]. domain-specific [Dev00, RM06a]. domain-to-range [WAK00]. domains [Bla00, LR01a, Mil00]. DoME [Ves00a]. Don [Fra04a]. Donald [WW06]. Dorset [Ba05a, Mar05b, Sau06b, Sau06c, Sch06e]. Dorsett [Tra06b]. Downey [Gve09b]. Dream [BLQ^{+}06]. driven [ARS05, BM09b, GZ05a, Ham05, JR05b, JC09, LLMY05, MBPRR01, Oqu06a, Rie07b, SNEC06, SBCO01, SN05, TR00f, TKA05, ZGL05]. driver [Ric00b]. driver-oracle [Ric00b]. Driving [GBRS06]. DSS [ZMC06]. DTD [BM08a]. during [AHW05, Egy04, RS07, SE04, ZMC06].

Dwight [Pet06]. Dybä [Tra06a]. DynaMetrics [SS08]. Dynamic [BFN^{+}05, BPSH05, EL04, PA05, Rei05b, Bar05b, CO04, DB05, Eng03, FCSM05, Lar06, MP05, MM09b, MACE02, Oqu04a, PE04, Pet05, QXCMW05, RCL05, RKG04, SSH^{+}05, WZX06, Whi00, WHW08, YYX06, YXP09]. Dynamically [SGY05, JAPM05]. Dynamics [BSM04, Fen07]. DynaMine [LZ05b]. dysfunction [Joh00b]. Dyson [Tra04d].

E-learning [ADD04]. Eamon [Fra04b]. Early [NWVO05, SNWV05, Tia00, CPK09, KY09, San09]. ease [JME05]. easing [WM00]. EasyWinWin [GB01]. eclipse [CT05, WW05a, WW05b, YWA05b, Bre06, Che05b]. economic [CEM02]. economical [CdOBT05]. Economics [KS01, KS02, SW05, WS05, Win07c]. ECSAM [Mar05b]. ECUA2001 [Kim02]. EDCS [Sti00]. Edit [Sau09a]. Edited [Ebe09, Cho09]. editing [VDM^{+}06]. Edition [Che05f, Che05g, Hei09b, Law05, Sau05d, BME^{+}08]. editor [Bla00, ACM05b, ACM06a]. Educating [Rei05a]. Education [Els06a, Els06b, Hen05a, Hen05c, Hen06d, Hen08b, Hen08c, Sau05b, Tra04c, Ali06, AH09a, AH09b, AH09c, Bag01, Haz07, Hen03, Hen04a, Hen04b, Hen04c, Hen05b, Hen05d, Hen06a, Hen06b, Hen06c, Hen07a, Hen07b, Hen07c, Hen08d, Hen09b, Hen09c, Kni05, PCP05]. EduLearn [PSK09]. Edwin [BM05d]. effect [CL05, JeMD04b, KON05, KL05, MRR02, SPC09, WBG05, Zel02]. Effective [Che05a, Fen07, Sau02, Ali06, BM01, CKZ05, Ell07, Gil05]. Effectively [ST02]. effectiveness [Ano00c, BCD06, Par05, Sil09, TE04]. Effects [LC05, BS06]. efficiency [HLL02]. Efficient [Att00, PE04, TH02, VJ05, BBG^{+}09, BDE^{+}03, CYD00, Kfo00, PD06, SMK09, TTOH01, XDS04, YCP04]. Effort [Tra04c, APS09, Asu05, HBZ^{+}05, KM06b, KM08]. EFIS [WJH^{+}04]. Egocentric [JR02]. Egon [Har04]. Eighth [TGF09]. EJB [EK01b, GM05]. Election [Ano05a]. electronic [EAG^{+}05, FSN05, Sca05]. element [Win08b]. Elements [Che05e, BM05e]. elicitation [LMG01, UKM02]. Eliciting [PAGDLN05a]. Eliminating [Ba05a, Tee09]. Ellen [WW04a]. Elsevier [Ber06a]. Embedded [BM05d, Mar05b, AMN^{+}06, BMSS09, CMV^{+}08, FLMW09, GR06, Jah01, LW04, Mar07, NI01, PGL^{+}05, PIÖ05, Ste00, TBM^{+}07]. emerging [Cla00b, ST05b]. emperor [dDLS02]. Empirical [BL04, LSH^{+}04, YW04, And04, BCD06, CM05a, CD01, CSTM06, EH04, Joh00b, JMV04, KSNM05, NE02b, SGS05a, SKG08, SCT01, Sil09, SKT05, ZM05, ZD05]. employer [Fra06]. empowering [Enc02]. Enabling [JAPM05, OACL05]. encoding [BDE^{+}03]. End [HBR05, LSBM09, BM05e, CKZ05, KM05a, PH05, RB05c, SSM05, Seg05, Sut05b]. End-User [LSBM09, HBR05, BM05e, CKZ05, KM05a, PH05, RB05c, Seg05, Sut05b]. Endeavors [TR00e]. Endgames [Ba05a]. enduring [Win05c]. energy [KPDdL00]. enforce [ZGL05]. enforcing [AAAG^{+}05, SAM05]. Engine [ACM05d, Joh00a, KDJ07, KDKJ08a, KDKJ08b, SMA05, Ves00b]. Engineering [Ano05a, BOR02, Ber06c, BÁ09a, BV00, BB07b, Cap04, Car07, Car09, CPK09, CSSW01, CRS^{+}05, DHH^{+}06, Doe00a, Doe00b, Doe01a, Doe01b, Doe01c, Doe01d, Doe02a, Doe02b, Doe02c, Doe02d, Doe03, Doe04a, Doe04b, Doe04c, Doe05a, Doe05b, Doe05c, Doe05d, Doe05e, Doe05f, Doe06b, Doe06c, Doe06d, Doe06e, Doe06f, Doe07a, Doe07f, Doe07b, Doe07c, Doe07d, Doe07e, Doe08a, Doe08b, Doe08c, Doe08d, Doe08e, Doe08f, Doe09a, Doe09b, Doe09c, Doe09d, Doe09e, Doe09f, Ebe09, GCL^{+}05a, GFF^{+}08, GWS04, Hei09b, Hen05a, Hen05c, Hen06d, Hen08b, Hen08c, Jah01, Kir01, KS01, KS02, KM05b, KH06, LSBM09, MZ01, Not00, PSS05, PSSS07, RPSS06, Rie05, RP01, Sau05d, Sau06a, TGF09, Tra04c, Tra05]. engineering [ASCP05a, ABF04, Ali05, Ali06, Ano05b, AH09a, AH09b, AH09c, Bag01, Bas07, BZG05, BÁ09b, BM00a, BDM05, BK05b, Cat09, CGL^{+}04, CC05, CLS02, DeL05, DH09, Ell07, GCL^{+}05b, GGR^{+}06, GHS^{+}06, GHP^{+}09, God01, GGZ00, HJ01, HDS05, Haz07, HD07, HW00, HS05, HSGPSF05, Hen03, Hen04a, Hen04b, Hen04c, Hen05b, Hen05d, Hen06a, Hen06b, Hen06c, Hen07a, Hen07b, Hen07c, Hen08d, Hen09b, Hen09c, JW01, JB05, KBBL05, Kni05, KGS09, Lan06, Law04c, LvL04, MK01, Mar05a, MS05, Mei00, MM08b, NTS07, Nam08, NI01, Oqu06a, OCE^{+}00, Ova05, PHP08, Pan09c, PPV09, PCP05, PSS04, Rac00, RJH05, Rec07a, Rec07b, Rho02, Rie06, Rom08, Ros04a, RVR06, SG05, Sau04e, SPH07, SSM05, Seg05, SKM07, SR05, Tem09, Tho01b, TS00, VJT08, WHS01]. engineering [WW00, WBG05, Wen01, Wol04, BFZ^{+}09, CSSW04, DM01, FFvdH01, FFHL05, FFH^{+}05, GZES04, GMAC09, JW01, JMT05, MEF01, SKK01b, THO^{+}01a, Win05d, WJH^{+}04, BM06, Ber06a]. engineers [DA05, Mar04, Rac06, RS07, Rei05a, SGS05a, Sau09b]. Enhanced [SYO09, KDJ07, KDKJ08a, KDKJ08b]. enhancement [BBG^{+}09]. enhancing [Ano00c, Joh00a]. Enough [Sau06b, HB05a]. ensure [XDS04]. ensuring [Ros00]. entailment [KK06]. enter [RB05b]. enterprise [Tra04d, Vog02b, XOWM06]. enterprise-resource-planning [Vog02b]. Entropy [BM08a, CCK04, YW04]. Environment [AMS05, BH05, CRL^{+}06, DGN05, HOST05, Hie04, HJS02, KTB05, LR01b, Man00a, MHB05, NDG05, PGL^{+}05, RCL05, ST02, YMI00]. Environments [GWS04, Fas07, JR02, Mil00, SBCO01]. ENVY [Pen02]. ENVY/Developer [Pen02]. epistemology [Sch09c]. equations [FLY03]. ERCIM [WWB09]. Eric [Bre06, Che05b]. Erich [WW05a]. Ericsson [Pec02]. ERP [ARG00, Dan01, Vog02b]. Error [GW05b, AE05a, DBK01, KAYE04, LZ05b, RSS^{+}04]. errors [CGB06, HJS02, Lev09, XE02]. Erwin [Rac01a]. Esprit [Rug00]. essence [Pie00]. essential [Rho02, WHS01, WLA05]. Essentials [Ana01a]. establishes [KM06b]. Establishing [LGA05, WXL06, WLA05]. estimate [ASCP05b, She05]. Estimating [Ano01b, PH07]. Estimation [BM07a, SKG08, APS09, Ano01c, AE05b, Asu05, BTA06, BM04b, CMPB05, Die06a, Die06b, KM08, MM08b, MM09a, NWVO05, SNWV05, Tra04e]. estimator [JR05b]. ETAPS [Mes01]. eternal [WAK00]. ethnocentric [Fas08]. Eugene [Coo01a]. European [AG01, Con00, Kim02]. EUSES [FR05]. evaluate [Hie04]. Evaluating [Dan01, HSP06, KS09b, LPH02, LPH06, Sut05b, CdOBT05, TE04]. Evaluation [ASCP05a, CG09, CFM05, DF04, GZES04, MM07, ANGRB05, CTZ02, HSGPSF05, KS09a, KDKJ08a, KDKJ08b, KP02a, KGK09, LSH^{+}04, NH09, NE02b, RRSW00, SE04, SAS09, SA09, SCS06, ZM05]. Evans [Ano07d]. Event [dH05, Abr01, HR06, KLS00, LKMU05, Luc00, TR00f, WV04]. Event-based [dH05, HR06, LKMU05]. Events [ACM05a, ACM05d, Tra05, ACM04d, BBM07c, ZMC06, ACM05b, ACM06a]. Everything [Sch07]. Evidence [BB07b, Joh05, BK05b, CM05a, KBBL05, LCLP05, RJH05, SGS05a, WDKM05]. Evidence-Based [BB07b, Joh05, BK05b, KBBL05, RJH05]. Evol [WWB09]. Evolution [GLM^{+}05, Leh00, WWB09, Win09a, BWKG05, BM00b, FORG05, KN05, NFH05, OLHL02, Per06, RRWB00, RK00, Ros01, RW00, Stu05, VPDJ05, WM00, WGS07, YML07, MW01]. Evolutionary [BBHK04, Ton04, GLB04, HFA^{+}01, Sti00]. Evolvability [WWB09, RFG05]. evolvable [BB07a, Bas08, SBCO01]. evolving [Ham05, JAPM05, KS09b, Min00, Rot00]. Ewatch [CG00]. EWSPT’7 [Con00]. examine [JeMD04b]. Examining [WBG05, HJS02]. example [ER05, Fra01, HWM05, Zir07, WW04b]. excellence [Bea04]. exception [FBR05]. exceptions [RM00a]. Exchange [Kie01, SK01b, God01]. exclusion [BBM07b]. execution [AM03, CPDGP01, GK05b, HKO^{+}05, KS06, Pon00, Ric00c, ZG05]. executions [OK05]. exhaustive [SYC^{+}04]. exion [LB05]. existing [LGO04, SMC05a, SWC05]. Expectation [TR00f, YML07]. Expectation-driven [TR00f]. Experience [Bow01, EAG^{+}05, HBW00, ARG00, CNS05, CD01, God01, KKB07, KKN07, LGO04, PJ05, WHS01]. experienced [CSTM06]. Experiences [MVT^{+}02, Pin00, TE04, FBC06]. experiment [LGO04, LS06, TDSCdlR08, YML07]. experimental [DBK01, KDKJ08a, KDKJ08b, OMK04, OYH04, SE04, SBZ00, WC00]. experimentation [DER04, FR05]. Experiments [MK04, CO02]. Expert [ST05a, ASCP05b]. experts [Pen03a]. Explaining [CGS04, MSA^{+}04]. explanatory [JM00]. explicit [LS06]. exploitable [ZLL04]. exploitation [KUST04a, KAYE04, SHRG05]. Exploiting [BQX02, FFQ04, Pan09a]. exploration [Cha05, GDC02, SSG05b, YWA05a, YWA05b]. Exploratory [RHK00, Miš06]. exploring [KMM05]. expressing [CYD00]. expression [SCS06]. expressive [Kfo00]. extensible [IMP05, VJT08]. extension [RGS06]. extensions [PM03]. Extracting [ST00, LZ05a]. extraction [Mar01, SM09b, WML02]. extractor [KN05]. Extreme [GP05, Che05a, Miš06, vD01]. extrinsic [RB05b]. eyes [Pic05].

F [Cha05]. F/OSS [Cha05]. facade

[Pan09b]. Facilitating [BWKG05]. facility [YXP09]. Factors [Gil05, JMT05, BKSM06, HW05, KM05a, LC05, MM09c, NvdBA05, SR05, SKR08]. Factory [Ros04b]. fail [Sch07]. failed [Vog02b]. Failure [Rie07b, WK05, AHW05, CZ02, DLP01, HZ00, WS05, Zei00]. Failure-driven [Rie07b]. failure-inducing [CZ02, HZ00]. failures [AHW05, DLP01, MSA^{+}04]. Fairness [Ban09, Ban07]. FAMILIAR [BA00]. families [BK05a, KFS05a, KFS05b]. family [FORG05, TTK04]. Fan [Sau04a]. FAQs [Not07a, Not07b, Not07c, Not07d, Not07e, Not08a, Not08b, Not09, WW05b]. far [Pin00]. Fascicle [WW06]. FASE [Bag01]. Faster [EJTY04, Sau05d]. Fault [BK09, NE05, BLS02, CL05, GFB05, Par05, PRSH07, Rom07, SKM09, SN05, WAK00, YCP04, Zei00]. fault-tolerant [GFB05]. faults [Mao09, OW02]. FDTs [GR06]. fear [Tra04f]. feasibility [YR08]. feasible [SGAY09]. Feature [CMPB05, MO04]. feature-oriented [MO04]. features [HA00, dCJL06, KUST04a, LKF02]. Feb [Con00]. Federated [WJH^{+}04]. federations [EVC01]. feedback [HDS05]. Fellow [Tra04c, Coo01b, Coo01a, DT06, Omi01]. Felsvlyi [Sau04a]. Fernando [Sau06a]. Few [GG04a, GKS07]. FFP [ZTSR05]. field [BRC^{+}05, Bow05, ELC^{+}02, RCJ^{+}08, RB05b]. fields [Lon04]. Fieldstone [Tra06b]. figures [Not07a, Not07b, Not07c, Not07d, Not07e, Not08a, Not08b, Not09]. file [Dro09, RC01, YMI00]. film [Ova05]. filtering [OOOiM05]. final [HDS05]. find [HSP06, XE02, Sau05b]. Finding [Fen09, JV00, LZ05b]. fine [CCWS02, Ste00]. fine-tuning [Ste00]. Finite [Cla00b, DH00, GGSV02]. finite-state [DH00]. first [CEHK07, Jac00, GZES04, KS01, LA06, MEF01, PTV01]. first-order [Jac00]. Fit [Els06a]. fix [Mar04]. fixes [ŚZZ05b]. flags [BBHK04]. FLAVERS [OC00a, TAC^{+}04]. flaw [WKP05]. flexible [BHLRM05, Hie04, SBCO01]. flight [Jiu07]. FLOSS [CHC05]. Flow [FLY03, Bha05a, DADY04, FBR05, FD00, Li02, MP05, Nau02, Nau03, RVR06, ZYL04]. flows [Fen09]. Fluent [LKMU05]. Focusing [Kni05, RSHJ01, XXJ05]. folklore [Ano05b]. forces [PH05]. foreword [ACM04e]. FORM [GBLM02]. Formal [BA00, BGK^{+}00, GGZ00, KMC^{+}06, Lea00, Sis00, Att00, BRV01, BCD06, Che00, CD01, Eme00, Ham06, HG05, LP05, LQV01, NSAK09, NRSS09, Oqu06a, PAGDLN05b, RC07, RM04b, SSH^{+}05, Wen01, MW01]. formalisms [Cle00]. formalization [BCD00]. Formally [Oqu04c, Oqu06b, MO06, Oqu04b]. format [CRL^{+}06, Kie01, LLL01, Sha00, SK01b]. format-free [CRL^{+}06]. formulas [Ezi04]. forum [Tra02]. Forward [Ano05a]. Fostering [Fas07]. foundation [GHP^{+}09, PHP08, SPH07]. Foundations [BOR02, LSBM09, MW01, Mai00]. four [FSN05]. fragments [YWA05a]. frame [WJS^{+}01]. Frames [CHR04]. Framework [Dev00, Els06a, BNL05, Bas07, BM09b, BCD06, BP05, BDE^{+}03, BX01, CCN^{+}05, CO04, CDF^{+}00, DB09, FBR05, FT05, GCS05, JS00, KJP05, KH05, MP09, Nad00, PV06, PSK09, QXCMW05, RSHJ01, RRCDHM05, RC07, SK01a, Shi04, SMK09, Väh05, WLA05]. frameworks [God01, Jan01, Lu04]. Franek [Fra05]. Frank [Man00b, Coo01b, Tra04c].

Frantisek [Fra05]. Franz [Law04b]. free [CRL^{+}06, IT01, Sca05]. free/open [Sca05]. Friedman [Coo01b]. Friendly [Sau04b]. Front [Sau06c, Tra03, Tra04a]. Front-Loaded [Sau06c]. Frontmatter [Ano05a, Ano05b, Ano07a, ACM05d, Tra04b, Tra04c, Tra05, ACM04e, ACM05c, ACM05b, ACM06a, ACM06b]. fully [Boo00]. Function [Boy01a, Mei04, Pet05, SS09, SBS09a]. Functional [BZG05, Din00, Hud00, Mei04, Ned00, Smi00, XZRA05]. Functionality [Bha05e]. functions [CP05]. Fundamental [Ran01]. Fundamentals [Mad01, Hor01]. fusion [Joh00a]. Future [ACM05a, ACM04d, AH01, AG01, Bha05c, BL04, Gri00, LLL01, Sha02, ACM06a]. fuzzy [ASCP05c, BTA06, LKS00, MB09, MBM09, SBS09b].

GA [BBG^{+}09]. Galen [Ba05a]. Gamma [OLHL02, WW05a]. Garson [BM08b]. Gary [Sau05c]. gcc [Bow05]. Geekonomics [Tra08]. Geeks [Win05a]. gems [Boy01b, Pet06]. genealogies [KSNM05]. genealogy [KN05]. General [Whi00, BDE^{+}03, KUST04b]. Generalized [KKG09]. Generalizing [KS06]. generated [SWC05]. Generating [GGSV02, OX04, RMLMG05, WW06, dFAM06, CJRZ01, MR06, Per06]. Generation [SD01, AMS05, BK09, BM00b, BJL^{+}05, DeM00, EK01a, JM00, Mil00, NE02a, Par05, Raj06, Rob05b, SM09a, SSG05b, Smi00, SPC09, SGAY09, VPK04, WW06]. Generative [But01, GST05, PS00]. Generic [Rob05a, BGPLHS05, CO04, EAN01, FLY03, GKK08]. genetic [BBG^{+}09, SGAY09, Zir06]. gentle [TV08]. Gerald [Tra06b]. Ghinwa [WW04b]. Gilb [Ber06a]. Glass [DT06]. Global [GFF^{+}08, HDLC04, CTZ02, MHK05, Tho01b]. Globus [KTB05]. GMD [Lak05]. go [Sau05a]. Goal [MBPRR01, LvL04, RM06a]. Goal-driven [MBPRR01]. Goals [PCP05, LvL02, RSHJ01]. God [Ada00]. goes [WLCP09]. Good [Sch09b, Rac02, Wol04, dSRC^{+}04]. Goodhack [Rac01a]. Google [Rec07a]. Gottesdiener [WW04a]. governance [DK09]. governed [Min05]. Grace [Law04c]. gradual [AKF06]. grained [CCWS02]. Grand [Sch07]. granularity [WZX06]. Graph [bReWgH^{+}09, NMS09, SPC09, WLF01, ZLX04]. graph-based [NMS09]. graphs [FLY03]. Gray [Che05e]. great [Win05c, Win09a]. greedy [TG06]. Gregory [Che05e]. GREW’07 [GFF^{+}08]. grey [Mao09]. grid [HBR05, KTB05, YEMT05]. grids [SL05]. Ground [Lak05]. Ground-Based [Lak05]. Group [Kim02, BZG05, DFMS09, Sut05a]. groupware [GB01]. groupware-supported [GB01]. Growing [Win06c]. growth [TvH00, Tra04f, Zei00]. GSwERC [Rie08b]. guarantee [BGP06]. GUI [Ano01a, MSP01]. guidance [SBCO01]. Guide [Bea05, Els06b, Sau06a, Sch09b, Ana01a, And00, BM04b, BM04c, Bie02, Boy00, Hub02, Pen01, Sau04a, SSM05, Tra04e]. guided [TAC^{+}04]. guidelines [BM04a, Ram05]. Guilds [Gri00]. Guiney [Fra04b]. GUIs [MPS00]. Gyurky [Fen07].

H [Ano07e, Coo01a]. hacker [Win06d]. Hacking [Win07b]. Half [WHS01]. Hall [Man00b, Sau04d, Sau06a, Sch06d]. Handbook [Tra06a, Law04b]. handling [Ols00, STG02]. hands [Ana01a]. hands-on [Ana01a]. Hard [Cho09, TBM^{+}07]. Hardback [Sau05d, Tra08]. hardcover [Cho09]. Hardgrave [Gla06]. hardware [Cla00a, CMV^{+}08, PC05b]. hardware-assisted [PC05b]. Harmonizing [Ehr05]. harnessing [GGT05]. Haskell [SNWV05]. HATARI [ŚZZ05a]. HCI [GSV05]. head [BM03]. headaches [God01]. healing [GK05b, GZ05b, QXCMW05, SC05b]. Heavyweight [PM03]. held [BK05b, BB07b]. Hell [Win08a]. here [Sau05a, Win06a]. heterogeneous [BFN^{+}05, Chi05, PGL^{+}05]. Heuristic [TAC^{+}04]. Heuristic-guided [TAC^{+}04]. Heuristics [Cle05, GV02]. Hibernate [Els06b]. Hidden [Gog00]. hiding [SGS^{+}05b]. Hierarchical [LWZ05]. hierarchically [COWL02]. hierarchy [YF00]. High [BB07a, Car07, You00, Har04, Hub01, MK04, Rom08, Sau04d, Tra04d, WBH^{+}05, Zir07]. high-capability [Tra04d]. High-level [BB07a, Har04, WBH^{+}05]. high-return [Sau04d]. higher [BJ05, Oqu04a, WC00]. higher-level [BJ05]. higher-order [Oqu04a, WC00]. Hiring [Sch06e]. historical [CR06b]. histories [HmL05, LZ05b, ZG05]. History [Chr08, WW06, RG06]. hoc [JR02, RM00b]. home [Mar04]. homogeneous [SPS^{+}08]. House [Ba05a, Mar05b, Sau06b, Sau06c, Sch06e, Tra06b]. HPC [Car07]. Huang [Sau04d]. Human [Cho09, KM05a, WW02, Aik04, Cap04, Joh05, KON05, Win08b, JMT05]. Human-Computer [Cho09, WW02, Joh05]. Humphrey [BM07b]. HW [CBB^{+}07]. HW/SW [CBB^{+}07]. hybrid [MLG05, RM06b]. hybridization [RRR04]. hypermedia [APS09].

ICSE

[CSSW01, CSSW04, DM01, KS01, KS02, MEF01, MZ01, PSS05, PSSS07, RPSS06, Ano05a, BFZ^{+}09, BB08a, BK05b, BB07b, CPK09, EL04, GK05a, KKS01, PA05, PSS04, Tra04c, dLGR05, dLR05, ACM05c]. ICSE2001 [Tho01b]. ICSM [ACM05c]. iDAVE [HB05b]. identification [Bra01, Cle05, RRCDHM05, RGB05, SYO09, WZX06]. identifiers [OG05a]. identifying [ZRWS05]. IDL [Bra01]. IDL/ADL [Bra01]. IEEE [BM06, BM08c, Ber06c, CLS02, Fen07, Sau06a]. Im [Ber04]. image [LN05, Shi04]. image-oriented [Shi04]. Impact [AK09, BRC^{+}05, ELC^{+}02, RCJ^{+}08, EH04, EAS07, JeMD04a, KB05, KM06a, LPH06, OCE^{+}00]. impacts [HDS05]. Imparting [Ali06]. implement [SD01]. implementation [ARG00, CLS00, Din00, GSX^{+}04, Gup00, Nau03, Ols00, Rey00, WJS^{+}01, XYH04]. implementations [JS01]. Implementing [GA01, WV04, FCSM05, FHK^{+}01]. implication [YWA05b]. Implications [KS01, KS02, CSR06]. Implicit [Chi05, WM00, LZ05a]. implied [UKM01, UKM02]. Importance [Gil06]. important [GG04a]. imprecise [NvdBA05]. improve [CMPB05, CKFS01, HBZ^{+}05, SKT05]. Improved [KM06b, SKM07]. Improvement [BM05a, Tra06a, US05, Boy00, Bro06, FD07, Joh00b, LSK^{+}05, PB05, Ric09, Tee09, Tia00, VGL05, Wen01, XN09, Law05, BM04a]. Improving [AAAG^{+}05, KK06, LE04, MACE02, PB05, RFG05, Red00, SA00, YR08, BS00, Sai09, Spr05]. in-car [BHLRM05]. in-process [NWVO05, SNWV05]. INCA [SA00]. incentives [RB05b]. incidents [MHK05]. Inconsistency [EC01, XC05]. incorporating [VJT08]. Incorrect [Sau05b]. increase [SHRG05]. Increasing [JS01, RM01]. Incremental [Ves00b, PE04, Väh05, Leh00]. incrementalism [GGT05]. Independent [MV05, SMCH06]. index [Zir06, ACM04c]. indicators [MLRW05]. induce [ŚZZ05b]. inducing [CZ02, HZ00]. inductive [Che02]. Industrial [Lan06, PJ05, Ber02, Enc02, Fra02b, FHK^{+}01, OW02, Pec02, Rom08, RCJ^{+}08, Sre02, Wil04, WLA05]. industrial-strength [Fra02b]. industry [Har02, JeMD04b, Lak05, LCS^{+}05, Rei05a]. inference [BA00, NE02b]. infinite [Che06]. influence [Haz07, HW05]. Influencing [BKSM06]. Information [ACM05d, SGS^{+}05b, Tra05, WJH^{+}04, AH01, ARV05, BMW01, CCK04, CO04, CP05, Eva00, Hee03, Kai00a, Kai00b, KM06b, KM06c, KM06d, Law01a, Li02, MP05, SV00, Wen01, YFR00, ACM06a]. information-flow [Li02]. informations [EAN01]. infrastructure [DER04, Lar06, LGA05]. inheritance [AK09]. inherited [YK09]. initial [PB05]. initiative [KAKM05, Rie08b]. inject [BMMG05]. injection [HO05]. inner [Tee09]. Innovations [Cam05]. input [HZ00, SK00, VPK04]. input-output [SK00]. inquiry [SYO09]. Ins [Bre06, WW05a, Che05b]. Insecure [Tra08]. INSERT [Leh00]. inspection [Ada00, Zir06]. inspections [RCJ^{+}08, SKT05]. Inspired [San09, TG06]. instability [JeMD04a]. Institute [Sau04a, Sau04b]. Institutional [Rie07c]. Instruction [AM03]. Instruction-level [AM03]. instrument [KP02d]. instrumentation [CO04, Ric00c, TH02]. intangibility [Haz07]. Integrated [Els06a, Law05, CTZ02, Kar00, MV01, RGB05, SMD^{+}05]. Integrating [Che00, JRHD00, PSK09, YFR00, BA00, Coo00, PSSA06]. Integration [Jiu07, BFN^{+}05, BHLRM05, BM04a, DPT05, FFPT05, GBLM02, HZ04, HIM00, MM09b, MG00, Mun00, Sch08a, Sch08b]. integrity [Egy00, Ros00]. Intel [Ano07b, Bea04]. Intellectual [KPDdL00]. intelligence [Ben00]. Intelligent [ADD04, JW01, Law02]. intensive [ASD03, CO00, GW05a]. intent [AAAG^{+}05]. inter [Ehr05, TK06]. inter-component [TK06]. inter-object [Ehr05]. interaction [BČVZ06, BC05b, GB08, Joh05, KAKM05, NS09, WW02, dFAM06, Cho09]. interactions [DGP^{+}01, HA00, RRR04, SHRG05]. interactive [Bre05, GJJL00, PP01, Pen03b, Red00]. interchange [HZ04, LLL01]. interconnection [GMS01]. interest [Chr08, DFMS09]. Interface [HLL02, dAH01, Fas08, GG04a, LBR06, Pen03b, SM00]. interfaces [HJM05, SMD^{+}05, SGS^{+}05b, WML02]. interleaving [RW00]. intermediate [Sha00]. internal [KS09b]. International [BOR02, BKW04, Car07, Car09, CFM05, CHR04, CRS^{+}05, DHH05, DMS09, EC01, EL04, GZES04, GFF^{+}08, HHM05, JW01, KS01, KS02, PTV01, PRRW04, PA05, WLK04, WWB09, DHH^{+}06, DF04, vdH01]. Internet [MZ01, Tra04d, Vog02a, Kai00a, Kai00b, TvH00, YML07]. Internet-based [Kai00a, Kai00b]. internetwares [WXL06]. Interoperability [SMC05b]. interoperation [MT00b, PWZ^{+}04]. interpretation [Sch00b]. Interpreting [BM05a]. interprocedural [Hin00]. interrelated [CRH01]. Interscience [BM07a, BM08c, BM09a, Ber06c, Hei09b, Hen09a]. Interviewers [Ano05b, Tra05]. interviews [Bro06]. Intl [PSS05, PSSS07, RPSS06, WGM06]. intra [Ehr05]. intra- [Ehr05]. Intractable [Vog02b]. intrinsic [RB05b]. introducing [EAG^{+}05, SCT01]. Introduction [Law05, Sau09b, NSAK09, TV08, Var00, FV09a]. Invariant [NE02b]. Invariants [Ham06, Min00, PE04]. investigate [JeMD04a]. Investigating [BLS02, NRSS09, RB05a]. Investigation [KUST04a, KL05, Rob05b]. invocation [Ols00]. involvement [vD01]. IOStreams [Pen01]. IP [Sch09a]. IPv6 [Law06]. irregularities [Cle05]. ISB [FV09b]. ISB-10 [FV09b]. ISBN [Bea05, BM08c, BM09a, Ba05a, Ber06a, Ber06b, Ber06c, Cho09, Ebe09, Els06a, Els06b, Hei09a, Hei09b, Law05, Man00b, Mar05b, Pet06, Sau05b, Sau05c, Sau05d, Sau06a, Sau06b, Sau06c, Sau08, Sau09a, Sau09b, Sch06c, Sch06d, Tra06a, WW05a, WW05b, WW06]. ISBN-13 [BM08c]. ISE3 [JW01]. ISO [BM08c, Bha05e]. Isolating [CZ02, Zel02]. isolation [BLS02]. ISSTA [Ber02, Enc02, Har02, Pec02, Sre02]. Issues [LG02, WBH^{+}05, AKL^{+}07, Bag01, BBM07b, JW01, MV05, MM08a, MM09a, Mor03, Pic05, PRSH07, ST05a, SMC05b, Ter04, XGY02]. IT-security [NKB05]. Italy [Pec02]. iterative [Väh05]. iterator [Lon04]. Iverson [Els06b]. IW [DHH^{+}06]. IW-SOSE06 [DHH^{+}06]. IWLWI01 [EC01].

J [Sau06a]. J2EE [Els06b, BG03, NC04]. Jackson [San09]. Jacky [FV09b]. Jakarta [Bat00]. Jalloul [WW04b]. James [BM08b]. Jan [Law04a]. Jane [Sau04d]. Jas [Sau05c]. Java [Pet06, Ana01a, BPSH05, BKM02, CR06a, Dil00a, EK01b, FRMW04, GV02, KLS00, Lar01, LBR06, Li02, LPH06, Mad01, MRR02, Nau02, OMK04, PV06, RM00a, RKG04, Sha01, VPK04, dH05, Gla06]. JavaBeans [XOWM06]. JavaScript [Doe06a]. Jayaswal [Sch06d]. Jeff [FV09a]. JIL [OC00b]. JIT [THL03]. JIT-compiler [THL03]. JML [LBR06, TE04]. JML-JUnit [TE04]. job [JeMD04b, Win05b]. jobs [KGS05]. Johanna [Sch06e]. John [BM08c, BM09a, Ber06c, Che05a, Hei09b, Law06, Sau04e, Sau05d, Tra04d, WW05b]. Johnson [BM05a, Gla06]. Joining [FJM05]. joinpoint [AX06]. Jon [Ano07b]. Jonah [Mar05b]. Jonathan [FV09b]. Joseph [Mar05b]. jRapture [SCFP00]. JTS [Bat00]. JUnit [Rie08a, TE04]. Jupiter [Are07]. Just [RM05, Wol04, Sau06b]. Just-in-time [RM05].

Kai [Sau09b]. Kaprun [Con00]. Keenan [Sau04a]. Kent [BM05a, WW05a]. KenyaEclipse [CT05]. kenyon [BWKG05]. key [Lon04]. Keynotes [ACM04d]. keywords [OG05a]. kind [DH09, GSX^{+}04, Rac00]. King [Sau09a]. kit [RN00]. Klop [Law04a]. Kluwer [Tra06a]. know [Ova05]. Knowledge [RM04a, Sch06e, TR00g, AKL^{+}07, ALK09, CD01, Far06, GL06, JMV04, OOOiM05, RRCDHM05, SHRG05, SvdVD06, dB06, LA06]. knowledge-based [RRCDHM05]. Knuth [WW06]. Konrad [BM04a]. Konstantina [Sch09a]. Korat [BKM02]. Krispies [Win07a]. Kudish [Mar05b]. Kulak [Fra04b]. Kulpa [BM05a]. Kutnic [Bea04].

L [Cho09, Coo01b, Hor01, Law04b]. lab [Pec02]. Laffra [WW05b]. Laird [BM07a]. Lan [Ber06c]. LANCET [VDM^{+}06]. Land [BM08c]. Language [Bil03, Oqu04a, Ana01a, CR06a, CO00, CL06, EAN01, FBHZ08, HG05, LBR06, Mai00, MO06, Ols00, Oqu04b, Ric00d, Sce00, SPT05, TS00, WLF01, YXP09]. language-based [Sce00]. Language-Dependent [Bil03]. languages [Ber06d, Bur00, Cla00d, Dev00, Din00, DH00, Kfo00, Nam08, NH09, NSAK09, PD06, Rey00, WC00]. Large [Fen07, GCL^{+}05a, Sch09a, CGL^{+}04, EV05, GCL^{+}05b, GGR^{+}06, GFB05, HRGB05, LZ05a, OG05a, OSH04, OW02, RB05a, Sch05a, Sch05b]. Large-Scale [GCL^{+}05a, Sch09a, CGL^{+}04, GCL^{+}05b, GGR^{+}06, GFB05, RB05a]. Latency [Ram00b]. Latest [Tra04c]. Lavi [Mar05b]. Law [Min05, Fas08]. Law-governed [Min05]. Lawrence [BM06]. layered [JR05a, KS09c, RDS05, XOWM06]. layout [MHB05]. LCMS [PSK09]. leadership [Rac06, Tra02, BB08a, BB09]. Leading [Tan02a]. leak [XA05]. Lean [Pen04]. LEAP [Joh00b]. learn [Dek05]. learned [LLL01]. Learning [SC05a, ARG00, ADD04, BRH04, CT05, Ell07, HmL05, KKB07, Pan09a, PSK09, RCJ^{+}08]. Learning/organizing [SC05a]. learnings [KKN07]. Lecture [Ebe09]. Leffingwell [Fra04a]. legacy [JRHD00, Law04c, MG00]. Legitimate [HmL05]. lemonade [PK01]. lemons [PK01]. lessons [LLL01]. Letter [ACM05b, ACM06a]. Letters [Ano05a, Ano05b, Ano07a, ACM05d, Tra04c, Tra05, ACM05c, ACM05b, ACM06a, Tra03, Tra04b, Tra04a]. level [AM03, BJ05, BB07a, Ber04, DSRJ04, Har04, RM09, WBH^{+}05]. Leveraging [BMMG05]. Lewis [Law04c]. Leymann [Man00b]. libraries [MSR02]. library [Cha05, KS06, NRSS09, ST00]. library-based [ST00]. libre [ABNR05, BRC^{+}05, HRGB05]. life [KS09d, Ova04]. life-cycle [Ova04]. lifecycle [Spr05]. Lightweight [BA05b, Joh00b, PJ05]. likely [PE04]. Limits [Ano01c, Sch05a]. Linda [BM07a]. Line [GMAC09, ABNR05, Cat09, CO00, DL05, GA01, GBLM02, MK01, NI01, PJ05, SG05]. Linear [ARV05, BD03]. Lines [KS02, ASCP05b, Ano05b, But01, ELC06, KJP05, KS01, KMM05, MT01, OG05b, SK01a]. Linguistic [Rie06]. Link [PD06]. Link-time [PD06]. Linux [SC05a, XMP08, YR08]. liquid [BD03]. Little [OC00b]. Little-JIL [OC00b]. live [AMN^{+}06, ZBPD01]. liveness [NC00]. Living [EC01]. Liz [Sau05c]. LMSA’08 [BB08a]. Loaded [Sau06c]. locales [Pen01]. localization [LYF^{+}05]. located [Dek05]. Locating [PAGDLN05b]. Logic [BM08b, Ezi04, GDC02, Hol02, Jac00, Law04b, LKMU05, MB09, MBM09, NR00b, RRSW00, Ram00a, Swi00]. logic-based [NR00b, RRSW00]. logical [BA00]. logics [Ehr05]. Lonesome [Win06d]. long [Mas05, Väh05, ZBPD01]. long-term [Väh05]. long-timescale [Mas05].

Longitudinal [Mas05, Not03]. Longshaw [Tra04d]. look [Fas08]. looming [Rom07]. loop [BBHK04]. loop-assigned [BBHK04]. LORs [PSK09]. love [Win07b]. Low [KCS06, Sau04d]. low-risk [Sau04d]. Lowe [Sau05c]. Ltd [Law06, Tra04d].

M [BM06, BM07a, Law05, Sau06b, Sch06c, Tra06b]. Machine [HLL02, FHNS02, Jan01, SKM09]. Machines [BE05, GK05a, WGM06, ARS05, GGSV02, Har04, Law02, Ler04]. MACS [Rob05c]. made [Sch07]. Madhur [Sau05c]. maintainability [HW05, MB09]. maintaining [GKS07, Man00a, Rot00]. Maintenance [BM09a, BSM04, BKSM06, Die06a, KH05, MM08a, MBM09, RS07, SG07, SBS09b]. Making [FFvdH01, LW04, Mye06, KWD05, OYH04, Sch06b, Wil04]. malicious [SYO09]. malware [CJ04]. MAME [YMI00]. man [Sch07, VS08]. man-made [Sch07]. man-rated [VS08]. manageability [MW05]. managed [AKF06, GK05b, KJP05]. Management [BB08a, BB09, Ben02, Fen06, Sau06b, Sch09a, vdH01, Abr01, APS09, BA00, Ber08, CCHS09, CCS00, CCWS02, DMNJ06, DB05, ELC^{+}02, FHK^{+}01, GvdHT05, Gil05, Kai00a, Kai00b, KS09c, KS09d, LSK^{+}05, MZY01, MO04, PSK09, Rie07c, RDS05, Sau05a, Sha01, TKA05, UT05, Var00, Ano07b, Ano07d, Ano07e, BM09a]. manager [Bie02]. Managing [BB01, CC05, DA05, Hen08a, Pan09b, VS08, CNPD06, EVC01, GBRS06, Ano07c, Fra04a, Tra04f]. MANETs [SHB05, SHRG05]. Manfred [Ebe09]. manifesto [Sch08a]. Manipulating [GK05b, FBS04]. manipulation [FLY03, HOST05]. manufacturing [Tra04f]. many [SGS03]. Mapping [AMN^{+}06, LLT05, ZYL04]. maps [Joh05]. Maraia [Fen06]. Marc [Law06]. Margaret [BM05a]. Margus [FV09b]. Mark [Law04a, Sau04d]. Market [PH05, ELC06, TvH00]. Marketing [Sau06b]. marking [YWA05a]. Marmoset [SSHP05]. Marquardt [Ebe09]. Mary [BM04a, Pen04]. MAS [BGPLHS05, BQX02]. Master [Fen06]. Mastering [Pen02]. Matching [GW05a, ZG05, NFH05]. material [ACM06b]. math [Fra06]. Mathematical [Sau09b]. Matlab [NE05]. matter [Tra03, Tra04a, ACM03, ACM04a]. maturity [Ber04, Sch09d]. Maude [MT00a]. maverick [SR05]. MBASE [BM00a]. MCDC [Bha07]. McGuinness [Law04b]. McMillan [Che05c]. MDA [BM09b]. means [KvB04]. measure [KFS05b, KKG09, KM06b, KM06c, KM06d, MM07, Mis07]. Measurement [BM07a, MK06, MK07, Sch09a, Bha05a, Boy01a, HBZ^{+}05, Joh00b, Pen03a, SCT04, Tia00, TWQ09, Sau04e]. measurements [Pan09c]. measures [CG09, KL05, MK06, MK07, ZLX04]. Measuring [DBK01, GKK08, Pet05, WGS07, WY04, YK09]. Mecano [GBLM02]. mechanism [BB07a, JSS01, SC05b]. mechanisms [BP05, FR05, Ros00, SSH^{+}05, VJT08]. mechatronic [GBSO04]. Mechatronics [Law02]. MediaDoc [JM00]. mediation [BMW01]. medical [CBB^{+}07]. meetings [Dek05]. Meets [Sau06b]. Meister [TS05b]. memory [KON05, Rie07c, WC00, XA05, XDS04, Fra05]. Merging [PC05b, UC04, DMNJ06, SNEC06]. message [BD03, BLQ^{+}06, SC02, UKM01]. message-oriented [BLQ^{+}06]. messages [AE05a]. MetaH [Ves00c]. metaheuristic [HJ01]. metal [CEH03]. metamodel [BGPLHS05, PM03]. metamodelling [HS05]. metamorphic [CTZ02]. metaphysical [Sch08a]. metaproblem [Sch05a]. Method [DRL^{+}05, HSGPSF05, Mar05b, Ros04b, CYD00, CTZ02, CKZ05, Har04, HS05, Jiu07, KGS05, Mar01, MT01, NRSS09, Oqu06a, SYO09, SMD^{+}05, Tee09, WZX06, Wor02, YK09, Oqu06a, Tra06b]. method-extraction [Mar01]. methodologies [BQX02, FLMW09, HS05, Rot04]. methodology [CBB^{+}07, CMV^{+}08, EKHG01, GB01, MS05, PAGDLN05a, Rob05a, SCT01, WBG05]. Methods [Cho09, Hen09a, Att00, BRC^{+}05, Eme00, Ham06, LP05, Lea00, RGB05, Rom08, San09, Sil09, Sis00, WHS01]. Metric [KM09, BM08a, GG04a, KM06b, SS08, SBS09a, YW04]. metric-based [SS08]. Metrics [BV08, CS09, KFS05a, SDF00, BM04c, Bha05e, Dan01, GB08, KS08c, KM06a, LRBH01, MM08b, NWVO05, RR09, SNWV05, YMI00, YWG04]. Metrics-based [BV08]. MIC [Kar00]. Michael [Ano07d, Che05c, Fen07, Sau04b]. micro [CEM02]. micro-economic [CEM02]. micromodularity [JSS01]. microprocessor [DBK01]. Microsoft [Che05d, Fen06, Das06]. Midcourse [Lak05]. middleware [BLQ^{+}06, DB09, EAS07, MG00, RC07, XC05]. Migrating [Law06]. migration [KLS00]. Mike [BM04a]. millennium [WW02]. Miller [Tra04f]. mindset [Tra04f]. Miner [LZ05a]. minimization [TG06, TGSP05]. Mining [DHH05, FORG05, Gve09a, HHM05, HmL05, MLRW05, AX06, ARV05, GCS05, HDS05, KCM05, LZ05b, OG05a, SSHP05, VGL05, YWA05b]. mirage [KS09a]. Miscellaneous [ACM06b]. Mischa [Sch06f]. Misfeldt [Che05e]. mishaps [MHK05]. mission [PH05]. mission-critical [PH05]. mitigation [VS08]. ML [Cla00c, Gun00]. Mobile [Sta01, BFN^{+}05, BH05, CEM02, Cap04, JAPM05, JR02, LMSF04, Lu04, Oqu04a, SMK09, TTOH01, YYX06, dFAM06, Ber06b, Sch06f]. mobility [LFW02, RP01]. Modal [BM08b]. mode [LE04, SN05]. mode-driven [SN05]. Model [BM00a, Dil00a, FBS04, GV02, GDC02, HZ04, Kar00, Lak05, NS09, OG05b, ASCP05b, ASCP05a, APS09, Amm00, BK09, BRV01, Bas08, BS00, BM09b, BMSS09, BQX02, Bul00, CRH01, DJP05, Eme00, FLMW09, FW06, GBLM02, GM05, GSX^{+}04, HCGGXFY00, HFH00, HN04, Hei00, HG05, dCJL06, KTB05, KM09, KGK09, KM06a, LLL01, LWZ05, LYF^{+}05, MZY01, MM09b, MM09a, NRSS09, NAD02, Oqu06a, PC05a, Par05, Raj00, RRSW00, RM06b, RGS06, RMLMG05, RRU05, Say05, SKM07, SSH^{+}05, SNWV05, SBS09a, SBS09b, SPS^{+}08, SM09b, UT05, UCKM04, VCST05, Wan05, WHW08, XN09, XYH04, ZGL05, ZD05, EK01b, FV09b]. Model-based [BM00a, Lak05, NS09, OG05b, BMSS09, DJP05, FLMW09, LWZ05, LYF^{+}05, NAD02, Say05, FV09b]. model-checkers [FW06]. model-driven [BM09b, Oqu06a, ZGL05]. Modeling [BFZ^{+}09, Ban07, Che05a, DS05, DB05, LN05, LJD05, Mar05b, MDC04, Rob05c, Sau09b, BB07a, BG09, BB08b, Boy02a, CR06a, DA01, Die06b, Die07a, FBHZ08, HOST05, KP05, KY09, LS06, Mil00, Pan00, Pen03b, RM05, RM04b, SS09, SvdVD06, XOWM06, Zei00, PRRW04]. Modelling [Ana00, EVC01, KvB04, JS00, LQV01, LMSF04, Oqu04b, Oqu06b, PGL^{+}05, Sch00b, WK05]. Models [Bag09, BE05, CFM05, DF04, GK05a, WGM06, AH01, ASCP05c, Asu05, AMS05, Aug09, Ban08, BMMG05, Bar05b, Che02, Coo00, EJT04, Egy00, EKHG01, GKKW05, HRGB05, KS08a, KJML05, KLS00, LM08, Ler04, LKMU05, LSH^{+}04, MR06, Mis00, PH07, SSG05b, UC04, UK06, Wen01]. Modern [Sau09a, Rom01]. Modernizing [Law04c]. modified [KM06c, Mis07]. Modular [FK01, GBSO04]. modularity [BS00, CS05, CKFS01, DNR06, SGCH01, YR08]. Modularization [GGH^{+}07]. modularly [KFG04]. module [ASC02, KL05]. Moe [Tra06a]. MonDe [CO06]. monitored [CO06]. Monitoring [BOH03, CO00, HO05, KCS06, TR00f]. monitors [PP00]. Monte [Sau04a]. Moose [DGN05, NDG05]. Moraine [YMI00]. MOST [DJP05]. MPEC’04 [DF04]. MPEC’05 [CFM05]. MSF [ZHG05]. MSR [DHH05, HHM05]. MTMOOD [KM09]. much [HB05a]. Mugridge [Els06a]. Multi [GCL^{+}05a, THO^{+}01a, BC05a, BBG^{+}09, CGL^{+}04, DL05, Ehr05, GCL^{+}05b, GGR^{+}06, GFB05, GM05, JR05a, KG05, LE04, Min05, OACL05, RDS05, WBSM00]. Multi-Agent [GCL^{+}05a, BC05a, CGL^{+}04, DL05, GCL^{+}05b, GGR^{+}06, GFB05, JR05a, KG05, Min05, OACL05]. Multi-Dimensional [THO^{+}01a, BBG^{+}09]. multi-layered [RDS05]. multi-mode [LE04]. multi-object [Ehr05]. multi-server [GM05]. multi-site [WBSM00]. multi-variable [BBG^{+}09]. Multiagent [DeL05, MHB05]. multicore [PPV09]. multimedia [JM00]. multimethod [Lea00]. multiple [TBM^{+}07, dSRC^{+}04]. multithreaded [BPSH05]. multithreading [PV06]. Munson [Sau04e]. MUS [PAGDLN05a]. Musa [Sau05d]. mutation [Amm00, OMK04]. mutual [BBM07b]. MyFaces [STB08].

N [Sau09a]. Nagl [Ebe09]. Naming [BM05e]. Nardi [Law04b]. Nathan [Sau06c]. National [Win05d]. Naveda [Sau06a]. near [Con00]. Nearest [Boe05b]. necessary [Pin00]. need [Asu05, KS09a, KAKM05, Ova05, Wil04]. Needed [Tra05, Ano05b]. Needham [Omi01]. needs [WW04a]. Negative [UKM02]. negotiation [GB01]. neighbor [Boe05b]. Nerds [Sch06e]. Net [BNL05, ASKS04, Doe00a, Doe00b, Doe01a, Doe01b, Doe01c, Doe01d, Doe02a, Doe02b, Doe02c, Doe02d, Doe03, Doe04a, Doe04b, Doe04c, Doe05a, Doe05b, Doe05c, Doe05d, Doe05e, Doe05f, Doe06b, Doe06c, Doe06d, Doe06e, Doe06f, Doe07a, Doe07f, Doe07b, Doe07c, Doe07d, Doe07e, Doe08a, Doe08b, Doe08c, Doe08d, Doe08e, Doe08f, Doe09a, Doe09b, Doe09c, Doe09d, Doe09e, Doe09f, Jah01, Not00]. net-centric [Jah01]. nets [AMN^{+}06]. network [ASCP05a, ASCP05c, BGK^{+}00, HR06, SAS09, XGY02, XOWM06]. network-centric [XGY02]. networked [MN05]. networks [DeM00, KUST04b, OOOiM05, Ric09, RM00b, Hei09a, Sch09a]. Neural [Hei09a, ASKS04, ASCP05a, ASCP05c, KUST04b]. news [ACM05c]. next [DeM00]. nifty [VDM^{+}06]. nightmares [Win08a]. Nils [Tra06a]. Non [CR06a, Ros04a, SPS^{+}08, BD03, Lon04, XZRA05]. Non-directive [Ros04a]. non-functional [XZRA05]. non-key [Lon04]. non-linear [BD03]. Non-null [CR06a]. nondeterministic [NVS^{+}04]. nonfunctional [DB05]. Norton [Sau09a]. notation [DGP^{+}01, KGS05]. notations [LQV01, NAD02]. Note [Ebe09]. Notenboom [BM05d]. Notes [Doe00a, Doe00b, Doe01a, Doe01b, Doe01c, Doe01d, Doe02a, Doe02b, Doe02c, Doe02d, Doe03, Doe04a, Doe04b, Doe04c, Doe05a, Doe05b, Doe05c, Doe05d, Doe05e, Doe05f, Doe06b, Doe06c, Doe06d, Doe06e, Doe06f, Doe07a, Doe07f, Doe07b, Doe07c, Doe07d, Doe07e, Doe08a, Doe08b, Doe08c, Doe08d, Doe08e, Doe08f, Doe09a, Doe09b, Doe09c, Doe09d, Doe09e, Doe09f, Not00]. notices [Tra03, Tra04b, Tra04a]. novel [BBG^{+}09]. novice [CSTM06]. NPIC [Hin00]. NSF [Eva00, TS00]. Nu [DNR06]. Nucci [Sch09a]. null [CR06a, HSP06]. numbers [Sau04d, Win06b]. numeric [CRH01]. Numerical [Hen09a]. Nurtures [Sau06c]. NYI [Bur00].

Object [BME^{+}08, Che05f, Che05g, CSTM06, DG00, KUST04a, KUST04b, Pen03b, WW00, BA05b, BQX02, BB08b, BLS02, BRBY00, Cac00, CYD00, Che00, CG09, DNR06, EJT04, Ehr05, EKHG01, HCGGXFY00, HFH00, JRHD00, KE01, KM09, Lee09, MRR02, MM08b, MM09a, NMS09, Nam08, Off00, PSK09, PND01, PD06, Ric00a, Ros00, SYO09, Sai09, SA09, SS08, SP00, ST00, Tai00, VJT08, WML02, Wor02, YF00, TDA07, Che05c, Che05d]. Object-Oriented [DG00, KUST04a, BME^{+}08, CSTM06, WW00, BQX02, Cac00, Che00, CG09, EKHG01, HFH00, MM09a, NMS09, Nam08, Off00, PND01, PD06, Ric00a, SS08, SP00, ST00, Tai00, VJT08, WML02, Wor02, YF00, Che05c]. Object-Z [TDA07]. objects [Ano01a, Ben00, Che06, CL06, LGO04, Pie00, ZXZ01]. observation [SCFP00]. observation-based [SCFP00]. Observations [SDD05, MM08b]. observer [Nau03]. observing [SM09b]. OCL [SAM05]. Off [CFM05, LCS^{+}05, RM01, Tan02b]. Off-the-Shelf [CFM05, LCS^{+}05, RM01, Tan02b]. office [Ova05]. Official [WW05b]. Offutt [FV09a]. Old [Pic05, BK01]. Oliveira [Sch09b]. OMEGA [CO00]. OMEN [SP00]. on-line [ABNR05, CO00]. On-Time [Ba05a]. one [GLB04, Van00b, NSAK09]. one-step-ahead [GLB04]. one-way [Van00b]. Online [VCST05, GBSO04]. online-reconfiguration [GBSO04]. only [Cla00b, Dil00a, Pin00, Wit00]. ontology [BMW01, GL06]. ontology-based [GL06]. OODPM [Dro09]. OOP [Mad01]. OOPSLA [SKK01b]. Open [DB02, FFvdH01, FFHL05, FFH^{+}05, KH05, RB05a, ACM05d, ZD05, Asu05, BB07a, Bas07, BG09, GGT05, GGH05b, JW01, LG02, LKF02, MV05, Mas05, MT00b, Rob05a, ST05b, Sca05, SM00, SDD05, Sto02, WGS07, ZE00, ZLL04, FFH^{+}05]. open-source [GGT05]. open-system [BG09]. OpenEC [Sca05]. OpenEC/B [Sca05]. operating [CKFS01]. operational [Bea04, Bis02, CJRZ01, FD00, LvL02]. operations [MDC04]. Operative [DRL^{+}05]. Opportunistic [SHRG05]. opportunities [KJML05, MV05, Sta01]. Optimal [NVS^{+}04, Din00, RMLMG05, SPS^{+}08, TGSP05]. optimization [ANGRB05, YF00, Zir06]. optimization-based [ANGRB05]. Optimizing [Suk01, WC00, Ezi04]. Options [BNL05]. Oracle [ASKS04, Ano01a, Ric00b]. oracles [CJRZ01, Dil00b, MPS00]. ORBIT [RK00]. ORBIT/VIRTUE [RK00]. orchestration [KvB04]. order [DA05, Jac00, Min05, Oqu04a, WC00]. Ordering [BBM07c, Ban08]. organisational [BC05a]. organization [DeL05, KON05, PND01, ACM04e]. organization-based [DeL05]. organizations [BQX02, Sch05a, Sch05b]. organizing [SC05a]. orientation [BB08b, Mei00]. Oriented [CPK09, DG00, DHH^{+}06, KUST04a, AX06, BV08, BB07a, Bas08, BG09, BLQ^{+}06, BQX02, BME^{+}08, BLS02, BČVZ06, Cac00, CS05, Che00, CG09, CSTM06, DS06, EKHG01, GB08, GKK08, HFH00, HS05, KUST04b, KM09, KS08c, KH01, KGS09, KKG09, KGK09, Lee09, MO04, MM08b, MM09a, NMS09, Nam08, Off00, Pan00, PND01, PD06, RR09, Ric00a, RSB04, RB05b, Sai09, SA09, Shi04, SS08, SP00, SGS^{+}05b, ST00, Tai00, VJT08, WZX06, WW00, WML02, Wor02, YF00, Che05c]. OSGi [TV08, WHW08]. Osiris [MR06]. OSS [Cha05, PGL^{+}05, ACM05d]. OSSmole [CHC05]. other [STB08]. OTS [CC05]. OTS-aware [CC05]. Our [ARG00, KKB07, KKN07, Win05b]. output [Mil00, SK00]. outsourced [BKSM06]. outsourcing [Kei05]. overflows [ZLL04]. overhead [KCS06]. Overview [SWC05].

P [Ano07e]. P-H [Ano07e]. P2P [PSK09]. packages [Bra00]. packaging [SBZ00]. page [BM05e, Shi04, ACM04e]. page-transition [Shi04]. pagination [STB08]. painful [Win06c]. pair [Aik04, KON05]. pairwise [BC05b]. PalME [VWW02]. Paltz [Hin00]. Pandian [BM04c]. Panel [Sre02, Ber02, OCE^{+}00, ZBPD01, Enc02]. Papagiannaki [Sch09a]. Paper [ACM05a, ACM04d]. Paperback [Hei09a, Sau05d, BM09a, Che05b, Che05f, Che05g, Cho09, FV09b, Fra05, Law05, Sau05c, Sau08, Sau09a, Sau09b, Sch09b]. papers [Law01b, Hor01]. paradox [Bha07]. Parallel [Gup00, Pon00, Chi05]. Parameterized [MRR02, TS05b, TS05a, Ler04]. Parnas [Hor01]. part [KP02c, KP02d, KP02a, KP02b, PK01]. partial [LvL04, UC04]. Participants [HmL05]. Participation [ACM04d, ER05]. particular [NTS07]. partitioning [Sto02]. Party [CNS05, DPT05]. passing [BD03]. PASTE [Das06, ACM06a]. Patel [Law04b]. Patel-Schneider [Law04b]. Patent [ACM05d]. path [AH03, CKFS01, DADY04, HYD06, VJ05, XA05]. path-sensitive [DADY04, HYD06, XA05]. path-specific [CKFS01]. PathFinder [VPK04]. paths [RRR05]. Pattern [RRR04, WBGM05, GSV05, HLL02, Lon04, Nau03, Pet05, Ros04b, XZRA05, dDLS02, Hei09a]. Patterns [WW05a, BJ05, Bil03, GW05a, GL06, GMS01, KWB05, LMSF04, LZ05b, MHK05, Mem05, NE05, PSSA06, SDD05, TTOH01, Tra04d, WV04]. Patton [Sch06d]. Paul [BM04b, Cho09, FV09a, Tra04d, Tra04e]. Pearson [Els06a, Els06b, Sau05b]. Peer [AH01, Hub02, Lee09, BD03, GWH05, PSK09]. Peer-to-peer [AH01, BD03, GWH05]. people [Bea04]. Perceptions [Miš06]. perfect [Sau04c]. perform [Sau04c]. Performance [Car07, MLG05, SAS09, SA09, SS09, SKK^{+}01a, XOWM06, Bil03, Hub01, KP05, MLRW05, MW05, SS05, Var00, WBH^{+}05]. period [SPS^{+}08]. Peripheral [HmL05]. Perl [Gve09a]. permissions [Nau02]. Permissive [HJM05]. Perplexed [Bea05]. Personality [YEMT05]. perspective [CR06b, Fra06, PND01, SS05, YK09]. Perspectives [KS01, KS02, Aik04]. perturbation [OX04]. pervasive [Cam05, FLMW09, SMC05a]. Peter [Law04b, Sch06d]. Petri [AMN^{+}06]. Petschenik [Sau06c]. phase [AK09]. phases [KS09d, Rei05b]. Philosophers [BBM07a, BM08b]. Philosophy [Ano05b]. phone [dFAM06]. physical [KG05]. PID [bReWgH^{+}09]. pipelined [Mil00]. Pithy [BÁ09b, BÁ09a]. plagiarism [ACM06a]. Plakosh [Law04c]. plan [Sau04c]. Planning [Abr01, BM04c, Väh05, Vog02b]. plans [PC05a]. platform [GFB05]. platforms [MV05]. Playscript [Rie08c]. Plug [Bre06, Che05b, WW05a]. Plug-Ins [Bre06, WW05a, Che05b]. Podcasts [Rec07b]. point [Boy01a, SS09, SBS09a]. pointcut [AX06]. pointer [HP00, HSP06]. pointers [KSWR06]. points [BC02, Die06b, MRR02, MACE02, MVT^{+}02]. points-to [BC02, MRR02, MACE02]. Poisson [SPS^{+}08]. Policies [SL05, SAM05]. Policy [TKA05, ACM06a]. Policy-driven [TKA05]. Pollice [Sau05c]. polymorphic [BD03]. polymorphism [Sai09]. Poppendieck [Pen04]. populations [KP02b]. portable [Joh00b]. portal [PJ05]. portfolios [CdOBT05]. Post [Car07, Hua07]. Post-development [Hua07]. Post-workshop [Car07]. postmortem [MSA^{+}04]. pp [BM08b, BM08c, BM09a, Cho09, FV09a, Hei09a, Hei09b, Man00b, Sau08, Tra08]. pp. [Ebe09, FV09b, Hen09a, Sau09a, Sau09b, Sch09a, Sch09b]. PR [LZ05a]. PR-Miner [LZ05a]. Practical [Ana02, BM07a, CLS00, God01, Law05, ZMC06, BM04b, Boy00, BL04, Fra01, Hub02, LW04, MT01, Pen03a, Ran01, RGS06, Tra04e, BM08c, Ber06c, Gve09a]. Practice [Che05a, TGF09, Tra06a, DMS09, Joh05, LRBH01, OCE^{+}00, RJH05, Rie05, RCJ^{+}08, Swi00, dSRC^{+}04]. practices [Boy01a, CNS05, Ell07, GGH05b, KJML05, LC05, Sau02, SR05, vdH01, Fen06, Bea04, Law04c]. Practitioners [Rac02]. pragmatic [GSV05]. Praise [Tra04c, Win05a]. Preaching [Tra05]. precision [LPH02, Pet05, SA00]. precondition [Ban08, BBM07a, BL06]. predicates [BKM02, FJM05]. predict [SKM09]. predictability [SHRG05]. Predicting [SBS09b, US05, SABS05, SBS09a]. prediction [Boe05b, CSSW04, GG04b, KUST04b, KL05, MSJ05, Mao09, RRU05, WS05, XOWM06, CSSW01]. predictive [ARV05]. predictor [HRGB05]. preferences [Jay00]. Preliminary [LBR06]. Prentice [Man00b, Sau04d, Sau06a, Sch06d]. prepare [Sau04c]. presence [BBHK04]. presentation [CM05a]. presentations [JM00]. presenting [CM05b]. preserving [DNR06]. Press [Ano07b, Bea04, BM04b, BM04c, BM06, BM08b, Ber06b, Che05e, Che05c, Che05f, Che05g, Che05d, Cho09, Fra05, Gla06, Gve09b, Hei09a, Law04b, Law04a, Pet06, Sch06f, Sch09a, Sch09b, Tra04e, WW04b]. preventing [SA00]. preventive [SG07]. primer [Rac06, Che05f, Che05g]. primitives [LFW02]. Principal [Kei05]. Principles [Hud00, KP02c, KP02d, KP02a, KP02b, PK01, Swi00, WW05a, Seg05, SvdVD06, Ber06b]. prioritization [BC05b, SW05, SKR08]. Prioritizing [EMR00, ASK04, ST02]. priority [SPS^{+}08]. privacy [SV00]. proactive [MM09b]. probabilistic [XN09]. Problem [CHR04, Amr05, BBM07a, Bas08, Doe07f, HLL02, HJ01, Mar04, Sri01]. problematic [DGP^{+}01]. Problems [Sau06a, CM05a, Fra02a, JW01, ST05b]. procedures [ACM06a]. Process [AG01, Bas08, BM05a, Con00, Ebe09, Law05, PRRW04, US05, BBM07b, BBM07c, BB07a, BG09, BM04a, BCD00, BB08b, Boy00, Bro06, CCO00, Coo00, Far06, FD07, GT07, GB09, HmL05, KS08b, KPCD00, KH05, KSM06, Ler04, LSK^{+}05, MHK05, MRE05, NWVO05, NKB05, PPLB07, PB05, Ric09, STG02, ST05a, Sch05a, SNWV05, SBCO01, SPS^{+}08, Suk01, Tan02b, VGL05, Wil04, ZRWS05, TLG01, Ano07c, Che05a, Tra06a]. process-based [NKB05]. process-driven [SBCO01]. Process-oriented [Bas08, BB07a, BG09]. Processes [CFM05, DF04, Boe05a, CO00, Law04c, Sau04c, SCT01]. processing [LN05, Luc00, MT01]. processor [WJS^{+}01]. processors [TBM^{+}07]. produce [Zir07]. produced [SGS05a]. Producing [PND01]. Product [GBLM02, GMAC09, KS02, NI01, RSHJ01, TTK04, Ano05b, BS06, BM04a, But01, Cat09, DL05, ELC06, FORG05, GA01, KJP05, KS01, KMM05, MK01, MT01, OG05b, PJ05, SG05, SK01a, Väh05]. product-line [DL05]. Production [Man01, LSH^{+}04, Man00b]. Productive [Ano07d]. Productivity [XN09, MBM09, Rom08, ACM05b]. products [Bal00]. profession [Win05c]. professional [Bag01, Sau06a]. professionalism [Sch06a, Tho01b]. professionals [Gri00]. Professor [Tom06]. profile [Bis02, Coo01b, Coo01a, DT06, DLP01, Omi01, Oqu06b, Tra04c]. profiles [CL05, Ham00]. profiling [AH03, EH04, KCS06, LPH02]. profit [RB05b]. profit-oriented [RB05b]. Program [PTV01, UK06, ASC02, BP05, CT05, DLP01, Eva00, Fel00, GL00, HKO^{+}05, JME05, JAPM05, Joh00a, KCS06, KM06b, Li02, LE04, MSA^{+}04, MACE02, Mye06, NE02a, Not03, OK05, Ram00a, RRR05, Rob05b, Sch00b, SD01, SMCH06, TS00, VS08, Wan00, XQZ^{+}05, ZG05]. program-based [BP05]. Programmer [ACM05b, Ova04, Pen01, Sch06a]. Programmers [Gla06, CKZ05, SSM05]. Programming [Ana01b, Che05a, Che05c, Fra05, GP05, Mai00, Sau09a, XMP08, Aik04, Ana01a, Ber06d, Bur00, Chi05, Cla00c, Cla00d, Din00, DH00, FLY03, Fra01, Fra02b, GST05, HBZ^{+}05, Hud00, JR02, Kfo00, KH01, KM05a, Lar01, LZ05a, Lie00, MO04, Miš06, Ned00, Ols00, PH05, Ram00a, Rey00, Shi04, SSHP05, Sre01, Swi00, Van00a, WC00, vD01, LSBM09]. programs [Att00, BL06, BJ05, Bas00, BS00, CF00, Car00, CJRZ01, Dil00a, Dil00b, DL04, GV02, HD07, KLS00, LPH06, NMS09, NR00a, Nau02, Pet00, Pon00, RKK04, RSB04, RM00a, VCST05, Wor02, XDS04, Zel02, dH05, Sau05b]. Progress [Wit00]. Project [BM08c, Ber06c, Joh00b, Sau05a, BBW05, BM04b, Ber08, CdOBT05, ELC^{+}02, Gup00, Inv02, JLG00, OOOiM05, OCE^{+}00, SSHP05, Tra04e, VWW02]. Projected [FHNS02]. projection [LSH^{+}04]. projects [Ano01b, Asu05, Boy01a, Dan01, HRGB05, MHK05, MV01, MM09a, OG05a, SDD05, Vog02b, XN09]. Prolog [Gup00]. prone [SKM09]. proof [Fas04]. propagation [AK09, HJS02, Raj00]. PROPANE [HJS02]. properties [CCO00, KM06d, MO06, NC00, SABS05, Ste00]. Property [FW06, BCD06]. property-based [BCD06]. proposal [FT05]. ProSim’04 [PRRW04]. protocol [RG06]. protocols [BRBY00, HR06, WV04]. prototype [Joh00a]. prototyping [Tai05]. provides [Zir06]. proving [CTZ02]. pruning [VJ05]. PSE [MSA^{+}04]. PSSM [XN09]. PTR [Man00b]. Public [Neu05a, RN00, Neu00a, Neu00b, Neu00c, Neu00d, Neu01a, Neu01b, Neu01c, Neu01d, Neu02a, Neu02b, Neu02c, Neu02d, Neu03, Neu04a, Neu04b, Neu04c, Neu05e, Neu05f, Neu05b, Neu05c, Neu05d, Neu06a, Neu06b, Neu06c, Neu06d, Neu06e, Neu06f, Neu07b, Neu07c, Neu07d, Neu07e, Neu07f, Neu07g, Neu07a, Neu08a, Neu08b, Neu08c, Neu08d, Neu08e, Neu08f, Neu09a, Neu09b, Neu09c, Neu09d, Neu09e, Neu09f, Rot04]. publication [Bag01]. Publications [BM04c, BM05a, BM05c, Sau04e]. published [BM09a, Cho09, FV09a, FV09b, Gve09a, Gve09b, Hei09a, Hei09b, Hen09a, Sau09a, Sau09b, Sch09a, Sch09b]. Publishers [Tra06a]. Publishing [Mar05b, Sau06b, Sau06c]. purity [FFQ04]. purpose [ARS05]. purpose-driven [ARS05]. Pursuing [DLP01]. Putting [LARSW00]. Python [Gve09b].

QFD [LSK^{+}05]. QoS [DDMW05]. qualitative [ZM05]. Quality [Bre06, Che05b, LP05, Tra04c, BS06, Bha05e, Bha09, Boe05a, Doe07f, Gil05, GHP^{+}09, JR05b, KS08a, KS09a, KUST04b, KKB07, KGK09, MS05, NWVO05, NvdBA05, PHP08, PB05, RB05a, RSHJ01, Sai09, SPH07, SKG08, WS05, WSH05, Wil04, Win09b, ZTSR05, ZE00, ZHG05, BOR02, Sau06c]. Quantification [Hee03]. Quantitative [Bar05b, BM06, KGK09, Sau04a]. Quantum [Law01a]. quasi [BG09]. quasi-agents [BG09]. queries [NB09]. query [GDC02, HG05]. Querying [Die09]. question [Hee03]. questionnaire [KP02a]. questions [Sch05a]. queuing [XOWM06]. Quotes [BÁ09a, BÁ09b].

R [Hei09b, Ben02]. radical [DA05]. Raging [GGT05]. RAIC [LR02]. raise [Chr08, NTS07]. raised [Hee03]. raising [ŚZZ05a]. range [WAK00]. ranking [KAYE04]. Rapid [RM00b, Tai05]. rated [VS08]. rates [Zei00]. ratio [WAK00]. Rational [Ben02, Sch06b]. rationale [AKL^{+}07, DBCP09, FBC06]. rationals [BA00]. ratios [Zir07]. Ravindranath [BM04c]. Ray [Ano07b]. RDBMS [KGS05]. Re [WLF01]. reactive [Alu00, Hud00, PAGDLN05b, PJ05, RKK04, Sto02]. reading [SKT05]. Real [FFPT05, Leh00, Sau06a, Tra08, Abr01, CO02, Dil00b, Fas07, JR05a, KFS05b, KD00, LQV01, PP00, Ste00, TBM^{+}07]. Real-Time [Leh00, FFPT05, CO02, Dil00b, JR05a, KD00, LQV01, PP00, Ste00, TBM^{+}07]. Real-World [Sau06a]. Realising [BK05b, KBBL05, BB07b]. reality [KS09a]. Realizations [BV00]. realizing [Nad00]. really [Ova05, Ros04b, Wol04]. Reasoning [LvL04, CSSW04, DYC^{+}05, KSWR06, RM04a]. REBSE [BB07b]. REBSE-2 [BB07b]. Rebuttal [Are07]. Recognition [Hei09a, Pet05]. recommendation [HWM05, WXL06]. Recompilation [THL03]. Reconciling [GKKW05]. reconfiguration [GBSO04, KJP05, MN05, RCL05, Whi00]. Reconn [LB05]. Reconn-exion [LB05]. reconstruction [GW05b]. record [Zir06]. Recovering [WH05, Man00a]. recovery [AHW05]. Reducing [Bin00]. reduction [RSS^{+}04, bReWgH^{+}09, SK00, VTK02]. redundancies [XE02]. redundancy [Tee09]. Reed [Gla06]. Reengineering [ASD03, DG00, LB05, DGN05, NDG05]. Refactoring [Bow05, Che02, DMNJ06, AX06, BX01, GW05b, HKKI05, JC09, MMvD05, Mar01, MB05, RFG05, YR08]. Refactoring-aware [DMNJ06]. refactorings [Per06]. reference [EAG^{+}05, FT05, LPH02, Pen01]. references [CR06a]. refinement [Bar05a, Oqu04b]. Refining [ZYL04, Oqu04c]. reflected [FSN05]. reflection [YML07]. Reflections [Let06]. reformulating [HJ01]. REFSQ [BOR02]. REFSQ’07 [SPH07]. REFSQ’08 [PHP08]. REFSQ’09 [GHP^{+}09]. regression [ASK04, Bin00, BMSS09, EMR00, KUST04b, MDR05, OSH04]. related [Neu00a, Neu00b, Neu00c, Neu00d, Neu01a, Neu01b, Neu01c, Neu01d, Neu02a, Neu02b, Neu02c, Neu02d, Neu03, Neu04a, Neu04b, Neu04c, Neu05e, Neu05f, RR09, YWA05a, ZE00]. relating [SGS03]. Relational [Bas00, KJ05, Jac00]. relations [LM02]. relationship [SNEC06]. relationship-driven [SNEC06]. relationships [GP05, WXL06]. Relativity [Pan09c, WT08]. released [EH04]. Relevancy [PWZ^{+}04]. Relevant [And04, Ber02, Enc02, Fra06, FW06, Har02, Pec02, Sre02]. Reliability [Ano05a, Sau05d, Tra04c, Zei00, BS00, Bis02, FD00, GHS^{+}06, KM06a, RRU05, Tia00, ZD05]. Reliable [Sau05d, Kfo00, KSM06, SC05b, Wil04]. remain [Rom01]. Reminiscences [Win06e]. remote [JS01, KMC^{+}06]. remotely [HKO^{+}05]. remotely-collected [HKO^{+}05]. repair [TGSP05]. repertory [YEMT05]. Replay [SCFP00, OK05]. replication [BB07a, GFB05]. Report [Car09, CPK09, CRS^{+}05, DHH^{+}06, DHH05, GZES04, GHP^{+}09, HHM05, LSBM09, LRBH01, Mes01, PHP08, PRRW04, RRWB00, ACM05a, WWB09, WJH^{+}04, ACM04d, AG01, BB08a, BB09, BK05b, BB07b, Car07, CGL^{+}04, CEHK07, DFMS09, DMS09, DK09, ELC^{+}02, FFHL05, GCL^{+}05a, GGR^{+}06, GGH^{+}07, GMAC09, HBW00, Inv02, Kim02, KS01, KS02, Lak05, MW01, SKK01b, WLK04, WLCP09, vD01]. reported [HBZ^{+}05]. repositories [ARV05, HG05, KCM05, MLRW05, PWZ^{+}04, PSK09, WH05, DHH05, HHM05]. Repository [VGL05, ABF04, CHC05, Mas05, SSHP05, YFR00, YWA05b]. Representation [SMCH06, GL06, NMS09]. Representation-independent [SMCH06]. Representations [Mun00]. Representing [MK05]. ReQuEst [JR05b]. requirement [VTK02]. Requirements [BOR02, BKW04, BJL^{+}05, CPK09, Die07a, GZES04, GFF^{+}08, JR05b, MK01, Mar05b, PP00, Sau06b, VPDJ05, Ano05b, Bas07, BC05a, CK01, CC05, Dan01, DL05, Die07b, Die09, DB05, DRBT02, Fra04a, Fra04b, GT07, GHP^{+}09, God01, GB01, HB05b, JeMD04a, KG01, KGS09, LMG01, LLMY05, LvL04, LJD05, MS05, Mei00, PHP08, PAGDLN05a, Raj06, RM06b, STG02, Sau02, SPH07, SW05, SKR08, SHD05, Vie05, WSH05, XZRA05, WW04a]. Requirements-based [MK01, PP00, Raj06, SW05]. Requirements-driven [JR05b, LLMY05]. Rescaling [Bis02]. Research [Ano05a, FD07, Jay00, LR02, PRSH07, RM09, TGF09, ABF04, And04, Ber02, BWKG05, BL04, EAS07, Enc02, ELC^{+}02, Eva00, GSX^{+}04, HDLC04, Har02, HW05, KP02c, KP02d, KP02a, KP02b, KJML05, LCLP05, Lu04, Mar05a, OCE^{+}00, Pec02, PK01, RJH05, RCJ^{+}08, SSM05, Seg05, SV00, Sre02, WLCP09, WLA05, Cho09]. researchers [Sau04a]. reservation [Vog02a]. resolution [CEM02, XC05]. RESOLVE [HBW00]. RESOLVE/C [HBW00]. Resolving [Egy04, HA00]. Resource [Tra05, FR05, PSPS05, RDS05, Vog02a, Vog02b]. resource-aware [PSPS05]. Resources [Ano05a, Pan09a]. Response [Jay06, Abr01]. responses [And04]. restructuring [CF00]. result [Lon04]. results [Ano05a, Boe05a, HDS05]. Rethinking [Ben00]. retrieval [BMW01]. Retrofitting [MN05]. retrospective [AR00, KS08a, WLCP09]. return [Sau04d, BM05b]. Reusability [SGK09b, Gil06]. Reusable [GMS01, Jan01, PND01]. Reuse [Ano05a, EV05, ABNR05, Dan01, DL05, GL06, LW04, Lon01, PWZ^{+}04, Ram05, TTOH01, WM00, YFR00, ZBPD01]. Reusing [LA06]. Reverse [BV00, AM03, Ali05, God01, RVR06, VJT08]. Review [Ano07b, Ano07c, Ano07d, Ano07e, Bea04, Bea05, BM04a, BM04b, BM04c, BM05a, BM05b, BM05c, BM05d, BM06, BM07a, BM07b, BM08b, BM08c, Ba05a, Ber06a, Ber06b, Ber06c, Bre06, Che05a, Che05b, Che05e, Che05c, Che05f, Che05g, Che05d, Cho09, Els06a, Els06b, Fen06, Fen07, Fra04a, Fra04b, Fra05, Gla06, Gve09a, Gve09b, Har04, Hei09a, Hei09b, Hen09a, Hor01, Law04b, Law04a, Law06, Law04c, Law05, Man00b, Mar05b, Pen04, Pet06, Sau04c, Sau04e, Sau04a, Sau04d, Sau04b, Sau05b, Sau05c, Sau05d, Sau06a, Sau06b, Sau06c, Sau08, Sau09a, Sch06c, Sch06d, Sch06e, Sch06f, Sch09a, Sch09b, Tra04d, Tra04e, Tra04f, Tra06a, Tra06b, Tra08, WW04a, WW04b, WW05a, WW05b, WW06, BM09a, Ebe09, FV09a, FV09b, Lee09, PPLB07, Sau09b]. reviewing [KON05]. Reviews [ACM05a, Hub02, KON05, RCJ^{+}08]. revision [Bra01, LZ05b]. revisited [Bha05b, DRL^{+}05, Gil05, GT07, JS00]. rewriting [Kfo00, Law04a, Liu00]. Rex [Ano07c, Sau04c]. Reza [Ber06b]. Rice [Tra08]. Richard [Bea05, Gla06, Law05]. Rick [Els06a]. right [Fra02a]. rigorous [Zün01]. Ripley [Hei09a]. rippling [RR09]. risk [CCHS09, CdOBT05, DPP05, FHK^{+}01, Rie07c, Sau04d, ŚZZ05a, SKR08, VS08]. risk-based [DPP05]. Risks [Neu00a, Neu00b, Neu00c, Neu00d, Neu01a, Neu01b, Neu01c, Neu01d, Neu02a, Neu02b, Neu02c, Neu02d, Neu03, Neu04a, Neu04b, Neu04c, Neu05e, Neu05f, Neu05a, Neu05b, Neu05c, Neu05d, Neu06a, Neu06b, Neu06c, Neu06d, Neu06e, Neu06f, Neu07b, Neu07c, Neu07d, Neu07e, Neu07f, Neu07g, Neu07a, Neu08a, Neu08b, Neu08c, Neu08d, Neu08e, Neu08f, Neu09a, Neu09b, Neu09c, Neu09d, Neu09e, Neu09f, Sch05b]. Risky [DPP05]. RMI [EK01b, Sha01]. Robert [Ba05a, Har04, Law04c, Sau08]. robotics [BRC^{+}05]. robust [RM00a]. Robustness [KM06d, FRMW04]. RODIN [ACM05c]. Roel [Law04a]. Roger [Gve09a, Omi01]. Role [KH06, HCGGXFY00, HFH00, NC04, Pan09b, Sha01]. role-based [NC04, Pan09b]. roles [BZG05, dSRC^{+}04]. Roller [Man00b, Win06f]. Rothman [Sch06e]. Round [SKK01b]. Round-Trip [SKK01b]. Roundtable [Tra02]. routines [CD01]. Roy [Tra04f]. Royer [Ano07e]. RRM [RN00]. RSL [NH09]. RTCORBA [RM04b]. RTCORBA-based [RM04b]. Rubel [Bre06, Che05b]. ruby [Ana01b]. rule [Bre05, Wan05]. rules [LZ05a, STG02, WH05]. run [Hor00]. run-time [Hor00]. runtime [CR06b, GvdHT05, HO05, KJP05, SS08, dH05]. runtimes [GK05b]. RUP [Sau05c, ZHG05]. RUP-Centric [Sau05c].

S [BM07b]. SAAGE [BM00b]. SABER [RSS^{+}04]. SableSpMT [PV06]. safe [CO06, DL05, LW04, MPCH05]. safeguarding [SWC05]. safety [Bha05c, Bha05d, CPDGP01, DL04, MP09, MK05, NC00, NR00b, Rie07b, WK05, XDS04]. safety-critical [CPDGP01, MP09, MK05]. safety-liveness [NC00]. Sakhinah [Hen09a]. SALADIN [Inv02]. Salleh [Hen09a]. Salzburg [Con00]. samples [KP02b]. sampling [Boe05b]. Sandy [BM04a]. SAS [Sau04a, Sau04b]. SASÒ [Sau04a]. satisfaction [BBW05, CGB06, JeMD04b, LvL04]. say [ZBPD01]. scalable [BG03, DADY04]. Scale [Fen07, GCL^{+}05a, Sch09a, CGL^{+}04, GCL^{+}05b, GGR^{+}06, GFB05, RB05a]. Scaling [OSH04]. Scenario [SKK01b, AMS05, GKKW05, RRU05, UKM02, UCKM04]. Scenario-Based [SKK01b, RRU05, UCKM04]. scenario-centered [GKKW05]. Scenarios [ARS05, GK05a, WGM06, KON05, KMM05, PAGDLN05a, UKM01, UKM02, BE05]. SCESM [WGM06]. schedules [CZ02]. scheduling [MM09a]. schemes [KD00]. Schneider [Law04b]. Schulte [FV09b]. Schwarz [Sch06f]. Science [Car09, Ebe09, Fen07, Ano05b, Haz07, LRBH01]. sciences [Tem09]. Scientific [Sch09b]. Scientists [Sau09b]. Scilab [XMP08]. SCM [vdH01].

SCM-10 [vdH01]. Scott [Che05a, Che05f, Che05g]. SCQL [HG05]. SCTL [PAGDLN05a]. SCTL-MUS [PAGDLN05a]. SDG [DK09]. SDT [GZ05a]. SE [Car07, Car09, Fra06, Win05b, Win05a]. SE-CSE [Car09]. SE-HPC [Car07]. SE2004 [Ano05a]. Seacord [Law04c, Sau08]. Sean [Sau04a]. Search [STB08, ACM05d, BBG^{+}09, HJ01, KDJ07, KDKJ08a, KDKJ08b, TAC^{+}04]. searches [Sch06a]. Searching [BC02]. SEC [KDJ07, KDKJ08a, KDKJ08b]. Second [Hei09b, Law05, BKW04, Car09, CFM05, DM01, EL04, KS02, WLK04]. Secure [TTOH01, VWW02, BDM05, Fra02a, Mye06, Sau08]. Securely [Bal00]. Security [MS05, BRV01, BP05, DPSJ05, DDMW05, Fra02a, GW05a, HR06, KTB05, NR00b, RM09, RTDR05, Sce00, Sch09c, TWQ09, Vie05, WKP05, Wol04, NKB05]. SEEd [AH09a, AH09b, AH09c, Hen03, Hen04a, Hen04b, Hen04c, Hen05a, Hen05b, Hen05c, Hen05d, Hen06a, Hen06b, Hen06c, Hen06d, Hen07a, Hen07b, Hen07c, Hen08d, Hen09b, Hen09c, Hen08b, Hen08c]. Seeing [Ano00b]. SEEWeb [OYH04]. SEI [HFA^{+}01]. SEI-CSE [HFA^{+}01]. Seidman [Sau06a]. Selected [Law01b]. selecting [CP05]. selection [BV08, BMSS09, CMPB05, DYC^{+}05, MRE05, MT01]. Selective [AH03, OK05]. Self [Sau06a, AKF06, GvdHT05, GK05b, GZ05b, HBZ^{+}05, KJP05, MM09b, QXCMW05, SC05b, TGSP05, Wan05]. self-adaptive [GvdHT05, Wan05]. self-healing [GK05b, GZ05b, QXCMW05, SC05b]. self-managed [AKF06, KJP05]. self-repair [TGSP05]. self-reported [HBZ^{+}05]. Self-Study [Sau06a]. SELMAS [GCL^{+}05a, GGR^{+}06, CGL^{+}04]. SELMAS’05 [GCL^{+}05b]. Semantic [MT00b, Mis00, Fas05, PWZ^{+}04]. semantics [GSV05, LvdBC00, NAD02, ZC05]. Semi [CTZ02, Mar07]. semi-automatic [Mar07]. Semi-proving [CTZ02]. SEMINAL [HJ01]. sense [FFvdH01]. sensitive [DADY04, HYD06, XA05]. Sensitivity [ASCP05c, RRU05, LPH06, MRR02]. sensor [HR06]. separate [PD06]. Separating [JME05]. Separation [Ham05, SG05, THO^{+}01a, WBGM05]. sequence [AMN^{+}06, RVR06, SC02, UKM01]. sequences [RMLMG05]. sequential [Ram00b]. SERPS’08 [TGF09]. server [GM05, MM08a, MM09c]. servers [Tan01]. Service [DHH^{+}06, BRV01, Jan01, KvB04, Lev09, LdSJR07, SMK09, WHW08]. services [CCN^{+}05, DPT05, FRMW04, GJJL00, GZ05b, HZ04, Kir01, KB05, LGA05, Mun00, OX04]. Session [Neu07a, OCE^{+}00, ZBPD01]. sessions [SSG^{+}05a]. set [Lon04]. sets [Jay00, MSJ05]. setting [GGH05b]. Seventh [BOR02]. Shaharuddin [Hen09a]. shape [Ras00, Tra02]. shaping [SMC05a]. shared [FR05, SGS03]. Sharing [LA06, Far06]. Shelf [CFM05, LCS^{+}05, RM01, Tan02b]. should [GHS^{+}06, HD07, HP00]. Shrum [BM04a]. shuttle [PAGDLN05a]. Siddhartha [Ric00b]. side [JS01, MRR02]. side-effect [MRR02]. Sigma [BM05c, VGL05]. signalling [MVT^{+}02]. signature [KWB05]. similarity [BJ05]. Simple [MPCH05]. Simplifying [HZ00]. Simulation [PRRW04, Sau09b, DBK01, GWH05, GM05, HYD06, KWD05, Hei09b]. Simulation-Based [Hei09b]. simulations [BGK^{+}00]. single [KON05]. Sistowicz [Ano07b]. site [WBSM00]. situation [FSN05]. Sivo [Sau04a]. Six [RB05c, VGL05, BM05c]. size [GG04b, KL05, WAK00]. skill [BS06]. SLEDE [HR06]. Slicing [LM08, NR00a, RKK04, SM09a, Li02, Mar01, MACE02, UK06, XQZ^{+}05].

Slicing-based [SM09a]. Small

[Sau05c, Sut05a]. Smalltalk [And00]. smart [Ben00, BJL^{+}05, RSS^{+}04]. SMIT [MV01]. snapshot [SSHP05]. SOA [TWQ09]. SoBelt [Ric00c]. SOBER [LYF^{+}05]. Social [JMT05, KG05, LC05, OOOiM05, SR05]. Society [BM06, BM08c, Ber06c, Fen07, Sau06a]. soft [LCLP05]. SofTware [BKW04, AG01, Ano05a, Ano07d, AH09a, AH09b, AH09c, Aug09, BFZ^{+}09, BTA06, BB08a, BB09, BM03, BM05c, BM06, BM08c, BOR02, Ben02, Ber06c, BÁ09a, Bha05b, Bha05c, BM00b, Bro06, BDM05, BB07b, Car07, Car09, CGL^{+}04, Con00, Coo00, CSSW01, CSSW04, CRS^{+}05, DeM00, DHH^{+}06, DHH05, Die06a, DADY04, DMS09, Ell07, Els06a, ELC06, FV09a, FFvdH01, FFHL05, FFH^{+}05, Fen06, Fen07, GCL^{+}05b, GG04b, Gre00, GWS04, Gve09b, HHM05, Hen03, Hen04a, Hen04b, Hen04c, Hen05a, Hen05b, Hen05c, Hen05d, Hen06a, Hen06b, Hen06c, Hen06d, Hen07a, Hen07b, Hen07c, Hen08b, Hen08c, Hen08d, Hen09b, Hen09c, JW01, Jay06, JMT05, KS02, KH06, KKS01, KWD05, KM08, Leh00, LSBM09, Lon01, Mao09, Mar05a, MEF01, MZ01, MW01, MB09, MBM09]. Software [NPSS01, Off00, PRRW04, PSS04, PSS05, PSSS07, RJH05, RB05a, Ram05, RPSS06, Rob05c, RP01, RW00, STG02, Sau05d, Sau06a, Sau06b, Sau06c, Sch09b, Sch09d, SSHP05, Sti00, SYC^{+}04, THO^{+}01a, TBM^{+}07, Tem09, TLG01, TGF09, Tra04c, US05, WS05, WLK04, WWB09, Win05d, Win07d, ZTSR05, ZHG05, vdH01, ABNR05, ASCP05a, ABF04, Ali06, Amr05, Ano00c, Ano01b, Ano01c, Ano05b, ARV05, AE05b, AKF06, Asu05, BV08, BB01, Bag09, Bag01, BMK^{+}05, BMMG05, Bar05a, BRC^{+}05, BBW05, BS06, BHP^{+}06, BZG05, BK01, Ber05, BÁ09b, Ber04, BWKG05, Bha05d, BSM04, BKSM06, Bha05e, BS01, BM00a, BKKS05, BJL^{+}05, BOH03, BRH04, Boy01a, Boy02a, BMW01, Bra00, BL04, BK05b, Bul00, CCHS09, CCK04, CK01, Cat09, CGB06, CS09, CMPB05, CNS05, CRH01, CCS00, CCWS02]. software [CO00, Cla00a, Cla00b, CR06b, CMV^{+}08, CdOBT05, CLS02, DJP05, DBCP09, Dek05, DA05, DH09, Die06b, DMNJ06, DER04, Doe07f, DRBT02, DPP05, EH04, EVC01, ELC^{+}02, EV05, FV09b, Fen09, FLMW09, FBR05, FK01, FW06, Fra02a, Fra04a, FHK^{+}01, FHNS02, FBS04, GHS^{+}06, GW05a, GvdHT05, GGT05, GCS05, Gil05, Gil06, GL06, GHP^{+}09, Gog00, GMS01, GL00, Gri00, GKS07, GGZ00, Ham02, HLL02, HDLC04, HJ01, HDS05, Haz07, HD07, HW00, Hei00, Hen08a, HRGB05, HJS02, HBW00, HB05a, Hua07, Hub01, Hub02, JB05, JeMD04b, JeMD04a, Jiu07, JAPM05, Joh00b, KS08a, KUST04b, KS08b, KS09b, KS09d, Kei05, KS08c, KDJ07, KDKJ08a, KDKJ08b, KFS05a, KFS05b, KBBL05, KS01, Kni05, KPCD00, KPDdL00, KJML05, KH05, KK06, KM05b, KGK09, KM06a, KSM06, KM06c, Lan06, Lar06]. software [LKS00, Law04c, LG02, Lea00, LvL02, LW04, LSH^{+}04, LZ05a, LWZ05, LLT05, LSK^{+}05, LZ05b, LRBH01, LJD05, MSR02, MHK05, Man00a, MV05, Mar07, MO06, MP09, Mei00, MZY01, MPCH05, MV01, MM08a, MM08b, MM09c, MM09a, MK06, MK07, MBPRR01, MDR05, Mun00, NWVO05, Nam08, NR00b, NI01, NvdBA05, OG05a, OG05b, Oqu04a, Oqu04b, Oqu04c, Oqu06b, Oqu06a, OLHL02, OSH04, OCE^{+}00, OW02, Ova05, PHP08, PC05a, Pan09a, Pan09b, Pan09c, PPV09, PCP05, PPLB07, Pen04, PB05, PV06, PRSH07, PH07, QXCMW05, Rac06, Raj00, RRWB00, RM01, RM04a, Ran01, Rec07a, Rec07b, RS07, Red00, RM09, Rei05a, Rei05b, RTDR05, Ric00d, Rie05, Rie06, Rie07a, Rie07b, Rob05a, Rom07, Ros04a, Rot00, SMC05a, Sai09, SG05, San09, SPH07, SA09, Sca05, SSM05, SGS05a, Seg05, SKM07]. software [SKG08, SGK09a, SGK09b, SR05, She05, SBZ00, SCT01, SG07, SS08, SBS09a, SBS09b, SKK^{+}01a, Smi00, SP00, Spr05, Sre01, SPS^{+}08, SKT05, SDD05, Suk01, SGCH01, SM09b, Tai00, Tan02a, Tan02b, Tho01b, Tia00, TvH00, TS00, Tra04f, Väh05, Var00, VS08, WM00, WHS01, Wan05, WGS07, WDKM05, WW00, WKP05, WBH^{+}05, WBG05, WBSM00, WH05, Win06b, Win07c, Win08a, Wit00, Wol04, WK05, XN09, XYH04, YYX06, YXP09, YML07, ZD05, Zün01, dSRC^{+}04, DK09, GCL^{+}05a, GGR^{+}06, GLM^{+}05, BM05b, BM05d, Sch06d, Tra08, BM04b, Tra04e, Ano07e, BM04c, BM07a, BM09a, Ba05a, Doe00a, Doe00b, Doe01a, Doe01b, Doe01c, Doe01d, Doe02a, Doe02b, Doe02c, Doe02d, Doe03, Doe04a, Doe04b, Doe04c, Doe05a, Doe05b, Doe05c, Doe05d, Doe05e]. Software [Doe05f, Doe06b, Doe06c, Doe06d, Doe06e, Doe06f, Doe07a, Doe07f, Doe07b, Doe07c, Doe07d, Doe07e, Doe08a, Doe08b, Doe08c, Doe08d, Doe08e, Doe08f, Doe09a, Doe09b, Doe09c, Doe09d, Doe09e, Doe09f, Hor01, Not00, Sau04e, Sau04d, Sau05c, Sau05d]. software-intensive [CO00, GW05a]. solid [JMV04]. solution [GGH05a]. solutions [Tra04d]. solve [HLL02]. solver [EK01a, JV00]. solving [Doe07f, EJTY04, Sri01]. Some [Fas05, Lu04, MM08b, MM09a, SR05, Tra04c, Fas07, God01, ST05a]. Sons [BM09a, Che05a, Law06, Tra04d]. soon [ZBPD01]. sort [STB08]. Sorting [Lon04]. SOSE06 [DHH^{+}06]. Source [FFvdH01, FFHL05, FFH^{+}05, RB05a, ACM05d, YWA05b, Asu05, DB02, GGT05, GGH05b, HG05, KCM05, KH05, LLL01, LG02, MV05, Mas05, MR06, NFH05, Rob05a, Sca05, SDD05, WGS07, ZE00, ZRWS05, ZD05, ZLL04]. sources [RGB05]. space [DLP01, LLT05, Ova05]. Spaces [FFH^{+}05, SC05a, YCP04]. Spafford [Coo01a]. spatial [Ban08, CG09]. special [DFMS09, ZMC06]. specialties [XXJ05]. specific [BLQ^{+}06, CEH03, CKFS01, Dev00, PND01, RM06a, WH05, Zir07]. Specification [Car00, Cle00, Mar05b, PSSA06, RRSW00, TK06, YKB02, BA05b, Bra01, BČVZ06, CL06, CLS00, DYC^{+}05, Ehr05, GR06, HCGGXFY00, HR06, KSWR06, Lak05, LBR06, OACL05, PP01, PAGDLN05b, Ric00a, RC07, SKK^{+}01a]. Specification-based [Car00, KSWR06, Ric00a]. specifications [Amm00, Che00, CJRZ01, LvL02, NE02a, UKM01, WY04, dFAM06]. specified [Bla00]. Specifying [ZC05, EKHG01, MO06, Oqu04a]. speculative [PV06]. Specware [McD00]. Spiral [HFA^{+}01]. spreadsheet [FR05]. spreadsheets [AE05a, Cle05]. Springer [Ebe09, Har04]. Springer-Verlag [Har04]. spurious [SA00]. SQL [HO05, SCT04]. SQL-injection [HO05]. SQUARE [MS05]. SRE [GHS^{+}06]. SS7 [MVT^{+}02]. stability [Ham05, WZX06, XN09]. stability-driven [Ham05]. staff [MM08b]. Staffing [BBW05]. Standard [SK01b, Bha05d, Gun00, KS09a, Lee09, PIÖ05, Pen01]. Standards [Ber06c]. standing [ZBPD01]. Stärk [Har04]. starting [WLCP09]. starvation [Ban07]. State [BE05, CCHS09, GK05a, SM09b, WGM06, AH01, ARS05, Cla00b, DH00, FHNS02, GKKW05, GGSV02, Ham06, Har04, Hie04, KPCD00, Ler04, Stu05, SGS03]. state-based [GKKW05, Hie04]. state-of-the-art [AH01]. statechart [Jiu07, LM02]. statecharts [LvdBC00]. Static [NC04, RKG04, RVR06, CEH03, Eng03, HO05, HSP06, KL05, Lar06, LARSW00, LPH02, MSA^{+}04, NE02b, Pet05, PD06, VJ05, ZLL04]. statistical [KP05, LYF^{+}05, SSG05b]. status [Bha05c, TvH00]. STCIM [WZX06]. steering [LE04]. stencil [DYC^{+}05]. step [Boy00, GLB04, SG07, BDE^{+}03]. step-by-step [Boy00]. Stephen [Sau04a, Sau06a]. steps [FBC06]. stepwise [Oqu04b]. stereotypes [SKT05]. Steve [BM05b]. Stewart [Sch09b]. STG [CJRZ01]. storage [CF00]. Story [GZ05a, NDG05]. Strategies [Ben02, Bag09, BL04, EH04, NVS^{+}04, NPSS01, RM04a, STG02, Whi00]. strategy [BHLRM05, SP00]. Strathcona [HWM05]. STRAW01 [CK01]. STRAW’03 [BKW04]. strawman [WLA05]. streams [Che06]. strength [Fra02b]. String [CKG06]. stringent [CMV^{+}08]. structural [GV02, Hee03, KM05b, MO06, Ric00c]. structure [Bow05, CCK04, CM05b, KS09c, Man00a, SGCH01]. structured [Boy00]. structures [Mad01]. Structuring [Die07b, BMK^{+}05, ELC06]. student [Chr08, Fra06, MLRW05, PCP05]. studies [BL04, Par05, Sau04a, Sil09, WC00]. Study [Sau06a, CS09, CD01, CSTM06, DS06, EAN01, EH04, FD00, FHK^{+}01, GGH05b, JeMD04a, KSNM05, KLS00, LARSW00, Miš06, NWVO05, Oqu04c, PAGDLN05a, SC05a, Ter04, WLA05, YEMT05, YR08, ZLX04, ZHG05]. study-based [WLA05]. style [MT01, PM03, Che05e, Sch09b]. styles [GWH05]. subdomains [Ham00]. subjectivity [CSTM06]. subset [CMPB05]. Substantial [Ebe09]. subtypes [EJTY04]. subtyping [FLF01]. successful [Boy01a, DOV^{+}01, Pin00, RCJ^{+}08]. SUDS [Lar06]. Suely [Sch09b]. sufficient [Pin00]. suggestions [Rob05b]. Suitability [LKS00]. Suite [Bat00, Rug00, Bha05e, LGO04, TG06]. suites [VTK02]. summaries [TS00, ACM05b]. Summarizing [SS05]. Summary [Con00, DM01, SKK01b, BOR02, BE05, CHR04, ELC^{+}02, EL04, GK05a, Inv02, JMT05, KB05, PA05, PSS04]. Superhack [Win07e]. Support [BM08c, Ber06c, TLG01, ANGRB05, Ale02, Bha05e, BK05a, Bow01, BLS02, BRBY00, Dil00b, DER04, GBLM02, GvdHT05, GK05b, HKKI05, MM09b, MRE05, Ols00, Per06, RS07, RK00, SMC05a, SKM09, SN05, THL03, VPDJ05, XC05, Ebe09]. supported [GB01]. Supporting [CCWS02, Dek05, ABNR05, FR05, KN05, MZY01, Ric09, RB05c, SSH^{+}05]. Surfing [Doe00a, Doe00b, Doe01a, Doe01b, Doe01c, Doe01d, Doe02a, Doe02b, Doe02c, Doe02d, Doe03, Doe04a, Doe04b, Doe04c, Doe05a, Doe05b, Doe05c, Doe05d, Doe05e, Doe05f, Doe06b, Doe06c, Doe06d, Doe06e, Doe06f, Doe07a, Doe07f, Doe07b, Doe07c, Doe07d, Doe07e, Doe08a, Doe08b, Doe08c, Doe08d, Doe08e, Doe08f, Doe09a, Doe09b, Doe09c, Doe09d, Doe09e, Doe09f, Not00]. surprise [Rie05]. survey [Ano05b, JeMD04b, KP02c, KP02d, KP02a, KP02b, PK01, TTK04, XQZ^{+}05, ZE00]. Surveying [HW05]. survival [And00]. Susan [BM08c, Ber06c]. SW [Ber06c, Ano05b, CBB^{+}07, Tra05]. Sweden [TGF09]. Swing [Gla06]. Symbolic [HYD06, CTZ02, CJRZ01, CPDGP01, KS06]. Symposium [CRS^{+}05]. symptoms [CS09]. synchronization [RC01]. synchronous [RKK04]. syntax [NFH05]. Synthesis [BGPLHS05, Att00, IT01, NS09, Ric00b, TBM^{+}07, UCKM04, YKB02]. System [Amm00, CSSW01, UCKM04, AAAG^{+}05, APS09, BBM07c, BG09, Bha09, BLQ^{+}06, BČVZ06, Cap04, CEH03, CO00, CKFS01, Dro09, EJT04, Egy00, GW05a, GKKW05, GM05, Har04, KFS05a, KFS05b, KMC^{+}06, Lak05, LvL02, LN05, Mao09, MM09b, NRSS09, NR00b, OMK04, OLHL02, OW02, PAGDLN05a, PSK09, RM06a, RM06b, RCL05, RSB04, Rom08, SAS09, SABS05, Sch06b, Sha01, SC05b, SSHP05, WH05, YMI00, Sau06c]. system-making [Sch06b]. system-specific [CEH03]. systematic [FBC06, GJJL00, PPLB07]. Systems [BM06, GCL^{+}05a, Rho02, Sch05c, WJH^{+}04, dLR05, AH01, ABNR05, Ale02, Alu00, AMN^{+}06, Ana00, Ano05b, AKF06, AHW05, BMK^{+}05, Ban07, Ban09, BMMG05, BB07a, Bas07, Bas08, BG09, BHLRM05, BHP^{+}06, Ber05, BM00a, Boy02a, BDM05, Bul00, Cac00, CO02, CRH01, CGL^{+}04, Cla00a, Cla00b, CLS00, Cle00, COWL02, CPDGP01, CLS02, DS05, DeL05, DL05, DB05, DPP05, EAN01, EAG^{+}05, GCL^{+}05b, GGR^{+}06, GBSO04, GG04a, GB08, GK05b, GKS07, GKK08, GR06, GFB05, Ham02, Ham05, Jah01, JR05a, JRHD00, KAKM05, KS09c, KE01, KS08c, KJP05, KG05, KD00, KKG09, Lan06, LLMY05, LQV01, Law02, Law04c, LKF02, LSH^{+}04, LE04, LWZ05, MSJ05, Man00a, MP09, MT00b, Min00, Min05, MT01, MK05, Mye06, NVS^{+}04, Nad00, Neu00a, Neu00b, Neu00c, Neu00d, Neu01a]. systems [Neu01b, Neu01c, Neu01d, Neu02a, Neu02b, Neu02c, Neu02d, Neu03, Neu04a, Neu04b, Neu04c, Neu05e, Neu05f, NI01, OACL05, OSH04, PP01, PAGDLN05b, Pen03b, PSSA06, PGL^{+}05, PP00, PIÖ05, PH05, PSS04, RK00, Ros00, Rot00, ST05a, Sai09, SA09, Sch05a, Sch05b, Sch06b, Sch07, Sch08a, Sch08b, SGK09a, SS08, Sis00, Sta00, Ste00, Sto02, Sut05a, TBM^{+}07, TKA05, VWW02, WGS07, WW00, Wen01, YFR00, dFAM06, CLS02, Leh00, PSS05, PSSS07, RPSS06, dLGR05, Hei09b, Mar05b, Law04a, Tra04d, Mar05b]. Szabolcs [Fen07].

table [SPC09]. Tabled [Ram00a, Swi00]. Tailoring [Far06, PPLB07]. Taking [Ano05b]. tale [Win06b]. talks [YWA05b]. Taming [Ros01]. task [Amr05, YWA05b]. taught [GHS^{+}06]. taxonomy [KCM05, WKP05, Wor02]. Tayntor [BM05c]. TBD [BÁ09b]. teach [Ali05, HD07]. Teaching [BK01, Bow01, Ell07, Pan09a]. teaching/learning [Pan09a]. Team [TLG01, BS06, JeMD04b, Tan02a, WBG05, YEMT05]. Team-Based [TLG01]. Teams [Sau05c, BZG05, Hub01, BM07b]. Techies [Sch06e]. Technical [Aik04, SR05, Tra02]. technique [ASK04, Li02, MB09]. Techniques [EAN01, GGH^{+}07, BL04, But01, CGB06, CLS00, DER04, Fra02b, HKO^{+}05, Hie04, JW01, JMV04, KMC^{+}06, Man01, MSC04, Par05, SCS06, SKT05, SWC05, WS05, Man00b]. Technologies [MEF01, BMW01, DH00, EK01b, Law04c, SBZ00, Sta01, You00]. Technology [Sha02, Ano05b, BG03, Cla00b, EAS07, Eva00, Fas08, Kai00a, Kai00b, Lu04, Sch00a, Van00b, WLA05, Win09c, AG01]. technology-dependent [WLA05]. telematics [BHLRM05]. Template [Dro09]. templates [Say05, XOWM06]. temporal [Ezi04, GDC02, LKMU05]. tension [DB02]. tentative [Väh05]. Term [Liu00, Väh05, Law04a]. terminological [Wen01]. terminology [BHP^{+}06]. Test [ASKS04, Ano07b, Ano07d, Boy00, BC05b, JC09, SKR08, VPK04, ASK04, BK09, BMSS09, BJL^{+}05, CO02, CJRZ01, CP05, EK01a, EMR00, FBHZ08, GT07, Gre00, Hie04, KM08, LGO04, MK04, MPS00, NMS09, NS09, OX04, Par05, Raj06, RCL05, Ric00b, Ric00d, SM09a, SSG05b, SK00, SW05, SGAY09, TG06, VTK02, dFAM06, DMS09]. test-case [BK09]. test-driven [JC09]. Testability [WAK00, BBHK04, KM09, MM09c]. tested [Sau06c]. tester [Enc02]. testers [Pin00]. Testing [Ano07c, Ano07e, BOP00, Cac00, CJ04, CSR06, DFW04, FRMW04, Ham00, Joh00a, PTV01, Rot00, Sai09, Sau06c, Tan02b, XXJ05, ZLL04, ASK04, ARS05, Amm00, And04, Ano00c, AMS05, Bag09, BBHK04, Ber04, BH05, Bin00, BS01, BGP06, BKM02, Boy00, BCD06, BP05, BL04, CL05, Car00, CDF^{+}00, CTZ02, CKZ05, DJP05, Dil00b, DER04, EMR00, EH04, FV09b, FD00, FW06, FHNS02, GZ05a, GT07, GLB04, GJJL00, Ham06, HN04, HIM00, JMV04, KP05, KB05, KKN07, Lak05, LM02, LN05, LGO04, LJD05, Mam00, Mei04, MSP01, MVT^{+}02, MSC04, MDR05, NVS^{+}04, NWVO05, Off00, OG05b, OSH04, Ric00a, RMHL06, SE04, Sau04c, Say05, SMA05, SCS06, SP00, SSHP05, Sre02, SPS^{+}08, SCFP00, SCT04, SYC^{+}04, Tai00, TE04, TTK04, TH02, Ton04, TDSCdlR08, Var00, VCST05, Wit00]. Testing [FV09a, BM05d]. tests [SMD^{+}05, ST02, TS05b, TS05a, Els06a]. TestTalk [Ric00d]. Text [Gve09a, HDS05]. textual [KK06]. Thailand [PB05]. their [EH04, KB05, WGS07, XXJ05, YWA05b]. theoretical [WC00]. Theories [Cla00d, LCS^{+}05]. Theory [Che06, Sch07, CS05, Fas05, JB05, Kei05, Kfo00, KPCD00, Mai00, Mao09, MK06, MK07, Sch06b, Sut05a, ZLX04]. theory-based [ZLX04]. things [Ras00]. Thinking [Che05d]. Third [JW01, BME^{+}08, DPT05, Car07, PA05]. third-party [DPT05]. Thomas [Ano07e]. Thoughts [Tra04c]. thread [CZ02, San09, SM00]. thread-aware [SM00]. threads [SGS03]. three [dDLS02]. thwarts [dSRC^{+}04]. Time [Ba05a, Leh00, Tra05, Abr01, ARV05, CO02, Dil00b, FFPT05, Hor00, JR05a, KFS05b, KD00, LQV01, LKMU05, PP00, PD06, RM05, Ste00, TBM^{+}07, TGSP05]. timescale [Mas05]. timing [Ste00]. tips [Pan09a]. title [ACM04e]. TML [FBHZ08]. TOC [Ano05a, Ano05b, Ano07a, ACM05d, Tra04c, Tra05, ACM05c, ACM05b, ACM06a, ACM06b]. Tockey [BM05b]. Today [Sau06a]. tolerance [Rom07, SN05]. tolerances [Rie07a]. tolerant [GFB05]. Tom [Ber06a, Pen04]. tomahawk [STB08]. Tomayko [Tom06]. tomography [BOH03]. too [Enc02, ZBPD01]. tookit [Pen04]. Tool [ANGRB05, Bat00, BV08, Ban08, Bow01, BRBY00, CNPD06, CJRZ01, FLY03, GST05, HKKI05, HWM05, IMP05, JS00, LMG01, Raj00, Ric00c, RN00, SS08, SCFP00, VDM^{+}06]. tool-assisted [JS00].

ToolBus [GSX^{+}04]. Tools [GK05a, WGM06, AR00, BRC^{+}05, GCS05, GGH05b, HN04, HBR05, KKB07, Mun00, PGL^{+}05, Pin00, RKK04, Spr05, WKP05, ZLL04, BE05]. toolset [Bul00]. Tore [Tra06a]. Torgeir [Tra06a]. TOSEM [Not08a, Not07a, Not07b, Not07c, Not08b]. trace [BDE^{+}03, Egy04]. traceability [BJL^{+}05, Hen08a, VPDJ05]. traces [Bow05, PC05b]. tracing [SHD05]. Tracking [MV01]. tradeoffs [MK05]. traffic [MDC04]. train [MSJ05]. training [ASCP05a, Bag01]. transfer [CD01, MVT^{+}02]. transformation [BBHK04, BM09b, GL00, JAPM05, Mye06, bReWgH^{+}09, WBGM05, XDS04]. transformations [SD01, Wan00]. transforming [KS09a]. transition [AKF06, Shi04]. Transparent [SMC05a, FFPT05]. transportation [ZMC06]. traumatica [Win07d]. treatment [RTDR05]. tree [NFH05]. Trees [WW06]. Trend [Rec07a]. trends [AKL^{+}07, Rec07a, ST05b]. Trevor [Che05e]. triangle [WAK00]. tricks [BK01]. Triggered [SC02]. Trip [SKK01b]. Triveni [JLG00]. trust [BMMG05, Cap04, DPP05, WXL06]. trustworthiness [LGA05]. Trustworthy [BHP^{+}06, Ber05, Jay06, Sch06d, BM06]. TSP [BM07b]. tuning [HSP06, Ste00]. Turner [Bea05, Law05]. turning [PK01]. tutorial [WHS01]. Tutorials [ACM04d]. Two [Seg05, BBM07b, BBM07c, Tem09, WBGM05]. two-dimensional [WBGM05]. two-process [BBM07b, BBM07c]. TXL [DL04]. TXM [CBB^{+}07]. Type [Kfo00, BLQ^{+}06, EJT04, Hor00, SGS05a]. Typed [Sha00, Oqu04a]. types [BCD00, BLQ^{+}06, MMvD05]. typestates [BA05b]. typing [TWQ09].

UAU [RCL05]. UML [Bar05b, Bha05e, Bha09, BV00, Egy00, HIM00, KG05, KKS01, LM08, LM02, LQV01, Mar07, NS09, Oqu06b, PM03, PH07, RVR06, SM09a, SAS09, SS09, SAM05, VJT08, YWG04, YW04, Zün01, TR00c, WW04b]. UML-Based [HIM00]. UML/Analyzer [Egy00]. uncertainties [Egy04]. uncut [Bro06]. Understanding [Ebe09, GGH05a, NFH05, GCS05, JME05, KN05]. Unified [Che05a, DS05, SSH^{+}05]. uniform [KS09c]. unifying [DPSJ05]. unit [AE05a, SMA05, TS05b, TS05a, TS05b]. units [COWL02]. Universal [Bha05d]. universe [BS00]. University [BM04b, BM08b, Ber06b, Che05e, Che05c, Che05f, Che05g, Cho09, Fra05, Gla06, Gve09b, Hei09a, Law04b, Law04a, Pet06, Sch06f, Sch09a, Sch09b, Tra04e, WW04b]. unmanned [Jiu07]. unsafe [MPCH05]. unstructured [BL06]. update [SSH^{+}05, WHW08]. updating [CO06]. UPEX [ER05]. upon [ELC06, OCE^{+}00]. usability [DFMS09, Mar04]. Usage [Say05, WC00]. Use [BV00, Die06b, KY09, KG01, SGAY09, BM01, Boy02a, BMW01, BLS02, FBC06, Fas07, Fra04a, HP00, MSR02, MV05, PIÖ05, Rie08c, ZXZ01, Zün01, dFAM06, Fra04b]. used [MSJ05, Wen01, YWA05a]. useful [GKS07]. User [Kim02, LSBM09, Bla00, BM05e, CKZ05, ER05, HBR05, KM05a, LMG01, Pen03b, PH05, RB05c, SSG^{+}05a, SSM05, SMD^{+}05, Seg05, SBCO01, Sut05b]. user-guidance [SBCO01]. users [Ber02, Enc02, Pec02, Sre02]. uses [RCJ^{+}08]. Using [Ber06c, CP05, CKFS01, CPDGP01, HB05b, KN05, LGO04, Mar05b, MP05, Nau03, SPT05, SvdVD06, SCT04, XE02, ASC02, ALK09, BNL05, BC05a, Bha09, BB08b, Bow05, BOH03, CHC05, DB09, Dan01, Dil00b, DL04, Dro09, GV02, GR06, GZ05b, Hen09a, HBW00, KUST04b, KP05, KG05, KPCD00, KL05, KK06, KWD05, KM06d, Lar06, LB05, Ler04, LPH02, LS06, Mar01, Mei04, MLG05, Mil00, MBM09, NWVO05, NTS07, NRSS09, NS09, NFH05, OX04, OOOiM05, PAGDLN05a, PJ05, Rie08a, RN00, SK00, SNWV05, SBS09b, SKT05, TTOH01, VJ05, WS05, XN09, ZLL04]. utility [CD01]. Utilizing [She05].

validate [MSJ05]. validated [RN00]. Validating [Mis07, Cla00b]. validation [BJL^{+}05, DADY04, FD07, JAPM05, LN05, MK06, MK07, She05, SBZ00]. valuation [NKB05]. Value [BNL05, Boe05a, Bre05, CS05, Dan01, DADY04, HB05a, JB05, SABS05, SGCH01]. value-added [Dan01]. Value-based [Boe05a, HB05a, JB05, SABS05]. value-oriented [CS05]. variabilities [GA01]. Variability [MO04, BB01, EV05, GBLM02, SvdVD06]. variable [BBG^{+}09]. variant [ARV05]. various [ASCP05a, RGB05]. VCH [Sau09b]. VDM [NH09]. Veanes [FV09b]. Vector [BBG^{+}09, SKM09]. vehicle [Jiu07, RSHJ01]. Velten [Sau09b]. Vendors [MV05]. verification [Alu00, AMN^{+}06, Ban08, BČVZ06, Che02, Cla00a, Cla00b, CLS00, CRL^{+}06, DH00, Ehr05, FD07, FK01, GBSO04, Hen00, Jiu07, KB05, LARSW00, LN05, MDC04, PP01, Ram00a, RM04b, She05, TK06, YKB02, dH05, PTV01]. verification-oriented [BČVZ06]. verified [BKKS05]. verify [HmL05]. Verifying [CCO00, DRBT02, KFG04, Ler04, LKF02, Ano01a, CPDGP01, KvB04, MO06]. Verisim [BGK^{+}00]. Verisoft [DL04]. Verlag [Ebe09, Har04]. Version [KS09d, Dro09, HmL05, KS09c, MV01]. versioning [Stu05, YMI00]. versions [CO06, ZG05]. versus [Mar05a, RJH05, TGSP05]. very [Zir07]. via [Amm00, DADY04, Hor00, Jan01, LE04, MSA^{+}04, VS08, WV04]. vice [ACM05b]. vice-chair [ACM05b]. view [AR00, DPSJ05, SNEC06]. Viewing [ZXZ01]. viewpoint [Rei05a]. viewpoints [GB09]. Views [SHB05]. Vincent [Fen06]. Violatility [SK01a]. violations [LZ05a]. Virtual [HLL02, Gri00, Jan01]. VIRTUE [RK00]. Visit [Kim02]. visitor [SD01]. Visual [Che05c, GST05, Hen09a, RS07, Bur00]. visualization [Eng03, RK00, Rei05b, RHK00, Van00b]. visualize [KGS05]. Visualizing [Sau04b]. Volume [Ano05a, MK04]. VORD [RGS06]. Vote [Tra05]. voting [KMC^{+}06]. Vrijer [Law04a]. vs [Pan09b, RB05b]. vulnerabilities [GW05a, RM09]. vulnerability [AK09].

W [Ano07d, BM08b, BM08c, Ber06c, Che05a, Che05f, Che05g, Sau09a]. WADS [dLR05]. walkthroughs [RCJ^{+}08]. Walz [BM08c, Ber06c]. WAN [CNS05]. WAP [Ana02]. Ward [Els06a]. warehouse [DA01, KGS09]. wars [Fas04]. Watts [BM07b]. way [Fra02a, Van00b]. Weakest [Ban08, BL06, BBM07a]. Weakest-precondition [BL06]. weakness [ASC02]. Web [DM01, Ana02, BM05e, Bow01, CCN^{+}05, CNPD06, DPT05, DFW04, Fas05, Fas07, FRMW04, GZ05b, HZ04, KvB04, KB05, KSM06, LdSJR07, MG00, NB09, OX04, Pan09a, PJ05, SSG^{+}05a, SMK09, XXJ05, APS09, Bie02, DB02, Kir01, Shi04]. web-based [CNPD06]. web-integration [MG00]. WebDAV [TR00h]. Weinberg [Tra06b, Tra06b]. Wesley [Bea04, Bea05, BM04a, BM05b, BM05d, BM07b, Bre06, Che05b, Fen06, Fra04a, Fra04b, Law04c, Law05, Pen04, Sau04c, Sau05c, Sau08, Sch06c, Tra04f, Tra08, WW04a, WW05a, WW05b, WW06]. West [Che05d]. Weyuker [KM06d]. Where [OWB04, Sau06b, SZPZ06, WLCP09, ZBPD01]. Which [HP00]. white [TDSCdlR08]. white-box [TDSCdlR08]. whole [Rac00]. WICSA [CNS05]. Wide [DB02]. widely [LSH^{+}04]. widely-deployed [LSH^{+}04]. Widrig [Fra04a]. Wiki [CNS05]. Wiley [Ano07c, Ano07d, BM07a, BM08c, BM09a, Ber06c, Che05a, Fen07, Gve09a, Hei09b, Hen09a, Law06, Sau09b, Tra04d]. Wiley-IEEE [Fen07]. Wiley-Interscience [BM07a, BM08c, Hei09b, Hen09a]. Wiley-VCH [Sau09b]. Will [Els06b]. Willem [Law04a]. Wireless [Bie02, Sch06f, Ana02, Sta01]. without [Tra04f]. WODA [EL04, PA05]. WoDiSEE [GWS04]. Wolfgang [Ebe09]. Wolfram [FV09b]. Wonderful [Are07]. WOOR’99 [DG00]. work [LLL01, LARSW00, Rac02]. workbench [KvB04]. Workers [Sch06e]. Workflow [Man00b, Man01, Sha01]. Workflow-Concepts [Man00b]. Working [Kim02, SPH07, GHP^{+}09, PHP08]. Workshop [AG01, BB09, BKW04, Car09, CFM05, Con00, CHR04, CSSW01, CSSW04, DHH^{+}06, DHH05, DMS09, DF04, EC01, FFvdH01, FFHL05, FFH^{+}05, GZES04, GMAC09, GK05a, GFF^{+}08, HHM05, JW01, Kim02, KS01, KS02, KKS01, LA06, MEF01, PTV01, PRRW04, PSS05, PSSS07, RPSS06, SK01b, ACM05d, SKK01b, Tra05, WLK04, WWB09, WGM06, dLGR05, dLR05, ACM05b, ACM06a, vdH01, BB08a, BOR02, BE05, BK05b, BB07b, Car07, CGL^{+}04, CEHK07, DOV^{+}01, DM01, DK09, EL04, GCL^{+}05a, GGR^{+}06, GGH^{+}07, GLM^{+}05, HFA^{+}01, HJ01, JMT05, MZ01, MW01, PA05, PSS04, RRWB00, Tho01b, WLCP09, WJH^{+}04, vD01, BFZ^{+}09, CPK09, DG00, GWS04, KS01, KS02, LSBM09, Rob05c, RP01, THO^{+}01a, TLG01]. Workshops [CLS02, ACM04d, WW04a]. World [Sau06a, Cha05, NPSS01, Rom01, DB02]. worlds [Tem09]. WoSEF [SK01b]. wrapping [Bal00]. WREN [LR01b]. write [CEH03]. Writing [BM01, Tra06b]. written [Hei09b, Sau09a]. x86 [CKG06]. XacT [RDS05]. Xitao [Sau04a]. XMI [PGL^{+}05]. XML [MEF01, BM08a, BM09b, BS01, FBHZ08, FBS04, GSV05, Kir01, WJS^{+}01]. XML-based [BM09b, BS01, FBHZ08]. XP [GMAC09, LR01a, YEMT05, ZHG05]. XP2001 [vD01]. XP2009 [WLCP09]. XSL [Kir01]. XSLT [DRBT02]. xTAO [OACL05].

year [WLCP09]. York [Sre01]. Yuhas

[BM06].

Z [Bow01, Mar05b, NH09, WY04, TDA07]. Zomaya [Hen09a].

References

Abi-Antoun:2005:ISD

[AAAG^{+}05] Marwan Abi-Antoun, Jonathan Aldrich, David Garlan, Bradley Schmerl, Nagi Nahas, and Tony Tseng. Improving system dependability by enforcing architectural intent. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alexander:2004:SER

[ABF04] Roger T. Alexander, James M. Bieman, and Robert B. France. A software engineering research repository. ACM SIGSOFT Software Engineering Notes, 29(5):1– 4, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Adams:2005:ARL

[ABNR05] Paul Adams, Cornelia Boldyreff, David Nutter, and Stephen Rank. Adaptive reuse of libre software systems for supporting on-line collaboration. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abrahams:2001:PRT

[Abr01] Alan Abrahams. Planning for real time event response management. ACM SIGSOFT Software Engineering Notes, 26(1):93–94, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2003:BM

[ACM03] ACM SIGSOFT Software Engineering Notes Staff. Back matter. ACM SIG-SOFT Software Engineering Notes, 28(3):22–37, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2004:BMA

[ACM04a] ACM SIGSOFT Software Engineering Notes Staff. Back matter (abstracts and calendar). ACM SIGSOFT Software Engineering Notes, 29(2):27–62, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2004:BAC

[ACM04b] ACM SIGSOFT Software Engineering Notes Staff. Backmatter (abstracts and calendar). ACM SIGSOFT Software Engineering Notes, 29(3):31–37, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2004:BAI

[ACM04c] ACM SIGSOFT Software Engineering Notes Staff. Backmatter (author index). ACM SIGSOFT Software Engineering Notes, 29(6): 273, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2004:BRA

[ACM04d] ACM SIGSOFT Software Engineering Notes Staff. Backmatter (report abstracts, paper abstracts, calendar of future events, call for participation, keynotes, workshops, tutorials). ACM SIGSOFT Software Engineering Notes, 29(5):42, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2004:FCT

[ACM04e] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (cover, title page, copyright, foreword, contents, organization). ACM SIGSOFT Software Engineering Notes, 29(6):0, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:2005:BRA

[ACM05a] ACM SIGSOFT Software Engineering Notes Staff. Backmatter (report abstracts, paper abstracts, book reviews, calendar of future events). ACM SIGSOFT Software Engineering Notes, 30(1):40, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2005:FTLc

[ACM05b] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (TOC, letter from the vice-chair, letters to the editor, programmer productivity, software assurance, workshop and conference summaries, calendar of events). ACM SIGSOFT Software Engineering Notes, 30(6):0, November 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2005:FTLa

[ACM05c] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (TOC, letters, ICSE news, ICSM, RODIN). ACM SIGSOFT Software Engineering Notes, 30(1):

0, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:2005:FTL

[ACM05d] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (TOC, letters, open source software (OSS) patent search engine, calendar of events, workshop and conference information). ACM SIGSOFT Software Engineering Notes, 30(2): 0, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2006:FTL

[ACM06a] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (TOC, letter from the chair, letter from the editor, letters to the editor, ACM policy and procedures on plagiarism, PASTE abstracts, calendar of future events, workshop and conference information). ACM SIGSOFT Software Engineering Notes, 31 (1):0, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

staff:2006:FTM

[ACM06b] ACM SIGSOFT Software Engineering Notes Staff. Frontmatter (TOC, miscellaneous material). ACM SIG-SOFT Software Engineering Notes, 31(6):0, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Adams:2000:GI

[Ada00] Tom Adams. The God of inspection. ACM SIGSOFT Software Engineering Notes, 25(2):30, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agarwal:2004:IAL

[ADD04] Rakesh Agarwal, Amrita Deo, and Swati Das. Intelligent agents in e-learning. ACM SIGSOFT Software Engineering Notes, 29(2): 1, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abraham:2005:HCU

[AE05a] Robin Abraham and Martin Erwig. How to communicate unit error messages in spreadsheets. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aranda:2005:AAS

[AE05b] Jorge Aranda and Steve Easterbrook. Anchoring and adjustment in software estimation. ACM SIGSOFT Software Engineering Notes, 30 (5):346–355, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ambriola:2001:EWS

[AG01] Vincenzo Ambriola and R. Mark Greenwood. 8th European Workshop on Software Process Technology: report and future directions. ACM SIGSOFT Software Engineering Notes, 26(5): 307–308, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aberer:2001:PPI

[AH01] Karl Aberer and Manfred Hauswirth. Peer-to-peer information systems: concepts and models, state-of-the-art, and future systems. ACM SIGSOFT Software Engineering Notes, 26(5):326– 327, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Apiwattanapong:2003:SPP

[AH03] Taweesup Apiwattanapong and Mary Jean Harrold. Selective path profiling. ACM SIGSOFT Software Engineering Notes, 28(1):35– 42, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2009:SEEa

[AH09a] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 34(4):8, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2009:SEEb

[AH09b] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 34(5): 7, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2009:SEEc

[AH09c] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 34(6):5– 8, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arshad:2005:DFD

[AHW05] Naveed Arshad, Dennis Heimbigner, and Alexander L. Wolf. Dealing with failures during failure recovery of distributed systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aiken:2004:THP

[Aik04]  Jason Aiken.  Technical

and human perspectives on pair programming. ACM SIGSOFT Software Engineering Notes, 29(5):1–14, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agrawal:2009:IIV

[AK09] A. Agrawal and R. A. Khan. Impact of inheritance on vulnerability propagation at design phase. ACM SIGSOFT Software Engineering Notes, 34(4):1–5, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arnautovic:2006:AGT

[AKF06] Edin Arnautovic, Hermann Kaindl, and Jürgen Falb. An architecture for gradual transition towards selfmanaged software systems. ACM SIGSOFT Software Engineering Notes, 31(6):1– 2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Avgeriou:2007:AKR

[AKL^{+}07] Paris Avgeriou, Philippe Kruchten, Patricia Lago, Paul Grisham, and Dewayne Perry. Architectural knowledge and rationale: issues, trends, challenges. ACM SIGSOFT Software Engineering Notes, 32 (4):41–46, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alexander:2002:DSS

[Ale02] Larry Alexander. Decision support systems in the 21st century. ACM SIGSOFT Software Engineering Notes, 27(5):104, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ali:2005:WTR

[Ali05] Muhammad Raza Ali. Why teach reverse engineering? ACM SIGSOFT Software Engineering Notes, 30(4): 1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ali:2006:IES

[Ali06] Muhammad Raza Ali. Imparting effective software engineering education. ACM SIGSOFT Software Engineering Notes, 31(4):1– 3, July 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Avgeriou:2009:TUA

[ALK09] Paris Avgeriou, Patricia Lago, and Philippe Kruchten.

Towards using architectural knowledge. ACM SIGSOFT Software Engineering Notes, 34(2):27–30, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alur:2000:CAV

[Alu00] Rajeev Alur. Computeraided verification of reactive systems. ACM SIGSOFT Software Engineering Notes, 25(1):32–33, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Akgul:2003:ILR

[AM03] Tankut Akgul and Vincent J. Mooney III. Instruction-level reverse execution for debugging. ACM SIGSOFT Software Engineering Notes, 28 (1):18–25, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ammann:2000:STM

[Amm00] Paul Ammann. System testing via mutation analysis of model checking specifications. ACM SIGSOFT Software Engineering Notes, 25 (1):33, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Amorim:2006:MLS

[AMN^{+}06] Leonardo Amorim, Paulo Maciel, Meuse Nogueira, Raimundo Barreto, and Eduardo Tavares. Mapping live sequence chart to coloured Petri nets for analysis and verification of embedded systems. ACM SIGSOFT Software Engineering Notes, 31 (3):1–25, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Amrit:2005:CSD

[Amr05] Chintan Amrit. Coordination in software development: the problem of task allocation. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Auguston:2005:EBM

[AMS05] Mikhail Auguston, James Bret Michael, and Man-Tak Shing. Environment behavior models for scenario generation and testing automation. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anantharam:2000:MS

[Ana00] Parasuram Anantharam. Modelling systems. ACM SIG-SOFT Software Engineering Notes, 25(3):58, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anantharam:2001:EJP

[Ana01a] Parasuram Anantharam. Essentials of the Java programming language: a hands-on guide. ACM SIGSOFT Software Engineering Notes, 26 (2):76, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anantharam:2001:PR

[Ana01b] Parasuram Anantharam. Programming ruby. ACM SIG-SOFT Software Engineering Notes, 26(4):89, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anantharam:2002:PWD

[Ana02] Parasuram Anantharam. Practical WAP: developing applications for the wireless web. ACM SIGSOFT Software Engineering Notes, 27 (2):69, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anderson:2000:DSS

[And00] Tom Anderson. The distributed Smalltalk survival guide. ACM SIGSOFT Software Engineering Notes, 25 (3):58, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Andrews:2004:RET

[And04] James H. Andrews. Relevant empirical testing research: challenges and responses. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Al-Naeem:2005:TSO

[ANGRB05] Tariq Al-Naeem, Ian Gorton, Fethi Rabhi, and Boualem Benatallah. Tool support for optimization-based architectural evaluation. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2000:AA

[Ano00a]  Anonymous.  Acme and

AcmeStudio. ACM SIG-SOFT Software Engineering Notes, 25(1):96, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2000:AAS

[Ano00b] Anonymous. ARGUS-1: “allseeing” architectural analysis. ACM SIGSOFT Software Engineering Notes, 25 (1):97, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2000:AES

[Ano00c] Anonymous. Assessing and enhancing software testing effectiveness. ACM SIGSOFT Software Engineering Notes, 25(1):50–51, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2001:AOV

[Ano01a] Anonymous. An automated oracle for verifying GUI objects. ACM SIGSOFT Software Engineering Notes, 26 (4):83–88, July 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2001:ESP

[Ano01b] Anonymous. Estimating software projects. ACM SIG-SOFT Software Engineering Notes, 26(4):60–67, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2001:LSE

[Ano01c] Anonymous. Limits to software estimation. ACM SIG-SOFT Software Engineering Notes, 26(4):54–59, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2005:FTLa

[Ano05a]  Anonymous.  Frontmatter

(TOC, letters, election results, software reliability resources!, computing curricula 2004 and the software engineering volume SE2004, software reuse research, ICSE 2005 forward). ACM SIG-SOFT Software Engineering Notes, 30(4):0, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2005:FTLb

[Ano05b]  Anonymous.  Frontmatter

(TOC, letters, philosophy of computer science, interviewers needed, taking software requirements creation from folklore to analysis, SW components and product lines: from business to systems and technology, software engineering survey). ACM SIGSOFT Software Engineering Notes, 30 (5):0, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2007:FTL

[Ano07a]  Anonymous.  Frontmat-

ter (TOC, letters, calendar, calls). ACM SIGSOFT Software Engineering Notes, 32 (1):0, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2007:RCB

[Ano07b]  Anonymous.  Review of

Change-based Test Management by Jon Sistowicz and Ray Arell, Intel Press 2001.

ACM SIGSOFT Software Engineering Notes, 32(2): 33–34, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2007:RMT

[Ano07c]  Anonymous.  Review of

Managing the Testing Process by Rex Black, Wiley 2002. ACM SIGSOFT Software Engineering Notes, 32 (2):28–32, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2007:RPS

[Ano07d]  Anonymous.  Review of

Productive Software Test Management by Michael W. Evans, Wiley 1984. ACM SIGSOFT Software Engineering Notes, 32(2): 32, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:2007:RST

[Ano07e] Anonymous. Review of Software Testing Management by Thomas C. Royer, P-H 1993. ACM SIGSOFT Software Engineering Notes, 32 (2):32–33, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:2009:WHC

[APS09] Naveen Aggarwal, Nupur Prakash, and Sanjeev Sofat. Web hypermedia content management system effort estimation model. ACM SIG-SOFT Software Engineering Notes, 34(2):1–7, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Albizuri-Romero:2000:RVC

[AR00] Miren Begoña Albizuri-Romero. A retrospective view of CASE tools adoption. ACM SIGSOFT Software Engineering Notes, 25 (2):46–50, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arell:2007:WJR

[Are07]  Ray Arell.  Wonderful

jupiter?: (a rebuttal). ACM SIGSOFT Software Engineering Notes, 32(3):26, May 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agarwal:2000:OEL

[ARG00] Rakesh Agarwal, Arup Ratan Raha, and Bhaskar Ghosh. Our experience and learning in ERP implementation. ACM SIGSOFT Software Engineering Notes, 25 (2):31–34, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alspaugh:2005:SSM

[ARS05] Thomas A. Alspaugh, Debra J. Richardson, and Thomas A. Standish. Scenarios, state machines and purpose-driven testing. ACM SIGSOFT Software Engineering Notes, 30(4):1– 5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Antoniol:2005:LPC

[ARV05] Giuliano Antoniol, Vincenzo Fabio Rollo, and Gabriele Venturi. Linear predictive coding and cepstrum coefficients for mining time variant information from software repositories. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:2002:CPW

[ASC02] K. K. Aggarwal, Yogesh Singh, and Jitender Kumar Chhabra. Computing program weakness using module coupling. ACM SIGSOFT Software Engineering Notes, 27(1):63–65, January 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:2005:EVT

[ASCP05a] K. K. Aggarwal, Yogesh Singh, Pravin Chandra, and Manimala Puri. Evaluation of various training algorithms in a neural network model for software engineering applications. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:2005:ECM

[ASCP05b] K. K. Aggarwal, Yogesh Singh, Pravin Chandra, and Manimala Puri. An expert committee model to estimate lines of code. ACM SIGSOFT Software Engineering Notes, 30(5):1–4, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:2005:SAF

[ASCP05c] K. K. Aggarwal, Yogesh Singh, Pravin Chandra, and Manimala Puri. Sensitivity analysis of fuzzy and neural network models. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agarwal:2003:RDI

[ASD03] Rakesh Agarwal, Ajit Sarangi, and Swati Das. Reengineering of database intensive application. ACM SIGSOFT Software Engineering Notes, 28(3):1, May 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggrawal:2004:CCB

[ASK04] K. K. Aggrawal, Yogesh Singh, and A. Kaur. Code coverage based technique for prioritizing test cases for regression testing. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:2004:NNB

[ASKS04] K. K. Aggarwal, Yogesh Singh, A. Kaur, and O. P. Sangwan. A neural net based approach to test oracle. ACM SIGSOFT Software Engineering Notes, 29 (3):1–6, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Asundi:2005:NEE

[Asu05]  Jai Asundi.  The need

for effort estimation models for open source software projects. ACM SIGSOFT Software Engineering Notes, 30(4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Attie:2000:EFM

[Att00] Paul C. Attie. Efficient formal methods for the synthesis of concurrent programs. ACM SIGSOFT Software Engineering Notes, 25 (1):34, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Auguston:2009:SAB

[Aug09] Mikhail Auguston. Software architecture built from behavior models. ACM SIGSOFT Software Engineering Notes, 34(5):1–15, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anbalagan:2006:CAJ

[AX06] Prasanth Anbalagan and Tao Xie. Clamp: automated joinpoint clustering and pointcut mining in aspect-oriented refactoring. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bailin:2000:FAM

[BA00] Sidney Bailin and Dean Allemang. Formal alternative management integrating logical inference and rationals (FAMILIAR). ACM SIG-SOFT Software Engineering Notes, 25(1):102, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-anachem:2005:RSE

[Ba05a] Mordechai Ben-anachem. Review of Software Endgames: Eliminating Defects, Controlling Change, and the Countdown to On-Time Delivery by Robert Galen, Dorset House, ISBN: 0-932633-62-5, 2005. ACM SIGSOFT Software Engineering Notes, 30 (4):82–83, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bierhoff:2005:LOS

[BA05b] Kevin Bierhoff and Jonathan Aldrich. Lightweight object specification with typestates. ACM SIGSOFT Software Engineering Notes, 30 (5):217–226, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2009:PSEb

[BÁ09a] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes new column. ACM SIGSOFT Software Engineering Notes, 34 (6):1, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2009:PSEa

[BÁ09b] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes new “TBD” column. ACM SIGSOFT Software Engineering Notes, 34(5):6, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bagert:2001:FSE

[Bag01] Donald J. Bagert. FASE:

a software engineering, education, training and professional issues publication. ACM SIGSOFT Software Engineering Notes, 26(1): 6, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bagchi:2009:MSD

[Bag09] Tapan P. Bagchi. Models for software defects and testing strategies. ACM SIGSOFT Software Engineering Notes, 34(2):1–4, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balzer:2000:SWC

[Bal00] Bob Balzer. Securely wrapping COTS products. ACM SIGSOFT Software Engineering Notes, 25(1):106– 107, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2007:MFS

[Ban07] Anup Kumar Bandyopadhyay. Modeling fairness and starvation in concurrent systems. ACM SIGSOFT Software Engineering Notes, 32 (6):2:1–2:??, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2008:WPB

[Ban08] Anup Kumar Bandyopadhyay. Weakest precondition based verification tool that models spatial ordering. ACM SIGSOFT Software Engineering Notes, 33 (2):2:1–2:??, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2009:FCC

[Ban09] Anup Kumar Bandyopadhyay. Fairness and conspiracy concepts in concurrent systems. ACM SIGSOFT Software Engineering Notes, 34 (2):1–8, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barbosa:2005:RCS

[Bar05a] Marco Antonio Barbosa. A refinement calculus for software components and architectures. ACM SIG-SOFT Software Engineering Notes, 30(5):377–380, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baroni:2005:QAU

[Bar05b] Aline Lúcia Baroni. Quantitative assessment of UML dynamic models. ACM SIGSOFT Software Engineering Notes, 30(5):366– 369, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastani:2000:RP

[Bas00] Farokh B. Bastani. Relational programs. ACM SIGSOFT Software Engineering Notes, 25(1):34– 35, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastani:2007:RAF

[Bas07] Behzad Bastani. A requirements analysis framework for open systems requirements engineering. ACM SIGSOFT Software Engineering Notes, 32(2):1–19, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastani:2008:POA

[Bas08] Behzad Bastani. Processoriented abstraction of the complex evolvable systems: problem model construction. ACM SIGSOFT Software Engineering Notes, 33(3): 1:1–1:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Batory:2000:JTS

[Bat00] Don Batory. Jakarta Tool Suite (JTS). ACM SIGSOFT Software Engineering Notes, 25(1):103–104, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bachmann:2001:MVS

[BB01] Felix Bachmann and Len Bass. Managing variability in software architectures. ACM SIGSOFT Software Engineering Notes, 26(3): 126–132, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastani:2007:HLO

[BB07a] Behzad Bastani and Hoda Bastani. High-level open evolvable systems design by process-oriented modeling: application to DNA replication mechanism. ACM SIGSOFT Software Engineering Notes, 32(6):3:1– 3:??, November 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Budgen:2007:REB

[BB07b] David Budgen and Pearl Brereton. Realising Evidence-Based Software Engineering (REBSE-2): a report from the workshop held at ICSE 2007. ACM SIGSOFT Software Engineering Notes, 32 (4):36–39, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bass:2008:LMS

[BB08a] Len Bass and Brian Berenbach. Leadership and Management in Software Architecture (LMSA’08): a report on an ICSE workshop. ACM SIGSOFT Software Engineering Notes, 33 (4):27–29, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borsoi:2008:DMP

[BB08b] Beatriz Terezinha Borsoi and Jorge Luis Risco Becerra. Definition and modeling of process using object orientation. ACM SIGSOFT Software Engineering Notes, 33 (3):2:1–2:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bass:2009:LMS

[BB09] Len Bass and Brian Berenbach. Leadership and Management in Software Architecture Workshop 2009 report. ACM SIGSOFT Software Engineering Notes, 34 (4):25–27, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhattacharyya:2009:VGN

[BBG^{+}09] Indrajit Bhattacharyya, Anup Kumar Bandypopadhyay, Bhaskar Gupta, Aloknath Chattopadhyay, Rajeswari Chattopadhyay, and Kiyotoshi Yasumoto. Vector GA: a novel enhancement of genetic algorithms for efficient multivariable or multi-dimensional search. ACM SIGSOFT Software Engineering Notes, 34 (6):1–5, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baresel:2004:ETP

[BBHK04] André Baresel, David Binkley, Mark Harman, and Bogdan Korel. Evolutionary testing in the presence of loopassigned flags: a testability transformation approach. ACM SIGSOFT Software Engineering Notes, 29(4): 108–118, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2007:ADW

[BBM07a] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. Application of Dijkstra’s weakest precondition calculus to Dining Philosophers problem. ACM SIGSOFT Software Engineering Notes, 32 (4):2:1–2:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2007:CIT

[BBM07b] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. On the correctness issues in two-process mutual exclusion algorithms. ACM SIGSOFT Software Engineering Notes, 32(6):1:1– 1:??, November 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2007:OET

[BBM07c] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. Ordering of events in twoprocess concurrent system. ACM SIGSOFT Software Engineering Notes, 32(4): 1:1–1:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barreto:2005:SSP

[BBW05] Ahilton Barreto, Márcio Barros, and Cláudia Werner. Staffing a software project: a constraint satisfaction approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bruns:2002:SPA

[BC02] Glenn Bruns and Satish Chandra. Searching for points-to analysis. ACM SIGSOFT Software Engineering Notes, 27(6):61– 70, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastos:2005:RMA

[BC05a] Lucia R. D. Bastos and Jaelson F. B. Castro. From requirements to multi-agent architecture using organisational concepts. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bryce:2005:TPP

[BC05b] Renée C. Bryce and Charles J. Colbourn. Test prioritization for pairwise interaction coverage. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernardo:2000:FAT

[BCD00] Marco Bernardo, Paolo Ciancarini, and Lorenzo Donatiello. On the formalization of architectural types with process algebras. ACM SIGSOFT Software Engineering Notes, 25(6):140– 148, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bradbury:2006:EFC

[BCD06] Jeremy S. Bradbury, James R.

Cordy, and Juergen Dingel. An empirical framework for comparing effectiveness of testing and property-based formal analysis. ACM SIGSOFT Software Engineering Notes, 31 (1):2–5, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brim:2006:CIA

[BČVZ06] Luboš Brim, Ivana Černá, Pavlína Vařeková, and Barbora Zimmerova. Componentinteraction automata as a verification-oriented componentbased system specification. ACM SIGSOFT Software Engineering Notes, 31(2): 4:1–4:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bayrak:2003:LAN

[BD03] Coskun Bayrak and Chad Davis. The liquid architecture: a non-linear peerto-peer distributed architecture with polymorphic message passing. ACM SIG-SOFT Software Engineering Notes, 28(3):2, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:2003:SFE

[BDE^{+}03] Rhodes Brown, Karel Driesen, David Eng, Laurie Hendren, John Jorgensen, Clark Verbrugge, and Qin Wang. STEP: a framework for the efficient encoding of general trace data. ACM SIGSOFT Software Engineering Notes, 28(1):27–34, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bruschi:2005:SES

[BDM05] Danilo Bruschi, Bart De Win, and Mattia Monga. Software engineering for secure systems. ACM SIGSOFT Software Engineering Notes, 30(4):1–2, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bontemps:2005:SSM

[BE05] Yves Bontemps and Alexander Egyed. Scenarios and State Machines: Models, Algorithms, and Tools: a summary of the 4th workshop. ACM SIGSOFT Software Engineering Notes, 30(5):1– 4, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beasley:2004:RBO

[Bea04] Charles Beasley. Review of Building operational excellence: IT people and best practices by Bruce Allen and Dale Kutnic. Intel Press and Addison-Wesley 2001 and 2002. ACM SIGSOFT Software Engineering Notes, 29 (5):35, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beasley:2005:RBA

[Bea05] Charles Beasley. Review of Balancing Agility and Discipline: a Guide for the Perplexed by Barry Boehm and Richard Turner; Addison Wesley, 2004, ISBN 0-321-18612-5. ACM SIGSOFT Software Engineering Notes, 30(2):28, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benedicenti:2000:RSO

[Ben00] Luigi Benedicenti. Rethinking smart objects: building artificial intelligence with objects. ACM SIGSOFT Software Engineering Notes, 25 (3):59, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benedicenti:2002:SCM

[Ben02]  Luigi Benedicenti.  Soft-

ware configuration management strategies and rational ClearCase(r). ACM SIG-SOFT Software Engineering Notes, 27(1):67, January 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bertolino:2002:IPI

[Ber02] Antonia Bertolino. ISSTA 2002 panel: is ISSTA research relevant to industrial users? ACM SIGSOFT Software Engineering Notes, 27 (4):201–202, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bertolino:2004:IML

[Ber04]  Antonia Bertolino.  The

(im)maturity level of software testing. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2005:TSS

[Ber05]  Larry Bernstein.  Trust-

worthy software systems. ACM SIGSOFT Software Engineering Notes, 30(1): 4–5, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2006:RCE

[Ber06a] Larry Bernstein. Review of Competitive Engineering by Tom Gilb, Elsevier, 2005. ISBN: 0-7506-6507-6. ACM SIGSOFT Software Engineering Notes, 31 (1):28, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2006:RMC

[Ber06b] Larry Bernstein. Review of Mobile Computing Principles by Reza B’Far, Cambridge University Press, 2005, ISBN: 0-521-81733-1. ACM SIGSOFT Software Engineering Notes, 31 (3):41–42, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2006:RPS

[Ber06c] Larry Bernstein. Review of Practical Support for CMMI-SW Software Project Documentation Using IEEE Software Engineering Standards by Susan K. Lan and John W. Walz, IEEE Computer Society Wiley Interscience, 2006, ISBN: 0-471-73849-2. ACM SIGSOFT Software Engineering Notes, 31 (3):41, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berzal:2006:CPL

[Ber06d] Fernando Berzal. A crossword on programming languages. ACM SIGSOFT Software Engineering Notes, 31(5):5, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berkun:2008:APM

[Ber08] Scott Berkun. The art of project management. ACM SIGSOFT Software Engineering Notes, 33(5):12:1– 12:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bartelt:2005:DIH

[BFN^{+}05] Christian Bartelt, Thomas Fischer, Dirk Niebuhr, Andreas Rausch, Franz Seidl, and Marcus Trapp. Dynamic integration of heterogeneous mobile devices. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baillargeon:2009:WMS

[BFZ^{+}09] Robert Baillargeon, Robert France, Steffen Zschaler, Bernhard Rumpe, Steven Völkel, and Geri Georg. Workshop on Modeling in Software Engineering at ICSE 2009. ACM SIGSOFT Software Engineering Notes, 34(4):34–37, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brebner:2003:HSJ

[BG03] Paul Brebner and Jeffrey Gosper. How scalable is J2EE technology? ACM SIGSOFT Software Engineering Notes, 28(3):4, May 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastani:2009:COS

[BG09] Behzad Bastani and David Greaves. Complex opensystem design by quasiagents: process-oriented modeling in agent-based systems. ACM SIGSOFT Software Engineering Notes, 34 (4):1–14, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhargavan:2000:VFA

[BGK^{+}00] Karthikeyan Bhargavan, Carl A.

Gunter, Moonjoo Kim, Insup Lee, Davor Obradovic, Oleg Sokolsky, and Mahesh Viswanathan. Verisim: Formal analysis of network simulations. ACM SIGSOFT Software Engineering Notes, 25(5):2–13, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blundell:2006:AGT

[BGP06] Colin Blundell, Dimitra Giannakopoulou, and Corina S. Păsăreanu. Assumeguarantee testing. ACM SIGSOFT Software Engineering Notes, 31(2):1:1– 1:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beydoun:2005:SGM

[BGPLHS05] Ghassan Beydoun, Cesar Gonzalez-Perez, Graham Low, and Brian Henderson-Sellers. Synthesis of a generic MAS metamodel. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Binder:2005:AMA

[BH05] Robert V. Binder and James E. Hanlon. The advanced mobile application testing environment. ACM SIGSOFT Software Engineering Notes, 30(4):1, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhansali:2005:CMD

[Bha05a] P. V. Bhansali. Complexity measurement of data and control flow. ACM SIGSOFT Software Engineering Notes, 30(1):1, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhansali:2005:SDD

[Bha05b] P. V. Bhansali. Software dissimilarity debate revisited. ACM SIGSOFT Software Engineering Notes, 30 (1):2, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhansali:2005:SSC

[Bha05c] P. V. Bhansali. Software safety: current status and future direction. ACM SIG-SOFT Software Engineering Notes, 30(1):3, January 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhansali:2005:USS

[Bha05d] P. V. Bhansali. Universal software safety standard. ACM SIGSOFT Software Engineering Notes, 30(5):1– 4, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatti:2005:WQI

[Bha05e] Shahid Nazir Bhatti. Why quality?: ISO 9126 software quality metrics (functionality) support by UML suite. ACM SIGSOFT Software Engineering Notes, 30 (2):1–5, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhansali:2007:MP

[Bha07] P. V. Bhansali. The MCDC paradox. ACM SIGSOFT Software Engineering Notes, 32(3):1–4, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatti:2009:DCQ

[Bha09] Shahid Nazir Bhatti. Deducing the complexity to quality of a system using UML. ACM SIGSOFT Software Engineering Notes, 34 (3):1–7, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bauer:2005:FIS

[BHLRM05] Thomas Bauer, Jens Herrmann, Peter Liggesmeyer, and Christopher Robinson-Mallett. A flexible integration strategy for in-car telematics systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Becker:2006:TSS

[BHP^{+}06] Steffen Becker, Wilhelm Hasselbring, Alexandra Paul, Marko Boskovic, Heiko Koziolek, Jan Ploski, Abhishek Dhama, Henrik Lipskoch, Matthias Rohr, Daniel Winteler, Simon Giesecke, Roland Meyer, Mani Swaminathan, Jens Happe, Margarete Muhle, and Timo Warns. Trustworthy software systems: a discussion of basic concepts and terminology. ACM SIGSOFT Software Engineering Notes, 31 (6):1–18, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bielikova:2002:WWM

[Bie02] Mária Bieliková. Wireless Web: a manager’s guide. ACM SIGSOFT Software Engineering Notes, 27(2): 69–70, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Billard:2003:LDP

[Bil03] Edward A. Billard. Languagedependent performance of design patterns. ACM SIG-SOFT Software Engineering Notes, 28(3):3, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Binkley:2000:RCR

[Bin00] David Binkley. Reducing the cost of regression testing. ACM SIGSOFT Software Engineering Notes, 25 (1):35–36, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bishop:2002:RRB

[Bis02] Peter G. Bishop. Rescaling reliability bounds for a new operational profile. ACM SIGSOFT Software Engineering Notes, 27(4):

180–190, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Basit:2005:DHL

[BJ05] Hamid Abdul Basit and Stan Jarzabek. Detecting higher-level similarity patterns in programs. ACM SIGSOFT Software Engineering Notes, 30(5):156– 165, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bouquet:2005:RTA

[BJL^{+}05] F. Bouquet, E. Jaffuel, B. Legeard, F. Peureux, and M. Utting. Requirements traceability in automated test generation: application to smart card software validation. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2001:TOS

[BK01] Lawrence Bernstein and David Klappholz. Teaching old software dogs, old tricks. ACM SIGSOFT Software Engineering Notes, 26 (2):33–34, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Biyani:2005:BCF

[BK05a] Karun N. Biyani and Sandeep S.

Kulkarni. Building component families to support adaptation. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Budgen:2005:REB

[BK05b] David Budgen and Barbara Kitchenham. Realising evidence-based software engineering: a report from the workshop held at ICSE 2005. ACM SIGSOFT Software Engineering Notes, 30(5):1– 5, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Babu:2009:FMT

[BK09] Chitra Babu and Harshini Ramnath Krishnan. Fault model and test-case generation for the composition of aspects. ACM SIGSOFT Software Engineering Notes, 34(1): 1–6, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Botaschanjan:2005:TVA

[BKKS05] J. Botaschanjan, L. Kof, C. Kühnel, and M. Spichkova. Towards verified automotive software. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyapati:2002:KAT

[BKM02] Chandrasekhar Boyapati, Sarfraz Khurshid, and Darko Marinov. Korat: automated testing based on Java predicates. ACM SIGSOFT Software Engineering Notes, 27 (4):123–133, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatt:2006:IFO

[BKSM06] Pankaj Bhatt, Williams K., Gautam Shroff, and Arun K. Misra. Influencing factors in outsourced software maintenance. ACM SIGSOFT Software Engineering Notes, 31 (3):1–6, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berry:2004:SIW

[BKW04] Daniel M. Berry, Rick Kazman, and Roel Wieringa. Second International Workshop on From SofTware Requirements to Architectures (STRAW’03). ACM SIG-SOFT Software Engineering Notes, 29(3):1–5, May 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Briand:2004:ESS

[BL04] Lionel Briand and Yvan Labiche. Empirical studies of software testing techniques: challenges, practical strategies, and future research. ACM SIGSOFT Software Engineering Notes, 29(5):1–3, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barnett:2006:WPU

[BL06] Mike Barnett and K. Rustan M. Leino. Weakestprecondition of unstructured programs. ACM SIGSOFT Software Engineering Notes, 31(1):82–87, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blazer:2000:CBD

[Bla00]  Bob Blazer.  A COTS-

based design editor for user specified domains. ACM SIGSOFT Software Engineering Notes, 25(1):100, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bidinger:2006:DTD

[BLQ^{+}06] Philippe Bidinger, Matthieu Leclercq, Vivien Quéma, Alan Schmitt, and Jean-Bernard Stefani. Dream types: a domain specific type system for componentbased message-oriented middleware. ACM SIGSOFT Software Engineering Notes, 31(2):2:1–2:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Briand:2002:IUA

[BLS02] L. C. Briand, Y. Labiche, and H. Sun. Investigating the use of analysis contracts to support fault isolation in object oriented code. ACM SIGSOFT Software Engineering Notes, 27 (4):70–80, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:2000:MBS

[BM00a] Barry Boehm and Neno Medvidovic. Model-based (systems) architecting and software engineering (MBASE). ACM SIGSOFT Software Engineering Notes, 25(1): 104–105, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:2000:SAA

[BM00b] Barry Boehm and Neno Medvidovic. Software architecture, analysis, generation, and evolution (SAAGE). ACM SIGSOFT Software Engineering Notes, 25(1): 106, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2001:WEU

[BM01] Mordechai Ben-Menachem.

Writing effective use cases. ACM SIGSOFT Software Engineering Notes, 26(1):94– 95, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2003:SYH

[BM03] Mordechai Ben-Menachem.

Software for your head. ACM SIGSOFT Software Engineering Notes, 28(3): 24, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2004:RCG

[BM04a] Mordechai Ben-Menachem.

Review of CMMI: guidelines for process integration and product improvement by Mary Beth Chrissis, Mike Konrad and Sandy Shrum. Addison Wesley 2003. ACM SIGSOFT Software Engineering Notes, 29(5):37– 38, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2004:RIP

[BM04b] Mordechai Ben-Menachem.

Review of IT project estimation: a practical guide to the costing of software by Paul Coombs. Cambridge University Press 2003. ACM SIG-SOFT Software Engineering Notes, 29(2):31, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2004:RSM

[BM04c] Mordechai Ben-Menachem. Review of Software metrics: a guide to planning, analysis and application by C. Ravindranath Pandian. Auerbach Publications, CRC Press 2004. ACM SIGSOFT Software Engineering Notes, 29(5):38–39, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2005:RIC

[BM05a] Mordechai Ben-Menachem. Review of Interpreting the CMMI: a Process Improvement Approach by Margaret K. Kulpa and Kent A. Johnson. Auerbach Publications 2003. ACM SIGSOFT Software Engineering Notes, 30 (1):33–34, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2005:RRS

[BM05b] Mordechai Ben-Menachem. Review of Return on Software by Steve Tockey; Addison Wesley, 2005. ACM SIGSOFT Software Engineering Notes, 30(3):39– 40, May 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2005:RSS

[BM05c] Mordechai Ben-Menachem.

Review of Six Sigma Software Development by Christine B. Tayntor. Auerbach Publications, 2003. ACM SIGSOFT Software Engineering Notes, 30(1):34, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2005:RTE

[BM05d] Mordechai Ben-Menachem.

Review of Testing Embedded Software by Bart Broekman and Edwin Notenboom; Addison Wesley.; 2003. ACM SIGSOFT Software Engineering Notes, 30 (2):30, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bolin:2005:NPE

[BM05e] Michael Bolin and Robert C.

Miller. Naming page elements in end-user web automation. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2006:RTS

[BM06] Mordechai Ben-Menachem. Review of Trustworthy Systems Through Quantitative Software Engineering by Lawrence Bernstein and C. M. Yuhas, IEEE Computer Society Press, 2005. ACM SIGSOFT Software Engineering Notes, 31(1):27– 28, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2007:RSM

[BM07a] Mordechai Ben-Menachem. Review of Software Measurement and Estimation: a Practical Approach by Linda M. Laird and M. Carol Brennan, Wiley-Interscience, 2006. ACM SIGSOFT Software Engineering Notes, 32 (1):47, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2007:RTC

[BM07b] Mordechai Ben-Menachem. Review of TSP: Coaching Development Teams by Watts S. Humphrey, Addison-Wesley, 2006. ACM SIGSOFT Software Engineering Notes, 32 (4):53, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Basci:2008:EMX

[BM08a] Dilek Basci and Sanjay Misra. Entropy metric for XML DTD documents. ACM SIGSOFT Software Engineering Notes, 33(4): 5:1–5:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2008:RML

[BM08b] Mordechai Ben-Menachem.

Review of Modal Logic for Philosophers by James W. Garson, Cambridge University Press, 2006, pp. 471. ACM SIGSOFT Software Engineering Notes, 33(3): 8:1–8:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2008:RPS

[BM08c] Mordechai Ben-Menachem.

Review of Practical Support for ISO 9001 Software Project Documentation by Susan K. Land and John W. Walz, IEEE Computer Society and Wiley-Interscience, 2006 (by IEEE CS), ISBN-13 978-0-471-76867-8, pp.

432. ACM SIGSOFT Software Engineering Notes, 33 (3):9:1–9:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2009:BRS

[BM09a] Mordechai Ben-Menachem. Book review: Software Maintenance Management by Alain April and Alain Abran, and published by Wiley Interscience (John Wiley and Sons, Inc.), 2008, (paperback), ISBN 978-0470-14707-8, pp. 334. ACM SIGSOFT Software Engineering Notes, 34(5):59–60, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatti:2009:XBF

[BM09b] Shahid Nazir Bhatti and Asif Muhammad Malik. An XML-based framework for bidirectional transformation in model-driven architecture (MDA). ACM SIGSOFT Software Engineering Notes, 34(3):1–5, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Booch:2008:OOA

[BME^{+}08] Grady Booch, Robert A. Maksimchuk, Michael W. Engle, Bobbi J. Young, Jim Connallen, and Kelli A. Houston. Object-oriented analysis and design with applications, third edition. ACM SIGSOFT Software Engineering Notes, 33(5): 11:1–11:??, September 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balasubramaniam:2005:SAA

[BMK^{+}05] Dharini Balasubramaniam, Ron Morrison, Graham Kirby, Kath Mickan, Brian Warboys, Ian Robertson, Bob Snowdon, R. Mark Greenwood, and Wykeen Seet. A software architecture approach for structuring autonomic systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2005:LAM

[BMMG05] Somo Banerjee, Chris A.

Mattmann, Nenad Medvidovic, and Leana Golubchik. Leveraging architectural models to inject trust into software systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Biswas:2009:MBR

[BMSS09] Swarnendu Biswas, Rajib Mall, Manoranjan Satpathy, and Srihari Sukumaran. A model-based regression test selection approach for embedded applications. ACM SIG-SOFT Software Engineering Notes, 34(4):1–9, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Braga:2001:UMO

[BMW01] Regina M. M. Braga, Marta Mattoso, and Cláudia M. L. Werner. The use of mediation and ontology technologies for software component information retrieval. ACM SIGSOFT Software Engineering Notes, 26(3): 19–28, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bajracharya:2005:UNO

[BNL05] Sushil Krishna Bajracharya, Trung Chi Ngo, and Cristina Videira Lopes. On using net options value as a value based design framework. ACM SIGSOFT Software Engineering Notes, 30(4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:2005:VBQ

[Boe05a] Barry Boehm. Value-based quality processes and results. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boetticher:2005:NNS

[Boe05b] Gary D. Boetticher. Nearest neighbor sampling for better defect prediction. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowring:2003:MDS

[BOH03] Jim Bowring, Alessandro Orso, and Mary Jean Harrold. Monitoring deployed software using software tomography. ACM SIGSOFT Software Engineering Notes, 28(1):2–9, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boothe:2000:FCB

[Boo00] Bob Boothe. A fully capable bidirectional debugger. ACM SIGSOFT Software Engineering Notes, 25 (1):36–37, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Buy:2000:ATC

[BOP00] Ugo Buy, Alessandro Orso, and Mauro Pezze. Automated testing of classes. ACM SIGSOFT Software Engineering Notes, 25(5):39– 48, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

BenAchour-Salinesi:2002:RWS

[BOR02] Camille Ben Achour-Salinesi, Andreas Opdahl, and Matti Rossi. REFSQ ’2001 workshop summary: Seventh International Workshop on Requirements Engineering: Foundations for Software Quality. ACM SIGSOFT Software Engineering Notes, 27(2):35–49, March 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowen:2001:ETZ

[Bow01] Jonathan P. Bowen. Experience teaching z with tool and web support. ACM SIGSOFT Software Engineering Notes, 26(2):69– 75, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowdidge:2005:RGU

[Bow05] Robert W. Bowdidge. Refactoring gcc using structure field access traces and concept analysis. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:2000:TPI

[Boy00] Kenneth W. Boyer, Jr. Test process improvement: a practical step-by-step guide to structured testing. ACM SIGSOFT Software Engineering Notes, 25(3):59– 60, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:2001:FPA

[Boy01a] Kenneth W. Boyer, Jr. Function point analysis: measurement practices for successful software projects. ACM SIG-SOFT Software Engineering Notes, 26(4):90, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:2001:MCG

[Boy01b] Kenneth W. Boyer, Jr. More C++ gems. ACM SIGSOFT Software Engineering Notes, 26(1):95, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:2002:AUC

[Boy02a] Kenneth W. Boyer, Jr. Advanced use case modeling: software systems. ACM SIG-SOFT Software Engineering Notes, 27(2):71, March 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:2002:CD

[Boy02b] Kenneth W. Boyer, Jr.

CMMI distilled. ACM SIG-SOFT Software Engineering Notes, 27(2):70, March 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Breech:2005:FTS

[BP05] Ben Breech and Lori Pollock.

A framework for testing security mechanisms for programbased attacks. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boroday:2005:DAJ

[BPSH05] S. Boroday, A. Petrenko, J. Singh, and H. Hallal. Dynamic analysis of Java applications for multithreaded antipatterns. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bo:2002:EOO

[BQX02] Gao Bo, Fei Qi, and Chen Xueguang. Exploiting objectoriented methodologies to model MAS on organizations. ACM SIGSOFT Software Engineering Notes, 27 (1):58–62, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brannum:2000:CPA

[Bra00] Garry Brannum. Capability packages for avionics software (CPAS). ACM SIGSOFT Software Engineering Notes, 25(1):98, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brada:2001:CRI

[Bra01] Přemysl Brada. Component revision identification based on IDL/ADL component specification. ACM SIGSOFT Software Engineering Notes, 26(5):297– 298, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Butkevich:2000:CTS

[BRBY00] Sergey Butkevich, Marco Renedo, Gerald Baumgartner, and Michal Young. Compiler and tool support for debugging object protocols. ACM SIGSOFT Software Engineering Notes, 25(6):50– 59, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barrera:2005:ILS

[BRC^{+}05] Pablo Barrera, Gregorio Robles, José M. Cañas, Francisco Martín, and Vicente Matellán. Impact of libre software tools and methods in the robotics field. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Breen:2005:CCR

[Bre05] Michael Breen. Combinatorial completion by rule definition with interactive value colouring. ACM SIGSOFT Software Engineering Notes, 30(2):1–6, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Breu:2006:REB

[Bre06]  Silvia Breu.  Review of

Eclipse — Building Commercial-Quality Plug-Ins by Eric Clayberg and Dan Rubel, Addison-Wesley, 2004, 0-321-22847-2. ACM SIGSOFT Software Engineering Notes, 31(4):39, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ren:2009:GTB

[bReWgH^{+}09] Sheng bing Ren, Xi e Wang, Zhi gang Hu, Ge Hu, and Guo jun Wang. Graph transformation based reduction analysis of PID. ACM SIGSOFT Software Engineering Notes, 34(4):1–5, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowring:2004:ALA

[BRH04] James F. Bowring, James M. Rehg, and Mary Jean Harrold. Active learning for automatic classification of software behavior. ACM SIGSOFT Software Engineering Notes, 29(4):195– 205, July 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brooks:2006:SPI

[Bro06] Andy Brooks. Software process improvement uncut: interviews. ACM SIGSOFT Software Engineering Notes, 31(6):1–11, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Basin:2001:FDM

[BRV01] David Basin, Frank Rittinger, and Luca Viganò. A formal data-model of the CORBA security service. ACM SIGSOFT Software Engineering Notes, 26 (5):303–304, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Behrends:2000:UMA

[BS00] Reimer Behrends and R. E. Kurt Stirewalt. The universe model: an approach for improving the modularity and reliability of concurrent programs. ACM SIGSOFT Software Engineering Notes, 25 (6):20–29, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bird:2001:XBA

[BS01] Colin Bird and Andrew Sermon. An XML-based approach to automated software testing. ACM SIGSOFT Software Engineering Notes, 26(2):64–65, March 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beaver:2006:EDT

[BS06] Justin M. Beaver and Guy A. Schiavone. The effects of development team skill on software product quality. ACM SIGSOFT Software Engineering Notes, 31 (3):1–5, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatt:2004:DSM

[BSM04] Pankaj Bhatt, Gautam Shroff, and Arun K. Misra. Dynamics of software maintenance. ACM SIGSOFT Software Engineering Notes, 29(5):1–5, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bajaj:2006:SEF

[BTA06] Nonika Bajaj, Alok Tyagi, and Rakesh Agarwal. Software estimation: a fuzzy approach. ACM SIGSOFT Software Engineering Notes, 31(3):1–5, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bultan:2000:CMC

[Bul00] Tevfik Bultan. A composite model checking toolset for analyzing software systems. ACM SIGSOFT Software Engineering Notes, 25 (1):37–38, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Burnett:2000:NAV

[Bur00] Margaret M. Burnett. NYI award: visual programming languages. ACM SIGSOFT Software Engineering Notes, 25(1):38, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Butler:2001:GTP

[But01] Greg Butler. Generative techniques for product lines. ACM SIGSOFT Software Engineering Notes, 26(6):74– 76, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bojic:2000:REU

[BV00] Dragan Bojic and Dusan Velasevic. Reverse engineering of use case realizations in UML. ACM SIGSOFT Software Engineering Notes, 25 (4):56–61, July 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Babu:2008:MBD

[BV08] Chitra Babu and R. Vijayalakshmi. Metrics-based design selection tool for aspect oriented software development. ACM SIGSOFT Software Engineering Notes, 33 (5):4:1–4:??, September 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bevan:2005:FSE

[BWKG05] Jennifer Bevan, E. James Whitehead, Jr., Sunghun Kim, and Michael Godfrey. Facilitating software evolution research with kenyon. ACM SIGSOFT Software Engineering Notes, 30(5): 177–186, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Butler:2001:CRF

[BX01] Greg Butler and Lugang Xu. Cascaded refactoring for framework. ACM SIGSOFT Software Engineering Notes, 26(3):51–57, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beranek:2005:FGR

[BZG05] Georgine Beranek, Wolfgang Zuser, and Thomas Grechenig. Functional group roles in software engineering teams. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Caci:2000:TOO

[Cac00] Claude Caci. Testing objectoriented systems. ACM SIG-SOFT Software Engineering Notes, 25(3):60–61, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Camara:2005:IPC

[Cam05] António S. Camara. Innovations in pervasive computing. ACM SIGSOFT Software Engineering Notes, 30(5):176, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Capra:2004:EHT

[Cap04] Licia Capra. Engineering human trust in mobile system collaborations. ACM SIGSOFT Software Engineering Notes, 29(6):107– 116, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carver:2000:SBT

[Car00] Richard H. Carver. Specificationbased testing of concurrent programs. ACM SIGSOFT Software Engineering Notes, 25(1):39–40, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carver:2007:PWR

[Car07]  Jeffrey Carver.  Post-

workshop report for the Third International Workshop on Software Engineering for High Performance Computing Applications (SE-HPC 07). ACM SIGSOFT Software Engineering Notes, 32(5):38–43, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carver:2009:RSI

[Car09] Jeffrey C. Carver. Report from the Second International Workshop on Software Engineering for Computational Science and Engineering (SE-CSE 09). ACM SIGSOFT Software Engineering Notes, 34(5):48– 51, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Catal:2009:BAS

[Cat09]  Cagatay Catal.  Barriers

to the adoption of software product line engineering. ACM SIGSOFT Software Engineering Notes, 34 (6):1–4, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cordeiro:2007:TAH

[CBB^{+}07] Lucas Cordeiro, Raimundo Barreto, Rafael Barcelos, Meuse Oliveira, Vicente Lucena, and Paulo Maciel. TXM: an agile HW/SW development methodology for building medical devices. ACM SIGSOFT Software Engineering Notes, 32(6): 4:1–4:??, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:2005:MCO

[CC05] Kendra Cooper and Lawrence Chung. Managing change in an OTS-aware requirements engineering approach. ACM SIGSOFT Software Engineering Notes, 30(4): 1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Calvo-ManzanoVillalon:2009:SAR

[CCHS09] Jose A. Calvo-Manzano Villalón, Gonzalo Cuevas Agustín, Gloria Gasca Hurtado, and Tomás San Feliu Gilabert. State of the art for risk management in software acquisition. ACM SIGSOFT Software Engineering Notes, 34(4):1–10, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cardoso:2004:ASS

[CCK04] Ana Isabel Cardoso, Rui Gustavo Crespo, and Peter Kokol. Assessing software structure by entropy and information density. ACM SIG-SOFT Software Engineering Notes, 29(2):2, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Canfora:2005:CCF

[CCN^{+}05] Gerardo Canfora, Piero Corte, Antonio De Nigro, Debora Desideri, Massimiliano Di Penta, Raffaele Esposito, Amedeo Falanga, Gloria Renna, Rita Scognamiglio, Francesco Torelli, Maria Luisa Villani, and Paolo Zampognaro. The C-Cube framework: developing autonomic applications through web services. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cobleigh:2000:VPP

[CCO00] Jamieson M. Cobleigh, Lori A. Clark, and Leon J. Osterweil. Verifying properties of process definitions. ACM SIGSOFT Software Engineering Notes, 25(5):96– 101, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chu-Carroll:2000:CBB

[CCS00] Mark C. Chu-Carroll and Sara Sprenkle. Coven: brewing better collaboration through software configuration management. ACM SIGSOFT Software Engineering Notes, 25(6):88– 97, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chu-Carroll:2002:SAF

[CCWS02] Mark C. Chu-Carroll, James Wright, and David Shields. Supporting aggregation in fine grained software configuration management. ACM SIGSOFT Software Engineering Notes, 27(6):99– 108, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Conradi:2001:ESU

[CD01] Reidar Conradi and Tore Dybå. An empirical study on the utility of formal routines to transfer knowledge and experience. ACM SIGSOFT Software Engineering Notes, 26(5):268– 276, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chays:2000:FTD

[CDF^{+}00] David Chays, Saikat Dan, Phyllis G. Frankl, Filippos I. Vokolos, and Elaine J. Weyuker. A framework for testing database applications. ACM SIGSOFT Software Engineering Notes, 25 (5):147–157, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Costa:2005:RBE

[CdOBT05] Hélio R. Costa, Marcio de O. Barros, and Guilherme H. Travassos. A risk based economical approach for evaluating software project portfolios. ACM SIGSOFT Software Engineering Notes, 30(4):1– 5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chelf:2003:HWS

[CEH03] Benjamin Chelf, Dawson Engler, and Seth Hallem. How to write system-specific, static checkers in metal. ACM SIGSOFT Software Engineering Notes, 28(1):51– 60, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clements:2007:AAD

[CEHK07] Paul Clements, David Emery, Rich Hilliard, and Philippe Kruchten. Aspects in Architectural Description: report on a first workshop at AOSD 2007. ACM SIGSOFT Software Engineering Notes, 32 (4):33–35, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Capra:2002:MEA

[CEM02] Licia Capra, Wolfgang Emmerich, and Cecilia Mascolo. A micro-economic approach to conflict resolution in mobile computing. ACM SIGSOFT Software Engineering Notes, 27 (6):31–40, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carter:2000:CCR

[CF00] Larry Carter and Jeanne Ferrante. CROPS: coordinated restructuring of programs and storage. ACM SIGSOFT Software Engineering Notes, 25(1):38–39, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chung:2005:SIW

[CFM05] Lawrence Chung, Xavier Franch, and Neil Maiden. Second International Workshop on Models and Processes for the Evaluation of Off-the-Shelf Components (MPEC’05). ACM SIGSOFT Software Engineering Notes, 30(5):1–3, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carlson:2000:E

[CG00] Bill Carlson and Chris Garrity. Ewatch. ACM SIG-SOFT Software Engineering Notes, 25(1):98, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chhabra:2009:EOO

[CG09] Jitender Kumar Chhabra and Varun Gupta. Evaluation of object-oriented spatial complexity measures. ACM SIG-SOFT Software Engineering Notes, 34(3):1–5, May 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ceballos:2006:CST

[CGB06] Rafael Ceballos, Rafael Martinez Gasca, and Diana Borrego. Constraint satisfaction techniques for diagnosing errors in design by contract software. ACM SIGSOFT Software Engineering Notes, 31(2):11:1– 11:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choren:2004:SEL

[CGL^{+}04] Ricardo Choren, Alessandro Garcia, Carlos Lucena, Martin Griss, David Kung, Naftaly Minsky, Alexander Romanovsky, Jaelson Castro, Rogério de Lemos, and Danny Weyns. Software engineering for large-scale multi-agent systems — SEL-MAS 2004: workshop report. ACM SIGSOFT Software Engineering Notes, 29 (5):1–10, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chaki:2004:EAC

[CGS04] Sagar Chaki, Alex Groce, and Ofer Strichman. Explaining abstract counterexamples. ACM SIGSOFT Software Engineering Notes, 29(6):73–82, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chawner:2005:FOL

[Cha05] Brenda Chawner. F/OSS in the library world: an exploration. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Conklin:2005:CUO

[CHC05] Megan Conklin, James Howison, and Kevin Crowston. Collaboration using OSSmole: a repository of FLOSS data and analyses. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheng:2000:IOO

[Che00] Betty H. C. Cheng. Integrating object-oriented analysis and formal specifications. ACM SIGSOFT Software Engineering Notes, 25 (1):40–41, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheng:2002:RDM

[Che02] Yung-Pin Cheng. Refactoring design models for inductive verification. ACM SIGSOFT Software Engineering Notes, 27(4):164–168, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:RAM

[Che05a] Haoyang Che. Review of Agile Modeling: Effective Practice for eXtreme Programming and the Unified Process by Scott W. Ambler, John Wiley and Sons, Inc, 2002, 0-471-20282-7. ACM SIGSOFT Software Engineering Notes, 30(4):83, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:REB

[Che05b] Haoyang Che. Review of Eclipse: Building Commercial Quality Plug-ins by Eric Clayberg and Dan Rubel. Addison Wesley, 2004, 0-321-22847-2 (paperback). ACM SIGSOFT Software Engineering Notes, 30(1): 36, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:ROO

[Che05c]  Haoyang Che.  Review

of Object-Oriented Programming with Visual Basic.NET by Michael McMillan; Cambridge University Press: Cambridge, 2004, 0-521-53983-8. ACM SIGSOFT Software Engineering Notes, 30(2):29–30, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:ROT

[Che05d] Haoyang Che. Review of Object Thinking by David West; Microsoft Press: Microsoft, 2004, 0-7356-1965-4. ACM SIGSOFT Software Engineering Notes, 30 (2):28–29, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:REC

[Che05e] Haoyang Che. Review of The Elements of C++ Style by Trevor Misfeldt, Gregory Bumgardner, and Andrew Gray; Cambridge University Press: Cambridge, 2004, 0-521-89308-9. ACM SIGSOFT Software Engineering Notes, 30(2):29, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:ROPa

[Che05f] Haoyang Che. Review of The Object Primer, 2nd Edition by Scott W. Ambler. Cambridge University Press: Cambridge, 2001, 0-521-785197 (paperback). ACM SIGSOFT Software Engineering Notes, 30(1): 35, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Che:2005:ROPb

[Che05g] Haoyang Che. Review of The Object Primer, 3rd Edition by Scott W. Ambler. Cambridge University Press: Cambridge, 2004, 0-521-54018-6 (paperback). ACM SIGSOFT Software Engineering Notes, 30(1):35– 36, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chekin:2006:TIS

[Che06] Konstantin Chekin. Theory of infinite streams and objects. ACM SIGSOFT Software Engineering Notes, 31 (2):15:1–15:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chiang:2005:IHP

[Chi05] Chia-Chu Chiang. Implicit heterogeneous and parallel programming. ACM SIG-SOFT Software Engineering Notes, 30(3):1–6, May 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chodkowski:2009:BRR

[Cho09] Steven Chodkowski. Book review: Research Methods for Human-Computer Interaction edited by Paul Cairns and Anna L. Cox, and published by Cambridge University Press, 2008, paperback (also available in Hard Cover), ISBN: 978-0-521-69031-7, 242 pp., (987-0-521-87012-2 hardcover). ACM SIGSOFT Software Engineering Notes, 34(5): 59, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cox:2004:IWA

[CHR04] Karl Cox, Jon G. Hall, and Lucia Rapanotti. 1st International Workshop on Advances and Applications of Problem Frames — summary. ACM SIGSOFT Software Engineering Notes, 29(5):1– 3, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chroust:2008:HCC

[Chr08] Gerhard Chroust. History of computers can raise student’s interest. ACM SIGSOFT Software Engineering Notes, 33(4):29, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Christodorescu:2004:TMD

[CJ04] Mihai Christodorescu and Somesh Jha. Testing malware detectors. ACM SIGSOFT Software Engineering Notes, 29(4):34–44, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clarke:2001:STG

[CJRZ01] Duncan Clarke, Thierry Jéron, Vlad Rusu, and Elena Zinovieva. STG: a tool for generating symbolic test programs and oracles from operational specifications. ACM SIGSOFT Software Engineering Notes, 26(5):301– 302, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Castro:2001:SRA

[CK01] Jaelson Castro and Jeff Kramer. From software requirements to architectures (STRAW01). ACM SIGSOFT Software Engineering Notes, 26(6):49– 51, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coady:2001:UAI

[CKFS01] Yvonne Coady, Gregor Kiczales, Mike Feeley, and Greg Smolyn. Using aspectC to improve the modularity of path-specific customization in operating system code. ACM SIGSOFT Software Engineering Notes, 26(5):88– 98, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Christodorescu:2006:SAX

[CKG06] Mihai Christodorescu, Nicholas Kidd, and Wen-Han Goh. String analysis for x86 binaries. ACM SIGSOFT Software Engineering Notes, 31 (1):88–95, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2005:ETM

[CKZ05] T. Y. Chen, F.-C. Kuo, and Zhi Quan Zhou. An effective testing method for end-user programmers. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cai:2005:ECC

[CL05] Xia Cai and Michael R.

Lyu. The effect of code coverage on fault detection under different testing profiles. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ciobanu:2006:SLC

[CL06] Gabriel Ciobanu and Dorel Lucanu. A specification language for coordinated objects. ACM SIGSOFT Software Engineering Notes, 31 (2):3:1–3:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clarke:2000:AVH

[Cla00a] Edmund M. Clarke. Automatic verification of hardware and software systems. ACM SIGSOFT Software Engineering Notes, 25(1):41– 42, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clarke:2000:FSV

[Cla00b]  Lori A. Clarke.  Finite

state verification (abstract only): an emerging technology for validating software systems. ACM SIGSOFT Software Engineering Notes, 25(5):146, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Claussen:2000:CPM

[Cla00c] Peter Claussen. Concurrent programming in ML. ACM SIGSOFT Software Engineering Notes, 25(3): 62–63, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Claussen:2000:TPL

[Cla00d] Peter Claussen. Theories of programming languages. ACM SIGSOFT Software Engineering Notes, 25(3): 61–62, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cleaveland:2000:SFC

[Cle00] Rance Cleaveland. Specification formalisms for componentbased concurrent systems. ACM SIGSOFT Software Engineering Notes, 25(1):42– 43, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clermont:2005:HAI

[Cle05] Markus Clermont. Heuristics for the automatic identification of irregularities in spreadsheets. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cleaveland:2000:PTD

[CLS00] Rance Cleaveland, Philip M.

Lewis, and Scott A. Smolka. Practical techniques for the design, specification, verification, and implementation of concurrent systems. ACM SIGSOFT Software Engineering Notes, 25(1):43– 44, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Crnkovic:2002:CBS

[CLS02] Ivica Crnkovic, Stig Larsson, and Judith Stafford. Component-based software engineering: building systems from components at 9th IEEE Conference and Workshops on Engineering of Computer-Based Systems. ACM SIGSOFT Software Engineering Notes, 27(3): 47–50, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ciolkowski:2005:APE

[CM05a] Marcus Ciolkowski and Jürgen Münch. Accumulation and presentation of empirical evidence: problems and challenges. ACM SIGSOFT Software Engineering Notes, 30 (4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coelho:2005:APC

[CM05b] Wesley Coelho and Gail C. Murphy. ActiveAspect: presenting crosscutting structure. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2005:FSS

[CMPB05] Zhihao Chen, Tim Menzies, Dan Port, and Barry Boehm. Feature subset selection can improve software cost estimation accuracy. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cordeiro:2008:ADM

[CMV^{+}08] Lucas Cordeiro, Carlos Mar, Eduardo Valentin, Fabiano Cruz, Daniel Patrick, Raimundo Barreto, and Vicente Lucena. An agile development methodology applied to embedded control software under stringent hardware constraints. ACM SIGSOFT Software Engineering Notes, 33(1):5:1–5:??, January 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Capilla:2006:WBT

[CNPD06] Rafael Capilla, Francisco Nava, Sandra Pérez, and Juan C. Dueñas. A webbased tool for managing architectural design decisions. ACM SIGSOFT Software Engineering Notes, 31(5): 4:1–4:??, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheng:2005:WWW

[CNS05] Shang-Wen Cheng, Robert L. Nord, and Judith A. Stafford. WICSA Wiki WAN party: capturing experience in software architecture best practices. ACM SIGSOFT Software Engineering Notes, 30 (1):1, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cimpan:2000:OLS

[CO00] Sorana Cîmpan and Flavio Oquendo. OMEGA: a language and system for online monitoring of softwareintensive processes. ACM SIGSOFT Software Engineering Notes, 25(4):62– 68, July 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cardell-Oliver:2002:CTE

[CO02] Rachel Cardell-Oliver. Conformance test experiments for distributed real-time systems. ACM SIGSOFT Software Engineering Notes, 27 (4):159–163, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chawla:2004:GIF

[CO04] Anil Chawla and Alessandro Orso. A generic instrumentation framework for collecting dynamic information.

ACM SIGSOFT Software Engineering Notes, 29(5):1– 4, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:2006:MSU

[CO06] Jonathan Cook and Alessandro Orso. MonDe: safe updating through monitored deployment of new component versions. ACM SIGSOFT Software Engineering Notes, 31(1):43–46, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Conradi:2000:SES

[Con00]  Reidar Conradi.  Sum-

mary from 7th European Software Process Workshop (EWSPT’7), Kaprun near Salzburg, 21–25 Feb. 2000. ACM SIGSOFT Software Engineering Notes, 25(3): 23, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:2000:SPA

[Coo00] Jonathan E. Cook. Software process analysis: integrating models and data. ACM SIGSOFT Software Engineering Notes, 25(1): 44, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:2001:AFPb

[Coo01a] Greg Cooper. ACM Fellow profile: Eugene H. Spafford. ACM SIGSOFT Software Engineering Notes, 26 (4):13–16, July 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:2001:AFPa

[Coo01b] Greg Cooper. ACM Fellow profile: Frank L. Friedman. ACM SIGSOFT Software Engineering Notes, 26 (2):12–13, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cobleigh:2002:CUH

[COWL02] Jamieson M. Cobleigh, Leon J. Osterweil, Alexander Wise, and Barbara Staudt Lerner. Containment units: a hierarchically composable architecture for adaptive systems. ACM SIGSOFT Software Engineering Notes, 27 (6):159–165, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clermont:2005:UIA

[CP05] Markus Clermont and David Parnas. Using information about functions in selecting test cases. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coen-Porisini:2001:USE

[CPDGP01] Alberto Coen-Porisini, Giovanni Denaro, Carlo Ghezzi, and Mauro Pezzé. Using symbolic execution for verifying safety-critical systems. ACM SIGSOFT Software Engineering Notes, 26 (5):142–151, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chitchyan:2009:REA

[CPK09] Ruzanna Chitchyan, Monica Pinto, and Safoora Shakil Khan. Report on early aspects at ICSE 2009: Workshop on Aspect-Oriented Requirements Engineering and Architecture Design. ACM SIGSOFT Software Engineering Notes, 34(5):30– 35, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chalin:2006:NNR

[CR06a] Patrice Chalin and Frédéric Rioux. Non-null references by default in the Java modeling language. ACM SIGSOFT Software Engineering Notes, 31(2):9:1–9:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clarke:2006:HPR

[CR06b] Lori A. Clarke and David S. Rosenblum. A historical perspective on runtime assertion checking in software development. ACM SIGSOFT Software Engineering Notes, 31 (3):25–37, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choi:2001:AAM

[CRH01] Yunja Choi, Sanjai Rayadurgam, and Mats P. E. Heimdahl. Automatic abstraction for model checking software systems with interrelated numeric constraints. ACM SIGSOFT Software Engineering Notes, 26(5):164– 174, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Collins:2006:ADV

[CRL^{+}06] Michael Collins, Charles Reynolds, Christine Le, Cihan Varol, and Coskun Bayrak. Automated data verification in a format-free environment. ACM SIGSOFT Software Engineering Notes, 31(2):1–4, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Crnkovic:2005:RIS

[CRS^{+}05] Ivica Crnkovic, Ralf Reussner, Heinz Schmidt, Kevin Simons, Judith Stafford, and Kurt Wallnau. Report of the international symposium on component-based software engineering. ACM SIGSOFT Software Engineering Notes, 30(3):1–9, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cai:2005:VOT

[CS05] Yuanfang Cai and Kevin J.

Sullivan. A value-oriented theory of modularity in design. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chahal:2009:MSS

[CS09] Kuljit Kaur Chahal and Hardeep Singh. Metrics to study symptoms of bad software designs. ACM SIG-SOFT Software Engineering Notes, 34(1):1–4, January 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cohen:2006:TAC

[CSR06] Myra B. Cohen, Joshua Snyder, and Gregg Rothermel. Testing across configurations: implications for combinatorial testing. ACM SIGSOFT Software Engineering Notes, 31(6):1–9, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Crnkovic:2001:IWC

[CSSW01] Ivica Crnkovic, Heinz Schmidt, Judith Stafford, and Kurt Wallnau. 4th ICSE Workshop on Component-Based Software Engineering: Component Certification and System Prediction. ACM SIGSOFT Software Engineering Notes, 26(6):33– 40, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Crnkovic:2004:IWC

[CSSW04] Ivica Crnkovic, Heinz Schmidt, Judith Stafford, and Kurt Wallnau. 6th ICSE Workshop on Component-Based Software Engineering: automated reasoning and prediction. ACM SIGSOFT Software Engineering Notes, 29 (3):1–7, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Counsell:2006:OOC

[CSTM06] Steve Counsell, Stephen Swift, Allan Tucker, and Emilia Mendes. Objectoriented cohesion subjectivity amongst experienced and novice developers: an empirical study. ACM SIGSOFT Software Engineering Notes, 31(5):1–10, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chatley:2005:KLP

[CT05] Robert Chatley and Thomas Timbul. KenyaEclipse: learning to program in eclipse. ACM SIGSOFT Software Engineering Notes, 30(5): 245–248, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2002:SPI

[CTZ02] T. Y. Chen, T. H. Tse, and Zhiquan Zhou. Semi-proving: an integrated method based on global symbolic evaluation and metamorphic testing. ACM SIGSOFT Software Engineering Notes, 27 (4):191–195, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2000:EME

[CYD00] Wang Chen, Zhou Ying, and Zhang Defu. An efficient method for expressing active object in C++. ACM SIGSOFT Software Engineering Notes, 25(3): 32–35, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choi:2002:IFI

[CZ02] Jong-Deok Choi and Andreas Zeller. Isolating failureinducing thread schedules. ACM SIGSOFT Software Engineering Notes, 27(4):

210–220, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dash:2001:DMD

[DA01] Ashiss Kumar Dash and Rakesh Agarwal. Dimensional modeling for a data warehouse. ACM SIGSOFT Software Engineering Notes, 26(6):83–84, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Desouza:2005:MRS

[DA05] Kevin C. Desouza and Yukika Awazu. Managing radical software engineers: between order and chaos. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dor:2004:SVS

[DADY04] Nurit Dor, Stephen Adams, Manuvir Das, and Zhe Yang. Software validation via scalable path-sensitive value flow analysis. ACM SIGSOFT Software Engineering Notes, 29(4):12–22, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deAlfaro:2001:IA

[dAH01] Luca de Alfaro and Thomas A. Henzinger. Interface automata. ACM SIGSOFT Software Engineering Notes, 26(5):109–120, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Daneva:2001:EVA

[Dan01]  Maya Daneva.  Evaluat-

ing the value-added benefits of using requirements reuse metrics in ERP projects. ACM SIGSOFT Software Engineering Notes, 26(3): 155–163, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Das:2006:PM

[Das06] Manuvir Das. PASTE at Microsoft. ACM SIGSOFT Software Engineering Notes, 31(1):1, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:2002:OSD

[DB02] Chad Davis and Coskun Bayrak. Open source development and the World Wide Web: a certain tension. ACM SIGSOFT Software Engineering Notes, 27 (5):93–97, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dinkel:2005:MNR

[DB05] Michael Dinkel and Uwe Baumgarten. Modeling nonfunctional requirements: a basis for dynamic systems management. ACM SIG-SOFT Software Engineering Notes, 30(4):1–8, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deBoer:2006:AKD

[dB06] Remco C. de Boer. Architectural knowledge discovery: why and how? ACM SIGSOFT Software Engineering Notes, 31(5):1:1– 1:??, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dahiya:2009:DMF

[DB09] Deepak Dahiya and Usha Batra. Distributed middleware framework using aspects. ACM SIGSOFT Software Engineering Notes, 34 (4):1–6, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Daughtry:2009:CRS

[DBCP09] John Daughtry, Janet Burge, John M. Carroll, and Colin Potts. Creativity and rationale in software design. ACM SIGSOFT Software Engineering Notes, 34(1):27– 29, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Desikan:2001:MEE

[DBK01] Rajagopalan Desikan, Doug Burger, and Stephen W. Keckler. Measuring experimental error in microprocessor simulation. ACM SIGSOFT Software Engineering Notes, 26(3):266– 277, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Junior:2006:CCC

[dCJL06] Francisco Heron de Carvalho Junior and Rafael Dueire Lins. A categorical characterization for the compositional features of the # component model. ACM SIGSOFT Software Engineering Notes, 31 (2):12:1–12:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

dAdderio:2002:PEC

[dDLS02] Luciana d’Adderio, Rick Dewar, Ashley Lloyd, and Perdita Stevens. Has the pattern emperor any clothes?: a controversy in three acts. ACM SIGSOFT Software Engineering Notes, 27(1):31– 35, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Demir:2005:DAA

[DDMW05] Omer Erdem Demir, Prem Devanbu, Nenad Medvidovic, and Eric Wohlstadter. DIS-COA: architectural adaptations for security and QoS. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dekel:2005:SDS

[Dek05]  Uri Dekel.  Supporting

distributed software design meetings: what can we learn from co-located meetings? ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeLoach:2005:MSE

[DeL05] Scott A. DeLoach. Multiagent systems engineering of organization-based multiagent systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeMillo:2000:SDN

[DeM00] Richard A. DeMillo. Software development for next generation communication networks. ACM SIGSOFT Software Engineering Notes, 25 (6):1, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Do:2004:BIS

[DER04] Hyunsook Do, Sebastian Elbaum, and Gregg Rothermel. Building an infrastructure to support experimentation with software testing techniques. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Devanbu:2000:FDD

[Dev00] Premkumar Devanbu. Framework for debugging domainspecific languages. ACM SIGSOFT Software Engineering Notes, 25(1):45, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubois:2004:IWM

[DF04] Eric Dubois and Xavier Franch. International Workshop on Models and Processes for the Evaluation of COTS Components (MPEC’04). ACM SIGSOFT Software Engineering Notes, 29(5):1– 3, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deFigueiredo:2006:GIT

[dFAM06] André L. L. de Figueiredo, Wilkerson L. Andrade, and Patrícia D. L. Machado. Generating interaction test cases for mobile phone systems from use case specifications. ACM SIGSOFT Software Engineering Notes, 31(6):1– 10, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Daughtry:2009:AUR

[DFMS09] John M. Daughtry, Umer Farooq, Brad A. Myers, and Jeffrey Stylos. API usability: report on special interest group at CHI. ACM SIGSOFT Software Engineering Notes, 34(4):27–29, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Deng:2004:TWD

[DFW04] Yuetang Deng, Phyllis Frankl, and Jiong Wang. Testing web database applications. ACM SIGSOFT Software Engineering Notes, 29 (5):1–10, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Demeyer:2000:WOO

[DG00] Serge Demeyer and Harald Gall. Workshop on Object-Oriented Reengineering (WOOR’99). ACM SIG-SOFT Software Engineering Notes, 25(1):27, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ducasse:2005:MAR

[DGN05] Stéphane Ducasse, Tudor Gîrba, and Oscar Nierstrasz. Moose: an agile reengineering environment. ACM SIGSOFT Software Engineering Notes, 30(5):99– 102, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:2001:NPA

[DGP^{+}01] L. Davis, R. Gamble, J. Payton, G. Jónsdóttir, and D. Underwood. A notation for problematic architecture interactions. ACM SIGSOFT Software Engineering Notes, 26(5):132– 141, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dwyer:2000:APL

[DH00] Matthew B. Dwyer and John Hatcliff. Adapting programming languages technologies for finite-state verification. ACM SIGSOFT Software Engineering Notes, 25 (1):46–49, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

dAmorim:2005:EBR

[dH05] Marcelo d’Amorim and Klaus Havelund. Event-based runtime verification of Java programs. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diaz-Herrera:2009:ESD

[DH09] Jorge L. Díaz-Herrera. The “engineering” of software, a different kind of engineering. ACM SIGSOFT Software Engineering Notes, 34(5):1– 3, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diehl:2005:RMI

[DHH05] Stephan Diehl, Ahmed E. Hassan, and Richard C. Holt. Report on MSR 2005: International Workshop on Mining Software Repositories. ACM SIGSOFT Software Engineering Notes, 30(5):1– 3, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DiNitto:2006:RIW

[DHH^{+}06] Elisabetta Di Nitto, Robert J. Hall, Jun Han, Yanbo Han, Andrea Polini, Kurt Sandkuhl, and Andrea Zisman. Report on the International Workshop on Service Oriented Software Engineering (IW-SOSE06). ACM SIGSOFT Software Engineering Notes, 31(5):36– 38, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diev:2006:SEM

[Die06a] Sergey Diev. Software estimation in the maintenance context. ACM SIGSOFT Software Engineering Notes, 31(2):1–8, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diev:2006:UCM

[Die06b]  Sergey Diev.  Use cases

modeling and software estimation: applying use case points. ACM SIGSOFT Software Engineering Notes, 31 (6):1–4, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diev:2007:RDM

[Die07a] Sergey Diev. Requirements development as a modeling activity. ACM SIGSOFT Software Engineering Notes, 32(2):1–3, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diev:2007:SCR

[Die07b]  Sergey Diev.  Structur-

ing complex requirements. ACM SIGSOFT Software Engineering Notes, 32(2): 1–5, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diev:2009:QCR

[Die09] Sergey Diev. Querying complex requirements. ACM SIGSOFT Software Engineering Notes, 34(1):1–7, January 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dill:2000:MCJ

[Dil00a] David Dill. Model checking Java programs (abstract only). ACM SIGSOFT Software Engineering Notes, 25 (5):179, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dillon:2000:AST

[Dil00b]  Laura K. Dillon.  Au-

tomated support for testing and debugging of realtime programs using oracles. ACM SIGSOFT Software Engineering Notes, 25 (1):45–46, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dinishak:2000:OIF

[Din00] Ron Dinishak. The optimal implementation of functional programming languages. ACM SIGSOFT Software Engineering Notes, 25(3):63, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dalal:2005:AMB

[DJP05] Siddhartha R. Dalal, Ashish Jain, and Jesse Poore. Advances in model-based software testing (A-MOST). ACM SIGSOFT Software Engineering Notes, 30(5):1– 3, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubinsky:2009:SDG

[DK09] Yael Dubinsky and Philippe Kruchten. Software development governance (SDG): report on 2nd workshop. ACM SIGSOFT Software Engineering Notes, 34(5):46– 47, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dingel:2004:ACS

[DL04] Juergen Dingel and Hongzhi Liang. Automating comprehensive safety analysis of concurrent programs using Verisoft and TXL. ACM SIGSOFT Software Engineering Notes, 29(6):13– 22, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dehlinger:2005:PLR

[DL05] Josh Dehlinger and Robyn R.

Lutz. A product-line requirements approach to safe reuse in multi-agent systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deLemos:2005:IWAa

[dLGR05] Rogério de Lemos, Cristina Gacek, and Alexander Romanovsky. ICSE 2004 Workshop on Architecting Dependable Systems. ACM SIGSOFT Software Engineering Notes, 30(1): 3, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dickinson:2001:PFD

[DLP01] William Dickinson, David Leon, and Andy Podgurski. Pursuing failure: the distribution of program failures in a profile space. ACM SIGSOFT Software Engineering Notes, 26(5):246– 255, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deLemos:2005:IWAb

[dLR05] Rogério de Lemos and Alexander Romanovsky. ICSE 2005 Workshop on Architecting Dependable Systems (WADS 2005). ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Deshpande:2001:SSI

[DM01] Yogesh Deshpande and San Murugesan. Summary of the Second ICSE workshop on Web Engineering. ACM SIGSOFT Software Engineering Notes, 26(1):76– 77, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dig:2006:RAS

[DMNJ06] Danny Dig, Kashif Manzoor, Tien N. Nguyen, and Ralph Johnson. Refactoringaware software merging and configuration management. ACM SIGSOFT Software Engineering Notes, 31(6):1– 2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dranidis:2009:CPI

[DMS09] Dimitris Dranidis, Stephen P.

Masticola, and Paul Strooper. Challenges in practice: 4th International Workshop on the Automation of Software Test report. ACM SIGSOFT Software Engineering Notes, 34(4):32–34, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dyer:2006:NPD

[DNR06] Robert Dyer, Harish Narayanappa, and Hridesh Rajan. Nu: preserving design modularity in object code. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2000:SNSa

[Doe00a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 25(2):8–16, March 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2000:SNSb

[Doe00b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 25(4):12–20, July 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2001:SNSa

[Doe01a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 26(1):11– 13, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2001:SNSb

[Doe01b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 26(2):14– 23, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2001:SNSc

[Doe01c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 26(4):17–26, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2001:SNSd

[Doe01d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 26(6):19– 28, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2002:SNSa

[Doe02a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 27(1):21– 29, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2002:SNSb

[Doe02b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 27(2):25– 34, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2002:SNSc

[Doe02c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 27(3):24–33, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2002:SNSd

[Doe02d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 27(5):21– 30, September 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2003:SNS

[Doe03] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 28(3):13–21, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2004:SNSa

[Doe04a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 29(2):17– 26, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2004:SNSb

[Doe04b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 29(3):15–24, May 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2004:SNSc

[Doe04c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 29(5):20– 28, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2005:SNSa

[Doe05a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 30(1):19– 27, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2005:SNSb

[Doe05b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 30(2):9–16, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2005:SNSc

[Doe05c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 30(3):12–21, May 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2005:SNSd

[Doe05d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 30(4):10–18, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2005:SNSe

[Doe05e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 30(5):8–15, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2005:SNSf

[Doe05f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 30(6):5–13, November 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2006:J

[Doe06a] Mark Doernhoefer. JavaScript.

ACM SIGSOFT Software Engineering Notes, 31(4): 16–24, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2006:SNSa

[Doe06b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 31(1):5–13, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2006:SNSb

[Doe06c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 31(2):17– 25, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2006:SNSc

[Doe06d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 31(3):8–16, May 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2006:SNSd

[Doe06e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 31(5):6–15, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2006:SNSe

[Doe06f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 31(6):11– 20, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2007:SNSa

[Doe07a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 32(1):6–14, January 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2007:SNSc

[Doe07b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 32(3):10–19, May 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2007:SNSd

[Doe07c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 32(4):10–19, July 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2007:SNSe

[Doe07d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 32(5):7–16, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2007:SNSf

[Doe07e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 32(6):10– 19, November 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2007:SNSb

[Doe07f] Mark Doernhoefer. Surfing the net for Software Engineering Notes: solving the software quality problem. ACM SIGSOFT Software Engineering Notes, 32 (2):8–17, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2008:SNSa

[Doe08a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 33(1):3:1– 3:??, January 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2008:SNSb

[Doe08b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 33(2):8–17, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2008:SNSc

[Doe08c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 33(3):7–16, May 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2008:SNSd

[Doe08d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 33(4):7–16, July 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2008:SNSe

[Doe08e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 33(5):1:1– 1:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2008:SNSf

[Doe08f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 33(6):5–14, November 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2009:SNSa

[Doe09a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 34(1):7–16, January 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2009:SNSb

[Doe09b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 34(2):5–14, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2009:SNSc

[Doe09c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 34(3):6–15, May 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2009:SNSd

[Doe09d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 34(4):9–16, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2009:SNSe

[Doe09e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 34(5):8–17, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2009:SNSf

[Doe09f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 34(6):11– 20, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dean:2001:CWC

[DOV^{+}01] John Dean, Patricia Oberndorf, Mark Vigder, Chris Abts, Hakan Erdogmus, Neil Maiden, Michael Looney, George Heineman, and Michael Guntersdorfer. COTS workshop: continuing collaborations for successful COTS development. ACM SIGSOFT Software Engineering Notes, 26(1):61–73, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dwaikat:2005:RTR

[DPP05] Zaid Dwaikat and Francesco Parisi-Presicce. Risky trust: risk-based analysis of software systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeWin:2005:TUV

[DPSJ05] Bart De Win, Frank Piessens, Jan Smans, and Wouter Joosen. Towards a unifying view on security contracts. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Denaro:2005:AIT

[DPT05] Giovanni Denaro, Mauro Pezzé, and Davide Tosi. Adaptive integration of thirdparty web services. ACM SIGSOFT Software Engineering Notes, 30(4):1– 6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duran:2002:VSR

[DRBT02] Amador Durán, Antonio Ruiz, Beatriz Bernárdez, and Miguel Toro. Verifying software requirements with XSLT. ACM SIGSOFT Software Engineering Notes, 27 (1):39–44, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dittrich:2005:COM

[DRL^{+}05] Yvonne Dittrich, Kari Rönkkö, Olle Lindeberg, Jeanette Erickson, and Christina Hansson. Co-operative method development revisited. ACM SIGSOFT Software Engineering Notes, 30 (4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drori:2009:TSD

[Dro09] Offer Drori. Template for a system design file using OODPM version 2010. ACM SIGSOFT Software Engineering Notes, 34(6):1– 4, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Darbari:2005:MBS

[DS05] Manuj Darbari and Vipin Saxena. Modeling biological systems: a unified approach. ACM SIGSOFT Software Engineering Notes, 30(5):1–4, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dahiya:2006:AAO

[DS06] Deepak Dahiya and Rajinder K. Sachdeva. Approaches to aspect oriented design: a study. ACM SIGSOFT Software Engineering Notes, 31(5):1–4, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deSouza:2004:HGS

[dSRC^{+}04] Cleidson R. B. de Souza, David Redmiles, Li-Te Cheng, David Millen, and John Patterson. How a good software practice thwarts collaboration: the multiple roles of APIs in software development. ACM SIGSOFT Software Engineering Notes, 29 (6):221–230, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dennis:2004:ACA

[DSRJ04] Greg Dennis, Robert Seater, Derek Rayside, and Daniel Jackson. Automating commutativity analysis at the design level. ACM SIGSOFT Software Engineering Notes, 29(4):165–174, July 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Darroch:2006:AFP

[DT06] Fiona Darroch and Mark Toleman. ACM Fellow profile: Bob Glass. ACM SIG-SOFT Software Engineering Notes, 31(4):2–6, July 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dong:2005:CAC

[DYC^{+}05] Jing Dong, Sheng Yang, Lawrence Chung, Paulo Alencar, and Donald Cowan. A COTS architectural component specification stencil for selection and reasoning. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eklund:2005:EIR

[EAG^{+}05] Ulrik Eklund, Örjan Askerdal, Johan Granholm, Anders Alminger, and Jakob Axelsson. Experience of introducing reference architectures in the development of automotive electronic systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eckstein:2001:TLC

[EAN01] Silke Eckstein, Peter Ahlbrecht, and Karl Neumann. Techniques and language constructs for developing generic informations systems: a case study. ACM SIGSOFT Software Engineering Notes, 26 (3):145–154, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Emmerich:2007:IRM

[EAS07] Wolfgang Emmerich, Mikio Aoyama, and Joe Sventek. The impact of research on middleware technology. ACM SIGSOFT Software Engineering Notes, 32 (1):21–46, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ebert:2009:BRC

[Ebe09] Jürgen Ebert. Book review: Collaborative and Distributed Chemical Engineering From Understanding to Substantial Design Process Support, Edited by Manfred Nagl, Wolfgang Marquardt (Springer Verlag, Berlin, 2008), as Lecture Note in Computer Science 4970, ISBN 3-540-70551-1, 849 pp. ACM SIGSOFT Software Engineering Notes, 34 (4):42–43, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Easterbrook:2001:IWL

[EC01] Steve Easterbrook and Marsha Chechik. 2nd International Workshop on Living with Inconsistency (IWLWI01). ACM SIGSOFT Software Engineering Notes, 26(6):76–78, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Egyed:2000:UAS

[Egy00] Alexander Egyed. UML/ Analyzer — a system for defining and analyzing the conceptual integrity of UML models. ACM SIGSOFT Software Engineering Notes, 25(1):108, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Egyed:2004:RUD

[Egy04] Alexander Egyed. Resolving uncertainties during trace analysis. ACM SIGSOFT Software Engineering Notes, 29(6):3–12, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elbaum:2004:ESP

[EH04] Sebastian Elbaum and Madeline Hardojo. An empirical study of profiling strategies for released software and their impact on testing activities. ACM SIGSOFT Software Engineering Notes, 29 (4):65–75, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ehrich:2005:HII

[Ehr05] Hans-Dieter Ehrich. Harmonizing intra- and interobject specification and verification: a multi-object logics approach. ACM SIGSOFT Software Engineering Notes, 30(4):1, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:2004:TSO

[EJT04] Jonathan Edwards, Daniel Jackson, and Emina Torlak. A type system for object models. ACM SIG-SOFT Software Engineering Notes, 29(6):189–199, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:2004:FCS

[EJTY04] Jonathan Edwards, Daniel Jackson, Emina Torlak, and Vincent Yeung. Faster constraint solving with subtypes. ACM SIGSOFT Software Engineering Notes, 29 (4):232–242, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edvardsson:2001:ACS

[EK01a] Jon Edvardsson and Mariam Kamkar. Analysis of the constraint solver in UNA based test data generation. ACM SIGSOFT Software Engineering Notes, 26(5):237– 245, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Emmerich:2001:CTJ

[EK01b] Wolfgang Emmerich and Nima Kaveh. Component technologies: Java Beans, COM, CORBA, RMI, EJB and the CORBA Component Model. ACM SIGSOFT Software Engineering Notes, 26 (5):311–312, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Engels:2001:MSA

[EKHG01] Gregor Engels, Jochem M. Küster, Reiko Heckel, and Luuk Groenewegen. A methodology for specifying and analyzing consistency of object-oriented behavioral models. ACM SIGSOFT Software Engineering Notes, 26(5):186– 195, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Evans:2004:IWS

[EL04] David Evans and Raimondas Lencevicius. ICSE 2004 workshop summary Second International Workshop on Dynamic Analysis (WODA 2004). ACM SIGSOFT Software Engineering Notes, 29(5):1–2, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Estublier:2002:IRC

[ELC^{+}02] Jacky Estublier, David Leblang, Geoff Clemm, Reidar Conradi, Walter Tichy, André van der Hoek, and Darcy Wiborg-Weber. Impact of the research community on the field of software configuration management: summary of an impact project report. ACM SIGSOFT Software Engineering Notes, 27 (5):31–39, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ereno:2006:SPL

[ELC06] Montse Ereño, Uxue Landa, and Dra. Rebeca Cortazar. Software product lines structuring based upon market demands. ACM SIGSOFT Software Engineering Notes, 31 (2):13:1–13:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ellis:2007:SEE

[Ell07] Heidi Ellis. Software engineering: effective teaching and learning approaches and practices. ACM SIGSOFT Software Engineering Notes, 32(1):4, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elssamadisy:2006:RFD

[Els06a] Amr Elssamadisy. Review of Fit for Developing Software: Framework for Integrated Tests by Rick Mugridge and Ward Cunningham, Pearson Education Inc., 2005, ISBN: 0-321-26934-9. ACM SIGSOFT Software Engineering Notes, 31 (3):43–44, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elssamadisy:2006:RHJ

[Els06b] Amr Elssamadisy. Review of Hibernate: a J2EE Developer’s Guide by Will Iverson, Pearson Education Inc., 2005, ISBN: 0-471-20282-7. ACM SIGSOFT Software Engineering Notes, 31 (3):42–43, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Emerson:2000:AFM

[Eme00]  E. Allen Emerson.  Au-

tomated formal methods: model checking and beyond. ACM SIGSOFT Software Engineering Notes, 25(1): 49, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elbaum:2000:PTC

[EMR00] Sebastian Elbaum, Alexey G.

Malishevsky, and Gregg Rothermel. Prioritizing test cases for regression testing. ACM SIGSOFT Software Engineering Notes, 25 (5):102–112, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Encontre:2002:IRR

[Enc02] Vincent Encontre. Is IS-STA research relevant to industrial users? Panel — ISSTA 2002: empowering the developer to be a tester too! ACM SIGSOFT Software Engineering Notes, 27 (4):203–204, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eng:2003:CSD

[Eng03] David Eng. Combining static and dynamic data in code visualization. ACM SIGSOFT Software Engineering Notes, 28(1):43–50, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Erfurth:2005:UUP

[ER05] Ivonne Erfurth and Wilhelm Rossak. UPEX: user participation by example. ACM SIGSOFT Software Engineering Notes, 30(5):374– 376, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Estublier:2005:RVL

[EV05] Jacky Estublier and German Vega. Reuse and variability in large software applications. ACM SIGSOFT Software Engineering Notes, 30 (5):316–325, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Evangelist:2000:NIT

[Eva00] Michael Evangelist. The NSF information technology research program. ACM SIG-SOFT Software Engineering Notes, 25(1):30, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Estublier:2001:MMS

[EVC01] J. Estublier, H. Verjus, and P. Y. Cunin. Modelling and managing software federations. ACM SIGSOFT Software Engineering Notes, 26 (5):299–300, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ezick:2004:OCB

[Ezi04]  James Ezick.  An opti-

mizing compiler for batches of temporal logic formulas. ACM SIGSOFT Software Engineering Notes, 29 (4):183–194, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farenhorst:2006:TKS

[Far06] Rik Farenhorst. Tailoring knowledge sharing to the architecting process. ACM SIGSOFT Software Engineering Notes, 31(5):3:1– 3:??, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fass:2004:AAP

[Fas04] Leona F. Fass. Approximations, anomalies and “the proof of correctness wars”. ACM SIGSOFT Software Engineering Notes, 29 (2):3, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fass:2005:SAT

[Fas05] Leona F. Fass. Some agent theory for the semantic web. ACM SIGSOFT Software Engineering Notes, 30(5):1– 6, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fass:2007:FWU

[Fas07] Leona F. Fass. Fostering web use in some real environments. ACM SIGSOFT Software Engineering Notes, 32(2):1–4, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fass:2008:ELL

[Fas08] Leona F. Fass. An ethnocentric look at the law and technology interface. ACM SIGSOFT Software Engineering Notes, 33(1):6:1– 6:??, January 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Falessi:2006:DDR

[FBC06] Davide Falessi, Martin Becker, and Giovanni Cantone. Design decision rationale: experiences and steps ahead towards systematic use. ACM SIGSOFT Software Engineering Notes, 31(5):2:1– 2:??, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Foos:2008:TXB

[FBHZ08] Rüdiger Foos, Christian Bunse, Hagen Höpfner, and Torsten Zimmermann. TML: an XML-based test modeling language. ACM SIGSOFT Software Engineering Notes, 33(2):1:1–1:??, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Filho:2005:FAE

[FBR05] Fernando Castor Filho, Patrick H. S. Brito, and Cecília Mary F. Rubira. A framework for analyzing exception flow in software architectures. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fu:2004:MCX

[FBS04] Xiang Fu, Tevfik Bultan, and Jianwen Su. Model checking XML manipulating software. ACM SIGSOFT Software Engineering Notes, 29 (4):252–262, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fleming:2005:AID

[FCSM05] Scott D. Fleming, Betty H. C.

Cheng, R. E. Kurt Stirewalt, and Philip K. McKinley. An approach to implementing dynamic adaptation in C++. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frankl:2000:CDR

[FD00] Phyllis G. Frankl and Yuetang Deng. Comparison of delivered reliability of branch, data flow and operational testing: a case study. ACM SIGSOFT Software Engineering Notes, 25 (5):124–134, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farooq:2007:RDV

[FD07] Ayaz Farooq and Reiner R. Dumke. Research directions in verification and validation process improvement. ACM SIGSOFT Software Engineering Notes, 32(4): 3:1–3:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Felleisen:2000:CPA

[Fel00] Matthias Felleisen. Components and program analyses. ACM SIGSOFT Software Engineering Notes, 25 (1):49–50, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feng:2006:RBM

[Fen06] Yu Feng. Review of The Build Master: Microsoft’s Software Configuration Management Best Practices by Vincent Maraia, Addison-Wesley, 2005, 0-321-33205-9. ACM SIGSOFT Software Engineering Notes, 31 (2):32, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feng:2007:RCD

[Fen07] Yu Feng. Review of The Cognitive Dynamics of Computer Science: Cost-Effective Large Scale Software Development by Szabolcs Michael de Gyurky, Wiley-IEEE Computer Society, 2006, 0-471-97047-6. ACM SIGSOFT Software Engineering Notes, 32(1):47–48, January 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feng:2009:FFS

[Fen09] Yu Feng. Finding flows in software development. ACM SIGSOFT Software Engineering Notes, 34(6):1– 3, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feller:2005:OSA

[FFH^{+}05] Joseph Feller, Brian Fitzgerald, Scott Hissam, Karim Lakhani, and Walt Scacchi. Open Source Application Spaces: 5th Workshop on Open Source Software Engineering. ACM SIGSOFT Software Engineering Notes, 30(6):1–2, November 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feller:2005:CCC

[FFHL05] Joseph Feller, Brian Fitzgerald, Scott Hissam, and Karim Lakhani. Collaboration, conflict and control: report on the 4th Workshop on Open Source Software Engineering. ACM SIGSOFT Software Engineering Notes, 30(3):1–2, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farcas:2005:RTC

[FFPT05] Emilia Farcas, Claudiu Farcas, Wolfgang Pree, and Josef Templ. Real-time component integration based on transparent distribution. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Flanagan:2004:EPA

[FFQ04] Cormac Flanagan, Stephen N. Freund, and Shaz Qadeer. Exploiting purity for atomicity. ACM SIGSOFT Software Engineering Notes, 29 (4):221–231, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feller:2001:MSB

[FFvdH01] Joseph Feller, Brian Fitzgerald, and André van der Hoek. Making sense of the bazaar: 1st Workshop on Open Source Software Engineering. ACM SIGSOFT Software Engineering Notes, 26(6):51–52, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Freimut:2001:ICS

[FHK^{+}01] Bernd Freimut, Susanne Hartkopf, Peter Kaiser, Jyrki Kontio, and Werner Kobitzsch. An industrial case study of implementing software risk management. ACM SIGSOFT Software Engineering Notes, 26(5):277– 287, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Friedman:2002:PSM

[FHNS02] G. Friedman, A. Hartman, K. Nagin, and T. Shiran. Projected state machine coverage for software testing. ACM SIGSOFT Software Engineering Notes, 27(4): 134–143, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fischer:2005:JDP

[FJM05] Jeffrey Fischer, Ranjit Jhala, and Rupak Majumdar. Joining dataflow with predicates. ACM SIGSOFT Software Engineering Notes, 30 (5):227–236, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fisler:2001:MVC

[FK01] Kathi Fisler and Shriram Krishnamurthi. Modular verification of collaborationbased software designs. ACM SIGSOFT Software Engineering Notes, 26(5):152– 163, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Findler:2001:BCB

[FLF01] Robert Bruce Findler, Mario Latendresse, and Matthias Felleisen. Behavioral contracts and behavioral subtyping. ACM SIGSOFT Software Engineering Notes, 26 (5):229–236, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fernandes:2009:RAM

[FLMW09] João M. Fernandes, Luis C. Lamb, Ricardo J. Machado, and Flávio R. Wagner. Recent advances in modelbased methodologies for pervasive and embedded software. ACM SIGSOFT Software Engineering Notes, 34 (5):37–39, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fiskio-Lasseter:2003:FEG

[FLY03] John Fiskio-Lasseter and Michal Young. Flow equations as a generic programming tool for manipulation of attributed graphs. ACM SIGSOFT Software Engineering Notes, 28(1):69– 76, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fischer:2005:MED

[FORG05] Michael Fischer, Johann Oberleitner, Jacek Ratzinger, and Harald Gall. Mining evolution data of a product family. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fisher:2005:ESC

[FR05] Marc Fisher and Gregg Rothermel. The EUSES spreadsheet corpus: a shared resource for supporting experimentation with spreadsheet dependability mechanisms. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frazer:2001:ACP

[Fra01]  Ken Frazer.  Accelerated

C++: practical programming by example. ACM SIGSOFT Software Engineering Notes, 26(1):95– 96, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frazer:2002:BSS

[Fra02a] Ken Frazer. Building secure software: how to avoid security problems the right way. ACM SIGSOFT Software Engineering Notes, 27 (2):71–72, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frazer:2002:CAI

[Fra02b] Ken Frazer. C++ in action: industrial-strength programming techniques. ACM SIGSOFT Software Engineering Notes, 27(5):104– 105, September 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frazer:2004:RMS

[Fra04a]  Ken Frazer.  Review of

Managing software requirements, a use case approach by Dean Leffingwell and Don Widrig. Addison-Wesley 2003. ACM SIGSOFT Software Engineering Notes, 29 (5):36, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frazer:2004:RUC

[Fra04b] Ken Frazer. Review of Use cases, requirements in context by Daryl Kulak and Eamon Guiney. Addison-Wesley 2004. ACM SIGSOFT Software Engineering Notes, 29 (5):36–37, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frazer:2005:RMP

[Fra05] Ken Frazer. Review of Memory as a Programming Concept in C and C++ by Frantisek Franek. Cambridge University Press, 2004, (paperback), 0-521-52043-6. ACM SIGSOFT Software Engineering Notes, 30(1):36– 37, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frailey:2006:WMR

[Fra06] Dennis J. Frailey. What math is relevant for a CS or SE student?: an employer’s perspective. ACM SIGSOFT Software Engineering Notes, 31(3):6–7, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fu:2004:TJW

[FRMW04] Chen Fu, Barbara G. Ryder, Ana Milanova, and David Wonnacott. Testing of Java web services for robustness. ACM SIGSOFT Software Engineering Notes, 29 (4):23–34, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Froberg:2005:BSR

[FSN05] Joakim Fröberg, Kristian Sandström, and Christer Norström. Business situation reflected in automotive electronic architectures: analysis of four commercial cases. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Franch:2005:TRF

[FT05] Xavier Franch and Marco Torchiano. Towards a reference framework for COTSbased development: a proposal. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farrell-Vinay:2009:BRI

[FV09a] Peter Farrell-Vinay. Book review: Introduction to Software Testing, by Paul Amman and Jeff Offutt, published by CUP, 2008, 978-0-521-88038 322 pp., 0-471-20282-7. ACM SIGSOFT Software Engineering Notes, 34(5):60–61, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farrell-Vinay:2009:BRM

[FV09b] Peter Farrell-Vinay. Book review: Model-based software testing and analysis in C# by Jonathan Jacky, Margus Veanes, Colin Campbell, and Wolfram Schulte, and published by CUP, 2008, (paperback) ISB-10: 0-521-68761-6, 349 pp. ACM SIGSOFT Software Engineering Notes, 34 (5):60, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:2006:PRS

[FW06] Gordon Fraser and Franz Wotawa. Property relevant software testing with modelcheckers. ACM SIGSOFT Software Engineering Notes, 31(6):1–10, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gacek:2001:IPL

[GA01] Critina Gacek and Michalis Anastasopoules. Implementing product line variabilities. ACM SIGSOFT Software Engineering Notes, 26 (3):109–117, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grunbacher:2001:EGS

[GB01] Paul Grünbacher and Barry Boehm. EasyWinWin: a groupware-supported methodology for requirements negotiation. ACM SIGSOFT Software Engineering Notes, 26(5):320–321, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2008:DIO

[GB08] Nasib S. Gill and Balkishan. Dependency and interaction oriented complexity metrics of component-based systems. ACM SIGSOFT Software Engineering Notes, 33(2):3:1–3:??, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2009:CAP

[GB09] Nasib S. Gill and Balkishan. Component adaptation process and adaptability viewpoints. ACM SIGSOFT Software Engineering Notes, 34 (6):1–3, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:2002:PLV

[GBLM02] Francisco José García, Juan-Antonio Barras, Miguel Ángel Laguna, and Jos´ e Manuel Marqués. Product line variability support by FORM and mecano model integration. ACM SIGSOFT Software Engineering Notes, 27 (1):35–38, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:2006:DMA

[GBRS06] Alessandro Garcia, Thais Batista, Awais Rashid, and Claudio Sant’Anna. Driving and managing architectural decisions with aspects. ACM SIGSOFT Software Engineering Notes, 31 (5):6:1–6:??, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giese:2004:MDV

[GBSO04] Holger Giese, Sven Burmester, Wilhelm Schäfer, and Oliver Oberschelp. Modular design and verification of component-based mechatronic systems with onlinereconfiguration. ACM SIG-SOFT Software Engineering Notes, 29(6):179–188, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:2005:SELb

[GCL^{+}05a] Alessandro Garcia, Ricardo Choren, Carlos Lucena, Alexander Romanovsky, Holger Giese, Danny Weyns, Tom Holvoet, and Paolo Giorgini. Software Engineering for Large-Scale Multi-Agent Systems — SEL-MAS 2005: workshop report. ACM SIGSOFT Software Engineering Notes, 30 (4):1–8, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:2005:SELa

[GCL^{+}05b] Alessandro Garcia, Ricardo Choren, Carlos Lucena, Alexander Romanovsky, Tom Holvoet, and Paolo Giorgini. Software engineering for large-scale multi-agent systems — SELMAS’05. ACM SIGSOFT Software Engineering Notes, 30(4):1– 2, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

German:2005:FDU

[GCS05] Daniel M. German, Davor Cubranić, and Margaret-Anne D. Storey. A framework for describing and understanding mining tools in software development. ACM SIGSOFT Software Engineering Notes, 30(4): 1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gurfinkel:2002:MET

[GDC02] Arie Gurfinkel, Benet Devereux, and Marsha Chechik. Model exploration with temporal logic query checking. ACM SIGSOFT Software Engineering Notes, 27 (6):139–148, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guessoum:2005:ARL

[GFB05] Zahia Guessoum, Nora Faci, and Jean-Pierre Briot. Adaptive replication of large-scale multi-agent systems: towards a fault-tolerant multi-agent platform. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gorschek:2008:IGR

[GFF^{+}08] Tony Gorschek, Samuel Fricker, Robert Felt, Claes Wohlin, and Michael Mattsson. 1st International Global Requirements Engineering Workshop — GREW’07. ACM SIGSOFT Software Engineering Notes, 33(2): 29–32, March 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2004:FIC

[GG04a] Nasib S. Gill and P. S. Grover. Few important considerations for deriving interface complexity metric for component-based systems. ACM SIGSOFT Software Engineering Notes, 29 (2):4, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2004:SSP

[GG04b] Nasib S. Gill and P. S. Grover. Software size prediction before coding. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gamble:2005:USA

[GGH05a] M. T. Gamble, R. Gamble, and M. Hepner. Understanding solution architecture concerns. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gurbani:2005:CSO

[GGH05b] Vijay K. Gurbani, Anita Garvert, and James D. Herbsleb. A case study of open source tools and practices in a commercial setting. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:2007:ACM

[GGH^{+}07] Alessandro Garcia, Phil Greenwood, George Heineman, Robert Walker, Yuanfang Cai, Hong Yul Yang, Elisa Baniassad, Cristina Videira Lopes, Christa Schwanninger, and Jianjun Zhao. Assessment of Contemporary Modularization Techniques — ACoM’07: workshop report. ACM SIGSOFT Software Engineering Notes, 32 (5):31–37, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:2006:SEL

[GGR^{+}06] Alessandro Garcia, Holger Giese, Alexander Romanovsky, Ricardo Choren, Ho fung Leung, Carlos Lucena, Florian Klein, and Eric Platon. Software engineering for large-scale multiagent systems — SELMAS 2006: workshop report. ACM SIGSOFT Software Engineering Notes, 31(5):24– 32, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grieskamp:2002:GFS

[GGSV02] Wolfgang Grieskamp, Yuri Gurevich, Wolfram Schulte, and Margus Veanes. Generating finite state machines from abstract state machines. ACM SIGSOFT Software Engineering Notes, 27(4): 112–122, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Georgas:2005:RIH

[GGT05] John C. Georgas, Michael M. Gorlick, and Richard N. Taylor. Raging incrementalism: harnessing change with opensource software. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gunter:2000:FSE

[GGZ00] Carl A. Gunter, Elsa L. Gunter, and Pamela Zave. Formal software engineering. ACM SIGSOFT Software Engineering Notes, 25 (1):54, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glinz:2009:RWC

[GHP^{+}09] Martin Glinz, Patrick Heymans, Anne Persson, Guttorm Sindre, Aybüke Aurum, Nazim Madhavji, Barbara Paech, Gil Regev, and Roel Wieringa. Report on the working conference on requirements engineering: foundation for software quality (REFSQ’09). ACM SIGSOFT Software Engineering Notes, 34(5):40– 45, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garzia:2006:HSS

[GHS^{+}06] Mario Garzia, John Hudepohl, Will Snipes, Michael Lyu, John Musa, Carol Smidts, and Laurie Williams. How should software reliability engineering (SRE) be taught? ACM SIGSOFT Software Engineering Notes, 31(4):1–5, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2005:FAE

[Gil05]  Nasib S. Gill.  Factors

affecting effective software quality management revisited. ACM SIGSOFT Software Engineering Notes, 30 (2):1–4, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2006:ISC

[Gil06] Nasib S. Gill. Importance of software component characterization for better software reusability. ACM SIGSOFT Software Engineering Notes, 31(1):1–3, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Godefroid:2000:AST

[GJJL00] Patrice Godefroid, Lalita J. Jagadeesan, Radha Jagadeesan, and Konstantin Läufer. Automated systematic testing for constraint-based interactive services. ACM SIGSOFT Software Engineering Notes, 25(6):40–49, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giese:2005:SIW

[GK05a] Holger Giese and Ingolf Krüger. A summary of the ICSE 2004 Workshop on “Scenarios and State Machines: Models, Algorithms, and Tools”. ACM SIGSOFT Software Engineering Notes, 30(1):2, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griffith:2005:MME

[GK05b] Rean Griffith and Gail Kaiser. Manipulating managed execution runtimes to support self-healing systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grover:2008:MCG

[GKK08] P. S. Grover, Rajesh Kumar, and Avadhesh Kumar. Measuring changeability for generic aspect-oriented systems. ACM SIGSOFT Software Engineering Notes, 33 (6):1–5, November 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giese:2005:RSC

[GKKW05] Holger Giese, Ekkart Kindler, Florian Klein, and Robert Wagner. Reconciling scenariocentered controller design with state-based system models. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grover:2007:FUC

[GKS07] P. S. Grover, Rajesh Kumar, and Arun Sharma. Few useful considerations for maintaining software components and component-based systems. ACM SIGSOFT Software Engineering Notes, 32(5):1–5, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goyal:2000:ADS

[GL00] Deepak Goyal and Y. Annie Liu. Automated development of software for program analysis and transformation. ACM SIGSOFT Software Engineering Notes, 25(1):52–53, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Girardi:2006:OBK

[GL06] Rosario Girardi and Alisson Neres Lindoso. An ontology-based knowledge base for the representation and reuse of software patterns. ACM SIGSOFT Software Engineering Notes, 31 (1):1–6, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:2006:RCP

[Gla06]  Robert L. Glass.  Re-

view of COBOL Programmers Swing With Java by E. Reed Doke, Bill C. Hardgrave and Richard A. Johnson, Cambridge University Press, 2005. ACM SIGSOFT Software Engineering Notes, 31(3):44–45, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gittens:2004:CEO

[GLB04] Mechelle Gittens, Hanan Lutfiyya, and Mike Bauer. Continuous evolutionary onestep-ahead testing. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:2005:DWD

[GLM^{+}05] David Garlan, Marin Litoiu, Hausi A. Müller, John Mylopoulos, Dennis B. Smith, and Kenny Wong. DEAS 2005: workshop on the Design and Evolution of Autonomic Application Software. ACM SIGSOFT Software Engineering Notes, 30 (4):1, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guinness:2005:SMM

[GM05] David Mc Guinness and Liam Murphy. A simulation model of a multi-server EJB system. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ghanam:2009:RXW

[GMAC09] Yaser Ghanam, Frank Maurer, Pekka Abrahamsson, and Kendra Cooper. A report on the XP Workshop on Agile Product Line Engineering. ACM SIGSOFT Software Engineering Notes, 34 (5):25–27, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gomaa:2001:RCI

[GMS01] Hassan Gomaa, Daniel A. Menascé, and Michael E. Shin. Reusable component interconnection patterns for distributed software architectures. ACM SIGSOFT Software Engineering Notes, 26 (3):69–77, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Godfrey:2001:PDE

[God01] Michael W. Godfrey. Practical data exchange for reverse engineering frameworks: some requirements, some experience, some headaches. ACM SIGSOFT Software Engineering Notes, 26(1):50– 52, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goguen:2000:HAC

[Gog00] Joseph Goguen. Hidden algebra and concurrent distributed software. ACM SIGSOFT Software Engineering Notes, 25(1):51– 52, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grisham:2005:CRE

[GP05] Paul S. Grisham and Dewayne E. Perry. Customer relationships and extreme programming. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guerrouat:2006:CBS

[GR06] Abdelaziz Guerrouat and Harald Richter. A componentbased specification approach for embedded systems using FDTs. ACM SIGSOFT Software Engineering Notes, 31 (2):14:1–14:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Green:2000:STA

[Gre00] Bradley S. Green. Software test automation. ACM SIG-SOFT Software Engineering Notes, 25(3):66, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griswold:2000:GVC

[Gri00] William G. Griswold. Guilds or virtual countries? the future of software professionals. ACM SIGSOFT Software Engineering Notes, 25 (2):2, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grigorenko:2005:VTG

[GST05] Pavel Grigorenko, Ando Saabas, and Enn Tyugu. Visual tool for generative programming. ACM SIG-SOFT Software Engineering Notes, 30(5):249–252, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gaffar:2005:HPS

[GSV05] Ashraf Gaffar, Ahmed Seffah, and John A. Van der Poll. HCI pattern semantics in XML: a pragmatic approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guo:2004:KNT

[GSX^{+}04] Bing Guo, Yan Shen, Jun Xie, Yong Wang, and Guang-Ze Xiong. A kind of new ToolBus model research and implementation. ACM SIG-SOFT Software Engineering Notes, 29(2):5, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2007:CTR

[GT07] Nasib S. Gill and Pradeep Tomar. CBS testing requirements and test case process documentation revisited. ACM SIGSOFT Software Engineering Notes, 32 (2):1–4, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gunter:2000:AAS

[Gun00]  Carl A. Gunter.  Auto-

mated analysis of standard ML. ACM SIGSOFT Software Engineering Notes, 25 (1):53, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gupta:2000:PIP

[Gup00]  Gopal Gupta.  Parallel

implementation of Prolog: the ACE project. ACM SIGSOFT Software Engineering Notes, 25(1):54– 55, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2002:MCJ

[GV02] Alex Groce and Willem Visser. Model checking Java programs using structural heuristics. ACM SIGSOFT Software Engineering Notes, 27(4):12–21, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Georgas:2005:ARC

[GvdHT05] John C. Georgas, André van der Hoek, and Richard N. Taylor. Architectural runtime configuration management in support of dependable self-adaptive software. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gvero:2009:BRPa

[Gve09a]  Igor Gvero.  Book re-

view: Practical Text Mining With Perl, by Roger Bilisoly, and published by Wiley, 2009, 978-0-470-1763-6, 295pp. ACM SIGSOFT Software Engineering Notes, 34 (6):31, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gvero:2009:BRPb

[Gve09b] Igor Gvero. Book review: Python for Software Design by Allen B. Downey, and published by Cambridge University Press, 2009, 978-0-521-72596-5, 251pp. ACM SIGSOFT Software Engineering Notes, 34(6):31– 32, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gegick:2005:MAP

[GW05a] Michael Gegick and Laurie Williams. Matching attack patterns to security vulnerabilities in software-intensive system designs. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gorg:2005:EDR

[GW05b] Carsten Görg and Peter Weißgerber. Error detection by refactoring reconstruction. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giesecke:2005:ASP

[GWH05] Simon Giesecke, Timo Warns, and Wilhelm Hasselbring. Availability simulation of peer-to-peer architectural styles. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grundy:2004:WDS

[GWS04] John Grundy, Ray Welland, and Hermann Stoeckle. Workshop on Directions in Software Engineering Environments (WoDiSEE). ACM SIGSOFT Software Engineering Notes, 29(5):1–3, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Geiger:2005:SDT

[GZ05a] Leif Geiger and Albert Zündorf. Story driven testing — SDT. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gurguis:2005:TAW

[GZ05b] Sherif A. Gurguis and Amir Zeid. Towards autonomic web services: achieving self-healing using web services. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gervasi:2004:RFI

[GZES04] Vincenzo Gervasi, Didar Zowghi, Steve Easterbrook, and Susan Elliott Sim. Report on the First International Workshop on Comparative Evaluation in Requirements Engineering. ACM SIGSOFT Software Engineering Notes, 29(2):1–4, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hay:2000:CFR

[HA00] Jonathan D. Hay and Joanne M. Atlee. Composing features and resolving interactions. ACM SIGSOFT Software Engineering Notes, 25(6): 110–119, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:2000:STP

[Ham00] Dick Hamlet. On subdomains: Testing, profiles, and components. ACM SIGSOFT Software Engineering Notes, 25(5):71–76, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:2002:CSS

[Ham02] Dick Hamlet. Continuity in software systems. ACM SIGSOFT Software Engineering Notes, 27(4):196– 200, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamza:2005:SCE

[Ham05] Haitham S. Hamza. Separation of concerns for evolving systems: a stability-driven approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:2006:IST

[Ham06] Dick Hamlet. Invariants and state in testing and formal methods. ACM SIGSOFT Software Engineering Notes, 31(1):48–51, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hartman:2002:IRR

[Har02]  A. Hartman.  Is ISSTA

research relevant to industry? ACM SIGSOFT Software Engineering Notes, 27 (4):205–206, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harris:2004:RAS

[Har04] Gregory H. Harris. Review of Abstract state machines: a method for high-level system design and analysis by Egon Börger and Robert Stärk. Springer-Verlag 2003. ACM SIGSOFT Software Engineering Notes, 29(2): 31–32, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hazzan:2007:ISI

[Haz07] Orit Hazzan. The influence of software intangibility on computer science and software engineering education. ACM SIGSOFT Software Engineering Notes, 32 (3):7–8, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:2005:DHM

[HB05a] LiGuo Huang and Barry Boehm. Determining how much software assurance is enough?: a value-based approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:2005:UID

[HB05b] LiGuo Huang and Barry Boehm. Using iDAVE to determine availability requirements. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hernandez:2005:EUT

[HBR05] Francisco Hernández, Purushotham Bangalore, and Kevin Reilly. End-user tools for grid computing. ACM SIGSOFT Software Engineering Notes, 30(4): 1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hollingsworth:2000:ERU

[HBW00] Joseph E. Hollingsworth, Lori Blankenship, and Bruce W. Weide. Experience report: using RESOLVE/C++ for commercial software. ACM SIGSOFT Software Engineering Notes, 25(6):11– 19, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hochstein:2005:CSR

[HBZ^{+}05] Lorin Hochstein, Victor R. Basili, Marvin V. Zelkowitz, Jeffrey K. Hollingsworth, and Jeff Carver. Combining selfreported and automatic data to improve programming effort measurement. ACM SIGSOFT Software Engineering Notes, 30(5):356– 365, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Han:2000:CSO

[HCGGXFY00] Yan Han, Xu Chun-Gen, Zhang Gong-Xuan, and Liu Feng-Yu. Constraint specification for object model of access control based on role. ACM SIGSOFT Software Engineering Notes, 25 (2):60–63, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hazzan:2007:WSE

[HD07] Orit Hazzan and Yael Dubinsky. Why software engineering programs should teach agile software development. ACM SIGSOFT Software Engineering Notes, 32 (2):1–3, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hargreaves:2004:GSD

[HDLC04] Elizabeth Hargreaves, Daniela Damian, Filippo Lanubile, and James Chisan. Global software development: building a research community. ACM SIGSOFT Software Engineering Notes, 29(5):1– 5, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hayes:2005:TMS

[HDS05] Jane Huffman Hayes, Alex Dekhtyar, and Senthil Sundaram. Text mining for software engineering: how analyst feedback impacts final results. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heering:2003:QSI

[Hee03] Jan Heering. Quantification of structural information: on a question raised by Brooks. ACM SIGSOFT Software Engineering Notes, 28(3):6, May 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heineman:2000:MDA

[Hei00] George T. Heineman. A model for designing adaptable software components. ACM SIGSOFT Software Engineering Notes, 25(1):55– 56, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heiman:2009:BRP

[Hei09a] Thomas Heiman. Book review: Pattern Recognition and Neural Networks by Brian D. Ripley, and published by Cambridge University Press, 2007, Paperback, ISBN 978-0521-71770-0, pp.

403. ACM SIGSOFT Software Engineering Notes, 34 (6):28, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heiman:2009:BRS

[Hei09b] Thomas Heiman. Book review: Simulation-Based Engineering of Complex Systems, Second Edition is written by John R. Clymer, and published by Wiley-Interscience, 2009, ISBN 978-00470-40129-3, pp. 503. ACM SIGSOFT Software Engineering Notes, 34(6):28– 29, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henzinger:2000:NDC

[Hen00] Thomas A. Henzinger. New directions in computer-aided verification. ACM SIGSOFT Software Engineering Notes, 25(1):56–57, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2003:SEE

[Hen03] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 28(3):11–12, May 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2004:SEEa

[Hen04a] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 29(2):6–8, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2004:SEEb

[Hen04b] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 29(3):5–6, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2004:SEEc

[Hen04c] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 29(5):12–13, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2005:SEEa

[Hen05a] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 30(1):8–9, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2005:SEEb

[Hen05b] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 30(2):3, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2005:SEEc

[Hen05c] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 30(4):6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2005:SEEd

[Hen05d] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 30(6):3–4, November 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2006:SEEa

[Hen06a] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 31(2):3, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2006:SEEb

[Hen06b] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 31(3):5, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2006:SEEc

[Hen06c] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 31(4):8, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2006:SEEd

[Hen06d] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 31(6):9, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2007:SEEa

[Hen07a] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 32(1):4, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2007:SEEb

[Hen07b] Peter B. Henderson. Software engineering education: (SEEd). ACM SIGSOFT Software Engineering Notes, 32(3):6–7, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2007:SEEc

[Hen07c] Peter B. Henderson. Software engineering education: (SEEd). ACM SIGSOFT Software Engineering Notes, 32(5):6, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2008:MSD

[Hen08a] Craig Henderson. Managing software defects: defect analysis and traceability. ACM SIGSOFT Software Engineering Notes, 33 (4):2:1–2:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2008:SEEa

[Hen08b] Peter B. Henderson. Software engineering education: SEEd. ACM SIGSOFT Software Engineering Notes, 33 (1):2:1–2:??, January 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2008:SEEb

[Hen08c] Peter B. Henderson. Software engineering education: SEEd. ACM SIGSOFT Software Engineering Notes, 33 (2):6, March 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2008:SEEc

[Hen08d] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 33(3):5, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2009:BRC

[Hen09a] Craig Henderson. Book review: Computing for Numerical Methods using Visual C++ by Shaharuddin Salleh, Albert Y. Zomaya, Sakhinah Abu Bakar, and published by Wiley-Interscience (December 14, 2007), 2007, 0-470-12795-3, 448 pp. ACM SIGSOFT Software Engineering Notes, 34(6):30– 31, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2009:SEEa

[Hen09b] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 34(2):3–4, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:2009:SEEb

[Hen09c] Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 34(3):4–5, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hansen:2001:SDE

[HFA^{+}01] W. J. Hansen, J. T. Foreman, C. C. Albert, E. Axelband, L. L. Brownsword, and E. C. Forrester. Spiral development and evolutionary acquisition: the SEI-CSE workshop. ACM SIGSOFT Software Engineering Notes, 26 (6):59, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Han:2000:OOM

[HFH00] Yan Han, Liu Fengyu, and Zhang Hong. An objectoriented model of access control based on role. ACM SIGSOFT Software Engineering Notes, 25(2):64– 68, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hindle:2005:SFM

[HG05] Abram Hindle and Daniel M.

German. SCQL: a formal model and a query language for source control repositories. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hassan:2005:RMI

[HHM05] Ahmed E. Hassan, Richard C.

Holt, and Audris Mockus. Report on MSR 2004: International Workshop on Mining Software Repositories. ACM SIGSOFT Software Engineering Notes, 30 (1):4, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hierons:2004:FEE

[Hie04] R. M. Hierons. A flexible environment to evaluate state-based test techniques. ACM SIGSOFT Software Engineering Notes, 29(5):1–3, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hartmann:2000:UBI

[HIM00] Jean Hartmann, Claudio Imoberdorf, and Michael Meisinger. UML-Based integration testing. ACM SIGSOFT Software Engineering Notes, 25(5):60– 70, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hind:2000:NNP

[Hin00] Michael Hind. NPIC — New Paltz interprocedural compiler. ACM SIGSOFT Software Engineering Notes, 25 (1):57–58, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harman:2001:SWR

[HJ01] Mark Harman and Bryan F. Jones. The SEMINAL workshop: reformulating software engineering as a metaheuristic search problem. ACM SIGSOFT Software Engineering Notes, 26(6):62– 66, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henzinger:2005:PI

[HJM05] Thomas A. Henzinger, Ranjit Jhala, and Rupak Majumdar. Permissive interfaces. ACM SIGSOFT Software Engineering Notes, 30 (5):31–40, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hiller:2002:PEE

[HJS02] Martin Hiller, Arshad Jhumka, and Neeraj Suri. PROPANE: an environment for examining the propagation of errors in software. ACM SIGSOFT Software Engineering Notes, 27(4):81–85, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Higo:2005:ARS

[HKKI05] Yoshiki Higo, Toshihiro Kamiya, Shinji Kusumoto, and Katsuro Inoue. ARIES: refactoring support tool for code clone. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Haran:2005:ACT

[HKO^{+}05] Murali Haran, Alan Karr, Alessandro Orso, Adam Porter, and Ashish Sanil. Applying classification techniques to remotely-collected program execution data. ACM SIGSOFT Software Engineering Notes, 30(5): 146–155, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Han:2002:VIM

[HLL02] Hong Han, Jun Lu, and Xianliang Lu. Virtual interface machine: a design pattern to solve the problem of coupling and efficiency in software design. ACM SIGSOFT Software Engineering Notes, 27 (3):88–92, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:2005:MVH

[HmL05] Shih-Kun Huang and Kang min Liu. Mining version histories to verify the learning process of legitimate peripheral participants. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hartman:2004:ATM

[HN04] A. Hartman and K. Nagin. The AGEDIS tools for model based testing. ACM SIGSOFT Software Engineering Notes, 29(4): 129–132, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Halfond:2005:CSA

[HO05] William G. J. Halfond and Alessandro Orso. Combining static analysis and runtime monitoring to counter SQLinjection attacks. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Holzmann:2002:LB

[Hol02] Gerard J. Holzmann. The logic of bugs. ACM SIG-SOFT Software Engineering Notes, 27(6):81–87, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horwitz:2000:DRT

[Hor00]  Susan Horwitz.  Debug-

ging via run-time type checking. ACM SIGSOFT Software Engineering Notes, 25 (1):58, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horning:2001:BRS

[Hor01] Jim Horning. Book review:

Software fundamentals: collected papers by David L. Parnas. ACM SIGSOFT Software Engineering Notes, 26(4):91, July 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrison:2005:CMC

[HOST05] William Harrison, Harold Ossher, Stanley Sutton, and Peri Tarr. Concern modeling in the concern manipulation environment. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hind:2000:WPA

[HP00] Michael Hind and Anthony Pioli. Which pointer analysis should i use? ACM SIGSOFT Software Engineering Notes, 25(5):113– 123, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hanna:2006:SEB

[HR06] Youssef Hanna and Hridesh Rajan. SLEDE: eventbased specification of sensor network security protocols. ACM SIGSOFT Software Engineering Notes, 31 (6):1–2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Herraiz:2005:TPM

[HRGB05] Israel Herraiz, Gregorio Robles, and Jesus M. Gonzalez-Barahona. Towards predictor models for large libre software projects. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson-Sellers:2005:AOM

[HS05] Brian Henderson-Sellers. Agentoriented methodologies: method engineering and metamodelling. ACM SIGSOFT Software Engineering Notes, 30(4):1, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson-Sellers:2005:MEC

[HSGPSF05] B. Henderson-Sellers, C. Gonzalez-Perez, M. K. Serour, and D. G. Firesmith. Method engineering and COTS evaluation. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hovemeyer:2006:ETS

[HSP06] David Hovemeyer, Jaime Spacco, and William Pugh. Evaluating and tuning a static analysis to find null pointer bugs. ACM SIG-SOFT Software Engineering Notes, 31(1):13–19, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:2007:PDS

[Hua07] Gang Huang. Post-development software architecture. ACM SIGSOFT Software Engineering Notes, 32(5):1–9, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huber:2001:CHP

[Hub01]  Andy Huber.  Creating

high performance software development teams. ACM SIGSOFT Software Engineering Notes, 26(1):96– 97, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huber:2002:PRS

[Hub02]  Andy Huber.  Peer re-

views in software: a practical guide. ACM SIGSOFT Software Engineering Notes, 27 (2):72–73, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hudak:2000:PFR

[Hud00] Paul Hudak. Principles of functional reactive programming. ACM SIGSOFT Software Engineering Notes, 25 (1):59, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heimbigner:2000:DSE

[HW00] Dennis Heimbigner and Alexander L. Wolf. Distributed software engineering. ACM SIGSOFT Software Engineering Notes, 25 (1):99, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hordijk:2005:SFI

[HW05] Wiebe Hordijk and Roel Wieringa. Surveying the factors that influence maintainability: research design. ACM SIGSOFT Software Engineering Notes, 30(5): 385–388, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Holmes:2005:SER

[HWM05] Reid Holmes, Robert J.

Walker, and Gail C. Murphy. Strathcona example recommendation tool. ACM SIGSOFT Software Engineering Notes, 30(5):237– 240, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hampapuram:2006:SPS

[HYD06] Hari Hampapuram, Yue Yang, and Manuvir Das. Symbolic path simulation in path-sensitive dataflow analysis. ACM SIGSOFT Software Engineering Notes, 31 (1):52–58, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hildebrandt:2000:SFI

[HZ00] Ralf Hildebrandt and Andreas Zeller. Simplifying failure-inducing input. ACM SIGSOFT Software Engineering Notes, 25(5):135– 145, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hall:2004:MII

[HZ04] Robert J. Hall and Andrea Zisman. Model interchange and integration for web services. ACM SIGSOFT Software Engineering Notes, 29 (5):1–11, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Inverardi:2005:CET

[IMP05] Paola Inverardi, Henry Muccini, and Patrizio Pelliccione. CHARMY: an extensible tool for architectural analysis. ACM SIGSOFT Software Engineering Notes, 30 (5):111–114, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Inverardi:2002:SPS

[Inv02] Paola Inverardi. The SAL-ADIN project: summary report. ACM SIGSOFT Software Engineering Notes, 27 (3):38–43, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Inverardi:2001:ASD

[IT01] Paola Inverardi and Massimo Tivoli. Automatic synthesis of deadlock free connectors for COM/DCOM applications. ACM SIGSOFT Software Engineering Notes, 26 (5):121–131, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:2000:AFO

[Jac00] Daniel Jackson. Automating first-order relational logic.

ACM SIGSOFT Software Engineering Notes, 25(6): 130–139, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jahnke:2001:ECB

[Jah01] Jens H. Jahnke. Engineering component-based netcentric systems for embedded applications. ACM SIGSOFT Software Engineering Notes, 26(5):218– 228, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jang:2001:ADR

[Jan01] Jun-Jang Jang. An approach to designing reusable service frameworks via virtual service machine. ACM SIGSOFT Software Engineering Notes, 26(3):58– 66, May 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jochen:2005:ECA

[JAPM05] Mike Jochen, Anteneh Addis Anteneh, Lori L. Pollock, and Lisa M. Marvel. Enabling control over adaptive program transformation for dynamically evolving mobile software validation. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jayaraman:2000:RSC

[Jay00] Bharat Jayaraman. Research on sets, constraints, and preferences. ACM SIGSOFT Software Engineering Notes, 25(1):60, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jayaswall:2006:RAD

[Jay06] Bijay Jayaswall. Response from authors of “Design For Trustworthy Software”. ACM SIGSOFT Software Engineering Notes, 31(6): 59, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Sch06d].

Jain:2005:DTV

[JB05] Apurva Jain and Barry Boehm. Developing a theory of value-based software engineering. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jiau:2009:TCD

[JC09] Hewijin Christine Jiau and Jinghong Cox Chen. Test code differencing for testdriven refactoring automation. ACM SIGSOFT Software Engineering Notes, 34 (1):1–10, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Javed:2004:SII

[JeMD04a] Talha Javed, Manzil e Maqsood, and Qaiser S. Durrani. A study to investigate the impact of requirements instability on software defects. ACM SIGSOFT Software Engineering Notes, 29 (3):1–7, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Javed:2004:SEE

[JeMD04b] Talha Javed, Manzil e Maqsood, and Qaiser S. Durrani. A survey to examine the effect of team communication on job satisfaction in software industry. ACM SIGSOFT Software Engineering Notes, 29(2):6, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jiufu:2007:ISB

[Jiu07] Liu Jiufu. Integration of statechart and b method based analysis and verification for flight control software of unmanned aerial vehicle. ACM SIGSOFT Software Engineering Notes, 32 (2):1–4, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jagadeesan:2000:TP

[JLG00] R. Jagadeesan, K. Läufer, and V. Gupta. The Triveni project. ACM SIGSOFT Software Engineering Notes, 25(1):59, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:2000:MAG

[JM00] Lewis Johnson and Stacy Marsella. MediaDoc: automated generation of multimedia explanatory presentations. ACM SIGSOFT Software Engineering Notes, 25 (1):105, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jakobac:2005:SAC

[JME05] Vladimir Jakobac, Nenad Medvidovic, and Alexander Egyed. Separating architectural concerns to ease program understanding. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

John:2005:HSF

[JMT05] Michael John, Frank Maurer, and Bjørnar Tessem. Human and Social Factors of Software Engineering: workshop summary. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Juristo:2004:TBS

[JMV04] N. Juristo, A. M. Moreno, and S. Vegas. Towards building a solid empirical body of knowledge in testing techniques. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johann:2000:TEP

[Joh00a]  Patricia Johann.  Test-

ing and enhancing a prototype program fusion engine. ACM SIGSOFT Software Engineering Notes, 25 (1):60–61, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:2000:PLL

[Joh00b] Philip M. Johnson. Project LEAP: lightweight, empirical, anti-measurement dysfunction, and portable software developer improvement. ACM SIGSOFT Software Engineering Notes, 25(1): 61, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

John:2005:EBP

[Joh05] Bonnie E. John. Evidencebased practice in humancomputer interaction and evidence maps. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Julien:2002:ECA

[JR02] Christine Julien and Gruia Catalin Roman. Egocentric contextaware programming in ad hoc mobile environments. ACM SIGSOFT Software Engineering Notes, 27(6):21– 30, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jamali:2005:LAR

[JR05a] Nadeem Jamali and Shangping Ren. A layered architecture for real-time distributed multi-agent systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–8, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Janakiram:2005:RRD

[JR05b] D. Janakiram and M. S. Rajasree. ReQuEst: Requirementsdriven quality estimator. ACM SIGSOFT Software Engineering Notes, 30(1): 4, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Juric:2000:ILS

[JRHD00] Matjaz B. Juric, Ivan Rozman, Marjan Hericko, and Tomaz Domajnko. Integrating legacy systems in distributed object architecture. ACM SIGSOFT Software Engineering Notes, 25 (2):35–39, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:2000:CRT

[JS00] Daniel Jackson and Kevin Sullivan. COM revisited: tool-assisted modelling of an architectural framework. ACM SIGSOFT Software Engineering Notes, 25(6): 149–158, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jagannathan:2001:ICS

[JS01] Ramesh Jagannathan and Paolo A. G. Sivilotti. Increasing client-side confidence in remote component implementations. ACM SIGSOFT Software Engineering Notes, 26(5):52–61, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:2001:MM

[JSS01] Daniel Jackson, Ilya Shlyakhter, and Manu Sridharan. A micromodularity mechanism. ACM SIGSOFT Software Engineering Notes, 26(5):62– 73, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:2000:FBC

[JV00] Daniel Jackson and Mandana Vaziri. Finding bugs with a constraint solver. ACM SIGSOFT Software Engineering Notes, 25(5):14– 25, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jahnke:2001:ITI

[JW01] Jens H. Jahnke and Joerg P. Wadsack. ISE3: the Third International Workshop on Intelligent Software Engineering open issues, new techniques, challenge problems in software engineering. ACM SIGSOFT Software Engineering Notes, 26 (1):43, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaiser:2000:IBIa

[Kai00a] Gail E. Kaiser. Internetbased information management technology. ACM SIG-SOFT Software Engineering Notes, 25(1):98, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaiser:2000:IBIb

[Kai00b] Gail E. Kaiser. Internetbased information management technology. ACM SIGSOFT Software Engineering Notes, 25(1):103, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaminski:2005:UNY

[KAKM05] Piotr Kaminski, Priyanka Agrawal, Holger Kienle, and Hausi Müller. < username >, I need you!: initiative and interaction in autonomic systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karsai:2000:MIC

[Kar00] Gabor Karsai. Model integrated computing (MIC). ACM SIGSOFT Software Engineering Notes, 25(1): 105, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kremenek:2004:CEE

[KAYE04] Ted Kremenek, Ken Ashcraft, Junfeng Yang, and Dawson Engler. Correlation exploitation in error ranking.

ACM SIGSOFT Software Engineering Notes, 29(6):83– 93, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnamurthi:2005:DSC

[KB05] Shriram Krishnamurthi and Tevfik Bultan. Discussion summary: characteristics of web services and their impact on testing, analysis and verification. ACM SIGSOFT Software Engineering Notes, 30(1):5, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:2005:REB

[KBBL05] Barbara Kitchenham, David Budgen, Pearl Brereton, and Step Hen Linkman. Realising evidence-based software engineering. ACM SIGSOFT Software Engineering Notes, 30(4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kagdi:2005:TTA

[KCM05] Huzefa Kagdi, Michael L. Collard, and Jonathan I. Maletic. Towards a taxonomy of approaches for mining of source code repositories. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2006:LOP

[KCS06] Naveen Kumar, Bruce R.

Childers, and Mary Lou Soffa. Low overhead program monitoring and profiling. ACM SIGSOFT Software Engineering Notes, 31 (1):28–34, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kolano:2000:CSA

[KD00] Paul Z. Kolano and Richard A.

Demmerer. Classification schemes to aid in the analysis of real-time systems. ACM SIGSOFT Software Engineering Notes, 25(5):86– 95, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khemakhem:2007:SES

[KDJ07] Sofien Khemakhem, Khalil Drira, and Mohamed Jmaiel. SEC+: an enhanced search engine for component-based software development. ACM SIGSOFT Software Engineering Notes, 32(4):4:1– 4:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khemakhem:2008:EESa

[KDKJ08a] Sofien Khemakhem, Khalil Drira, Emna Khemakhem, and Mohamed Jmaiel. An experimental evaluation of SEC+, an enhanced search engine for component-based software development. ACM SIGSOFT Software Engineering Notes, 33(3):3:1– 3:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khemakhem:2008:EESb

[KDKJ08b] Sofien Khemakhem, Khalil Drira, Emna Khemakhem, and Mohamed Jmaiel. An experimental evaluation of SEC+, an enhanced search engine for component-based software development. ACM SIGSOFT Software Engineering Notes, 33(4):3:1– 3:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaveh:2001:DDD

[KE01] Nima Kaveh and Wolfgang Emmerich. Deadlock detection in distribution object systems. ACM SIGSOFT Software Engineering Notes, 26(5):44–51, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Keil:2005:PAT

[Kei05]  Patrick Keil.  Principal

agent theory and its application to analyze outsourcing of software development. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnamurthi:2004:VAA

[KFG04] Shriram Krishnamurthi, Kathi Fisler, and Michael Greenberg. Verifying aspect advice modularly. ACM SIG-SOFT Software Engineering Notes, 29(6):137–146, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kfoury:2000:TTR

[Kfo00] A. J. Kfoury. Type theory and rewriting theory for expressive, efficient and reliable programming languages. ACM SIGSOFT Software Engineering Notes, 25(1):61– 62, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kiebusch:2005:MSS

[KFS05a] Sebastian Kiebusch, Bogdan Franczyk, and Andreas Speck. Metrics for software system families. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kiebusch:2005:RTM

[KFS05b] Sebastian Kiebusch, Bogdan Franczyk, and Andreas Speck. A real time measure of software system families. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kulak:2001:UCR

[KG01] Daryl Kulak and Eamonn Guiney. Use cases: requirements in context. ACM SIGSOFT Software Engineering Notes, 26(1):101, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Klein:2005:ADP

[KG05] Florian Klein and Holger Giese. Analysis and design of physical and social contexts in multi-agent systems using UML. ACM SIGSOFT Software Engineering Notes, 30(4):1–8, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2009:QEA

[KGK09] Avadhesh Kumar, P. S. Grover, and Rajesh Kumar. A quantitative evaluation of aspect-oriented software quality model (AOSQUAMO). ACM SIGSOFT Software Engineering Notes, 34(5):1– 9, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kendapadi:2005:BMN

[KGS05] Ananth Kendapadi, Matt Gandolfo, and Amitosh Shukla. BatchFlow: a method and notation to visualize RDBMS batch jobs. ACM SIGSOFT Software Engineering Notes, 30(3): 1–5, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2009:AOR

[KGS09] Manoj Kumar, Anjana Gosain, and Yogesh Singh. Agent oriented requirements engineering for a data warehouse. ACM SIGSOFT Software Engineering Notes, 34(5):1–4, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kiczales:2001:AOP

[KH01] Gregor Kiczales and Erik Hilsdale. Aspect-oriented programming. ACM SIG-SOFT Software Engineering Notes, 26(5):313, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koponen:2005:OSS

[KH05] Timo Koponen and Virpi Hotti. Open source software maintenance process framework. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kramer:2006:RAS

[KH06] Jeff Kramer and Orit Hazzan. The role of abstraction in software engineering. ACM SIGSOFT Software Engineering Notes, 31 (6):38–39, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kienle:2001:EFB

[Kie01]  Holger M. Kienle.  Ex-

change format bibliography. ACM SIGSOFT Software Engineering Notes, 26(1):56– 60, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:2002:VRE

[Kim02] Hyoseob Kim. Visit report: 2nd European COTS User Working Group Workshop (ECUA2001). ACM SIG-SOFT Software Engineering Notes, 27(1):30, January 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kirda:2001:EWS

[Kir01] Engin Kirda. Engineering of Web services with XML and XSL. ACM SIGSOFT Software Engineering Notes, 26 (5):318–319, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kuncak:2005:RAA

[KJ05] Viktor Kuncak and Daniel Jackson. Relational analysis of algebraic datatypes. ACM SIGSOFT Software Engineering Notes, 30(5):207– 216, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kontio:2005:CPR

[KJML05] Jyrki Kontio, Jani-Pekka Jokinen, Markus M. Mäkelä, and Virve Leino. Current practices and research opportunities in software business models. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:2005:PLS

[KJP05] Minseong Kim, Jaemin Jeong, and Sooyong Park. From product lines to self-managed systems: an architecturebased runtime reconfiguration framework. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kotb:2006:ISD

[KK06] Yasser Kotb and Takuya Katayama. Improving software documentation using textual entailment. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnan:2007:CQT

[KKB07] R. Krishnan, S. Murali Krishna, and Nishil Bharill. Code quality tools: learning from our experience. ACM SIGSOFT Software Engineering Notes, 32(4): 5:1–5:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2009:GCM

[KKG09] Avadhesh Kumar, Rajesh Kumar, and P. S. Grover. Generalized coupling measure for aspect-oriented systems. ACM SIGSOFT Software Engineering Notes, 34 (3):1–6, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnan:2007:CTL

[KKN07] R. Krishnan, S. Murali Krishna, and P. Siva Nandhan. Combinatorial testing: learnings from our experience. ACM SIGSOFT Software Engineering Notes, 32(3):1–8, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kruchten:2001:IWD

[KKS01] Philippe Kruchten, Wojtek Kozaczynski, and Bran Selic. ICSE 2001 Workshop on Describing Software Architecture with UML. ACM SIGSOFT Software Engineering Notes, 26(6):78– 79, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koru:2005:IEM

[KL05] A. Günes Koru and Hongfang Liu. An investigation of the effect of module size on defect prediction using static measures. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2000:SAM

[KLS00] Bharath M. Kumar, R. Lakshminarayanan, and Y. N. Srikant. A study of automatic migration of programs across the Java event models. ACM SIGSOFT Software Engineering Notes, 25 (3):24–29, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ko:2005:HFA

[KM05a] Andrew J. Ko and Brad A. Myers. Human factors affecting dependability in end-user programming. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kramer:2005:EDS

[KM05b] Jeff Kramer and Jeff Magee. Engineering distributed software: a structural discipline. ACM SIGSOFT Software Engineering Notes, 30 (5):283–285, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kushwaha:2006:CCM

[KM06a] Dharmender Singh Kushwaha and A. K. Misra. Cognitive complexity metrics and its impact on software reliability based on cognitive software development model. ACM SIGSOFT Software Engineering Notes, 31 (2):1–6, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kushwaha:2006:ICI

[KM06b] Dharmender Singh Kushwaha and A. K. Misra. Improved cognitive information complexity measure: a metric that establishes program comprehension effort. ACM SIGSOFT Software Engineering Notes, 31(5):1– 7, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kushwaha:2006:MCI

[KM06c] Dharmender Singh Kushwaha and A. K. Misra. A modified cognitive information complexity measure of software. ACM SIGSOFT Software Engineering Notes, 31(1):1–4, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kushwaha:2006:RAC

[KM06d] Dharmender Singh Kushwaha and A. K. Misra. Robustness analysis of cognitive information complexity measure using Weyuker properties. ACM SIGSOFT Software Engineering Notes, 31 (1):1–6, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kushwaha:2008:STE

[KM08] Dharmender Singh Kushwaha and A. K. Misra.

Software test effort estimation. ACM SIGSOFT Software Engineering Notes, 33 (3):6:1–6:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khan:2009:MBT

[KM09] R. A. Khan and K. Mustafa. Metric based testability model for object oriented design (MTMOOD). ACM SIGSOFT Software Engineering Notes, 34(2):1–6, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kiniry:2006:FTR

[KMC^{+}06] Joseph R. Kiniry, Alan E. Morkan, Dermot Cochran, Martijn Oostdijk, and Engelbert Hubbers. Formal techniques in a remote voting system. ACM SIGSOFT Software Engineering Notes, 31 (6):1–2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kruger:2005:SAE

[KMM05] Ingolf H. Krüger, Reena Mathew, and Michael Meisinger. From scenarios to aspects: exploring product lines. ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:2005:UCG

[KN05] Miryung Kim and David Notkin. Using a clone genealogy extractor for understanding and supporting evolution of code clones. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knight:2005:FSE

[Kni05] John C. Knight. Focusing software education on engineering. ACM SIGSOFT Software Engineering Notes, 30(2):3–5, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koru:2005:EHM

[KON05] A. Günes Koru, A. Ant Ozok, and Anthony F. Norcio. The effect of human memory organization on code reviews under different single and pair code reviewing scenarios. ACM SIGSOFT Software Engineering Notes, 30 (4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:2002:PSRc

[KP02a] Barbara Kitchenham and Shari Lawrence Pfleeger. Principles of survey research part 4: questionnaire evaluation. ACM SIGSOFT Software Engineering Notes, 27 (3):20–23, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:2002:PSRd

[KP02b] Barbara Kitchenham and Shari Lawrence Pfleeger. Principles of survey research: part 5: populations and samples. ACM SIGSOFT Software Engineering Notes, 27 (5):17–20, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:2002:PSRa

[KP02c] Barbara A. Kitchenham and Shari Lawrence Pfleeger. Principles of survey research part 2: designing a survey. ACM SIGSOFT Software Engineering Notes, 27 (1):18–20, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:2002:PSRb

[KP02d] Barbara A. Kitchenham and Shari Lawrence Pfleeger. Principles of survey research: part 3: constructing a survey instrument. ACM SIGSOFT Software Engineering Notes, 27(2):20–24, March 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karr:2005:DPT

[KP05] Alan F. Karr and Adam A.

Porter. Distributed performance testing using statistical modeling. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:2000:ASS

[KPCD00] Peter Kokol, Vili Podgorelec, Ana Isabel Cardoso, and Francis Dion. Assessing the state of the software process development using the chaos theory. ACM SIGSOFT Software Engineering Notes, 25 (3):41–43, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:2000:IES

[KPDdL00] Peter Kokol, Vili Podgorelec, Francis Dion, and Rich de Loach. Intellectual energy in software design. ACM SIGSOFT Software Engineering Notes, 25 (3):44–45, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knauber:2001:PSP

[KS01] Peter Knauber and Giancarlo Succi. Perspectives on software product lines: report on First International Workshop on Software Product Lines: Economics, Architectures, and Implications Workshop #15 at 22nd International Conference on Software Engineering (ICSE). ACM SIGSOFT Software Engineering Notes, 26(2): 29–33, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knauber:2002:PSP

[KS02] Peter Knauber and Giancarlo Succi. Perspectives on software product lines: report on Second International Workshop on Software Product Lines: Economics, Architectures, and Implications Workshop at 23rd International Conference on Software Engineering (ICSE). ACM SIGSOFT Software Engineering Notes, 27(2): 40–45, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khurshid:2006:GSE

[KS06] Sarfraz Khurshid and Yuk Lai Suen. Generalizing symbolic execution to library classes. ACM SIGSOFT Software Engineering Notes, 31 (1):103–110, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kalaimagal:2008:RSC

[KS08a] Sivamuni Kalaimagal and Rengaramanujam Srinivasan. A retrospective on software component quality models. ACM SIGSOFT Software Engineering Notes, 33(6):1– 10, November 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2008:CPS

[KS08b] Parminder Kaur and Hardeep Singh. Certification process of software components. ACM SIGSOFT Software Engineering Notes, 33 (4):4:1–4:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kharb:2008:CMC

[KS08c] Latika Kharb and Rajender Singh. Complexity metrics for component-oriented software systems. ACM SIGSOFT Software Engineering Notes, 33(2):4:1– 4:??, March 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kalaimagal:2009:NTC

[KS09a] Sivamuni Kalaimagal and Rengaramanujam Srinivasan. The need for transforming the COTS component quality evaluation standard mirage to reality. ACM SIGSOFT Software Engineering Notes, 34(5):1–4, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2009:EES

[KS09b] Kuljit Kaur and Hardeep Singh. Evaluating an evolving software component: case of internal design. ACM SIG-SOFT Software Engineering Notes, 34(4):1–4, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2009:LSU

[KS09c] Parminder Kaur and Hardeep Singh. A layered structure for uniform version management in component based systems. ACM SIGSOFT Software Engineering Notes, 34(6):1–7, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2009:VMC

[KS09d] Parminder Kaur and Hardeep Singh. Version management and composition of software components in different phases of software development life cycle. ACM SIG-SOFT Software Engineering Notes, 34(4):1–9, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kushwaha:2006:CWB

[KSM06] Dharmender Singh Kushwaha, R. K. Singh, and A. K. Misra. Cognitive web based software development process: towards a more reliable approach. ACM SIGSOFT Software Engineering Notes, 31(4):1–6, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:2005:ESC

[KSNM05] Miryung Kim, Vibha Sazawal, David Notkin, and Gail Murphy. An empirical study of code clone genealogies. ACM SIGSOFT Software Engineering Notes, 30 (5):187–196, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kulczycki:2006:SBA

[KSWR06] Gregory Kulczycki, Murali Sitaraman, Bruce W. Weide, and Atanas Rountev. A specification-based approach to reasoning about pointers. ACM SIGSOFT Software Engineering Notes, 31(2): 7:1–7:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kanaskar:2005:GSM

[KTB05] Nitin V. Kanaskar, Umit Topaloglu, and Coskun Bayrak. Globus security model for grid environment. ACM SIGSOFT Software Engineering Notes, 30(6):1– 9, November 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kanmani:2004:IEO

[KUST04a] S. Kanmani, V. Rhymend Uthariaraj, V. Sankaranarayanan, and P. Thambidurai. Investigation into the exploitation of objectoriented features. ACM SIG-SOFT Software Engineering Notes, 29(2):7, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kanmani:2004:OOS

[KUST04b] S. Kanmani, V. Rhymend Uthariaraj, V. Sankaranarayanan, and P. Thambidurai. Object oriented software quality prediction using general regression neural networks. ACM SIGSOFT Software Engineering Notes, 29(5):1–6, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koshkina:2004:MVW

[KvB04] Mariya Koshkina and Franck van Breugel. Modelling and verifying web service orchestration by means of the concurrency workbench. ACM SIGSOFT Software Engineering Notes, 29(5):1– 10, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:2005:ASC

[KWB05] Sunghun Kim, E. James Whitehead, and Jennifer Bevan. Analysis of signature change patterns. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kuehne:2005:SAU

[KWD05] Ralph Kuehne, Cornelius Wille, and Reiner Dumke. Software agents using simulation for decision-making. ACM SIGSOFT Software Engineering Notes, 30(1): 5, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kong:2009:UCM

[KY09] Leilei Kong and Tao Yuan.

Use case modeling approach for early aspect acquisition. ACM SIGSOFT Software Engineering Notes, 34 (4):1–6, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lago:2006:FWS

[LA06] Patricia Lago and Paris Avgeriou. First Workshop on Sharing and Reusing Architectural Knowledge. ACM SIGSOFT Software Engineering Notes, 31(5):32– 36, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lakey:2005:MBS

[Lak05] Peter B. Lakey. Model-based specification and testing applied to the ground-based midcourse defense (GMD) system: an industry report. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lanman:2006:ISE

[Lan06] Jeremy T. Lanman. Industrial software engineering: developing software systems. ACM SIGSOFT Software Engineering Notes, 31 (2):3–5, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Larsen:2001:JPB

[Lar01] Albert L. Larsen. Java programming: from the beginning. ACM SIGSOFT Software Engineering Notes, 26 (1):97–98, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Larson:2006:SID

[Lar06] Eric Larson. SUDS: an infrastructure for dynamic software bug detection using static analysis. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lev-Ami:2000:PSA

[LARSW00] Tal Lev-Ami, Thomas Reps, Mooly Sagiv, and Reinhard Wilhelm. Putting static analysis to work for verification: a case study. ACM SIGSOFT Software Engineering Notes, 25(5):26–38, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2001:QCQ

[Law01a] Jim Law. Quantum computation and quantum information. ACM SIGSOFT Software Engineering Notes, 26(4):91, July 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2001:SPA

[Law01b]  Jim Law.  Selected pa-

pers on analysis of algorithms. ACM SIGSOFT Software Engineering Notes, 26 (1):98–99, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2002:MDI

[Law02]  James Law.  Mechatron-

ics and the design of intelligent machines and systems. ACM SIGSOFT Software Engineering Notes, 27 (3):93, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2004:RTR

[Law04a]  James Law.  Review of

Term rewriting systems by Mark Bezem, Jan Willem Klop, and Roel de Vrijer. Cambridge University Press 2003. ACM SIGSOFT Software Engineering Notes, 29 (2):33, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2004:RDL

[Law04b] James Law. Review of The description logic handbook by Franz Bader, Diego Calvanese, Deborah L. McGuinness, Daniele Nardi, Peter Patel-Schneider. Cambridge University Press 2003. ACM SIGSOFT Software Engineering Notes, 29(2): 32–33, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lawler:2004:RML

[Law04c] Brian Lawler. Review of Modernizing legacy systems: software technologies, engineering processes and business practices by Robert C. Seacord, Daniel Plakosh and Grace A. Lewis. Addison Wesley 2003. ACM SIG-SOFT Software Engineering Notes, 29(5):37, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lawler:2005:RCD

[Law05] Brian Lawler. Review of CMMI Distilled: a Practical Introduction to Integrated Process Improvement, Second Edition by Dennis M. Ahern, Aaron Clouse and Richard Turner. Addison Wesley, 2004, paperback, ISBN 0-321-18613-3. ACM SIGSOFT Software Engineering Notes, 30(1):37– 38, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2006:RMI

[Law06] James Law. Review of Migrating to IPv6 by Marc Blanchet, John Wiley and Sons, Ltd., 2006, 0-471-49892-0. ACM SIGSOFT Software Engineering Notes, 31(4):39–40, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

LeGear:2005:RTC

[LB05] Andrew Le Gear and Jim Buckley. Reengineering towards components using “Reconn-exion”. ACM SIG-SOFT Software Engineering Notes, 30(5):370–373, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leavens:2006:PDJ

[LBR06] Gary T. Leavens, Albert L. Baker, and Clyde Ruby. Preliminary design of JML: a behavioral interface specification language for Java. ACM SIGSOFT Software Engineering Notes, 31(3): 1–38, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:2005:EAP

[LC05] Amy Law and Raylene Charron. Effects of agile practices on social factors. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2005:DEB

[LCLP05] WenQian Liu, Charles L. Chen, Vidya Lakshminarayanan, and Dewayne E. Perry. A design for evidence — based soft research. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2005:BDS

[LCS^{+}05] Jingyue Li, Reidar Conradi, Odd Petter N. Slyngstad, Christian Bunse, Umair Khan, Maurizio Morisio, and Marco Torchiano. Barriers to disseminating off-the-shelf based development theories to IT industry. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lins:2007:AWS

[LdSJR07] Fernando Antônio Aires Lins, José Carlos dos Santos Júnior, and Nelson Souto Rosa. Adaptive web service composition. ACM SIGSOFT Software Engineering Notes, 32(4):6:1– 6:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lin:2004:IAM

[LE04] Lee Lin and Michael D.

Ernst. Improving the adaptability of multi-mode systems via program steering. ACM SIGSOFT Software Engineering Notes, 29(4): 206–216, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leavens:2000:FMM

[Lea00] Gary T. Leavens. Formal methods for multimethod software components. ACM SIGSOFT Software Engineering Notes, 25(1):62– 63, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:2009:PRA

[Lee09] Tony Lee. Peer review assessment standard for object oriented analysis and design. ACM SIGSOFT Software Engineering Notes, 34 (1):1–3, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lehoczky:2000:IIS

[Leh00] John Lehoczky. INSERT — Incremental Software Evolution for Real-Time Systems. ACM SIGSOFT Software Engineering Notes, 25 (1):103, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lerner:2004:VPM

[Ler04] Barbara Staudt Lerner. Verifying process models built using parameterized state machines. ACM SIGSOFT Software Engineering Notes, 29 (4):274–284, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lethbridge:2006:RC

[Let06] Timothy C. Lethbridge. Reflections on CSEE&T 2006. ACM SIGSOFT Software Engineering Notes, 31(4): 8–9, July 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Levine:2009:DDE

[Lev09] Gertrude Neuman Levine.

Defining defects, errors, and service degradations. ACM SIGSOFT Software Engineering Notes, 34(2): 1–14, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lopes:2002:APD

[LFW02] Antónia Lopes, José Luiz Fiadeiro, and Michel Wermelinger. Architectural primitives for distribution and mobility. ACM SIGSOFT Software Engineering Notes, 27(6):41–50, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lawrie:2002:IDO

[LG02] Tony Lawrie and Cristina Gacek. Issues of dependability in open source software development. ACM SIGSOFT Software Engineering Notes, 27(3):34– 37, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:2005:ETS

[LGA05] Seok Won Lee, Robin A.

Gandhi, and Gail-Joon Ahn. Establishing trustworthiness in services of the critical infrastructure through certification and accreditation. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lindstrom:2004:UES

[LGO04] B. Lindström, M. Grindal, and J. Offutt. Using an existing suite of test objects: experience from a testing experiment. ACM SIGSOFT Software Engineering Notes, 29(5):1–3, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2002:AIF

[Li02] Bixin Li. Analyzing informationflow in Java program based on slicing technique. ACM SIGSOFT Software Engineering Notes, 27(5):98– 103, September 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lieberherr:2000:DAP

[Lie00] Karl J. Lieberherr. Demeter/ adaptive programming. ACM SIGSOFT Software Engineering Notes, 25(1):100– 101, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2000:TRA

[Liu00] Chang Liu. Term rewriting and all that. ACM SIGSOFT Software Engineering Notes, 25(3):63, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lott:2005:MRC

[LJD05] C. Lott, A. Jain, and S. Dalal. Modeling requirements for combinatorial software testing. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2002:VCC

[LKF02] Harry Li, Shriram Krishnamurthi, and Kathi Fisler. Verifying cross-cutting features as open systems. ACM SIGSOFT Software Engineering Notes, 27(6):89– 98, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Letier:2005:FTL

[LKMU05] Emmanuel Letier, Jeff Kramer, Jeff Magee, and Sebastian Uchitel. Fluent temporal logic for discrete-time event-based models. ACM SIGSOFT Software Engineering Notes, 30(5):70– 79, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lather:2000:SAS

[LKS00] Anu Singh Lather, Shakti Kumar, and Yogesh Singh. Suitability assessment of software developers: a fuzzy approach. ACM SIGSOFT Software Engineering Notes, 25(3):30–31, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lapierre:2001:DTS

[LLL01] Sébastien Lapierre, Bruno Laguë, and Charles Leduc. Datrix ^{T M} source code model and its interchange format: lessons learned and considerations for future work. ACM SIGSOFT Software Engineering Notes, 26(1):53– 56, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lapouchnian:2005:TRD

[LLMY05] Alexei Lapouchnian, Sotirios Liaskos, John Mylopoulos, and Yijun Yu. Towards requirements-driven autonomic systems design. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2005:MCS

[LLT05] Jing (Janet) Liu, Robyn R.

Lutz, and Jeffrey M. Thompson. Mapping concern space to software architecture: a connector-based approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Latella:2002:TCR

[LM02] Diego Latella and Mieke Massink. On testing and conformance relations for UML statechart diagrams behaviours. ACM SIGSOFT Software Engineering Notes, 27(4):144–153, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lallchandani:2008:SUA

[LM08] Jaiprakash T. Lallchandani and R. Mall. Slicing UML architectural models. ACM SIGSOFT Software Engineering Notes, 33(3):4:1– 4:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laguna:2001:URE

[LMG01] Miguel Angel Laguna, José Manuel Marqués, and Francisco José García. A user requirements elicitation tool. ACM SIGSOFT Software Engineering Notes, 26(2):35– 37, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lima:2004:AMA

[LMSF04] Emerson F. A. Lima, Patrícia D. L. Machado, Fláavio R. Sampaio, and Jorge C. A. Figueiredo. An approach to modelling and applying mobile agent design patterns. ACM SIGSOFT Software Engineering Notes, 29 (3):1–8, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2005:MIP

[LN05] Xing Li and Ramesh Nagarajan. Modeling for image processing system validation, verification and testing. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Long:2001:SRA

[Lon01] John Long. Software reuse antipatterns. ACM SIG-SOFT Software Engineering Notes, 26(4):68–76, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Long:2004:SNK

[Lon04] Brad Long. Sorting nonkey fields in the distributed result set iterator pattern. ACM SIGSOFT Software Engineering Notes, 29(5):1– 2, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Langari:2005:QCF

[LP05] Zarrin Langari and Anne Banks Pidduck. Quality, cleanroom and formal methods. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liang:2002:EPS

[LPH02] Donglin Liang, Maikel Pennings, and Mary Jean Harrold. Evaluating the precision of static reference analysis using profiling. ACM SIGSOFT Software Engineering Notes, 27(4):22–32, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liang:2006:EIC

[LPH06] Donglin Liang, Maikel Pennings, and Mary Jean Harrold. Evaluating the impact of context-sensitivity on Andersen’s algorithm for Java programs. ACM SIGSOFT Software Engineering Notes, 31(1):6–12, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lavazza:2001:CUF

[LQV01] Luigi Lavazza, Gabriele Quaroni, and Matteo Venturelli. Combining UML and formal notations for modelling real-time systems. ACM SIGSOFT Software Engineering Notes, 26(5):196– 206, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lippert:2001:AXC

[LR01a] Martin Lippert and Stefan Roock. Adapting XP to complex application domains. ACM SIGSOFT Software Engineering Notes, 26 (5):316–317, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Luer:2001:WEC

[LR01b] Chris Lüer and David S. Rosenblum. WREN — an environment for componentbased development. ACM SIGSOFT Software Engineering Notes, 26(5):207– 217, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2002:RDR

[LR02] Chang Liu and Debra J.

Richardson. Research directions in RAIC. ACM SIGSOFT Software Engineering Notes, 27(3):43– 46, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loconsole:2001:RMS

[LRBH01] Annabella Loconsole, Daniel Rodriguez, Jürgen Börstler, and Rachel Harrison. Report on metrics 2001: the science and practice of software metrics conference. ACM SIGSOFT Software Engineering Notes, 26(6):52– 57, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lumpe:2006:CEM

[LS06] Markus Lumpe and Jean-Guy Schneider. Classboxes: an experiment in modeling compositional abstractions using explicit contexts. ACM SIGSOFT Software Engineering Notes, 31(2): 6:1–6:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lewis:2009:RWS

[LSBM09] Grace Lewis, Dennis Smith, Len Bass, and Brad Myers. Report of the Workshop on Software Engineering Foundations for End-User Programming. ACM SIGSOFT Software Engineering Notes, 34(5):51–54, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2004:EED

[LSH^{+}04] Paul Luo Li, Mary Shaw, Jim Herbsleb, Bonnie Ray, and P. Santhanam. Empirical evaluation of defect projection models for widely-deployed production software systems. ACM SIGSOFT Software Engineering Notes, 29(6):263– 272, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2005:QAS

[LSK^{+}05] Xiaoqing (Frank) Liu, Yan Sun, Gautam Kane, Yuji Kyoya, and Kunio Noguchi. QFD application in software process management and improvement based on CMM. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lu:2004:SRC

[Lu04] Jian Lu. Some research on componentware frameworks based on mobile agent technology. ACM SIGSOFT Software Engineering Notes, 29 (2):8, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Luckham:2000:CEP

[Luc00]  David Luckham.  Com-

plex event processing (CEP). ACM SIGSOFT Software Engineering Notes, 25(1): 99, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Luttgen:2000:CAS

[LvdBC00] Gerald Lüttgen, Michael von der Beeck, and Rance Cleaveland. A compositional approach to statecharts semantics. ACM SIGSOFT Software Engineering Notes, 25(6):120– 129, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Letier:2002:DOS

[LvL02] Emmanuel Letier and Axel van Lamsweerde. Deriving operational software specifications from system goals. ACM SIGSOFT Software Engineering Notes, 27(6): 119–128, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Letier:2004:RAP

[LvL04] Emmanuel Letier and Axel van Lamsweerde. Reasoning about partial goal satisfaction for requirements and design engineering. ACM SIGSOFT Software Engineering Notes, 29(6):53– 62, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leveson:2004:MES

[LW04] Nancy G. Leveson and Kathryn Anne Weiss. Making embedded software reuse practical and safe. ACM SIGSOFT Software Engineering Notes, 29(6):171– 178, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Litoiu:2005:HMB

[LWZ05] Marin Litoiu, Murray Woodside, and Tao Zheng. Hierarchical model-based autonomic control of software systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2005:SSM

[LYF^{+}05] Chao Liu, Xifeng Yan, Long Fei, Jiawei Han, and Samuel P. Midkiff. SOBER: statistical model-based bug localization. ACM SIG-SOFT Software Engineering Notes, 30(5):286–295, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2005:PMA

[LZ05a] Zhenmin Li and Yuanyuan Zhou. PR-Miner: automatically extracting implicit programming rules and detecting violations in large software code. ACM SIGSOFT Software Engineering Notes, 30 (5):306–315, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Livshits:2005:DFC

[LZ05b] Benjamin Livshits and Thomas Zimmermann. DynaMine: finding common error patterns by mining software revision histories. ACM SIGSOFT Software Engineering Notes, 30(5):296– 305, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mock:2002:IPS

[MACE02] Markus Mock, Darren C.

Atkinson, Craig Chambers, and Susan J. Eggers. Improving program slicing with dynamic points-to data. ACM SIGSOFT Software Engineering Notes, 27(6):71– 80, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Madrigal:2001:FOD

[Mad01] Victor M. Madrigal III. Fundamentals of OOP and data structures in Java. ACM SIGSOFT Software Engineering Notes, 26(1): 99, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mairson:2000:PLF

[Mai00]  Harry Mairson.  Pro-

gramming language foundations of computation theory. ACM SIGSOFT Software Engineering Notes, 25 (1):63, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mamone:2000:DT

[Mam00] Salvatore Mamone. Documentation testing. ACM SIGSOFT Software Engineering Notes, 25(2):26– 29, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mancoridis:2000:TDE

[Man00a] Spiros Mancoridis. Toward a design environment for recovering and maintaining the structure of software systems. ACM SIGSOFT Software Engineering Notes, 25 (1):63–64, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Manolescu:2000:BRP

[Man00b] Dragos Manolescu. Book review: Production Workflow-Concepts and Techniques by Frank Leymann and Dieter Roller (Prentice Hall PTR, 2000, ISBN 0-13-02175-0, 479 pp.). ACM SIGSOFT Software Engineering Notes, 25 (4):69, July 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Manolescu:2001:PWC

[Man01] Dragos Manolescu. Production workflow: concepts and techniques. ACM SIGSOFT Software Engineering Notes, 26(1):99–100, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mao:2009:SFP

[Mao09] Chengying Mao. Software faults prediction based on grey system theory. ACM SIGSOFT Software Engineering Notes, 34(2):1–6, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maruyama:2001:AME

[Mar01] Katsuhisa Maruyama. Automated method-extraction refactoring by using blockbased slicing. ACM SIG-SOFT Software Engineering Notes, 26(3):31–40, May 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marks:2004:UPH

[Mar04] Joe Marks. The usability problem for home appliances: engineers caused it, engineers can fix it! ACM SIGSOFT Software Engineering Notes, 29(6):95, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marcos:2005:SER

[Mar05a] Esperanza Marcos. Software engineering research versus software development. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marliss:2005:RSM

[Mar05b] G. S. Marliss. Review of Systems Modeling and Requirements Specification Using ECSAM: an Analysis Method for Embedded and Computer-Based Systems by Jonah Z. Lavi and Joseph Kudish, Dorset House Publishing Co. Inc., 2005, ISBN 0-932633-45-5. ACM SIGSOFT Software Engineering Notes, 30 (4):83–84, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marovac:2007:UBE

[Mar07] Nenad Marovac. UML based embedded documentation for semi-automatic software development. ACM SIGSOFT Software Engineering Notes, 32(5):1–3, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Massey:2005:LAL

[Mas05] Bart Massey. Longitudinal analysis of long-timescale open source repository data. ACM SIGSOFT Software Engineering Notes, 30(4): 1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McCloskey:2005:ANA

[MB05] Bill McCloskey and Eric Brewer. ASTEC: a new approach to refactoring C. ACM SIGSOFT Software Engineering Notes, 30(5):21– 30, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mittal:2009:SMA

[MB09] Harish Mittal and Pradeep Bhatia. Software maintainability assessment based on fuzzy logic technique. ACM SIGSOFT Software Engineering Notes, 34(3): 1–5, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mittal:2009:SMP

[MBM09] J. P. Mittal, Pradeep Bhatia, and Harish Mittal. Software maintenance productivity assessment using fuzzy logic. ACM SIGSOFT Software Engineering Notes, 34(5):1– 4, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mittermeir:2001:GDC

[MBPRR01] Roland T. Mittermeir, Andreas Bollin, Heinz Pozewaunig, and Dominik Rauner-Reithmayer. Goal-driven combination of software comprehension approaches for component based development. ACM SIGSOFT Software Engineering Notes, 26 (3):95–102, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McDonald:2000:S

[McD00] Jim McDonald. Specware. ACM SIGSOFT Software Engineering Notes, 25(1): 107, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Munoz:2004:MVA

[MDC04] César A. Muñoz, Gilles Dowek, and Víctor Carreño. Modeling and verification of an air traffic concept of operations. ACM SIGSOFT Software Engineering Notes, 29 (4):175–182, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Muccini:2005:TSA

[MDR05] Henry Muccini, Marcio S.

Dias, and Debra J. Richardson. Towards software architecture-based regression testing. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mascolo:2001:FIW

[MEF01] Cecilia Mascolo, Wolfgang Emmerich, and Anthony Finkelsteiin. First ICSE 2001 Workshop on XML Technologies and Software Engineering. ACM SIGSOFT Software Engineering Notes, 26 (6):59–61, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mei:2000:CAR

[Mei00] Hong Mei. A complementary approach to requirements engineering—software architecture orientation. ACM SIGSOFT Software Engineering Notes, 25(2):40– 45, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meinke:2004:ABB

[Mei04] Karl Meinke. Automated black-box testing of functional correctness using function approximation. ACM SIGSOFT Software Engineering Notes, 29(4):143– 153, July 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Memmert:2005:CPA

[Mem05]  Juri Memmert.  Con-

cern patterns and analysis. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meseguer:2001:RE

[Mes01] José Meseguer. Report on ETAPS 2000. ACM SIG-SOFT Software Engineering Notes, 26(1):39, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mondal:2000:CMW

[MG00] Sakib Abdul Mondal and Kingshuk Das Gupta. Choosing a middleware for webintegration of a legacy application. ACM SIGSOFT Software Engineering Notes, 25 (3):50–53, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mertens:2005:CAD

[MHB05] Koenraad Mertens, Tom Holvoet, and Yolande Berbers. A case for adaptation of the distributed environment layout in multiagent applications. ACM SIGSOFT Software Engineering Notes, 30 (4):1–8, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

MacGregor:2005:CPS

[MHK05] Eve MacGregor, Yvonne Hsieh, and Philippe Kruchten. Cultural patterns in software process mishaps: incidents in global projects. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Milicev:2000:COG

[Mil00]  Dragan Milicev.  Cus-

tomizable output generation in modeling environments using pipelined domains. ACM SIGSOFT Software Engineering Notes, 25 (3):46–50, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Minsky:2000:TAI

[Min00] Naftaly H. Minsky. Towards architectural invariants of evolving systems. ACM SIGSOFT Software Engineering Notes, 25(1): 65, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Minsky:2005:LGM

[Min05]  Naftaly Minsky.  Law-

governed multi-agent systems: from anarchy to order. ACM SIGSOFT Software Engineering Notes, 30 (4):1, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mislove:2000:SMC

[Mis00] Michael W. Mislove. Semantic models for concurrency. ACM SIGSOFT Software Engineering Notes, 25 (1):65–66, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misic:2006:PEP

[Miš06]  Vojislav B. Mišić.  Per-

ceptions of extreme programming: an exploratory study. ACM SIGSOFT Software Engineering Notes, 31 (2):1–8, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2007:VMC

[Mis07] Sanjay Misra. Validating modified cognitive complexity measure. ACM SIGSOFT Software Engineering Notes, 32(3):1–5, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mannion:2001:RBP

[MK01] Mike Mannion and Hermann Kaindl. Requirementsbased product line engineering. ACM SIGSOFT Software Engineering Notes, 26 (5):322–323, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McGee:2004:EHV

[MK04] Pat McGee and Cem Kaner.

Experiments with high volume test automation. ACM SIGSOFT Software Engineering Notes, 29(5):1–3, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morris:2005:RDT

[MK05] Jennifer Morris and Philip Koopman. Representing design tradeoffs in safetycritical systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2006:MTV

[MK06] Sanjay Misra and Hurevren Kilic. Measurement theory and validation criteria for software complexity measures. ACM SIGSOFT Software Engineering Notes, 31 (6):1–3, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2007:MTV

[MK07] Sanjay Misra and Hurevren Kilic. Measurement theory and validation criteria for software complexity measures. ACM SIGSOFT Software Engineering Notes, 32 (2):1–3, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Metz:2005:PDC

[MLG05] Edu Metz, Raimondas Lencevicius, and Teofilo F. Gonzalez. Performance data collection using a hybrid approach. ACM SIGSOFT Software Engineering Notes, 30(5): 126–135, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mierle:2005:MSC

[MLRW05] Keir Mierle, Kevin Laven, Sam Roweis, and Greg Wilson. Mining student CVS repositories for performance indicators. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2007:ECC

[MM07] Sanjay Misra and A. K.

Misra. Evaluation and comparison of cognitive complexity measure. ACM SIGSOFT Software Engineering Notes, 32(2):1–5, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2008:DIC

[MM08a] Deepti Mishra and Alok Mishra. Design issues in client-server software maintenance. ACM SIGSOFT Software Engineering Notes, 33 (5):6:1–6:??, September 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2008:SOS

[MM08b] Deepti Mishra and Alok Mishra. Some observations on staff estimation metrics for object: oriented software engineering. ACM SIGSOFT Software Engineering Notes, 33(5):5:1– 5:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2009:SIS

[MM09a] Alok Mishra and Deepti Mishra. Some issues on scheduling estimation model for object-oriented software projects. ACM SIGSOFT Software Engineering Notes, 34(4):1–4, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2009:CAP

[MM09b] Arun Mishra and A. K. Misra. Component assessment and proactive model for support of dynamic integration in self adaptive system. ACM SIGSOFT Software Engineering Notes, 34 (4):1–9, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2009:DDF

[MM09c] Deepti Mishra and Alok Mishra. A discussion on design factors of clientserver software testability. ACM SIGSOFT Software Engineering Notes, 34(5):1– 5, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marin:2005:AAR

[MMvD05] Marius Marin, Leon Moonen, and Arie van Deursen. An approach to aspect refactoring based on crosscutting concern types. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Merideth:2005:RNA

[MN05] Michael G. Merideth and Priya Narasimhan. Retrofitting networked applications to add autonomic reconfiguration. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mezini:2004:VMF

[MO04] Mira Mezini and Klaus Ostermann. Variability management with feature-oriented programming and aspects. ACM SIGSOFT Software Engineering Notes, 29(6): 127–136, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mateescu:2006:AAA

[MO06] Radu Mateescu and Flavio Oquendo. π-AAL: an architecture analysis language for formally specifying and verifying structural and behavioural properties of software architectures. ACM SIGSOFT Software Engineering Notes, 31 (2):1–19, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morrisett:2003:AIC

[Mor03] Greg Morrisett. Analysis issues for cyclone. ACM SIG-SOFT Software Engineering Notes, 28(1):26, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Masri:2005:UDI

[MP05] Wes Masri and Andy Podgurski. Using dynamic information flow analysis to detect attacks against applications. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Medikonda:2009:FSS

[MP09] Ben Swarup Medikonda and Seetha Ramaiah Panchumarthy. A framework for software safety in safetycritical systems. ACM SIG-SOFT Software Engineering Notes, 34(2):1–9, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Menzies:2005:SSC

[MPCH05] Tim Menzies, Dan Port, Zhihao Chen, and Jairus Hihn. Simple software cost analysis: safe or unsafe? ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Memon:2000:ATO

[MPS00] Atif M. Memon, Martha E.

Pollack, and Mary Lou Soffa. Automated test oracles for GUIs. ACM SIGSOFT Software Engineering Notes, 25 (6):30–39, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Memmert:2006:OGC

[MR06] Juri Memmert and Hridesh Rajan. Osiris: generating concern models from source code. ACM SIGSOFT Software Engineering Notes, 31 (6):1–2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mohamed:2005:DSC

[MRE05] Abdallah Mohamed, Guenther Ruhe, and Armin Eberlein. Decision support for customization of the COTS selection process. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Milanova:2002:POS

[MRR02] Ana Milanova, Atanas Rountev, and Barbara G. Ryder. Parameterized object sensitivity for points-to and side-effect analyses for Java. ACM SIGSOFT Software Engineering Notes, 27 (4):1–11, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mead:2005:SQR

[MS05] Nancy R. Mead and Ted Stehney. Security quality requirements engineering (SQUARE) methodology. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Manevich:2004:PEP

[MSA^{+}04] Roman Manevich, Manu Sridharan, Stephen Adams, Manuvir Das, and Zhe Yang. PSE: explaining program failures via postmortem static analysis. ACM SIGSOFT Software Engineering Notes, 29(6):63–72, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morasca:2004:ACT

[MSC04] Sandro Morasca and Stefano Serra-Capizzano. On the analytical comparison of testing techniques. ACM SIGSOFT Software Engineering Notes, 29(4):154–164, July 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mair:2005:ADS

[MSJ05] Carolyn Mair, Martin Shepperd, and Magne Jørgensen. An analysis of data sets used to train and validate cost prediction systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Memon:2001:CCG

[MSP01] Atif M. Memon, Mary Lou Soffa, and Martha E. Pollack. Coverage criteria for GUI testing. ACM SIGSOFT Software Engineering Notes, 26(5):256– 267, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

M:2002:UCL

[MSR02] Bharath Kumar M., Y. N.

Srikant, and Lakshminarayanan R. On the use of connector libraries in distributed software architectures. ACM SIGSOFT Software Engineering Notes, 27(1):45– 52, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meseguer:2000:M

[MT00a] Jose Meseguer and Carolyn Talcott. Maude. ACM SIGSOFT Software Engineering Notes, 25(1):104, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mesenguer:2000:SIO

[MT00b] José Mesenguer and Carolyn Talcott. Semantic interoperation of open systems. ACM SIGSOFT Software Engineering Notes, 25 (1):64–65, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morisawa:2001:ASP

[MT01] Yoshitomi Morisawa and Koji Torii. An architectural style of product lines for distributed processing systems, and practical selection method. ACM SIGSOFT Software Engineering Notes, 26(5):11–20, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Munson:2000:RTS

[Mun00] Ethan V. Munson. Representations, tools, and services for the complete integration of software development documents. ACM SIGSOFT Software Engineering Notes, 25 (2):25, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mihajlovic:2001:TSP

[MV01] Zorica Mihajlovic and Dusan Velasevic. Tracking software projects with the integrated version control in SMIT. ACM SIGSOFT Software Engineering Notes, 26 (2):38–43, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mannaert:2005:UOS

[MV05] Herwig Mannaert and Kris Ven. The use of open source software platforms by independent software vendors: issues and opportunities. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moors:2002:EAT

[MVT^{+}02] Tim Moors, Malathi Veeraraghavan, Zhifeng Tao, Xuan Zheng, and Ramesh Badri. Experiences in automating the testing of SS7 signalling transfer points. ACM SIGSOFT Software Engineering Notes, 27(4): 154–158, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mens:2001:FFS

[MW01] Tom Mens and Michel Wermelinger. Formal Foundations of Software Evolution: workshop report. ACM SIGSOFT Software Engineering Notes, 26(4): 27–29, July 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Munawar:2005:BPB

[MW05] Mohammad A. Munawar and Paul A. S. Ward. Better performance or better manageability? ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Myers:2006:MDS

[Mye06] Andrew Myers. Making distributed systems secure with program analysis and transformation. ACM SIGSOFT Software Engineering Notes, 31(1):47, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maurer:2001:IWS

[MZ01] Frank Maurer and Carmen Zannier. 4th ICSE workshop on “Software Engineering over the Internet”. ACM SIGSOFT Software Engineering Notes, 26(6):29– 31, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mei:2001:SCM

[MZY01] Hong Mei, Lu Zhang, and Fuqing Yang. A software configuration management model for supporting component-based software development. ACM SIGSOFT Software Engineering Notes, 26(2):53– 58, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nadathur:2000:FRD

[Nad00] Gopalan Nadathur. A framework for realizing derivation systems. ACM SIGSOFT Software Engineering Notes, 25(1):66–67, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Niu:2002:CSM

[NAD02] Jianwei Niu, Joanne M.

Atlee, and Nancy A. Day. Composable semantics for model-based notations. ACM SIGSOFT Software Engineering Notes, 27(6):149– 158, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nami:2008:COO

[Nam08] Mohammad Reza Nami.

A comparison of objectoriented languages in software engineering. ACM SIGSOFT Software Engineering Notes, 33(4):6:1– 6:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:2002:CAC

[Nau02] Gleb Naumovich. A conservative algorithm for computing the flow of permissions in Java programs. ACM SIGSOFT Software Engineering Notes, 27(4): 33–43, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:2003:UOD

[Nau03] Gleb Naumovich. Using the observer design pattern for implementation of data flow analyses. ACM SIGSOFT Software Engineering Notes, 28(1):61–68, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nelson:2009:CWQ

[NB09] Stephen Nelson and Coşkun Bayrak. Categorizing web queries. ACM SIGSOFT Software Engineering Notes, 34(2):1–4, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:2000:CPA

[NC00] Gleb Naumovich and Lori A. Clarke. Classifying properties: an alternative to the safety-liveness classification. ACM SIGSOFT Software Engineering Notes, 25 (6):159–168, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:2004:SAR

[NC04] Gleb Naumovich and Paolina Centonze. Static analysis of role-based access control in J2EE applications. ACM SIGSOFT Software Engineering Notes, 29(5):1– 10, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nierstrasz:2005:SMA

[NDG05] Oscar Nierstrasz, Stéphane Ducasse, and Tudor Gĭrba. The story of Moose: an agile reengineering environment. ACM SIGSOFT Software Engineering Notes, 30 (5):1–10, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nimmer:2002:AGP

[NE02a] Jeremy W. Nimmer and Michael D. Ernst. Automatic generation of program specifications. ACM SIGSOFT Software Engineering Notes, 27(4):229–239, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nimmer:2002:IIS

[NE02b] Jeremy W. Nimmer and Michael D. Ernst. Invariant inference for static checking: an empirical evaluation. ACM SIGSOFT Software Engineering Notes, 27 (6):11–20, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nkwocha:2005:FPM

[NE05] Fidel Nkwocha and Sebastian Elbaum. Fault patterns in Matlab. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nedunuri:2000:FAP

[Ned00] Srinivas Nedunuri. The functional approach to programming. ACM SIGSOFT Software Engineering Notes, 25 (3):63–65, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2000:RPCa

[Neu00a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 25(1):22– 26, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2000:RPCb

[Neu00b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 25(2): 17–21, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2000:RPCc

[Neu00c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 25(3): 15–23, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2000:RPCd

[Neu00d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIG-SOFT Software Engineering Notes, 25(4):7–11, July 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2001:RPCa

[Neu01a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 26(1):14– 38, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2001:RPCb

[Neu01b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 26(2): 5–11, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2001:RPCc

[Neu01c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIG-SOFT Software Engineering Notes, 26(4):3–12, July 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2001:RPCd

[Neu01d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 26(6):6– 15, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2002:RPCa

[Neu02a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 27(1):7– 17, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2002:RPCb

[Neu02b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 27(2): 5–19, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2002:RPCc

[Neu02c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIG-SOFT Software Engineering Notes, 27(3):5–19, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2002:RPCd

[Neu02d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 27(5):6– 16, September 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2003:RPC

[Neu03] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIG-SOFT Software Engineering Notes, 28(3):5–9, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2004:RPCa

[Neu04a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 29(2): 8–16, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2004:RPCb

[Neu04b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIG-SOFT Software Engineering Notes, 29(3):7–14, May 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2004:RPCc

[Neu04c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 29(5):13– 18, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2005:RPa

[Neu05a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 30(2):17–25, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2005:RPb

[Neu05b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 30(3):22–37, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2005:RPd

[Neu05c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 30(4):19–35, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2005:RPe

[Neu05d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 30(6):14–25, November 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2005:RPCa

[Neu05e] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIG-SOFT Software Engineering Notes, 30(1):9, January 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2005:RPCb

[Neu05f] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 30(1):9– 18, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2006:RPa

[Neu06a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 31(1):14–19, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2006:RPb

[Neu06b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 31(2):6–16, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2006:RPc

[Neu06c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 31(3):17–24, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2006:RPd

[Neu06d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 31(4):25–37, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2006:RPe

[Neu06e] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 31(5):16–21, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2006:RPf

[Neu06f] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 31(6):21–37, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:SDR

[Neu07a] Peter Neumann. Session details: Risks to the public. ACM SIGSOFT Software Engineering Notes, 32 (1):??, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:RPa

[Neu07b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 32(1):15–20, January 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:RPb

[Neu07c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 32(2):18–23, March 2007.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:RPc

[Neu07d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 32(3):20–24, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:RPd

[Neu07e] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 32(4):20–29, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:RPe

[Neu07f] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 32(5):17–25, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2007:RPf

[Neu07g] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 32(6):20–27, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2008:RPa

[Neu08a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 33(1):4:1–4:??, January 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2008:RPb

[Neu08b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 33(2):18–28, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2008:RPc

[Neu08c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 33(3):17–28, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2008:RPd

[Neu08d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 33(4):17–26, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2008:RPe

[Neu08e] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 33(5):2:1–2:??, September 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2008:RPf

[Neu08f] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 33(6):15–24, November 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2009:RPa

[Neu09a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 34(1):17–26, January 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2009:RPb

[Neu09b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 34(2):15–24, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2009:RPc

[Neu09c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 34(3):16–29, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2009:RPd

[Neu09d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 34(4):17–24, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2009:RPe

[Neu09e] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 34(5):18–24, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2009:RPf

[Neu09f] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 34(6):21–24, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neamtiu:2005:USC

[NFH05] Iulian Neamtiu, Jeffrey S.

Foster, and Michael Hicks. Understanding source code evolution using abstract syntax tree matching. ACM SIGSOFT Software Engineering Notes, 30(4):1– 5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nami:2009:CEZ

[NH09] Mohammad Reza Nami and Fatemeh Hassani. A comparative evaluation of the z, CSP, RSL, and VDM languages. ACM SIGSOFT Software Engineering Notes, 34(3):1–4, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Niemela:2001:PLS

[NI01] Eila Niemelä and Tuomas Ihme. Product line software engineering of embedded systems. ACM SIGSOFT Software Engineering Notes, 26 (3):118–125, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neubauer:2005:BPB

[NKB05] Thomas Neubauer, Markus Klemen, and Stefan Biffl. Business process-based valuation of IT-security. ACM SIGSOFT Software Engineering Notes, 30(4):1– 5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Najumudheen:2009:DGB

[NMS09] ESF Najumudheen, Rajib Mall, and Debasis Samanta. A dependence graph-based representation for test coverage analysis of objectoriented programs. ACM SIGSOFT Software Engineering Notes, 34(2):1–8, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2000:SNS

[Not00] David Notkin. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 25 (3):5–10, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2003:LPA

[Not03] David Notkin. Longitudinal program analysis. ACM SIG-SOFT Software Engineering Notes, 28(1):1, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2007:ATFa

[Not07a]  David Notkin.  ACM

TOSEM: FAQs and figures. ACM SIGSOFT Software Engineering Notes, 32 (3):6, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2007:ATFb

[Not07b]  David Notkin.  ACM

TOSEM: FAQs and figures. ACM SIGSOFT Software Engineering Notes, 32 (4):8, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2007:ATFc

[Not07c]  David Notkin.  ACM

TOSEM: FAQs and figures. ACM SIGSOFT Software Engineering Notes, 32(5):5– 6, September 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2007:FFa

[Not07d] David Notkin. FAQs and figures. ACM SIGSOFT Software Engineering Notes, 32 (1):3, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2007:FFb

[Not07e] David Notkin. FAQs and figures. ACM SIGSOFT Software Engineering Notes, 32 (2):6–7, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2008:ATFa

[Not08a] David Notkin. ACM TOSEM FAQs and figures. ACM SIG-SOFT Software Engineering Notes, 33(2):5–6, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2008:ATFb

[Not08b]  David Notkin.  ACM

TOSEM: FAQs and figures. ACM SIGSOFT Software Engineering Notes, 33 (4):6, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2009:FF

[Not09] David Notkin. FAQs and figures. ACM SIGSOFT Software Engineering Notes, 34(4):6, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nord:2001:SAC

[NPSS01] Robert L. Nord, Daniel J.

Paulish, Robert W. Schwanke, and Dilip Soni. Software architecture in a changing world: developing design strategies that anticipate change. ACM SIG-SOFT Software Engineering Notes, 26(5):309–310, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nanda:2000:SCP

[NR00a] Mangala Gowri Nanda and S. Ramesh. Slicing concurrent programs. ACM SIGSOFT Software Engineering Notes, 25(5):180– 190, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Necula:2000:LBA

[NR00b] George C. Necula and Shree Rahul. A logic-based approach to software system safety and security. ACM SIGSOFT Software Engineering Notes, 25(1):67– 68, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nami:2009:INF

[NRSS09] Mohammad Reza Nami, Hamid Razavi, Milad Saeedi, and Navid Shahidi. Investigating a new formal model for a library system using B method. ACM SIGSOFT Software Engineering Notes, 34(4):1–4, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nayak:2009:MBT

[NS09] Ashalatha Nayak and Debasis Samanta. Modelbased test cases synthesis using UML interaction diagrams. ACM SIGSOFT Software Engineering Notes, 34 (2):1–10, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nami:2009:CIC

[NSAK09] Mohammad Reza Nami, Bahareh Sheikh-Abbasi, and Marjaneh Khoshandam. A comparative introduction to CSP and ACT-ONE formal languages. ACM SIGSOFT Software Engineering Notes, 34(6):1–5, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nami:2007:ADE

[NTS07] Mohammad Reza Nami, Mehdi Sagheb Tehrani, and Mohsen Sharifi. Applying domain engineering using RAISE into a particular banking domain. ACM SIG-SOFT Software Engineering Notes, 32(2):1–6, March 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Noppen:2005:DIQ

[NvdBA05] Joost Noppen, Pim van den Broek, and Mehmet Aksit. Dealing with imprecise quality factors in software design. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nachmanson:2004:OST

[NVS^{+}04] Lev Nachmanson, Margus Veanes, Wolfram Schulte, Nikolai Tillmann, and Wolfgang Grieskamp. Optimal strategies for testing nondeterministic systems. ACM SIGSOFT Software Engineering Notes, 29(4): 55–64, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nagappan:2005:EES

[NWVO05] Nachiappan Nagappan, Laurie Williams, Mladen Vouk, and Jason Osborne. Early estimation of software quality using in-process testing metrics: a controlled case study. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Oliveira:2005:XED

[OACL05] Toacy C. Oliveira, Paulo Alencar, Don Cowan, and Carlos Lucena. xTAO: enabling a declarative approach to the specification of multiagent systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2000:F

[OC00a] Leon Osterweil and Lori A. Clarke. FLAVERS. ACM SIGSOFT Software Engineering Notes, 25(1):101– 102, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2000:LJ

[OC00b] Leon Osterweil and Lori A. Clarke. Little-JIL. ACM SIGSOFT Software Engineering Notes, 25(1):104, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2000:IPP

[OCE^{+}00] Leon J. Osterweil, Lori A. Clarke, Michael Evangelist, Jeffrey Kramer, Dieter Rombach, and Alexander L. Wolf. The impact project (panel session): determining the impact of software engineering research upon practice. ACM SIGSOFT Software Engineering Notes, 25(6):108– 109, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:2000:STA

[Off00] Jeff Offutt. Software testing and analysis of objectoriented software. ACM SIG-SOFT Software Engineering Notes, 25(1):68, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ohba:2005:TMC

[OG05a] Masaru Ohba and Katsuhiko Gondow. Toward mining “concept keywords” from identifiers in large software projects. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Olimpiew:2005:MBT

[OG05b] Erika Mir Olimpiew and Hassan Gomaa. Model-based testing for applications derived from software product lines. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Orso:2005:SCR

[OK05] Alessandro Orso and Bryan Kennedy. Selective capture and replay of program executions. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Orso:2002:GSC

[OLHL02] Alessandro Orso, Donglin Liang, Mary Jean Harrold, and Richard Lipton. Gamma system: continuous evolution of software after deployment. ACM SIGSOFT Software Engineering Notes, 27 (4):65–69, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Olsson:2000:CPL

[Ols00] Ronald A. Olsson. Concurrent programming language support for invocation handling: design and implementation. ACM SIGSOFT Software Engineering Notes, 25 (1):68–69, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Omitol:2001:AFP

[Omi01] Tope Omitol. ACM Fellow profile: Roger Needham. ACM SIGSOFT Software Engineering Notes, 26 (1):7–10, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:2004:EMS

[OMK04] Jeff Offutt, Yu-Seung Ma, and Yong-Rae Kwon. An experimental mutation system for Java. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ohira:2005:ACP

[OOOiM05] Masao Ohira, Naoki Ohsugi, Tetsuya Ohoka, and Ken ichi Matsumoto. Accelerating cross-project knowledge collaboration using collaborative filtering and social networks. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Oquendo:2004:AAD

[Oqu04a] Flavio Oquendo. π-ADL: an architecture description language based on the higherorder typed π-calculus for specifying dynamic and mobile software architectures. ACM SIGSOFT Software Engineering Notes, 29(3): 1–14, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Oquendo:2004:AAR

[Oqu04b] Flavio Oquendo. π-ARL: an architecture refinement language for formally modelling the stepwise refinement of software architectures. ACM SIGSOFT Software Engineering Notes, 29 (5):1–20, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Oquendo:2004:FRS

[Oqu04c] Flavio Oquendo. Formally refining software architectures with π-ARL: a case study. ACM SIGSOFT Software Engineering Notes, 29 (5):1–26, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Oquendo:2006:MMD

[Oqu06a] Flavio Oquendo. π-Method: a model-driven formal method for architecture-centric software engineering. ACM SIG-SOFT Software Engineering Notes, 31(3):1–13, May 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Oquendo:2006:FMS

[Oqu06b] Flavio Oquendo. Formally modelling software architectures with the UML 2.0 profile for π-ADL. ACM SIGSOFT Software Engineering Notes, 31(1):1–13, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Orso:2004:SRT

[OSH04] Alessandro Orso, Nanjuan Shi, and Mary Jean Harrold. Scaling regression testing to large software systems. ACM SIGSOFT Software Engineering Notes, 29 (6):241–251, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ovans:2004:PLC

[Ova04] Russell Ovans. The programmer life-cycle. ACM SIGSOFT Software Engineering Notes, 29(3):25– 26, May 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ovans:2005:AWR

[Ova05] Russell Ovans. All we really need to know about software engineering is in the film office space. ACM SIGSOFT Software Engineering Notes, 30(4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ostrand:2002:DFL

[OW02] Thomas J. Ostrand and Elaine J. Weyuker. The distribution of faults in a large industrial software system. ACM SIGSOFT Software Engineering Notes, 27 (4):55–64, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ostrand:2004:WB

[OWB04] Thomas J. Ostrand, Elaine J. Weyuker, and Robert M. Bell. Where the bugs are. ACM SIGSOFT Software Engineering Notes, 29(4): 86–96, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:2004:GTC

[OX04] Jeff Offutt and Wuzhi Xu. Generating test cases for web services using data perturbation. ACM SIGSOFT Software Engineering Notes, 29(5):1–10, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:2004:SME

[OYH04] Jeff Offutt, Yuan Yang, and Jane Huffman Hayes. SEEWeb: making experimental artifacts available. ACM SIGSOFT Software Engineering Notes, 29(5):1– 3, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pollock:2005:IWS

[PA05] Lori Pollock and James Andrews. ICSE 2005 workshop summary Third International Workshop on Dynamic Analysis (WODA 2005). ACM SIGSOFT Software Engineering Notes, 30 (4):1–2, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pazos-Arias:2005:ERS

[PAGDLN05a] José J. Pazos-Arias, Jorge García-Duque, and Martín López-Nores. Eliciting requirements and scenarios using the SCTL-MUS methodology: the shuttle system case study. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pazos-Arias:2005:LCC

[PAGDLN05b] José J. Pazos-Arias, Jorge García-Duque, and Martín López-Nores. Locating crosscutting concerns in the formal specification of distributed reactive systems. ACM SIGSOFT Software Engineering Notes, 30(4): 1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

PanYunhe:2000:AOA

[Pan00] Zhou Yonglin PanYunhe. Agent-oriented analysis and modeling. ACM SIGSOFT Software Engineering Notes, 25(3):36–40, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pandey:2009:EWR

[Pan09a] R. K. Pandey. Exploiting web resources for teaching/ learning best software design tips. ACM SIGSOFT Software Engineering Notes, 34 (6):1–7, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pandey:2009:MSD

[Pan09b]  R. K. Pandey.  Manag-

ing software design complexity: facade vs role-based design. ACM SIGSOFT Software Engineering Notes, 34 (1):1–4, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pandey:2009:RSE

[Pan09c] R. K. Pandey. Relativity in software engineering measurements. ACM SIGSOFT Software Engineering Notes, 34(2):1–3, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paradkar:2005:CSF

[Par05] Amit Paradkar. Case studies on fault detection effectiveness of model based test generation techniques. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Phongpaibul:2005:IQT

[PB05] Monvorath Phongpaibul and Barry Boehm. Improving quality through software process improvement in Thailand: initial analysis. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Painter:2005:MSP

[PC05a] Robert R. Painter and David Coppit. A model for software plans. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Palankar:2005:MTH

[PC05b] Mayur Palankar and Jonathan E.

Cook. Merging traces of hardware-assisted data breakpoints. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:2005:GSE

[PCP05] David Parnas and Lillian Chik-Parnas. Goals for software engineering student education. ACM SIGSOFT Software Engineering Notes, 30(4):6–8, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Privat:2006:LTS

[PD06] Jean Privat and Roland Ducournau. Link-time static analysis for efficient separate compilation of objectoriented languages. ACM SIGSOFT Software Engineering Notes, 31(1):20– 27, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perkins:2004:EIA

[PE04] Jeff H. Perkins and Michael D. Ernst. Efficient incremental algorithms for dynamic detection of likely invariants. ACM SIGSOFT Software Engineering Notes, 29 (6):23–32, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Peciola:2002:ELI

[Pec02]  Emilia Peciola.  Ericsson

lab Italy: is ISSTA research relevant to industrial users? ACM SIGSOFT Software Engineering Notes, 27(4):207, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pentinmaki:2001:SCI

[Pen01]  Isaac Pentinmaki.  Stan-

dard C++ IOStreams and locales: advanced programmer’s guide and reference. ACM SIGSOFT Software Engineering Notes, 26(2): 76–77, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pentinmaki:2002:MED

[Pen02]  Isaac Pentinmaki.  Mas-

tering ENVY/Developer. ACM SIGSOFT Software Engineering Notes, 27(2): 73, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pentinmaki:2003:IMP

[Pen03a] Isaac Pentinmaki. IT measurement: practical advice from the experts. ACM SIG-SOFT Software Engineering Notes, 28(3):25, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pentinmaki:2003:OMU

[Pen03b] Isaac Pentinmaki. Object modeling and user interface design designing interactive systems. ACM SIGSOFT Software Engineering Notes, 28(3):24, May 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pentinmaki:2004:RLS

[Pen04] Isaac Pentinmaki. Review of Lean software development: an agile tookit by Mary and Tom Poppendieck. Addison Wesley 2003. ACM SIG-SOFT Software Engineering Notes, 29(2):33, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perkins:2006:AGR

[Per06]  Jeff H. Perkins.  Auto-

matically generating refactorings to support API evolution. ACM SIGSOFT Software Engineering Notes, 31 (1):111–114, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pethia:2000:BP

[Pet00] Richard D. Pethia. Bugs in the programs. ACM SIG-SOFT Software Engineering Notes, 25(6):79, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pettersson:2005:MPS

[Pet05] Niklas Pettersson. Measuring precision for static and dynamic design pattern recognition as a function of coverage. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Petro:2006:RMJ

[Pet06] Andrew Petro. Review of More Java Gems by Dwight Deugo, Cambridge University Press, 2000, ISBN: 0-521-77477-2. ACM SIGSOFT Software Engineering Notes, 31(3):42, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Persson:2005:OTH

[PGL^{+}05] Anna Persson, Henrik Gustavsson, Brian Lings, Björn Lundell, Anders Mattsson, and Ulf Ärlig. OSS tools in a heterogeneous environment for embedded systems modelling: an analysis of adoptions of XMI. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prechelt:2005:MFE

[PH05] Lutz Prechelt and Daniel J.

Hutzel. Market forces and end-user programming for mission-critical systems. ACM SIGSOFT Software Engineering Notes, 30(4):

1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Podgorelec:2007:ESC

[PH07] Vili Podgorelec and Marjan Heričko. Estimating software complexity from UML models. ACM SIGSOFT Software Engineering Notes, 32 (2):1–5, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paech:2008:RWC

[PHP08] Barbara Paech, Patrick Heymans, and Anne Persson. Report working conference on requirements engineering: foundation for software quality (REFSQ’08). ACM SIGSOFT Software Engineering Notes, 33(5):3:1– 3:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pickard:2005:OIN

[Pic05] Michael M. Pickard. Old issues, new eyes. ACM SIG-SOFT Software Engineering Notes, 30(4):1–3, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pierce:2000:EO

[Pie00] Benjamin C. Pierce. The essence of objects. ACM SIGSOFT Software Engineering Notes, 25(1):69– 71, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pincus:2000:ANF

[Pin00] Jon Pincus. Analysis is necessary, but far from sufficient (abstract only): Experiences building and deploying successful tools for developers and testers. ACM SIGSOFT Software Engineering Notes, 25(5):1, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pettersson:2005:AUC

[PIÖ05] Fredrik Pettersson, Martin Ivarsson, and Peter Öhman. Automotive use case standard for embedded systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pettersson:2005:IEB

[PJ05] Ulf Pettersson and Stan Jarzabek. Industrial experience with building a web portal product line using a lightweight, reactive approach. ACM SIG-SOFT Software Engineering Notes, 30(5):326–335, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfleeger:2001:PSR

[PK01] Shari Lawrence Pfleeger and Barbara A. Kitchenham. Principles of survey research: part 1: turning lemons into lemonade. ACM SIGSOFT Software Engineering Notes, 26(6):16–18, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perez-Martinez:2003:HEU

[PM03] Jorge Enrique Pérez-Martínez. Heavyweight extensions to the UML metamodel to describe the C3 architectural style. ACM SIGSOFT Software Engineering Notes, 28 (3):5, May 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Price:2001:PRO

[PND01] Margaretha W. Price, Donald M. Needham, and Steven A. Demurjian, Sr. Producing reusable objectoriented components: a domain-and-organization-specific perspective. ACM SIGSOFT Software Engineering Notes, 26(3):41–50, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pontelli:2000:PDE

[Pon00]  Enrico Pontelli.  Paral-

lel and distributed execution of constraint programs. ACM SIGSOFT Software Engineering Notes, 25(1):71– 72, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Peters:2000:RBM

[PP00] Dennis K. Peters and David L.

Parnas. Requirements-based monitors for real-time systems. ACM SIGSOFT Software Engineering Notes, 25 (5):77–85, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Palanque:2001:DSV

[PP01] Philippe Palanque and Fabio Paternò. Design, specification, and verification of interactive systems. ACM SIGSOFT Software Engineering Notes, 26(1):74– 75, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pedreira:2007:SRS

[PPLB07] Oscar Pedreira, Mario Piattini, Miguel R. Luaces, and Nieves R. Brisaboa. A systematic review of software process tailoring. ACM SIG-SOFT Software Engineering Notes, 32(3):1–6, May 2007.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pankratius:2009:AMS

[PPV09] Victor Pankratius, Adam Porter, and Larry Votta. Advances in multicore software engineering. ACM SIGSOFT Software Engineering Notes, 34(4):30–31, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfahl:2004:RPI

[PRRW04] Dietmar Pfahl, David Raffo, Ioana Rus, and Paul Wernick. Report on ProSim’04: the 5th International Workshop on Software Process Simulation and Modeling. ACM SIGSOFT Software Engineering Notes, 29(5):1– 6, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ploski:2007:RIS

[PRSH07] Jan Ploski, Matthias Rohr, Peter Schwenkenberg, and Wilhelm Hasselbring. Research issues in software fault categorization. ACM SIGSOFT Software Engineering Notes, 32(6):6:1– 6:??, November 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pulvermuller:2000:TGC

[PS00] Elke Pulvermüller and Andreas Speck. Towards generative components. ACM SIGSOFT Software Engineering Notes, 25(2):22– 24, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prakash:2009:IEL

[PSK09] Lakshmi Sunil Prakash, Dinesh Kumar Saini, and N. S. Kutti. Integrating EduLearn learning content management system (LCMS) with cooperating learning object repositories (LORs) in a peer to peer (p2p) architectural framework. ACM SIGSOFT Software Engineering Notes, 34(3):1–7, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poladian:2005:ACR

[PSPS05] Vahe Poladian, João Sousa, Frank Padberg, and Mary Shaw. Anticipatory configuration of resource-aware applications. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pretschner:2004:SEA

[PSS04] Alexander Pretschner, Christian Salzmann, and Thomas Stauner. Software engineering for automotive systems at ICSE 2004 workshop summary. ACM SIGSOFT Software Engineering Notes, 29(5):1–2, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pretschner:2005:IIW

[PSS05] Alexander Pretschner, Christian Salzmann, and Thomas Stauner. 2nd Intl. ICSE Workshop on Software Engineering for Automotive Systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–2, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perez-Serrano:2006:SDC

[PSSA06] Pedro L. Pérez-Serrano and Marisol Sánchez-Alonso. Specification and design of component-based coordination systems by integrating coordination patterns. ACM SIGSOFT Software Engineering Notes, 31(2):10:1– 10:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pretschner:2007:IIW

[PSSS07] Alexander Pretschner, Christian Salzmann, Bernhard Schätz, and Thomas Stauner. 4th Intl. ICSE Workshop on Software Engineering for Automotive Systems. ACM SIGSOFT Software Engineering Notes, 32(4):40, July 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Penix:2001:FIW

[PTV01] John Penix, Nigel Tracey, and Willem Visser. The First International Workshop on Automated Program Analysis, Testing and Verification. ACM SIGSOFT Software Engineering Notes, 26 (1):40, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pickett:2006:SSF

[PV06] Christopher J. F. Pickett and Clark Verbrugge. SableSpMT: a software framework for analysing speculative multithreading in Java. ACM SIGSOFT Software Engineering Notes, 31(1):59– 66, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pan:2004:RBS

[PWZ^{+}04] Ying Pan, Lei Wang, Lu Zhang, Bing Xie, and Fuqing Yang. Relevancy based semantic interoperation of reuse repositories. ACM SIGSOFT Software Engineering Notes, 29 (6):211–220, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Qun:2005:FDS

[QXCMW05] Yang Qun, Yang Xian-Chun, and Xu Man-Wu. A framework for dynamic software architecture-based selfhealing. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:2000:WNK

[Rac00]  L. B. S. Raccoon.  A

whole new kind of engineering. ACM SIGSOFT Software Engineering Notes, 25 (1):109–113, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:2001:CEG

[Rac01a] L. B. S. Raccoon. A conversation with erwin goodhack. ACM SIGSOFT Software Engineering Notes, 26 (1):92, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:2001:DD

[Rac01b] L. B. S. Raccoon. Definitions and demographics. ACM SIGSOFT Software Engineering Notes, 26(1):82– 91, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:2002:PDG

[Rac02] L. B. S. Raccoon. Practitioners do good work. ACM SIGSOFT Software Engineering Notes, 27(2): 46–52, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:2006:LPS

[Rac06] L. B. S. Raccoon. A leadership primer for software engineers. ACM SIGSOFT Software Engineering Notes, 31 (4):10–15, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rajlich:2000:MTC

[Raj00] Václav Rajlich. A model and a tool for change propagation in software. ACM SIGSOFT Software Engineering Notes, 25(1):72, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rajan:2006:ARB

[Raj06] Ajitha Rajan. Automated requirements-based test case generation. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramakrishnan:2000:TLP

[Ram00a] C. R. Ramakrishnan. Tabled logic programming for verification and program analysis. ACM SIGSOFT Software Engineering Notes, 25 (1):73, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramaswamy:2000:LDS

[Ram00b] Ramkumar Ramaswamy. Latency in distributed, sequential application designs. ACM SIGSOFT Software Engineering Notes, 25(2): 51–55, March 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramachandran:2005:SRG

[Ram05] Muthu Ramachandran. Software reuse guidelines. ACM SIGSOFT Software Engineering Notes, 30(3):1– 8, May 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ran:2001:FCP

[Ran01] Alexander Ran. Fundamental concepts for practical software architecture. ACM SIGSOFT Software Engineering Notes, 26(5):328– 329, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rashid:2000:STC

[Ras00] Richard F. Rashid. The shape of things to come? ACM SIGSOFT Software Engineering Notes, 25(6): 169, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raja:2005:IQL

[RB05a] Uzma Raja and Evelyn Barry. Investigating quality in large-scale open source software. ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rossi:2005:WPO

[RB05b] Cristina Rossi and Andrea Bonaccorsi. Why profitoriented companies enter the OS field?: intrinsic vs. extrinsic incentives. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ruthruff:2005:SCS

[RB05c] Joseph R. Ruthruff and Margaret Burnett. Six challenges in supporting end-user debugging. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramsey:2001:AAF

[RC01] Norman Ramsey and Előd Csirmaz. An algebraic approach to file synchronization. ACM SIGSOFT Software Engineering Notes, 26 (5):175–185, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosa:2007:FFM

[RC07] Nelson Souto Rosa and Paulo Roberto Freire Cunha. A formal framework for middleware behavioural specification. ACM SIGSOFT Software Engineering Notes, 32 (2):1–7, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rombach:2008:IRP

[RCJ^{+}08] Dieter Rombach, Marcus Ciolkowski, Ross Jeffery, Oliver Laitenberger, Frank McGarry, and Forrest Shull. Impact of research on practice in the field of inspections, reviews and walkthroughs: learning from successful industrial uses. ACM SIGSOFT Software Engineering Notes, 33(6):26– 35, November 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rawashdeh:2005:UTD

[RCL05] Osamah A. Rawashdeh, Garrett D. Chandler, and James E. Lumpp. A UAU test and development environment based on dynamic system reconfiguration. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rits:2005:XBB

[RDS05] Maarten Rits, Benjamin De Boe, and Andreas Schaad. XacT: a bridge between resource management and access control in multi-layered applications. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rech:2007:DTS

[Rec07a]  Jörg Rech.  Discovering

trends in software engineering with Google Trend. ACM SIGSOFT Software Engineering Notes, 32(2): 1–2, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rech:2007:PAS

[Rec07b] Jörg Rech. Podcasts about software engineering. ACM SIGSOFT Software Engineering Notes, 32(2):1–2, March 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Redmiles:2000:IDI

[Red00] David F. Redmiles. Improving the design of interactive software. ACM SIGSOFT Software Engineering Notes, 25(1):74–75, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reifer:2005:ESE

[Rei05a] Donald J. Reifer. Educating software engineers: an industry viewpoint. ACM SIG-SOFT Software Engineering Notes, 30(3):8–9, May 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reiss:2005:DDV

[Rei05b] Steven P. Reiss. Dynamic detection and visualization of software phases. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reynolds:2000:DDI

[Rey00] John C. Reynolds. The design, definition and implementation of programming languages. ACM SIGSOFT Software Engineering Notes, 25(1):75, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ratzinger:2005:IET

[RFG05] Jacek Ratzinger, Michael Fischer, and Harald Gall. Improving evolvability through refactoring. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rotondi:2006:CHA

[RG06] Guido Rotondi and Gianpiero Guerrera. A consistent history authentication protocol. ACM SIGSOFT Software Engineering Notes, 31 (3):1–7, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robles:2005:DIM

[RGB05] Gregorio Robles and Jesus M.

Gonzalez-Barahona. Developer identification methods for integrated data from various sources. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rekoske:2006:PEV

[RGS06] Jeffrey Rekoske, Hamid Ghasemiyeh, and Ahmed Salem. A practical extension to the VORD model. ACM SIGSOFT Software Engineering Notes, 31(5):1– 5, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roman:2000:EV

[RHK00] Gruia-Catalin Roman, Delbert Hart, and Eileen Kraemer. Exploratory visualization. ACM SIGSOFT Software Engineering Notes, 25 (1):76–77, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rhodes:2002:SEE

[Rho02] Donna H. Rhodes. Systems engineering: an essential discipline for the 21st century. ACM SIGSOFT Software Engineering Notes, 27(5):40–45, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:2000:BDA

[Ric00a] Debra Richardson. DAS-BOOT: design-, architectureand specification-based approaches to object-oriented testing. ACM SIGSOFT Software Engineering Notes, 25(1):100, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:2000:SAT

[Ric00b]  Debra Richardson.  Sid-

dhartha — automated test driver-oracle synthesis. ACM SIGSOFT Software Engineering Notes, 25(1):106– 107, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:2000:SSB

[Ric00c] Debra Richardson. SoBelt:

structural and behavioral execution instrumentation tool. ACM SIGSOFT Software Engineering Notes, 25(1): 107, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:2000:TST

[Ric00d] Debra Richardson. TestTalk:

software test description language. ACM SIGSOFT Software Engineering Notes, 25 (1):108, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richter:2009:CBN

[Ric09] Karsten Richter. Causalbased networks supporting process improvement. ACM SIGSOFT Software Engineering Notes, 34(1): 1–9, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2005:ESC

[Rie05] Richard Riehle. Engineering on the surprise continuum: as applied to software practice. ACM SIGSOFT Software Engineering Notes, 30(5):1–6, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2006:LCS

[Rie06] Richard Riehle. Linguistic continuity in software engineering. ACM SIGSOFT Software Engineering Notes, 31(1):1–5, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2007:DSC

[Rie07a] Richard Riehle. Designing software components to tolerances. ACM SIGSOFT Software Engineering Notes, 32 (4):7:1–7:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2007:FDS

[Rie07b]  Richard Riehle.  Failure-

driven software safety. ACM SIGSOFT Software Engineering Notes, 32(5):1–4, September 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2007:IMR

[Rie07c] Richard Riehle. Institutional memory and risk management. ACM SIGSOFT Software Engineering Notes, 32 (6):5:1–5:??, November 2007.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2008:JDU

[Rie08a]  Dirk Riehle.  JUnit 3.8

documented using collaborations. ACM SIGSOFT Software Engineering Notes, 33(2):5:1–5:??, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2008:GI

[Rie08b] Richard Riehle. The GSwERC initiative. ACM SIGSOFT Software Engineering Notes, 33(6):25, November 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riehle:2008:PUC

[Rie08c] Richard Riehle. Playscript for use cases. ACM SIG-SOFT Software Engineering Notes, 33(5):7:1–7:??, September 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rainer:2005:SEP

[RJH05] Austen Rainer, Dorota Jagielska, and Tracy Hall. Software engineering practice versus evidence-based software engineering research. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reed:2000:OVC

[RK00] Daniel Reed and Simon Kaplan. ORBIT/VIRTUE — collaboration and visualization support for complex systems evolution. ACM SIGSOFT Software Engineering Notes, 25(1):105– 106, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rountev:2004:SDA

[RKG04] Atanas Rountev, Scott Kagan, and Michael Gibas. Static and dynamic analysis of call chains in Java. ACM SIGSOFT Software Engineering Notes, 29(4): 1–11, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramesh:2004:STS

[RKK04] S. Ramesh, A. Kulkarni, and V. Kamat. Slicing tools for synchronous reactive programs. ACM SIGSOFT Software Engineering Notes, 29 (4):217–220, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robillard:2000:DRJ

[RM00a] Martin P. Robillard and Gail C. Murphy. Designing robust Java programs with exceptions. ACM SIGSOFT Software Engineering Notes, 25(6):2–10, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roman:2000:RDD

[RM00b] Gruia-Catalin Roman and Amy L. Murphy. Rapid development of dependable applications over ad hoc networks. ACM SIGSOFT Software Engineering Notes, 25 (1):77–78, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rakic:2001:ICS

[RM01] Marija Rakic and Nenad Medvidovic. Increasing the confidence in off-the-shelf components: a software connector-based approach. ACM SIGSOFT Software Engineering Notes, 26(3):11–18, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramachandran:2004:KBR

[RM04a] Muthu Ramachandran and Domenic Mangano. Knowledge based reasoning for software architectural design strategies. ACM SIGSOFT Software Engineering Notes, 29(3):1–4, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rossi:2004:FAM

[RM04b] Matteo Rossi and Dino Mandrioli. A formal approach for modeling and verification of RTCORBA-based applications. ACM SIGSOFT Software Engineering Notes, 29 (4):263–273, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robillard:2005:JTC

[RM05] Martin P. Robillard and Gail C. Murphy. Just-in-time concern modeling. ACM SIG-SOFT Software Engineering Notes, 30(4):1–3, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ranjan:2006:ABS

[RM06a] Prabhat Ranjan and A. K. Misra. Agent based system development: a domainspecific goal approach. ACM SIGSOFT Software Engineering Notes, 31(6):1–6, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ranjan:2006:HMA

[RM06b] Prabhat Ranjan and A. K. Misra. A hybrid model for agent based system requirements analysis. ACM SIG-SOFT Software Engineering Notes, 31(3):1–7, May 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rehman:2009:RSD

[RM09] S. Rehman and K. Mustafa.

Research on software design level security vulnerabilities. ACM SIGSOFT Software Engineering Notes, 34 (6):1–5, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robinson-Mallett:2006:ACC

[RMHL06] Christopher Robinson-Mallett, Robert M. Hierons, and Peter Liggesmeyer. Achieving communication coverage in testing. ACM SIGSOFT Software Engineering Notes, 31(6):1–10, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robinson-Mallett:2005:GOD

[RMLMG05] Christopher Robinson-Mallett, Peter Liggesmeyer, Tilo Mücke, and Ursula Goltz. Generating optimal distinguishing sequences with a model checker. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:2000:PCB

[RN00] D. Rine and N. Nada. Public case-base and tool kit using a validated RRM.

ACM SIGSOFT Software Engineering Notes, 25(1): 76, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roberts:2005:GMO

[Rob05a] Keith A. Roberts. Generic methodology for open source software development. ACM SIGSOFT Software Engineering Notes, 30(2):1–5, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robillard:2005:AGS

[Rob05b] Martin P. Robillard. Automatic generation of suggestions for program investigation. ACM SIGSOFT Software Engineering Notes, 30 (5):11–20, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robillard:2005:WMA

[Rob05c] Martin P. Robillard. Workshop on the Modeling and Analysis of Concerns in Software (MACS 2005). ACM SIGSOFT Software Engineering Notes, 30(4):1– 3, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Romanovsky:2001:CAA

[Rom01] Alexander Romanovsky. Coordinated atomic actions: how to remain ACID in the modern world. ACM SIGSOFT Software Engineering Notes, 26(2):66– 68, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Romanovsky:2007:LFT

[Rom07] Alexander Romanovsky. A looming fault tolerance software crisis? ACM SIGSOFT Software Engineering Notes, 32(2):1–4, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Romanovsky:2008:DID

[Rom08] Alexander Romanovsky. DE-PLOY: industrial deployment of advanced system engineering methods for high productivity and dependability. ACM SIGSOFT Software Engineering Notes, 33 (6):1–3, November 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:2000:CME

[Ros00] David S. Rosenblum. CA-REER: mechanisms for ensuring the integrity of distributed object systems. ACM SIGSOFT Software Engineering Notes, 25(1):78– 79, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:2001:TAE

[Ros01] David Rosenblum. Taming architectural evolution. ACM SIGSOFT Software Engineering Notes, 26(5):1– 10, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosen:2004:NDS

[Ros04a] Clive Rosen. Non-directive software engineering. ACM SIGSOFT Software Engineering Notes, 29(2):9, March 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rost:2004:FMR

[Ros04b]  Johann Rost.  Is “Fac-

tory Method” really a pattern? ACM SIGSOFT Software Engineering Notes, 29 (5):1, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rothermel:2000:TME

[Rot00]  Gregg Rothermel.  Test-

ing and maintaining evolving software systems. ACM SIGSOFT Software Engineering Notes, 25(1):79– 80, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rotondi:2004:AMP

[Rot04] Guido Rotondi. Assessment methodologies for public contractors. ACM SIGSOFT Software Engineering Notes, 29(5):1–6, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roman:2001:WSE

[RP01] Gruia-Catalin Roman and Gian Pietro Picco. Workshop on Software Engineering and Mobility. ACM SIGSOFT Software Engineering Notes, 26(6):67–70, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rappl:2006:IIW

[RPSS06] Martin Rappl, Alexander Pretschner, Christian Salzmann, and Thomas Stauner. 3rd Intl. ICSE Workshop on Software Engineering for Automotive Systems. ACM SIGSOFT Software Engineering Notes, 31(5):22– 23, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reddy:2009:DOC

[RR09] K. Reddy Reddy and A. Ananda Rao. Dependency oriented complexity metrics to detect rippling related design defects. ACM SIGSOFT Software Engineering Notes, 34 (4):1–7, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Requile-Romanczuk:2005:TKB

[RRCDHM05] Annya Réquilé-Romanczuk, Alejandra Cechich, Anne Dourgnon-Hanoune, and Jean-Christophe Mielnik. Towards a knowledge-based framework for COTS component identification. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ram:2004:PHB

[RRR04] D. Janaki Ram, P. Jithendra Kumar Reddy, and M. S. Rajasree. Pattern hybridization: breeding new designs out of pattern interactions. ACM SIGSOFT Software Engineering Notes, 29 (3):1–10, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Renieris:2005:APP

[RRR05] Manos Renieris, Shashank Ramaprasad, and Steven P. Reiss. Arithmetic program paths. ACM SIGSOFT Software Engineering Notes, 30 (5):90–98, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramakrishnan:2000:SEL

[RRSW00] C. R. Ramakrishnan, I. V.

Ramakrishnan, Scott A. Smolka, and David S. Warren. Specification and evaluation of logic-based model checking. ACM SIGSOFT Software Engineering Notes, 25(1):73–74, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rodrigues:2005:SAS

[RRU05] Genaína N. Rodrigues, David S.

Rosenblum, and Sebastian Uchitel. Sensitivity analysis for a scenario-based reliability prediction model. ACM SIGSOFT Software Engineering Notes, 30(4): 1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rajlich:2000:RWS

[RRWB00] V. T. Rajlich, S. Rank, N. Wilde, and K. H. Bennett. Report on a workshop on software change and evolution. ACM SIGSOFT Software Engineering Notes, 25 (1):28–29, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rech:2007:VSS

[RS07] Jörg Rech and Waldemar Schäfer. Visual support of software engineers during development and maintenance. ACM SIGSOFT Software Engineering Notes, 32 (2):1–3, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rinard:2004:CSA

[RSB04] Martin Rinard, Alexandru Salcianu, and Suhabe Bugrara. A classification system and analysis for aspectoriented programs. ACM SIGSOFT Software Engineering Notes, 29(6):147– 158, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rangarajan:2001:PQF

[RSHJ01] Krishnan Rangarajan, N. Swaminathan, Vishu Hegde, and Jacob Jacob. Product quality framework: a vehicle for focusing on product quality goals. ACM SIGSOFT Software Engineering Notes, 26 (4):77–82, July 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reimer:2004:SSA

[RSS^{+}04] Darrell Reimer, Edith Schonberg, Kavitha Srinivas, Harini Srinivasan, Bowen Alpern, Robert D. Johnson, Aaron Kershenbaum, and Larry Koved. SABER: smart analysis based error reduction. ACM SIGSOFT Software Engineering Notes, 29(4): 243–251, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ren:2005:TAT

[RTDR05] Jie Ren, Richard Taylor, Paul Dourish, and David Redmiles. Towards an architectural treatment of software security: a connector-centric approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rugaber:2000:ECS

[Rug00] Spencer Rugaber. Esprit de Corps suite. ACM SIGSOFT Software Engineering Notes, 25(1):101, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rountev:2006:SCF

[RVR06] Atanas Rountev, Olga Volgin, and Miriam Reddoch. Static control-flow analysis for reverse engineering of UML sequence diagrams. ACM SIGSOFT Software Engineering Notes, 31(1):96– 102, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rugaber:2000:SEI

[RW00] Spencer Rugaber and Linda Wills. Software evolution and interleaving. ACM SIGSOFT Software Engineering Notes, 25(1):80, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Siegel:2000:IPI

[SA00] Stephen F. Siegel and George S. Avrunin. Improving the precision of INCA by preventing spurious cycles. ACM SIGSOFT Software Engineering Notes, 25(5):191– 200, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saxena:2009:PEO

[SA09] Vipin Saxena and Deepak Arora. Performance evaluation for object oriented software systems. ACM SIG-SOFT Software Engineering Notes, 34(2):1–5, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scaffidi:2005:VBA

[SABS05] Chris Scaffidi, Ashish Arora, Shawn Butler, and Mary Shaw. A value-based approach to predicting system properties from design. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saini:2009:TPO

[Sai09] Dinesh Kumar Saini. Testing polymorphism in object oriented systems for improving software quality. ACM SIGSOFT Software Engineering Notes, 34 (2):1–5, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sohr:2005:AEA

[SAM05] Karsten Sohr, Gail-Joon Ahn, and Lars Migge. Articulating and enforcing authorisation policies with UML and OCL. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sanden:2009:ISD

[San09] Bo Sandén. Inspired software design early Jackson methods to thread architectures. ACM SIGSOFT Software Engineering Notes, 34 (4):1–6, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saxena:2009:PEN

[SAS09] Vipin Saxena, Deepak Arora, and Manish Shrivastava. Performance evaluation of network system through UML. ACM SIGSOFT Software Engineering Notes, 34(5):1– 3, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2002:ERP

[Sau02] Joe Saur. Effective requirements practices. ACM SIG-SOFT Software Engineering Notes, 27(3):93–94, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2004:RSM

[Sau04a] Joe Saur. Review of SASÒ for Monte Carlo studies: a guide for quantitative researchers by Xitao Fan, Ákos Felsvlyi, Stephen A. Sivo, and Sean C. Keenan. SAS Institute, Inc. 2002. ACM SIGSOFT Software Engineering Notes, 29(5): 39, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2004:RVC

[Sau04b] Joe Saur. Review of Visualizing categorical data by Michael Friendly. SAS Institute, Inc. 2000. ACM SIGSOFT Software Engineering Notes, 29(5):39– 40, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2004:RCT

[Sau04c] Joseph M. Saur. Review of Critical testing processes: plan, prepare, perform, perfect by Rex Black. Addison-Wesley 2004. ACM SIG-SOFT Software Engineering Notes, 29(3):30, May 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2004:RSN

[Sau04d] Joseph M. Saur. Review of Software by numbers: lowrisk, high-return development by Mark Denne and Jane Cleland-Huang. Prentice Hall 2004. ACM SIGSOFT Software Engineering Notes, 29 (3):29–30, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2004:RSE

[Sau04e] Joseph M. Saur. Review of Software engineering measurement by John C. Munson. Auerbach Publications 2003. ACM SIGSOFT Software Engineering Notes, 29 (3):30, May 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2005:PMH

[Sau05a] Joe Saur. Project management: here we go again! ACM SIGSOFT Software Engineering Notes, 30(4): 8, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2005:RFB

[Sau05b] Joe Saur. Review of Find the Bug: a Book of Incorrect Programs by Adam Barr; Pearson Education, 2004, ISBN 0-321-22391-8. ACM SIGSOFT Software Engineering Notes, 30(2):30–31, March 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2005:RSD

[Sau05c] Joe Saur. Review of Software Development for Small Teams: a RUP-Centric Approach by Gary Pollice, Liz Augustine, Chris Lowe, and Jas Madhur. Addison-Wesley, 2004, paperback, ISBN 0-321-19950-2. ACM SIGSOFT Software Engineering Notes, 30(1):38, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2005:RSR

[Sau05d] Joe Saur. Review of Software Reliability Engineering: More Reliable Software Faster and Cheaper — 2nd Edition by John Musa. AuthorHouse, Inc., 2004, Hardback, ISBN 1-4184-9388-0, Paperback, ISBN 1-4184-9387-2. ACM SIGSOFT Software Engineering Notes, 30 (1):38–39, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2006:RIC

[Sau06a]  Joe Saur.  Review of

IEEE Computer Society Real-World Software Engineering Problems: a Self-Study Guide for Today’s Software Professional by J. Fernando Naveda and Stephen B. Seidman, Prentice Hall, 2006, ISBN 0-471-71051-2. ACM SIGSOFT Software Engineering Notes, 31(6): 58, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2006:RJE

[Sau06b] Joe Saur. Review of Just Enough Requirements Management: Where Software Development Meets Marketing by Alan M. Davis, Dorset House Publishing, 2005, ISBN 0-932633-64-1. ACM SIGSOFT Software Engineering Notes, 31 (2):33, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2006:RST

[Sau06c] Joe Saur. Review of System Testing with an Attitude: an Approach That Nurtures Front-Loaded (Designed and built in … not tested in!) Software Quality by Nathan Petschenik, Dorset House Publishing, 2005, ISBN 0-932633-46-3. ACM SIGSOFT Software Engineering Notes, 31(2):32–33, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2008:RSC

[Sau08] Joe Saur. Review of Secure Coding in C and C++ by Robert C. Seacord, Addison Wesley, 2006, paperback, ISBN 0-321-33572-4, 341 pp., $39.99. ACM SIGSOFT Software Engineering Notes, 33 (3):7:1–7:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2009:BRC

[Sau09a] Joe Saur. Book review: C Programming: a Modern Approach (2nd Edit.) written by K. N. King, and published by W. W. Norton and Company, paperback, ISBN 978-0-393-97950-3, 832 pp. ACM SIGSOFT Software Engineering Notes, 34(5):57–58, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2009:BRM

[Sau09b] Joe Saur. Book review:

Mathematical Modeling and Simulation: Introduction for Scientists and Engineers by Kai Velten, and published by Wiley-VCH, 2009, paperback, ISBN 978-3-527-40758-3, 348 pp. ACM SIGSOFT Software Engineering Notes, 34(5):58, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sayre:2005:UMB

[Say05] Kirk Sayre. Usage modelbased automated testing of C++ templates. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sliski:2001:AFE

[SBCO01] Timothy J. Sliski, Matthew P.

Billmers, Lori A. Clarke, and Leon J. Osterweil. An architecture for flexible, evolvable process-driven userguidance environments. ACM SIGSOFT Software Engineering Notes, 26(5):33– 43, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2009:AMP

[SBS09a] Yogesh Singh, Pradeep Kumar Bhatia, and Omprakash Sangwan. ANN model for predicting software function point metric. ACM SIG-SOFT Software Engineering Notes, 34(1):1–4, January 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2009:PSM

[SBS09b] Yogesh Singh, Pradeep Kumar Bhatia, and Omprakash Sangwan. Predicting software maintenance using fuzzy model. ACM SIGSOFT Software Engineering Notes, 34 (4):1–6, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shull:2000:EVP

[SBZ00] Forrest Shull, Victor R. Basili, and Marvin Zelkowitz. The experimental validation and packaging of software technologies. ACM SIGSOFT Software Engineering Notes, 25(1):82– 83, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sengupta:2002:TMS

[SC02] Bikram Sengupta and Rance Cleaveland. Triggered message sequence charts. ACM SIGSOFT Software Engineering Notes, 27(6):167– 176, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaikh:2005:LOL

[SC05a] Maha Shaikh and Tony Cornford. Learning/organizing in Linux: a study of the ‘spaces in between’. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shin:2005:CBS

[SC05b] Michael E. Shin and Daniel Cooke. Connector-based selfhealing mechanism for components of a reliable system. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scacchi:2005:OBE

[Sca05]  Walt Scacchi.  OpenEC/

B: electronic commerce and free/open source software development. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scedrov:2000:LBS

[Sce00] Andre Scedrov. A languagebased security analysis. ACM SIGSOFT Software Engineering Notes, 25(1):80– 81, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Steven:2000:JCR

[SCFP00] John Steven, Pravir Chandra, Bob Fleck, and Andy Podgurski. jRapture: a capture/replay tool for observationbased testing. ACM SIG-SOFT Software Engineering Notes, 25(5):158–167, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scherlis:2000:ACT

[Sch00a] William L. Scherlis. Adaptation and commitment technology (ACT). ACM SIG-SOFT Software Engineering Notes, 25(1):96, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schmidt:2000:AIP

[Sch00b] David Schmidt. Abstract interpretation and program modelling. ACM SIGSOFT Software Engineering Notes, 25(1):81–82, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2005:DQM

[Sch05a] Robert Schaefer. Deeper questions: the metaproblem of large organizations developing complex systems and the limits of process.

ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2005:RLO

[Sch05b] Robert Schaefer. The risks of large organizations in developing complex systems. ACM SIGSOFT Software Engineering Notes, 30(5):1– 3, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2005:SSC

[Sch05c] Robert Schaefer. Systems of systems and coordinated atomic actions. ACM SIG-SOFT Software Engineering Notes, 30(1):6, January 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2006:CPS

[Sch06a] Robert Schaefer. A critical programmer searches for professionalism. ACM SIGSOFT Software Engineering Notes, 31(4):1–17, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2006:RTS

[Sch06b] Robert Schaefer. A rational theory of system-making systems. ACM SIGSOFT Software Engineering Notes, 31 (2):1–20, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2006:RCD

[Sch06c] Robert Schaefer. Review of Catastrophe Disentanglement by E. M. Bennatan, Addison Wesley, 2006, ISBN 0-321-33662-0. ACM SIGSOFT Software Engineering Notes, 31(6):59–60, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2006:RDT

[Sch06d] Robert Schaefer. Review of Design For Trustworthy Software by Bijay K. Jayaswal and Peter C. Patton, Prentice Hall, 2007, ISBN 0-13-187250-8. ACM SIGSOFT Software Engineering Notes, 31(6):58–59, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See response [Jay06].

Schaefer:2006:RHB

[Sch06e] Robert Schaefer. Review of Hiring The Best Knowledge Workers, Techies and Nerds by Johanna Rothman, Dorset House, 2004, 0-932633-59-5. ACM SIGSOFT Software Engineering Notes, 31 (2):34, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2006:RMW

[Sch06f]  Robert Schaefer.  Re-

view of Mobile Wireless Communications by Mischa Schwarz, Cambridge University Press, 2005, 0-521-84347-2. ACM SIGSOFT Software Engineering Notes, 31 (2):34, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2007:GTE

[Sch07] Robert Schaefer. The grand theory of everything: what man-made systems are, and why they fail. ACM SIG-SOFT Software Engineering Notes, 32(4):8:1–8:??, July 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2008:DDM

[Sch08a] Robert Schaefer. Debugging debugged, a metaphysical manifesto of systems integration. ACM SIGSOFT Software Engineering Notes, 33 (3):5:1–5:??, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2008:SAS

[Sch08b] Robert Schaefer. A systems analysis of systems integration. ACM SIGSOFT Software Engineering Notes, 33 (1):7:1–7:??, January 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2009:BRD

[Sch09a] Robert Schaefer. Book review: Design, Measurement and Management of Large-Scale IP Networks by Antonio Nucci and Konstantina Papagiannaki, and published by Cambridge University Press, 2009, 978-0-521-88069-5, 394 pp. ACM SIGSOFT Software Engineering Notes, 34(6):29, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2009:BRS

[Sch09b] Robert Schaefer. Book review: Scientific Software: a Guide to Good Style by Suely Oliveira and David Stewart, and published by Cambridge University Press, 2006, (paperback), 0-521-67595-2, 303 pp. ACM SIGSOFT Software Engineering Notes, 34 (6):30, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2009:ECS

[Sch09c] Robert Schaefer. The epistemology of computer security. ACM SIGSOFT Software Engineering Notes, 34(6):8–10, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2009:SMD

[Sch09d] Robert Schaefer. Software maturity: design as dark art. ACM SIGSOFT Software Engineering Notes, 34 (1):1–36, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2006:EBE

[SCS06] R. K. Singh, Pravin Chandra, and Yogesh Singh. An evaluation of Boolean expression testing techniques. ACM SIGSOFT Software Engineering Notes, 31(5):1– 6, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shull:2001:EMI

[SCT01] Forrest Shull, Jeffrey Carver, and Guilherme H. Travassos. An empirical methodology for introducing software processes. ACM SIGSOFT Software Engineering Notes, 26 (5):288–296, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Suarez-Cabal:2004:USC

[SCT04] María José Suárez-Cabal and Javier Tuya. Using an SQL coverage measurement for testing database applications. ACM SIGSOFT Software Engineering Notes, 29 (6):253–262, November 2004.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stirewalt:2001:GVC

[SD01] R. E. Kurt Stirewalt and Laura K. Dillon. Generation of visitor components that implement program transformations. ACM SIGSOFT Software Engineering Notes, 26(3):86–94, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stewart:2005:OPD

[SDD05] Katherine J. Stewart, David P. Darcy, and Sherae L. Daniel. Observations on patterns of development in open source software projects. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schmietendorf:2000:MBA

[SDF00] Andreas Schmietendorf, Reiner Dumke, and Erik Foltin. Metrics based asset assessment. ACM SIGSOFT Software Engineering Notes, 25 (4):51–55, July 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saff:2004:EEC

[SE04] David Saff and Michael D. Ernst. An experimental evaluation of continuous testing during development. ACM SIGSOFT Software Engineering Notes, 29 (4):76–85, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Segal:2005:TPE

[Seg05] Judith Segal. Two principles of end-user software engineering research. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saleh:2005:SCS

[SG05] Mazen Saleh and Hassan Gomaa. Separation of concerns in software product line engineering. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2007:STS

[SG07] Yogesh Singh and Bindu Goel. A step towards software preventive maintenance. ACM SIGSOFT Software Engineering Notes, 32(4): 10:1–10:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srivastava:2009:UGA

[SGAY09] Praveen Ranjan Srivastava, Priyanka Gupta, Yogita Arrawatia, and Suman Yadav. Use of genetic algorithm in generation of feasible test data. ACM SIGSOFT Software Engineering Notes, 34 (2):1–4, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sullivan:2001:SVM

[SGCH01] Kevin J. Sullivan, William G. Griswold, Yuanfang Cai, and Ben Hallen. The structure and value of modularity in software design. ACM SIGSOFT Software Engineering Notes, 26(5):99– 108, September 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2009:DAC

[SGK09a] Arun Sharma, P. S. Grover, and Rajesh Kumar. Dependency analysis for componentbased software systems. ACM SIGSOFT Software Engineering Notes, 34(4): 1–6, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2009:RAS

[SGK09b] Arun Sharma, P. S. Grover, and Rajesh Kumar. Reusability assessment for software components. ACM SIGSOFT Software Engineering Notes, 34(2):1–6, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutherland:2003:CMC

[SGS03] Dean F. Sutherland, Aaron Greenhouse, and William L. Scherlis. The code of many colors: relating threads to code and shared state. ACM SIGSOFT Software Engineering Notes, 28(1):77– 83, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Segal:2005:TEP

[SGS05a] Judith Segal, Antony Grinyer, and Helen Sharp. The type of evidence produced by empirical software engineers. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sullivan:2005:IHI

[SGS^{+}05b] Kevin Sullivan, William G.

Griswold, Yuanyuan Song, Yuanfang Cai, Macneil Shonle, Nishit Tewari, and Hridesh Rajan. Information hiding interfaces for aspect-oriented design. ACM SIGSOFT Software Engineering Notes, 30 (5):166–175, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schmerl:2005:DDA

[SGY05] Bradley Schmerl, David Garlan, and Hong Yan. Dynamically discovering architectures with DiscoTect. ACM SIGSOFT Software Engineering Notes, 30(5):103– 106, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shao:2000:TCI

[Sha00] Zhong Shao. Typed common intermediate format. ACM SIGSOFT Software Engineering Notes, 25(1): 82, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaofeng:2001:RJR

[Sha01] Wang Shaofeng. The role of Java RMI in designing workflow management system. ACM SIGSOFT Software Engineering Notes, 26 (2):49–52, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shakir:2002:TAB

[Sha02]  Maha Shakir.  Technol-

ogy acquisition: buying the future of your business. ACM SIGSOFT Software Engineering Notes, 27 (2):73–74, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schelfthout:2005:VCA

[SHB05] Kurt Schelfthout, Tom Holvoet, and Yolande Berbers. Views: customizable abstractions for contextaware applications in MANETs. ACM SIGSOFT Software Engineering Notes, 30(4):1–8, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sundaram:2005:BRT

[SHD05] Senthil Karthikeyan Sundaram, Jane Huffman Hayes, and Alexander Dekhtyar. Baselines in requirements tracing. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sherriff:2005:UVV

[She05] Mark Sherriff. Utilizing verification and validation certificates to estimate software defect density. ACM SIGSOFT Software Engineering Notes, 30(5):381– 384, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shimomura:2004:PTF

[Shi04] Takao Shimomura. A pagetransition framework for image-oriented Web programming. ACM SIGSOFT Software Engineering Notes, 29(2):10, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sen:2005:OEK

[SHRG05] Rohan Sen, Gregory Hackmann, Gruia-Catalin Roman, and Christopher Gill. Opportunistic exploitation of knowledge to increase predictability of agent interactions in MANETs. ACM SIGSOFT Software Engineering Notes, 30(4):1– 7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sillitti:2009:DES

[Sil09]  Alberto Sillitti.  Design-

ing empirical studies: assessing the effectiveness of agile methods. ACM SIGSOFT Software Engineering Notes, 34(5):35–37, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sistla:2000:FMC

[Sis00]  A. Prasad Sistla.  For-

mal methods in concurrent and distributed systems. ACM SIGSOFT Software Engineering Notes, 25 (1):84, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schroeder:2000:BBT

[SK00] Patrick J. Schroeder and Bogdan Korel. Black-box test reduction using inputoutput analysis. ACM SIGSOFT Software Engineering Notes, 25(5):173– 177, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Savolainen:2001:VAF

[SK01a] Juha Savolainen and Juha Kuusela. Violatility analysis framework for product lines. ACM SIGSOFT Software Engineering Notes, 26 (3):133–141, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sim:2001:WWS

[SK01b] Susan Elliot Sim and Rainer Koschke. WoSEF: Workshop on Standard Exchange Format. ACM SIGSOFT Software Engineering Notes, 26 (1):44–49, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2008:EQS

[SKG08] Arun Sharma, Rajesh Kumar, and P. S. Grover. Estimation of quality for software components: an empirical approach. ACM SIGSOFT Software Engineering Notes, 33(6):1–10, November 2008.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sitaraman:2001:PSS

[SKK^{+}01a] Murali Sitaraman, Greg Kulczycki, Joan Krone, William F. Ogden, and A. L. N. Reddy. Performance specification of software components. ACM SIGSOFT Software Engineering Notes, 26(3):3–10, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Systa:2001:SRO

[SKK01b] Tarja Systä, Rudolf K. Keller, and Kai Koskimies. Summary report of the OOP-SLA 2000 Workshop on Scenario-Based Round-Trip Engineering. ACM SIGSOFT Software Engineering Notes, 26(2):24–28, March 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Senthil:2007:ICM

[SKM07] R. Senthil, D. S. Kushwaha, and A. K. Misra. An improved component model for component based software engineering. ACM SIGSOFT Software Engineering Notes, 32(4):9:1–9:??, July 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2009:ASV

[SKM09] Yogesh Singh, Arvinder Kaur, and Ruchika Malhotra. Application of support vector machine to predict fault prone classes. ACM SIGSOFT Software Engineering Notes, 34(1): 1–6, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srivastva:2008:TCP

[SKR08] Praveen Ranjan Srivastva, Krishan Kumar, and G. Raghurama. Test case prioritization based on requirements and risk factors. ACM SIGSOFT Software Engineering Notes, 33(4):7:1–7:??, July 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staron:2005:EAU

[SKT05] Miroslaw Staron, Ludwik Kuzniarz, and Christian Thurn. An empirical assessment of using stereotypes to improve reading techniques in software inspections. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Simmons:2005:PGA

[SL05] Bradley Simmons and Hanan Lutfiyya. Policies, grids and autonomic computing. ACM SIGSOFT Software Engineering Notes, 30(4): 1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schulz:2000:TAD

[SM00] Daniel Schulz and Frank Mueller. A thread-aware debugger with an open interface. ACM SIGSOFT Software Engineering Notes, 25 (5):201–211, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Samuel:2009:SBT

[SM09a] Philip Samuel and Rajib Mall. Slicing-based test case generation from UML activity diagrams. ACM SIGSOFT Software Engineering Notes, 34(6):1–14, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Suman:2009:SME

[SM09b] Rajiv Ranjan Suman and Rajib Mall. State model extraction of a software component by observing its behavior. ACM SIGSOFT Software Engineering Notes, 34 (1):1–7, January 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sen:2005:CCU

[SMA05] Koushik Sen, Darko Marinov, and Gul Agha. CUTE: a concolic unit testing engine for C. ACM SIGSOFT Software Engineering Notes, 30 (5):263–272, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sadjadi:2005:TSE

[SMC05a] S. Masoud Sadjadi, Philip K.

McKinley, and Betty H. C. Cheng. Transparent shaping of existing software to support pervasive and autonomic computing. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:2005:IIA

[SMC05b] Dennis Smith, Edwin Morris, and David Carney. Interoperability issues affecting autonomic computing. ACM SIG-SOFT Software Engineering Notes, 30(4):1–3, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Strout:2006:RIP

[SMCH06] Michelle Mills Strout, John Mellor-Crummey, and Paul Hovland. Representationindependent program analysis. ACM SIGSOFT Software Engineering Notes, 31 (1):67–74, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schilling:2005:IMD

[SMD^{+}05] Albert Schilling, Kelma Madeira, Paula Donegan, Kênia Sousa, Elizabeth Furtado, and Vasco Furtado. An integrated method for designing user interfaces based on tests. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smidts:2000:AAG

[Smi00] Carol S. Smidts. An approach to the automatic generation of software functional architecture. ACM SIGSOFT Software Engineering Notes, 25(1):84–85, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2009:EAM

[SMK09] Ranjit Singh, Shakti Mishra, and D. S. Kushwaha. An efficient asynchronous mobile web service framework. ACM SIGSOFT Software Engineering Notes, 34(6):1– 7, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srivastava:2005:ASM

[SN05] Deepti Srivastava and Priya Narasimhan. Architectural support for mode-driven fault tolerance in distributed applications. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sabetzadeh:2006:RDA

[SNEC06] Mehrdad Sabetzadeh, Shiva Nejati, Steve Easterbrook, and Marsha Chechik. A relationship-driven approach to view merging. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sherriff:2005:EED

[SNWV05] Mark Sherriff, Nachiappan Nagappan, Laurie Williams, and Mladen Vouk. Early estimation of defect density using an in-process Haskell metrics model. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Souter:2000:OST

[SP00] Amie L. Souter and Lori L.

Pollock. OMEN: a strategy for testing object-oriented software. ACM SIGSOFT Software Engineering Notes, 25(5):49–59, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srivastava:2009:CEG

[SPC09] Praveen Ranjan Srivastava, Parshad Patel, and Siddharth Chatrola. Cause effect graph to decision table generation. ACM SIGSOFT Software Engineering Notes, 34 (2):1–4, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sawyer:2007:WCR

[SPH07] Pete Sawyer, Barbara Paech, and Patrick Heymans. Working conference on requirements engineering: foundation for software quality (REFSQ’07). ACM SIG-SOFT Software Engineering Notes, 32(4):47–53, July 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Spraragen:2005:CCT

[Spr05] Susan L. Spraragen. The challenges in creating tools for improving the software development lifecycle. ACM SIGSOFT Software Engineering Notes, 30 (4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srivastava:2008:NHP

[SPS^{+}08] Praveen R. Srivastava, Deepak Pareek, Kailash Sati, Dinesh C. Pujari, and G. Raghurama. Non homogeneous Poisson process based cumulative priority model for determining optimal software testing period. ACM SIGSOFT Software Engineering Notes, 33(2):6:1– 6:??, March 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shepherd:2005:ULC

[SPT05] David Shepherd, Lori Pollock, and Tom Tourwé. Using language clues to discover crosscutting concerns. ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharp:2005:SSF

[SR05] Helen Sharp and Hugh Robinson. Some social factors of software engineering: the maverick, community and technical practices. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sreedhar:2001:YPS

[Sre01] Vugranam C. Sreedhar. York: programming software components. ACM SIGSOFT Software Engineering Notes, 26(5):305–306, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sreenivas:2002:PDI

[Sre02] Ashok Sreenivas. Panel discussion: is ISSTA testing research relevant to industrial users? ACM SIGSOFT Software Engineering Notes, 27 (4):208–209, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srinivasa:2001:CPS

[Sri01]  Srinath Srinivasa.  Con-

notations of problem solving. ACM SIGSOFT Software Engineering Notes, 26 (6):80–82, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srinivas:2005:SAP

[SS05] Kavitha Srinivas and Harini Srinivasan. Summarizing application performance from a components perspective. ACM SIGSOFT Software Engineering Notes, 30(5): 136–145, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2008:DRM

[SS08] Paramvir Singh and Hardeep Singh. DynaMetrics: a runtime metric-based analysis tool for object-oriented software systems. ACM SIGSOFT Software Engineering Notes, 33(6):1–6, November 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saxena:2009:PFP

[SS09] Vipin Saxena and Manish Shrivastava. Performance of function point analysis through UML modeling. ACM SIGSOFT Software Engineering Notes, 34 (2):1–4, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sampath:2005:ACW

[SSG^{+}05a] Sreedevi Sampath, Sara Sprenkle, Emily Gibson, Lori Pollock, and Amie Souter. Analyzing clusters of web application user sessions. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sant:2005:ESM

[SSG05b] Jessica Sant, Amie Souter, and Lloyd Greenwald. An exploration of statistical models for automated test case generation. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shen:2005:TUF

[SSH^{+}05] Junrong Shen, Xi Sun, Gang Huang, Wenpin Jiao, Yanchun Sun, and Hong Mei. Towards a unified formal model for supporting mechanisms of dynamic component update. ACM SIGSOFT Software Engineering Notes, 30(5):80–89, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Spacco:2005:SRM

[SSHP05] Jaime Spacco, Jaymie Strecker, David Hovemeyer, and William Pugh. Software repository mining with marmoset: an automated programming project snapshot and testing system. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scaffidi:2005:ACE

[SSM05] Christopher Scaffidi, Mary Shaw, and Brad Myers. An approach for categorizing end user programmers to guide software engineering research. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sweeney:2000:ELB

[ST00] Peter F. Sweeney and Frank Tip. Extracting library-based object-oriented applications. ACM SIGSOFT Software Engineering Notes, 25(6):98– 107, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srivastava:2002:EPT

[ST02] Amitabh Srivastava and Jay Thiagarajan. Effectively prioritizing tests in development environment. ACM SIGSOFT Software Engineering Notes, 27(4):97– 106, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sagheb-Tehrani:2005:ESD

[ST05a] Mehdi Sagheb-Tehrani. Expert systems development: some issues of design process. ACM SIGSOFT Software Engineering Notes, 30 (2):1–5, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Salehie:2005:ACE

[ST05b] Mazeiar Salehie and Ladan Tahvildari. Autonomic computing: emerging trends and open problems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stavridou:2000:AAC

[Sta00] Victoria Stavridou. Architectural analysis of componentbased systems. ACM SIG-SOFT Software Engineering Notes, 25(1):96–97, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stalica:2001:MCO

[Sta01] Mark Stalica. Mobile commerce: opportunities, applications, and technologies of wireless business. ACM SIGSOFT Software Engineering Notes, 26(6): 85, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Serfass:2008:SSP

[STB08] Doug Serfass, Kalyan S. Thiparthi, and Coskun Bayrak. Search, sort, pagination, association and (other) CRUD: a comparison of ActiveScaffold and MyFaces tomahawk. ACM SIGSOFT Software Engineering Notes, 33 (4):1:1–1:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stewart:2000:AAF

[Ste00]  David B. Stewart.  Au-

tomated analysis and finetuning of timing properties in embedded real-time systems. ACM SIGSOFT Software Engineering Notes, 25 (1):86–87, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sagheb-Tehrani:2002:SDP

[STG02] Mehdi Sagheb-Tehrani and Arbi Ghazarian. Software development process: strategies for handling business rules and requirements. ACM SIGSOFT Software Engineering Notes, 27(2): 58–62, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stidolph:2000:EDC

[Sti00] Wayne Stidolph. Evolutionary Design of Complex Software (EDCS) demonstration days 1999. ACM SIGSOFT Software Engineering Notes, 25(1):95, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stoller:2002:DPO

[Sto02] Scott D. Stoller. Domain partitioning for open reactive systems. ACM SIGSOFT Software Engineering Notes, 27(4):44–54, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stuckenholz:2005:CEV

[Stu05] Alexander Stuckenholz. Component evolution and versioning state of the art. ACM SIGSOFT Software Engineering Notes, 30(1): 7, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sukhodolsky:2001:OSP

[Suk01] Jacob Sukhodolsky. Optimizing software process control. ACM SIGSOFT Software Engineering Notes, 26 (2):59–63, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutcliffe:2005:ASG

[Sut05a] Alistair Sutcliffe. Applying small group theory to analysis and design of CSCW systems. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutcliffe:2005:ECB

[Sut05b] Alistair Sutcliffe. Evaluating the costs and benefits of enduser development. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:2000:BRI

[SV00] Geoffrey Smith and Dennis Volpano. Basic research in information privacy. ACM SIGSOFT Software Engineering Notes, 25 (1):86, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sinnema:2006:UVM

[SvdVD06] Marco Sinnema, Jan Salvador van der Ven, and Sybren Deelstra. Using variability modeling principles to capture architectural knowledge. ACM SIGSOFT Software Engineering Notes, 31 (5):5:1–5:??, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srikanth:2005:ERB

[SW05] Hema Srikanth and Laurie Williams. On the economics of requirementsbased test case prioritization. ACM SIGSOFT Software Engineering Notes, 30 (4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sturmer:2005:OES

[SWC05] Ingo Stürmer, Daniela Weinberg, and Mirko Conrad.

Overview of existing safeguarding techniques for automatically generated code. ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Swift:2000:PPA

[Swi00]  Terrance Swift.  Princi-

ples, practice, and applications of tabled logic programming. ACM SIGSOFT Software Engineering Notes, 25 (1):87–88, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sullivan:2004:SAB

[SYC^{+}04] Kevin Sullivan, Jinlin Yang, David Coppit, Sarfraz Khurshid, and Daniel Jackson. Software assurance by bounded exhaustive testing. ACM SIGSOFT Software Engineering Notes, 29 (4):133–142, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saini:2009:EIM

[SYO09] Dinesh Kumar Saini, Jabar H. Yousif, and Wail M. Omar. Enhanced inquiry method for malicious object identification. ACM SIGSOFT Software Engineering Notes, 34 (3):1–5, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schroter:2006:WDB

[SZPZ06] Adrian Schröter, Thomas Zimmermann, Rahul Premraj, and Andreas Zeller. Where do bugs come from? ACM SIGSOFT Software Engineering Notes, 31(6):1– 2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sliwerski:2005:HRR

[ŚZZ05a] Jacek Śliwerski, Thomas Zimmermann, and Andreas Zeller. HATARI: raising risk awareness. ACM SIGSOFT Software Engineering Notes, 30(5):107– 110, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sliwerski:2005:WDC

[ŚZZ05b] Jacek Śliwerski, Thomas Zimmermann, and Andreas Zeller. When do changes induce fixes? ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tan:2004:HGC

[TAC^{+}04] Jianbin Tan, George S.

Avrunin, Lori A. Clarke, Shlomo Zilberstein, and Stefan Leue. Heuristic-guided counterexample search in FLAVERS. ACM SIG-SOFT Software Engineering Notes, 29(6):201–210, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tai:2000:ATC

[Tai00] K. C. Tai. Analysis and testing of concurrent objectoriented software. ACM SIG-SOFT Software Engineering Notes, 25(1):88, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tai:2005:CAR

[Tai05] Grace Tai. A communication architecture from rapid prototyping. ACM SIGSOFT Software Engineering Notes, 30(4):1–3, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tanuan:2001:BAS

[Tan01] Meyer Tanuan. Building application servers. ACM SIG-SOFT Software Engineering Notes, 26(2):77, March 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tanuan:2002:LSD

[Tan02a]  Meyer Tanuan.  Lead-

ing a software development team. ACM SIGSOFT Software Engineering Notes, 27 (3):94, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tanuan:2002:TIS

[Tan02b] Meyer Tanuan. Testing it:

an off-the-shelf software testing process. ACM SIGSOFT Software Engineering Notes, 27(5):105, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tavares:2007:SSH

[TBM^{+}07] Eduardo Tavares, Raimundo Barreto, Paulo Maciel, Meuse Oliveira, Jr., Leonardo Amorim, Fernando Rocha, and Ricardo Lima. Software synthesis for hard real-time embedded systems with multiple processors. ACM SIGSOFT Software Engineering Notes, 32(2):1–10, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taibi:2007:CCO

[TDA07] Fathi Taibi, Jacob K. Daniel, and Fouad Mohammed Abbou. On checking the consistency of Object-Z classes. ACM SIGSOFT Software Engineering Notes, 32(4): 11:1–11:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tuya:2008:CEW

[TDSCdlR08] Javier Tuya, Javier Dolado, Ma Jose Suarez-Cabal, and Claudio de la Riva. A controlled experiment on whitebox database testing. ACM SIGSOFT Software Engineering Notes, 33(1):8:1– 8:??, January 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tan:2004:EEE

[TE04] Roy Patrick Tan and Stephen H. Edwards. Experiences evaluating the effectiveness of JML-JUnit testing. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tee:2009:EMR

[Tee09] Sim Hui Tee. Eliminating method redundancy for the improvement of inner class design. ACM SIGSOFT Software Engineering Notes, 34 (6):1–3, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teitelbaum:2000:C

[Tei00] Tim Teitelbaum. CodeSurfer. ACM SIGSOFT Software Engineering Notes, 25(1): 99, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tempero:2009:SEC

[Tem09] Ewan Tempero. Software engineering and computer sciences: two worlds. ACM SIG-SOFT Software Engineering Notes, 34(3):6, May 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terekhov:2004:DAI

[Ter04] Andrey A. Terekhov. Dealing with architectural issues: a case study. ACM SIGSOFT Software Engineering Notes, 29(2):11, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tallam:2006:CAI

[TG06] Sriraman Tallam and Neelam Gupta. A concept analysis inspired greedy algorithm for test suite minimization. ACM SIGSOFT Software Engineering Notes, 31 (1):35–42, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Torkar:2009:ECS

[TGF09] Richard Torkar, Tony Gorschek, and Robert Feldt. Eighth Conference on Software Engineering Research and Practice in Sweden (SERPS’08). ACM SIGSOFT Software Engineering Notes, 34(2):

31–33, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tichy:2005:COS

[TGSP05] Matthias Tichy, Holger Giese, Daniela Schilling, and Wladimir Pauls. Computing optimal self-repair actions: damage minimization versus repair time. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tikir:2002:EIC

[TH02] Mustafa M. Tikir and Jeffrey K. Hollingsworth. Efficient instrumentation for code coverage testing. ACM SIGSOFT Software Engineering Notes, 27(4):86– 96, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tikir:2003:RDS

[THL03] Mustafa M. Tikir, Jeffrey K. Hollingsworth, and Guei-Yuan Lueh. Recompilation for debugging support in a JIT-compiler. ACM SIGSOFT Software Engineering Notes, 28(1):10– 17, January 2003. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tarr:2001:WMD

[THO^{+}01a] Peri Tarr, William Harrison, Harold Ossher, Anthony Finkelsteiin, Bashar Nuseibeh, and Dewayne Perry. Workshop on Multi-Dimensional Separation of Concerns in Software Engineering. ACM SIGSOFT Software Engineering Notes, 26(1):78–81, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thompson:2001:IWC

[Tho01b] J. Barrie Thompson. ICSE2001 workshop to consider global aspects of software engineering professionalism. ACM SIGSOFT Software Engineering Notes, 26(6):40– 44, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tian:2000:EMI

[Tia00] Jeff Tian. Early measurement and improvement of software reliability. ACM SIGSOFT Software Engineering Notes, 25(1):89, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thang:2006:SVI

[TK06] Nguyen Truong Thang and Takuya Katayama. Specification and verification of inter-component constraints in CTL. ACM SIGSOFT Software Engineering Notes, 31(2):8:1–8:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tripathi:2005:PDC

[TKA05] Anand Tripathi, Devdatta Kulkarni, and Tanvir Ahmed. Policy-driven configuration and management of agent based distributed systems. ACM SIGSOFT Software Engineering Notes, 30(4): 1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tiako:2001:PSD

[TLG01] Pierre F. Tiako, Tim Lindquist, and Volker Gruhn. Process Support for Distributed Team-Based Software Development Workshop. ACM SIGSOFT Software Engineering Notes, 26(6):31– 33, November 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tomayko:2006:PJT

[Tom06] James E. Tomayko. Professor James E. Tomayko. ACM SIGSOFT Software Engineering Notes, 31(2): 6, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tonella:2004:ETC

[Ton04] Paolo Tonella. Evolutionary testing of classes. ACM SIGSOFT Software Engineering Notes, 29(4):119– 128, July 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:Aa

[TR00a] Richard Taylor and David Redmiles. Arabica. ACM SIGSOFT Software Engineering Notes, 25(1):96, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:Ab

[TR00b] Richard Taylor and David Redmiles. ArchStudio. ACM SIGSOFT Software Engineering Notes, 25(1): 97, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:AU

[TR00c] Richard Taylor and David Redmiles. Argo/UML. ACM SIGSOFT Software Engineering Notes, 25(1): 97, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:C

[TR00d] Richard Taylor and David Redmiles. Chimera. ACM SIGSOFT Software Engineering Notes, 25(1):98, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:E

[TR00e] Richard Taylor and David Redmiles. Endeavors. ACM SIGSOFT Software Engineering Notes, 25(1):101, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:EDE

[TR00f] Richard Taylor and David Redmiles. Expectationdriven event monitoring. ACM SIGSOFT Software Engineering Notes, 25(1): 101, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:KD

[TR00g] Richard Taylor and David Redmiles. Knowledge depot. ACM SIGSOFT Software Engineering Notes, 25 (1):104, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:2000:W

[TR00h] Richard Taylor and David Redmiles. WebDAV. ACM SIGSOFT Software Engineering Notes, 25(1):108, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2002:RTL

[Tra02]  Will Tracz.  Roundtable

on technical leadership: a SHAPE forum dialog. ACM SIGSOFT Software Engineering Notes, 27(5):106– 107, September 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2003:FML

[Tra03] Will Tracz. Front matter (letters and notices). ACM SIG-SOFT Software Engineering Notes, 28(3):1–8, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2004:FML

[Tra04a]  Will Tracz.  Front mat-

ter (letters and notices). ACM SIGSOFT Software Engineering Notes, 29(2): 0–6, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2004:FLN

[Tra04b] Will Tracz. Frontmatter (letters and notices). ACM SIG-SOFT Software Engineering Notes, 29(3):0–4, May 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2004:FTL

[Tra04c]  Will Tracz.  Frontmatter

(TOC, letters, Frank David Anger, ACM Fellow Profile, ICSE 2005, praise and criticism, some thoughts on software reliability, latest DoD effort to achieve quality in software, software engineering education). ACM SIGSOFT Software Engineering Notes, 29(5): 0, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2004:RAE

[Tra04d] Will Tracz. Review of Architecting enterprise solutions: patterns for high-capability Internet based systems by Paul Dyson and Andy Longshaw. John Wiley and Sons, Ltd. 2004. ACM SIGSOFT Software Engineering Notes, 29(5):41, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2004:RIP

[Tra04e]  Will Tracz.  Review of

IT project estimation: a practical guide to costing of software by Paul Coombs. Cambridge University Press 2003. ACM SIGSOFT Software Engineering Notes, 29 (5):41, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2004:RMS

[Tra04f] Will Tracz. Review of Managing software for growth: without fear, control, and the manufacturing mindset by Roy Miller. Addison-Wesley 2003. ACM SIGSOFT Software Engineering Notes, 29 (5):40–41, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2005:FTL

[Tra05]  Will Tracz.  Frontmatter

(TOC, letters, preaching to the choir!, interviewers needed, time to vote!, new resource on SW engineering, calendar of events, workshop and conference information). ACM SIGSOFT Software Engineering Notes, 30(3):0, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2006:RPI

[Tra06a] Will Tracz. Review of Process Improvement in Practice — A Handbook for IT Companies by Tore Dybä, Torgeir Dingsøyr, and Nils Brede Moe, Kluwer Academic Publishers, 2004, ISBN 1-4020-7869-2. ACM SIGSOFT Software Engineering Notes, 31 (2):35, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2006:RWW

[Tra06b]  Will Tracz.  Review of

Weinberg on Writing: The Fieldstone Method by Gerald M. Weinberg, Dorsett House, 2006. ACM SIGSOFT Software Engineering Notes, 31(4):40, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2008:RGR

[Tra08] Will Tracz. Review of Geekonomics: The Real Cost of Insecure Software by David Rice, Addison-Wesley, 2008, 362 pp., 0-321-47789-8 (hardback), $29.99. ACM SIG-SOFT Software Engineering Notes, 33(1):9:1–9:??, January 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2000:NSE

[TS00] Will Tracz and Wayne Stidolph. NSF software engineering and language program summaries. ACM SIGSOFT Software Engineering Notes, 25(1):30– 32, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tillmann:2005:PUTb

[TS05a] Nikolai Tillmann and Wolfram Schulte. Parameterized unit tests. ACM SIGSOFT Software Engineering Notes, 30(5):253– 262, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tillmann:2005:PUTa

[TS05b] Nikolai Tillmann and Wolfram Schulte. Parameterized unit tests with Unit Meister. ACM SIGSOFT Software Engineering Notes, 30 (5):241–244, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tevanlinna:2004:PFT

[TTK04] Antti Tevanlinna, Juha Taina, and Raine Kauppinen. Product family testing: a survey. ACM SIGSOFT Software Engineering Notes, 29 (2):12, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tahara:2001:SEM

[TTOH01] Yasuyuki Tahara, Nobukazu Toshiba, Akihiko Ohsuga, and Shinichi Honiden. Secure and efficient mobile agent application reuse using patterns. ACM SIGSOFT Software Engineering Notes, 26 (3):78–85, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tavares:2008:GIO

[TV08] Andre L. C. Tavares and Marco Tulio Valente. A gentle introduction to OSGi. ACM SIGSOFT Software Engineering Notes, 33(5): 8:1–8:??, September 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Traas:2000:SCM

[TvH00] Vincent Traas and Jos van Hillegersberg. The software component market on the Internet: current status and conditions for growth. ACM SIGSOFT Software Engineering Notes, 25(1): 114, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tsui:2009:DST

[TWQ09] Frank Tsui, Andy Wang, and Kai Qian. A discussion on security typing and measurement for SOA. ACM SIG-SOFT Software Engineering Notes, 34(3):1–5, May 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Uchitel:2004:MPB

[UC04] Sebastian Uchitel and Marsha Chechik. Merging partial behavioural models. ACM SIGSOFT Software Engineering Notes, 29(6):43– 52, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Uchitel:2004:SAC

[UCKM04] Sebastian Uchitel, Robert Chatley, Jeff Kramer, and Jeff Magee. System architecture: the context for scenario-based model synthesis. ACM SIGSOFT Software Engineering Notes, 29(6):33– 42, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Uzuncaova:2006:PSD

[UK06] Engin Uzuncaova and Sarfraz Khurshid. Program slicing for declarative models. ACM SIGSOFT Software Engineering Notes, 31 (6):1–2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Uchitel:2001:DIS

[UKM01] Sebastian Uchitel, Jeff Kramer, and Jeff Magee. Detecting implied scenarios in message sequence chart specifications. ACM SIGSOFT Software Engineering Notes, 26 (5):74–82, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Uchitel:2002:NSI

[UKM02] Sebastian Uchitel, Jeff Kramer, and Jeff Magee. Negative scenarios for implied scenario elicitation. ACM SIGSOFT Software Engineering Notes, 27(6):109– 118, November 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Umarji:2005:PAS

[US05] Medha Umarji and Carolyn Seaman. Predicting acceptance of software process improvement. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ubayashi:2005:CMC

[UT05] Naoyasu Ubayashi and Tetsuo Tamai. Concern management for constructing model compilers. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vahaniitty:2005:TFC

[Väh05] Jarno Vähäniitty. A tentative framework for connecting long-term business and product planning with iterative and incremental software product development. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanHentenryck:2000:CP

[Van00a] Pascal Van Hentenryck. Constraint programming. ACM SIGSOFT Software Engineering Notes, 25(1):89– 90, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanderZanden:2000:NVD

[Van00b] Bradley T. Vander Zanden.

New visualization and debugging technology for oneway dataflow constraints. ACM SIGSOFT Software Engineering Notes, 25(1):90– 91, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Varma:2000:AST

[Var00] Tathagat Varma. Automated software testing: introduction, management and performance. ACM SIGSOFT Software Engineering Notes, 25(3):65, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Veanes:2005:OTM

[VCST05] Margus Veanes, Colin Campbell, Wolfram Schulte, and Nikolai Tillmann. Online testing with model programs. ACM SIGSOFT Software Engineering Notes, 30(5):

273–282, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanDeursen:2001:CIE

[vD01] Arie van Deursen. Customer involvement in extreme programming: XP2001 workshop report. ACM SIGSOFT Software Engineering Notes, 26(6):70–73, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanderHoek:2001:IWS

[vdH01] André van der Hoek. International Workshop on Software Configuration Management (SCM-10): new practices, new challenges, and new boundaries. ACM SIGSOFT Software Engineering Notes, 26(6):57–58, November 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanPut:2006:LNC

[VDM^{+}06] Ludo Van Put, Bjorn De Sutter, Matias Madou, Bruno De Bus, Dominique Chanet, Kristof Smits, and Koen De Bosschere. LANCET: a nifty code editing tool. ACM SIGSOFT Software Engineering Notes, 31(1):75– 81, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vestal:2000:D

[Ves00a] Steve Vestal. DoME. ACM SIGSOFT Software Engineering Notes, 25(1):101, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vestal:2000:ICE

[Ves00b] Steve Vestal. Incremental constraint engine. ACM SIGSOFT Software Engineering Notes, 25(1):102– 103, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vestal:2000:M

[Ves00c] Steve Vestal. MetaH. ACM SIGSOFT Software Engineering Notes, 25(1):105, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanHilst:2005:RMS

[VGL05] Michael VanHilst, Pankaj K.

Garg, and Christopher Lo. Repository mining and six sigma for process improvement. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Viega:2005:BSR

[Vie05] John Viega. Building security requirements with CLASP. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vipindeep:2005:ESA

[VJ05] V. Vipindeep and Pankaj Jalote. Efficient static analysis with path pruning using coverage data. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vinita:2008:REO

[VJT08] Vinita, Amita Jain, and Devendra K. Tayal. On reverse engineering an objectoriented code into UML class diagrams incorporating extensible mechanisms. ACM SIGSOFT Software Engineering Notes, 33(5):9:1– 9:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vogt:2002:ACR

[Vog02a]  Christian Vogt.  Admis-

sion control and resource reservation on the internet. ACM SIGSOFT Software Engineering Notes, 27 (3):80–87, May 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vogt:2002:IEC

[Vog02b] Christian Vogt. Intractable ERP: a comprehensive analysis of failed enterpriseresource-planning projects. ACM SIGSOFT Software Engineering Notes, 27(2): 62–68, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Verhanneman:2005:RTS

[VPDJ05] Tine Verhanneman, Frank Piessens, Bart De Win, and Wouter Joosen. Requirements traceability to support evolution of access control. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Visser:2004:TIG

[VPK04] Willem Visser, Corina S.

Păsăreanu, and Sarfraz Khurshid. Test input generation with Java PathFinder. ACM SIGSOFT Software Engineering Notes, 29(4): 97–107, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Varnell-Sarjeant:2008:MMR

[VS08] Julia F. Varnell-Sarjeant.

Managing a man-rated software development program via risk mitigation. ACM SIGSOFT Software Engineering Notes, 33(4):8:1– 8:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vaysburg:2002:DAR

[VTK02] Boris Vaysburg, Luay H. Tahat, and Bogdan Korel. Dependence analysis in reduction of requirement based test suites. ACM SIGSOFT Software Engineering Notes, 27(4):107–111, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vetterling:2002:SSD

[VWW02] Monika Vetterling, Guido Wimmel, and Alexander Wisspeintner. Secure systems development based on the common criteria: the PalME project. ACM SIG-SOFT Software Engineering Notes, 27(6):129–138, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Woodward:2000:TFS

[WAK00] Martin R. Woodward and Zuhoor A. Al-Khanjari. Testability, fault size and the domain-to-range ratio: an eternal triangle. ACM SIGSOFT Software Engineering Notes, 25(5):168– 172, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wand:2000:ABP

[Wan00] Mitchell Wand. Analysisbased program transformations. ACM SIGSOFT Software Engineering Notes, 25 (1):91–92, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2005:TRM

[Wan05] Qianxiang Wang. Towards a rule model for self-adaptive software. ACM SIGSOFT Software Engineering Notes, 30(1):8, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wellington:2005:ETC

[WBG05] Carol A. Wellington, Thomas Briggs, and C. Dudley Girard. Examining team cohesion as an effect of software engineering methodology. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wu:2005:PTT

[WBGM05] Xiaoqing Wu, Barrett R.

Bryant, Jeff Gray, and Marjan Mernik. Pattern transformation for twodimensional separation of concerns. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weide:2005:IPC

[WBH^{+}05] Bruce W. Weide, Paolo Bucci, Wayne D. Heym, Murali Sitaraman, and Giorgio Rizzoni. Issues in performance certification for highlevel automotive control software. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wiederhold:2000:CMS

[WBSM00] Gio Wiederhold, Dorothea Beringer, Neal Sample, and Laurence Melloul. Composition of multi-site software (CHAIMS). ACM SIGSOFT Software Engineering Notes, 25(1):99–100, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wand:2000:OMU

[WC00] Mitchell Wand and William D. Clinger. Optimizing memory usage in higher-order programming languages: theoretical and experimental studies. ACM SIGSOFT Software Engineering Notes, 25(1):92, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weaver:2005:CSE

[WDKM05] Rob Weaver, Georgios Despotou, Tim Kelly, and John Mc-Dermid. Combining software evidence: arguments and assurance. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wendorff:2001:FAA

[Wen01] Peter Wendorff. A formal approach to the assessment and improvement of terminological models used in information systems engineering. ACM SIGSOFT Software Engineering Notes, 26 (5):83–87, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whittle:2006:IWS

[WGM06] Jon Whittle, Leif Geiger, and Michael Meisinger. 5th Intl. Workshop on Scenarios and State Machines: Models — Algorithms — and Tools (SCESM). ACM SIGSOFT Software Engineering Notes, 31(5):39–40, September 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2007:MEO

[WGS07] Yi Wang, Defeng Guo, and Huihui Shi. Measuring the evolution of open source software systems with their communities. ACM SIG-SOFT Software Engineering Notes, 32(6):7:1–7:??, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:2005:RSS

[WH05] Chadd C. Williams and Jeffrey K. Hollingsworth. Recovering system specific rules from software repositories. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

White:2000:GSD

[Whi00] Elizabeth L. White. General strategies for dynamic reconfiguration. ACM SIGSOFT Software Engineering Notes, 25(1):93, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wallnau:2001:HDT

[WHS01] Kurt Wallnau, Scott Hissam, and Robert C. Seacord. Half day tutorial in methods of component-based software engineering essential concepts and classroom experience. ACM SIGSOFT Software Engineering Notes, 26 (5):314–315, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wu:2008:ABM

[WHW08] Jiankun Wu, Linpeng Huang, and Dejun Wang. ASMbased model of dynamic service update in OSGi. ACM SIGSOFT Software Engineering Notes, 33(2):8:1– 8:??, March 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:2004:NPM

[Wil04] Laurie Williams. On the need for a process for making reliable quality comparisons with industrial data. ACM SIGSOFT Software Engineering Notes, 29(5):1– 4, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2005:PSG

[Win05a] Michael Wing. In praise of SE geeks. ACM SIGSOFT Software Engineering Notes, 30(4):8–9, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2005:OJC

[Win05b] Michael Wing. It’s our job to care for SE. ACM SIGSOFT Software Engineering Notes, 30(2):6–8, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2005:TGE

[Win05c] Michael Wing. Toward a great and enduring profession. ACM SIGSOFT Software Engineering Notes, 30 (3):9–11, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2005:TNA

[Win05d] Michael Wing. Towards a National Academy of Software Engineering. ACM SIGSOFT Software Engineering Notes, 30(1):6–7, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2006:HN

[Win06a]  Michael Wing.  Be here

now. ACM SIGSOFT Software Engineering Notes, 31 (3):7, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2006:BAN

[Win06b] Michael Wing. Burt add numbers: a tale of software. ACM SIGSOFT Software Engineering Notes, 31 (6):10, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2006:GP

[Win06c] Michael Wing. Growing up is painful. ACM SIGSOFT Software Engineering Notes, 31(4):9, July 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2006:LHB

[Win06d] Michael Wing. Lonesome hacker blues. ACM SIG-SOFT Software Engineering Notes, 31(2):5, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2006:R

[Win06e]  Michael Wing.  Reminis-

cences. ACM SIGSOFT Software Engineering Notes, 31 (1):4, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2006:RC

[Win06f] Michael Wing. Roller coaster.

ACM SIGSOFT Software Engineering Notes, 31(5): 4, September 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2007:CK

[Win07a]  Michael Wing.  Cocoa

krispies. ACM SIGSOFT Software Engineering Notes, 32(2):7, March 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2007:HL

[Win07b] Michael Wing. Hacking with love. ACM SIGSOFT Software Engineering Notes, 32 (1):4–5, January 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2007:ES

[Win07c] Michael Wing. On the economics of software. ACM SIGSOFT Software Engineering Notes, 32(6):8–9, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2007:ST

[Win07d] Michael Wing. Software traumatica. ACM SIGSOFT Software Engineering Notes, 32(3):8–9, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2007:S

[Win07e]  Michael Wing.  Super-

hack. ACM SIGSOFT Software Engineering Notes, 32 (4):9, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2008:HSN

[Win08a] Michael Wing. Hell’s software nightmares. ACM SIGSOFT Software Engineering Notes, 33(3):6, May 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2008:HE

[Win08b] Michael Wing. The human element. ACM SIGSOFT Software Engineering Notes, 33(1):1:1–1:??, January 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2009:EGA

[Win09a] Michael Wing. Evolution and great apps. ACM SIGSOFT Software Engineering Notes, 34(5):6–7, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2009:Q

[Win09b] Michael Wing. On quality. ACM SIGSOFT Software Engineering Notes, 34 (4):7–8, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2009:TCC

[Win09c] Michael Wing. On technology and cultural change. ACM SIGSOFT Software Engineering Notes, 34(2): 4, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wittenberg:2000:PTC

[Wit00] Craig H. Wittenberg. Progress in testing component-based software (abstract only). ACM SIGSOFT Software Engineering Notes, 25(5): 178, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wyss:2004:REF

[WJH^{+}04] C. M. Wyss, A. James, W. Hasselbring, S. Conrad, and Hagen Höpfner. Report on the Engineering Federated Information Systems 2003 workshop (EFIS 2003). ACM SIGSOFT Software Engineering Notes, 29 (2):1–3, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wong:2001:XIF

[WJS^{+}01] Tak Wong, Stan Jarzabek, Soe Myat Swe, Ru Shen, and Hongyu Zhang. XML implementation of frame processor. ACM SIGSOFT Software Engineering Notes, 26 (3):164–172, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wu:2005:FMS

[WK05] Weihang Wu and Tim Kelly.

Failure modelling in software architecture design for safety. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber:2005:SFT

[WKP05] Sam Weber, Paul A. Karger, and Amit Paradkar. A software flaw taxonomy: aiming tools at security. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:2005:EEC

[WLA05] Laurie Williams, Lucas Layman, and Pekka Abrahamsson. On establishing the essential components of a technology-dependent framework: a strawman framework for industrial case studybased research. ACM SIG-SOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2009:WAR

[WLCP09] Xiaofeng Wang, Michael Lane, Kieran Conboy, and Minna Pikkarainen. Where agile research goes: starting from a 7-year retrospective (report on agile research workshop at XP2009). ACM SIGSOFT Software Engineering Notes, 34(5):28– 30, September 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wermelinger:2001:GBA

[WLF01] Michel Wermelinger, Antónia Lopes, and José Luiz Fiadeiro. A graph based architectural (re)configuration language. ACM SIGSOFT Software Engineering Notes, 26(5):21–32, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walenstein:2004:SIW

[WLK04] Andrew Walenstein, Arun Lakhotia, and Rainer Koschke. The Second International Workshop on Detection of Software Clones: workshop report. ACM SIGSOFT Software Engineering Notes, 29 (2):1–5, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walker:2000:ICE

[WM00] Robert J. Walker and Gail C. Murphy. Implicit context: easing software evolution and reuse. ACM SIGSOFT Software Engineering Notes, 25 (6):69–78, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whaley:2002:AEO

[WML02] John Whaley, Michael C.

Martin, and Monica S. Lam. Automatic extraction of object-oriented component interfaces. ACM SIGSOFT Software Engineering Notes, 27(4):218–228, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wolf:2004:SER

[Wol04] Alexander L. Wolf. Is security engineering really just good software engineering? ACM SIGSOFT Software Engineering Notes, 29(6): 1, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Workman:2002:CMT

[Wor02]  David A. Workman.  A

class and method taxonomy for object-oriented programs. ACM SIGSOFT Software Engineering Notes, 27 (2):53–58, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wagner:2005:SQE

[WS05] Stefan Wagner and Tilman Seifert. Software quality economics for defect-detection techniques using failure prediction. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weyns:2005:ADD

[WSH05] Danny Weyns, Kurt Schelfthout, and Tom Holvoet. Architectural design of a distributed application with autonomic quality requirements. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wei:2008:CBR

[WT08] Wang Wei and Li Tong. Component behavior relativity analysis. ACM SIGSOFT Software Engineering Notes, 33(2):7:1–7:??, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walker:2004:IPD

[WV04] Robert J. Walker and Kevin Viggers. Implementing protocols via declarative event patterns. ACM SIGSOFT Software Engineering Notes, 29 (6):159–169, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2000:OOS

[WW00] Debora Weber-Wulff. Objectoriented software engineering: conquering complex and changing systems. ACM SIGSOFT Software Engineering Notes, 25(3):65– 66, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2002:HCI

[WW02] Debora Weber-Wulff. Humancomputer interaction in the new millennium. ACM SIG-SOFT Software Engineering Notes, 27(3):94–95, May 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2004:RRC

[WW04a] Debora Weber-Wulff. Review of Requirements by collaboration — workshops for defining needs by Ellen Gottesdiener Addison–Wesley 2002. ACM SIGSOFT Software Engineering Notes, 29(5): 40, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2004:RUE

[WW04b] Debora Weber-Wulff. Review of UML by example by Ghinwa Jalloul. Cambridge University Press 2004. ACM SIGSOFT Software Engineering Notes, 29(5): 40, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2005:RCE

[WW05a] Debora Weber-Wulff. Review of Contributing to eclipse: Principles, Patterns, and Plug-Ins by Erich Gamma and Kent Beck, Addison-Wesley, 2004, ISBN 0-321-20575-8. ACM SIGSOFT Software Engineering Notes, 30(3):40–41, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2005:ROE

[WW05b] Debora Weber-Wulff. Review of Official eclipse 3.0 FAQs by John Arthorne and Chris Laffra; Addison-Wesley, 2004, ISBN 0-321-26838-5. ACM SIGSOFT Software Engineering Notes, 30(3):40, May 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber-Wulff:2006:RFG

[WW06] Debora Weber-Wulff. Review of Fascicle 4: Generating All Trees — History of Combinatorial Generation by Donald E. Knuth, Addison-Wesley, 2006, ISBN 0-321-33570-8. ACM SIGSOFT Software Engineering Notes, 31(6):60, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wermelinger:2009:RIE

[WWB09] Michel Wermelinger, Paul Wernick, and Ciarán Bryce. Report on the 4th International ERCIM Workshop on Software Evolution and Evolvability (Evol ’08). ACM SIGSOFT Software Engineering Notes, 34 (2):25–27, March 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2006:ERT

[WXL06] Yuan Wang, Feng Xu, and Jian Lü. Establishing recommendation trust relationships for internetwares. ACM SIGSOFT Software Engineering Notes, 31(1): 1–5, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wu:2004:MZS

[WY04] Fangjun Wu and Tong Yi.

Measuring z specifications. ACM SIGSOFT Software Engineering Notes, 29(5):1– 5, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2006:SDG

[WZX06] Zhong-Jie Wang, De-Chen Zhan, and Xiao-Fei Xu. STCIM: a dynamic granularity oriented and stability based component identification method. ACM SIG-SOFT Software Engineering Notes, 31(3):1–14, May 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xie:2005:CPS

[XA05] Yichen Xie and Alex Aiken. Context- and path-sensitive memory leak detection. ACM SIGSOFT Software Engineering Notes, 30(5):115– 125, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2005:IDR

[XC05] Chang Xu and S. C. Cheung. Inconsistency detection and resolution for context-aware middleware support. ACM SIGSOFT Software Engineering Notes, 30(5):336– 345, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2004:EBC

[XDS04] Wei Xu, Daniel C. DuVarney, and R. Sekar. An efficient and backwards-compatible transformation to ensure memory safety of c programs. ACM SIGSOFT Software Engineering Notes, 29(6):117– 126, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xie:2002:URF

[XE02] Yichen Xie and Dawson Engler. Using redundancies to find errors. ACM SIGSOFT Software Engineering Notes, 27(6):51–60, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xiaochun:2002:AIN

[XGY02] Xu Xiaochun, Xu Guanghui, and Xu Yongsen. Architectural issues in networkcentric computing. ACM SIGSOFT Software Engineering Notes, 27(1):53– 57, January 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xia:2008:PSA

[XMP08] Feng Xia, Longhua Ma, and Zhe Peng. Programming Scilab in ARM Linux. ACM SIGSOFT Software Engineering Notes, 33(5):10:1– 10:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xavier:2009:PIS

[XN09] P. Eugene Xavier and E. R.

Naganathan. Productivity improvement in software projects using 2-dimensional probabilistic software stability model (PSSM). ACM SIGSOFT Software Engineering Notes, 34(5):1–3, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2006:PMP

[XOWM06] Jing Xu, Alexandre Oufimtsev, Murray Woodside, and Liam Murphy. Performance modeling and prediction of enterprise JavaBeans with layered queuing network templates. ACM SIGSOFT Software Engineering Notes, 31(2):5:1–5:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2005:BSP

[XQZ^{+}05] Baowen Xu, Ju Qian, Xiaofang Zhang, Zhongqiang Wu, and Lin Chen. A brief survey of program slicing. ACM SIGSOFT Software Engineering Notes, 30 (2):1–36, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2005:TWA

[XXJ05] Lei Xu, Baowen Xu, and Jixiang Jiang. Testing web applications focusing on their specialties. ACM SIGSOFT Software Engineering Notes, 30(1):10, January 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2004:BMC

[XYH04] Guoqing Xu, Zongyuan Yang, and Haitao Huang. A basic model for components implementation of software architecture. ACM SIGSOFT Software Engineering Notes, 29(5):1–11, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:2005:APN

[XZRA05] Lihua Xu, Hadar Ziv, Debra Richardson, and Thomas A. Alspaugh. An architectural pattern for non-functional dependability requirements. ACM SIGSOFT Software Engineering Notes, 30(4): 1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yilmaz:2004:CAE

[YCP04] Cemal Yilmaz, Myra B.

Cohen, and Adam Porter. Covering arrays for efficient fault characterization in complex configuration spaces. ACM SIGSOFT Software Engineering Notes, 29(4): 45–54, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:2005:PCX

[YEMT05] S. Michelle Young, Helen M.

Edwards, Sharon McDonald, and J. Barrie Thompson.

Personality characteristics in an XP team: a repertory grid study. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yuan:2000:OCH

[YF00] Xiaodong Yuan and Ying Feng. The optimization of class hierarchy in objectoriented design. ACM SIG-SOFT Software Engineering Notes, 25(2):56–59, March 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ye:2000:IAI

[YFR00] Yunwen Ye, Gerhard Fischer, and Brent Reeves. Integrating active information delivery and reuse repository systems. ACM SIGSOFT Software Engineering Notes, 25 (6):60–68, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yadav:2009:MDC

[YK09] A. Yadav and R. A. Khan. Measuring design complexity: an inherited method perspective. ACM SIGSOFT Software Engineering Notes, 34 (4):1–5, July 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yavuz-Kahveci:2002:SVS

[YKB02] Tuba Yavuz-Kahveci and Tevfik Bultan. Specification, verification, and synthesis of concurrency control components. ACM SIGSOFT Software Engineering Notes, 27 (4):169–179, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yamamoto:2000:AVF

[YMI00] Tetsuo Yamamoto, Makoto Matsushita, and Katsuro Inoue. Accumulative versioning file system moraine and its application to metrics environment MAME. ACM SIGSOFT Software Engineering Notes, 25(6):80– 87, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yu:2007:EER

[YML07] Ping Yu, Xiaoxing Ma, and Jian Lu. Expectation, experiment and reflection on Internet software evolution. ACM SIGSOFT Software Engineering Notes, 32 (3):1–5, May 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:2000:HAT

[You00] Michal Young. High assurance technologies. ACM SIGSOFT Software Engineering Notes, 25(1):106, January 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yu:2008:IMR

[YR08] Liguo Yu and Srini Ramaswamy. Improving modularity by refactoring code clones: a feasibility study on Linux. ACM SIGSOFT Software Engineering Notes, 33(2):9:1–9:??, March 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yi:2004:EAE

[YW04] Tong Yi and Fangjun Wu. Empirical analysis of entropy distance metric for UML class diagrams. ACM SIGSOFT Software Engineering Notes, 29(5):1–6, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ying:2005:EHC

[YWA05a] Annie T. T. Ying, James L. Wright, and Steven Abrams. An exploration of how comments are used for marking related code fragments. ACM SIGSOFT Software Engineering Notes, 30(4): 1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ying:2005:SCT

[YWA05b] Annie T. T. Ying, James L.

Wright, and Steven Abrams. Source code that talks: an exploration of Eclipse task comments and their implication to repository mining. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yi:2004:CMU

[YWG04] Tong Yi, Fangjun Wu, and Chengzhi Gan. A comparison of metrics for UML class diagrams. ACM SIGSOFT Software Engineering Notes, 29(5):1–6, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yang:2009:DLF

[YXP09] Qun Yang, Man-Wu Xu, and De-Chang Pi. Delegation: a language facility for dynamic software adaptation. ACM SIGSOFT Software Engineering Notes, 34 (3):1–5, May 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yang:2006:MAA

[YYX06] Qun Yang, Xianchun Yang, and Manwu Xu. A mobile agent approach to dynamic architecture-based software adaptation. ACM SIGSOFT Software Engineering Notes, 31(3):1–7, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zand:2001:CPP

[ZBPD01] Manzour Zand, Paul Bassett, and Rubén Prieto-Díaz. Closing panel (panel session): where are we standing? can we say “reuse is dead, long live reuse” or is it too soon? ACM SIGSOFT Software Engineering Notes, 26 (3):173–175, May 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2005:SAS

[ZC05] Ji Zhang and Betty H. C. Cheng. Specifying adaptation semantics. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhou:2005:OSS

[ZD05] Ying Zhou and Joseph Davis. Open source software reliability model: an empirical approach. ACM SIGSOFT Software Engineering Notes, 30(4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhao:2000:SQR

[ZE00] Luyin Zhao and Sebastian Elbaum. A survey on quality related activities in open source. ACM SIGSOFT Software Engineering Notes, 25 (3):54–57, May 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zeil:2000:RGM

[Zei00] Steven J. Zeil. Reliability growth modeling from fault failure rates. ACM SIGSOFT Software Engineering Notes, 25(1):94, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zeller:2002:ICE

[Zel02]  Andreas Zeller.  Isolat-

ing cause-effect chains from computer programs. ACM SIGSOFT Software Engineering Notes, 27(6):1–10, November 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2005:MEH

[ZG05] Xiangyu Zhang and Rajiv Gupta. Matching execution histories of program versions. ACM SIGSOFT Software Engineering Notes, 30 (5):197–206, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2005:MDA

[ZGL05] Jing Zhang, Jeff Gray, and Yuehua Lin. A modeldriven approach to enforce crosscutting assertion checking. ACM SIGSOFT Software Engineering Notes, 30 (4):1–5, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zuser:2005:SQD

[ZHG05] Wolfgang Zuser, Stefan Heil, and Thomas Grechenig. Software quality development and assurance in RUP, MSF and XP: a comparative study. ACM SIGSOFT Software Engineering Notes, 30 (4):1–6, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zirkind:2006:HDI

[Zir06] Givon Zirkind. How data inspection and consideration, provides for index compression and record access optimization of genetic databases. ACM SIGSOFT Software Engineering Notes, 31(1):1–4, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zirkind:2007:ADC

[Zir07] Givon Zirkind. AFIS data compression: an example of how domain specific compression algorithms can produce very high compression ratios. ACM SIGSOFT Software Engineering Notes, 32 (6):8:1–8:??, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zitser:2004:TSA

[ZLL04] Misha Zitser, Richard Lippmann, and Tim Leek. Testing static analysis tools using exploitable buffer overflows from open source code. ACM SIGSOFT Software Engineering Notes, 29(6):97– 106, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhou:2004:CSG

[ZLX04] Yuming Zhou, Jiangtao Lu, and Hongmin Lu Baowen Xu. A comparative study of graph theory-based class cohesion measures. ACM SIGSOFT Software Engineering Notes, 29(2):13, March 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zannier:2005:QEE

[ZM05] Carmen Zannier and Frank Maurer. A qualitative empirical evaluation of design decisions. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zacharias:2006:PAD

[ZMC06] Manousaridis Zacharias, Sagheb-Tehrani Mehdi, and Mamaloukas Christos. Practical aspects of DSS design for commodities transportation during special events. ACM SIGSOFT Software Engineering Notes, 31(1): 1–6, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zheng:2005:PIC

[ZRWS05] Jiang Zheng, Brian Robinson, Laurie Williams, and Karen Smiley. A process for identifying changes when source code is not available. ACM SIGSOFT Software Engineering Notes, 30 (4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zayaraz:2005:SQA

[ZTSR05] G. Zayaraz, P. Thambidurai, Madhu Srinivasan, and Paul Rodrigues. Software quality assurance through COS-MIC FFP. ACM SIGSOFT Software Engineering Notes, 30(5):1–5, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zundorf:2001:UCC

[Zün01] Albert Zündorf. From use cases to code — rigorous software development with UML. ACM SIGSOFT Software Engineering Notes, 26 (5):324–325, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2001:VUC

[ZXZ01] Lu Zhang, Dan Xie, and Wei Zou. Viewing use cases as active objects. ACM SIGSOFT Software Engineering Notes, 26(2):44–48, March 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2004:RCD

[ZYL04] Xiaofang Zhang, Michal Young, and John H. E. F. Lasseter. Refining codedesign mapping with flow analysis. ACM SIGSOFT Software Engineering Notes, 29(6):231–240, November 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).